A Complete Bibliography of Publications in Technology and Culture: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 August 2024 Version 1.01

Title word cross-reference

#FoodPeople [CG16].

e [But12, Gil10, Ora15, Pan18b].

/To [Kra11].

10-14 [ERST13]. 150th [Mül16a]. 1600 [Ben14]. 1700 [Bra15a]. 1750 [San11]. 1810 [Rab13a]. 1820 [Cas13, Cro18, TS12]. 1830 [Hea12, Rey12]. 1840 [Str11]. 1840s [LeC11, Wan10]. 1850 [Hah19, Hep16, Jaf11]. 1860s [Wan10]. 1862 [Gam11]. 1865 [Pre10b]. 1868 [Cut13]. 1870 [Lak13, Pan18a]. 1870s [Hir17]. 1880 [Ung15]. 1880s [Hir10]. 1880s-1930s [Hir10]. 1884 [Rhe14]. 1890 [Man15, Cas10]. 1898 [Koe18]. 18th [Ros10]. 18th-Century [Ros10]. 1900 [War14]. 1903 [Joh16]. 1904 [Gil11a]. 1910 [Cah14, Coc17b, Gro15, For10]. 1914 [Bhi13, Fre15, Gol12]. 1915 [Bal11a].

1917 [Sch14c]. 1918 [Ket16, Tuc16, Wol14]. 1919 [Amm11, Tuc12]. 1920 [Arc13b, Jør13, Ken19]. 1923 [Sla14]. 1925 [Eis10]. 1929 [Wri12a]. 1930 [Ham16a, Wil17a]. 1930s [Hir10, Lig11, Mir14, Nye11]. 1932 [Hul16]. 1933 [Vol10]. 1935 [Swa18]. 1937 [Swa13b, Yan19]. 1939 [Dic10]. 1940 [Lif18, Rie16]. 1940s [Bar14b]. 1941 [Kai10]. 1942 [Gan17b]. 1943 [Gro19]. 1945 [Bro14b, Hay18, Hea13a, Jør17, Moo16b, MS16, Ens11, Rol10a]. 1947 [Sat10b]. 1948 [Rob10a]. 1950 [Fru17, Pan10, The10a]. 1950s [Kri19b, Par16]. 1960 [Gol14b, Rom10, Sch12e]. 1960s [Abb18, Boy16, Par16, Ste16a]. 1963 [Moo16a]. 1964 [Moo15, Pat15, Joh12]. 1967 [Bur13, Hic12]. 1968 [Gai15]. 1970 [Jon15, Tob13, Mor12a, Sjö11, Str10]. 1970s [Boy16, RS14, Wet15, Gor17, Hir17]. 1973 [Cas16a]. 1980 [Gai19]. 1980s [Ste16a]. 1989 [Jen11]. 1990s [Pre19, LeC11]. 1992 [Wes15]. 19th [Hur15, Lac16, SH15]. 19th- [Lac16]. 19th-Century [Hur15].

2000 [Bar16a, Jac16, Wei13, Dio11, Wil10, Wu11]. 2003 [New18]. 2004 [McC14a]. 2005 [Die12]. 2006 [Nie12]. 2008 [Mey15]. 2009 [Ano10e, OS12, Pra13]. 2010 [Ano11c, Mis11b, Wil11]. 2010s [Abb18]. 2011 [Ber12, HHH12, Soc12]. 2012 [Ano13b, ERST13, Seg14, Col14]. 2013 [And14, Ano14b, KLŁ14]. 2014 [Ano15b, BBŁT15, Sta14b]. 2015 [Ano16a, KNSZ16, Gan19]. 2016 [Ano17b, BGW^{+}17]. 2017 [Ano18b, GKMV18]. 2018 [Ano19c, BBB^{+}19]. 20th [Lac16]. 20th-Century [Lac16]. 21st [Fer16, For11].

4-7 [Ano13b]. 41 [Ste14].

500-year [DL18]. 50th [Wor10].

60 [Ros13].

9 [Tay14].

âge [Cas19, Pan18a, Lev17].

Aaron [Bro14b, Ket16, Wil17a]. Abbate [Hay14]. Abbott [Pan18d]. Abby [Sto13a]. Abigail [Boo17]. Above [Jac14, Ste19]. Abyss [Her11]. Academic [Ken14, Lin11, Tei11]. Accelerated [Gan17a, Luc14]. Acceleration [Jør17]. Accelerator [Wes10b]. Accepting [Moo11]. Access [Bor10, Lig11]. Accessible [Alp16]. Accident [Wel17, Luc10]. Accidents [Gan13, Gan17a, Gan19, Tay14]. Across [vR12, Lan18b, McC13, Vee10, vdV14]. Acting [Jas12]. Action [De 17]. Activism [Boy16, Pre19]. Actor [Gro18]. Actor-Network [Gro18]. Actors [Hog14b]. Adam [Bon16, Coc18, Sch13a]. Adams [Fro14, SH15]. Adapting [Gai15]. Adaptive [Bau11]. Addiction [Lee14]. Adelman [Hri16]. Adinkra [Got12]. Administrative [FR18]. Adopted [Wel15]. Adoption [Gif15b]. Adopts [Gam11]. Adrian [Sla15, Kra11]. Adrienne [Aid19, Hen19, Wes16]. Adrift [Hog13a]. Advertising [Pop13a]. Aerial [Day11, Hay18, Sim17b, Ste19]. Aermotor [Cou17b]. Aeromotive [Bug13]. Aeronautical [Rol18]. Aeronautics [Mey15]. Aerospace [Kin14, Mey17]. Aerostatic [Kim13]. Aesthetic [Vai13]. Aesthetics [Dow10, Hui16, Low14]. Affect [Wea16]. Affective [Hal12]. Africa [Bhi13, Edw18, Rob10a, Sol11, LL15, Mav10]. African [Mai16, Rob10a, San17, Zyl11a, Ded18, Mac16a]. Africans [Abr13]. After [Bro12b, RV18, Bra15a, Buh17, Kru16, Mur12, Oka17b, Oka18, Ona18, Rud12, Ven19, Vin15, Wad16, Wil17c, Ham18a, Hog14b, Shi17]. Aftermath [Med18]. Aftermaths [Ste19]. Agadir [Wil17c]. Again [Ber16]. Against [Ihd16]. Agamben [Str13]. Age [Bat11, Bre18, Col19a, Coo15, Dow17, Dus11, Edw12a, Est10, Hac15, Hah19, Her11, Hic12, Kin10, Las18, Low10, Luc18a, Mey12, Nee17, Nof18, Olw14, Pop19, Pop11, Pre10b, Rob10c, Smi11b, Str10, Ves19, Wal18, Whi13a, Wor10, Wos17, Acl17, Cos17, Fin17, Hey10, Rab13b, SH17, Wil14c, Zia17, Bow14, De 19, De 16, Dob16, Fer13, Joh15b, Ret16, Sla19, Bes18, Bij12, Luc10, Mou10, Vin12]. Agency [Rob13, FR18]. Agent [Ham14a]. Ages [Mor10b]. Ago [Ber14]. Agrarian [Bea19b]. Agribusiness [Ham14b, Kle10]. Agriculture [Fin10, Sah13b, Wil17b, Ham18c]. Agro [vdHP13]. Agro-Industrialism [vdHP13]. Ahuja [Dec18]. Aidoo [Pur17]. AIDS [Hog18b]. Air [Bor14, Cor14, Gai19, Her16a, Kou10, Rot16, Sim18, Gai13]. Air-to-Air [Gai19]. Aircraft [Sim18]. Airline [vR12]. Airplane [Rol18]. Airport [McD17]. Airpower [Mey18, Mey12]. Airs [Bar12]. Airship [Cro14a]. Airspace [Sim18]. Akarnan [Muk19]. Akhil [Oak19]. al [Hea15]. al. [Fer13, Gan17b, Gud18, Lau17, Pla17, Rhe14, Sah13a, Sch13c]. Alan [Deu17, McC17, See15a, Swa13a]. Albert [Joh11a, Sch14c]. Alberts [Paj16]. Albuquerque [Ano16a]. Alchemy [Sea11, Ram10, Ram11]. Alcohol [Coo11a]. Aldrich [Gan19]. Alex [Mir19]. Alexander [DB13, Hen18a, Ihd16, Jos13, Las18, Lau13, Mir14, Mil13b]. Alexandra [Cah14]. Alferov [Sid12]. Alford [Vol17]. Alfred [Pet16]. Alger [Aga16]. Algorithms [Nof18]. Alic [Gai14]. Alive [LL15]. alla [Cam16]. Allan [Hau15]. Alleghenies [Chu15]. Allen [Car13a]. Allied [Boc16]. Allow [Pre19]. Allure [Kim19]. Alois [Gif17a]. Alone [Kit12]. Along [Mei13]. Alphabet [Mul18a]. Alpi [Mor18]. alpino [Mor18]. Alps [Mor18]. Altar [For11]. Alternative [Abb18, Vol17]. Altshuler [Hea15]. Alvin [Joh18b]. Always [Kat14]. Amateur [Tuc16]. Amazing [Bus19b]. Ambassador [Fre12]. Ambient [Dow17]. Ambition [Lau11, Wri12a]. Ambivalence [Wal18]. America [Arc13b, Bac15, Bar10, Bow15, Cas10, De 16, Gai18, Gam11, Gua13, Hay18, Hea11, Hri16, Hus10, Jac16, Kle10, Lai12, Las18, Lau11, Luc18b, Mei13, Nye11, Rol10b, Str11, Wil10, Wri12a, vR12, Aik19, Bar14b, Cro16, Edw12b, Gai13, Mas17, McK13b, Mer12, Pen18, Pla16, Tay14, App14, App17, Ber18, Bra13, Bro15, Car13c, Kas17, Kru14, Lan15, Luc13, MF17, McC17, Mer17b, Olw14, Roh19, Tel13, Whi13b, Wil14a, Bro11, Col12c, Dru12, Fei12, Hoo10, Lov11, Mis11a, Osl10, Pos15, Rol16, Sca11, Sch10c, Sha10, Wak11, Wel10, Wro10]. American [Amm11, Amm14, And10, Ang17, Arc12, Bar16b, Ben17, Bow15, Bow16, Bre10b, Bye19, Car14a, Cla11, Coc17b, Cor14, Cor12a, Cry11, Cut11, Cut13, Fin11, Fre18, Gan10, Gan19, Ger13, Gor11, Gor12, Gru16b, Gud18, Ham16a, Ham18c, Han13, Her17a, Hig12, Hin10, Hor14, Kno16, Kru12, Lan11, Lie18, Man18, Mey19, Mil13b, Moh12, Nee18, Nem14, Ore13, Ran11, Ree18, Rol10a, Rol10b, Rom10, Ros11, Rub18, San17, Sch13b, Sch15a, Sch14b, Sch13d, Sch12e, Seg14, Shi17, Sut12, Swa13b, Tar15, Van14, Vet19, Wil13, Zyl11a, Acl17, BH17, Día16, Goo19a, Kri19b, Lin10, Mor14a, Pat19, Pre13, Rud12, Wil13, Fai16, Gam15, Gor17, Her11, Ken19, Mei10, Ste14, Tel11]. Americanist [KL19]. Americanization [Jør10]. Americans [Cas16a, SH15]. Americas [Joh16]. amerikanischen [Lin13]. amid [Nan12, Pre10a]. Amiga [Nov14]. Amit [DiM15]. Ammon [Van18]. Amplification [Wij14]. Amy [McC14a, Rae16, Sch15a, Ver15]. Anahid [Kat14]. Analog [Ves19]. Analysis [Hag15b, Hir10, Wol11b, God16a]. Analyze [ZT18]. Analyzing [Hog13a, Muk12, Low14]. Anand [Oak19]. Anatomical [Mae13]. Anatomizing [Mae13]. Anatomy [Ham16b]. Anchor [OL17]. Ancient [Hen18a, Hen18b, Hen19, KO14, Nye15, Sea11]. Andean [Clo11]. Andersen [Bhi13]. Andes [Hos13]. Andrea [Bou19]. Andreas [Anb18, DB13, Qui17]. Andrew [Hle19, Ihd18, Man15, ZT16]. Andrews [Smi13]. Angeles [Est10, Hir12]. Angelina [Neu14a]. Angus [Bjö13, Blo13]. Animal [Deu17, Kir12]. Animals [SH17, Sop15]. Anique [Moh15]. Ankerson [Cor19]. Ann [Eva16, Mod17, Sop15]. Anna [Mar17a]. Annales [Gui16]. Anne [Cut13, Kra17, Lin13, Low14, Pop19, Yan19]. Anniversary [Mül16a, Wor10]. Antebellum [Mul16b, Reg18]. Anthony [Bow16, Mer17b, Bye19]. Anthropocene [Jør17, JJ16]. Anthropological [Mil19, Ros12b]. Anthropology [KL19, Ros12b]. Anti [LeC17]. Anti-Landscape [LeC17]. Antiaircraft [Sla12a]. Antibiotic [Li16, SH17]. Antibiotics [Ven19]. Antibodies [Tim17]. Anticipatory [MV15]. Antikythera [Hen18a]. Antiquity [KO13, Mor10b]. Antivirus [Ens10]. Antonio [Mor18, Ber17]. Antwerp [De 10]. Apartment [Pat15]. Apollo [Shi17, Wei15, Col19a]. Apparatus [Cum19]. Appeals [Jac15]. Appel [Oak19]. Apple [Cro14b]. Appletopia [Fis15]. Appliances [Par16, Wil14a]. Applied [DK18, Tka19, Fri16]. Applying [OF18]. Appraisals [Lig11]. Apprenticeship [Jaf10]. Approach [Par18, ZT18]. April [Rhe14, Sta14b, Ano10h, Ano11f, Ano12f, Ano13e, Ano14e, Ano15e, Ano16e, Ano17e, Ano18e, Ano19f, Ano19g]. Apt [Smi10c]. Aquarius [Col19a]. Aqueducts [KO12]. Arabia [Cit12]. Arabs [Cas16a]. Araldo [Bal10]. Arapostathis [Joh15a]. Arbuckle [MM11]. Arcanum [Kle14]. Archaeological [Kui19]. Archaeology [Amm11, Eva16]. Archer [Sum17]. architecte [Lan10]. Architectonica [Eds10]. Architects [Hor11]. Architectural [Hen18b]. Architecture [Aid19, Arc13a, Bar18, For17, Gar18, Gru16a, Gru16b, Kar17, McL17, Mei13, Moh12, Mor18, San11]. Archiv [Kur19b]. Archival [Tka19]. Archive [BT19, Buh17, Riz10, She14a, SW15, Wal10, Sza15]. Archives [Boo17, KL19, dlP10]. Archiving [Bir19, Fea19, KL19]. Arctic [McN14, Hög15]. Argentina [Mur12, Cas16b]. Argonne [Wes15]. Arguments [Col14]. Ariel [Jac15]. Arithmetic [Ros17]. Armin [Gla17]. Arming [Jen15]. Armory [Sim10]. Armoured [Dou16]. Arms [Mac18, Rol10d, Sla12a, TAF^{+}18, Zim10]. Armstrong [Cas14]. Army [Dou16, Joh18a, Kni18]. Arne [Gif17b, Sam13, Smi14b]. Arnold [Cer16, Cha14, Mat17b]. Arnoldi [Sta17]. ARPANET [RS14]. Arrangement [Kar18]. Arresting [Deu17]. Arrive [LL15]. Arsenault [Jin19]. Arsenic [Blu12]. Art [Arc18, Bai18, Bra17, Bro17a, Cla13, Dob16, Fig10, Gal15, Gla17, Gop15, Hey10, Hin11, Low14, Low16b, Mar18, Pac18, Rat12, Riz10, Rob10c, San12, See15a, Seg14, Tho10a, Wos17, Kan16, Pop19, Tay17]. Arthur [Mar17a]. Artifacts [Ran14]. Artificial [Ham16b, Sch11b]. Artillerie [Dou16]. Artillery [Ans17]. Artisan [Ben14]. Artisan/Practitioners [Ben14]. Artisanal [Haf18]. Artisans [Ham18d, Ora15, Gue13]. Arts [ERST13, Smi11b, Pan18b, Sta17, Pan12]. Arun [Set16]. Arwen [Kno16]. Asbestos [Mai10]. Ascendency [Cor14]. Ascending [Kim13]. Ashley [Car16]. Ashok [Neu14a]. Ashworth [Hor18]. Asia [Bro14b, vdS19, Leu14, McC13, McD17, Set16]. Asian [Gol15, Kui19]. Asif [Smi13]. Aspects [Eds10]. Assa [Sin15]. Assault [Cry11]. Assembly [Pos15]. Assessing [De 10]. Astounding [Tel13]. Astride [Ang17]. Astroculture [Lau13]. Astrolabe [Eds10]. Astronaut [McC14a, Ger13]. Astronauts [Kil15]. Astronomers [McC14b]. Astronomical [CR19]. Astronomy [Rus11a, Kra14]. Asymmetrical [McD15]. Atari [Jin17]. Ate [Mun19]. atelier [But12]. Athens [Lak13]. Atkinson [Cer15]. Atlanta [Har16]. Atlantic [Col13, Gor10, Sol11, Mül16a, Ros17, Kai12, Rie16]. Atlas [Gai13]. Atmosphere [Ber18, Gil10]. Atmospheric [Mun14, Wea18, Kim19, Kru10]. Atomic [Bre10b, Car14a, DiM10, McK13b, Oka17a, BF19, Olw10, Olw14, Wel12, Zia17, Car13a, Kru12, Ram14]. Atoms [DiM10]. Atomtheoretiker [Kai10]. Attack [Ald16a, Jos19]. Attacks [Yos16]. Attempt [Dav17]. Attendant [Tya12]. Attention [Dow17]. Attwood [De 12]. Audio [Rus12]. Audiovisual [Moh16]. Auditory [Lac16]. Audra [Olw14]. Aufbruch [Ole17b]. August [BBŁT15, Ber12, KNSZ16, Wil11]. Aurality [Bor14, Bra12]. Aurora [Arc13a]. Austin [Hle19]. Austria [BF19]. Authentication [Hus19]. Authoritarianism [Cam19]. Authority [Gil15, Hir14, Maa16, Ram10, Ram11]. Authors [Cum19]. Auto [De 16]. Autobahn [Sjö11]. Automata [Kan13]. Automating [BM14, Gai19]. Automation [Hic12]. Automaton [Dup15, Rei12]. Automatons [Lie18]. Automobile [Gan10, Luc14, Mom14, Mor14a, Vin15, Wet15, Zim11, Kre14a, Har12].

Automobiles [Bor14]. Automobilism [Rie16]. Automobility [De 16, Wel10, Zar13, Ros12a]. Automotive [Mor17, Wet15]. Autopia [Est10]. aux [Har13, Lev11]. AV [Acl17]. Avert [Rub18]. Aviation [Bye13, Cla11, Cor14, Hoo10, McC17, Vol10, Bye19, Uzi19]. Aviv [KNSZ16]. Awards [Ano10a, Ano11a, Ano12a, Ano13a, Ano14a, Ano15a, Ano16b, Ano17a, Ano18a, Ano19a]. Ax [Sah13a]. Axel [Mil13b]. Axis [Moh12]. Aya [Ona18]. Aylen [Wel15]. Ayurveda [Muk19].

B [Aga16, Bye13, Cas16b, Gal17, Hor17, Joh15b, Rhe14, Smi14b, Wil17a]. Ba [Pur17]. Babbage [Nof15]. Babel [Mor14c]. Babies [Bjö13]. Bache [Mil13b]. Back [Ber16, Cow10, O’D12, Pos10, Hal16, App17, Gan19]. Background [Ber17]. Backwater [San17]. Backyard [BH17]. Bad [Jos19]. Badash [Olw10]. Badenoch [DB13]. Baekeland [Mer12]. Baggott [Eva15]. Bailey [Wis17]. Bakelite [Mer12]. Baker [Ole17a]. Bala [Set16]. Balbi [Col17b, Cam16]. Ballet [Lae15]. Ballistic [Gai13]. Balloon [Sul19]. Ballooning [TS12]. Ballots [Pfa13]. Baltic [Gor10]. BAM [Smi10a]. Bamboo [Wu11]. Banchoff [Col13]. Banff [Cas14]. Bang [Mor14c]. Banham [Ber11]. Banks [Bra15a]. Barbara [McC15, Mol15, Ole17a, Pfa13, Swa13b]. Barber [Kar17]. Barcelona [ERST13]. Bards [Jin19]. Barefoot [Hai15]. Barnstorming [Sim17b]. Baroque [KO12, HV15]. Barricades [Lev13b]. Barry [Mat17c]. Barton [Ang17]. Base [Lev13a]. Based [Pat19, CC18, See15b, Wea18]. Basic [DK18]. Basistexte [Die11]. Bastard [Cor13a]. Bates [Kar18]. Bathroom [Mol15]. Baton [Moo11, Sta10b]. Battery [Eis17, Pet19]. Battle [Dus18, Hög13b, Sat13b]. Battles [Mul16b]. Bau [Pop13b]. Bay [Fel14]. Bayonet [Sto12]. BBC [Jon15, Bru19]. be [Kra11]. Beatty [Vit17]. Beaudouin [Sch19]. Became [Car18, Cas14, Ken14, NPA14, Sch12b]. Beckert [Coc16]. Become [Cer10]. Becoming [Col15, Seg10, Wis11]. Beer [Sch10b]. Beetle [Lad14]. Before [Bro12b, Gar18, Lif18, Bal10, De 10, Hir14, SDJ19, Tom15, Ven19, Wor16b, SH19a]. Began [Nof15]. Beginner [The10b]. Beginnings [Gai13, Ona14]. Behaviour [Sin12]. behind [Bra15b, Ham18b]. Being [Boy16, Ore12, Abr13, Ore12]. Belgian [De 11a, DH11, Web15]. Bell [Tay17]. Belong [Pat19]. Belonging [Sha17]. Below [Car15]. Belville [Ste10b]. Benesch [Tar15]. Bengal [Hea13b, Muk19]. Benjamin [Hir16, Hor17, Nik17, Wol15, vdH17]. Benny [Amm12]. Berenice [Pan18d]. Berghoff [Coc18]. Berkeley [Mun19]. Berlage [Swa18]. Berlin [Sop14]. Berliner [Kur19b]. Berman [Ken14]. Bernaldo [Lev17]. Bernard [Bow14, Moh16, Sla11b]. Berner [Åbe15]. Bernhard [Lad14]. Berridge [Sta14a]. Bessner [Kri19a]. Best [Fei12, Reg18]. Beth [Car13c, Dup15]. Better [All11, Gat18, God18]. Between [DH11, Haf11, Hog14b, Mar18, Mil11, GKMV18, Pre10a, Whi11b, Ihd11, Ver12]. Betz [Sim17a]. Beverage [Sam13]. Beyond [DK18, Fin10, Lan18a, Lev13a, PR16, Bat11, Lev13b, Mor14c, Car16, God16a, Hri16]. Bicycle [MF17, Epp15]. Bicycles [Lai12, MF11, Mor15, Vol17]. Bicycling [Ros12a]. Big [Car16, Hag15b, Kum19a, Mod11, Mor14c, Ram14, Riz10, Ste16b]. Bigelow [Reg18]. Bijker [Moh15]. Bijsterveld [Bra15b, Mil13a, Sca14]. Biles [Jac16]. Bill [Gor15, Ziz19]. Biodefense [Fai16]. Biography [Fre18]. Bioinformatics [Hag15a]. Bioinsecurities [Dec18]. Biological [Cur13, Fai16, Nor15a]. Biology [Rud12, Scr12, Ste15, Wad16, Wes11]. Biomass [Mit18]. Biomedical [Sta10a, Sch13a]. Biomedicine [Moo16b, Hey10]. Biometrics [Col12b]. Biopolitics [Hog14b, Str13]. Biosensing [Col17a]. Biotech [Fea16, Ona14]. Biotechnology [Low16a, Cha18a]. Birdwomen [Cla11]. Birth [Jen15, KO12, OL17, Pre19, Sat10b]. bis [Kai10, Pop13b]. Bitches [ZT18]. Bitcoin [Zim17]. BITS [Sha17]. Bitter [Eva18]. Bix [Sch15a]. Black [Aid19, Eis17, Mer17b, Pra19, Lae16]. Blackouts [Mis11a]. Blainey [Dow18]. Blake [McA14]. Blandford [Oka18]. Blaszczyk [Kal14]. Blimp [SS13]. Blind [Bra10]. Bloc [Łot16, De 12, Zar13]. Blocking [Jon10]. Blood [Car11, Jos13, Mer17b]. Bloom [Kru13a]. Blowout [Col12c]. Blue [Ded18, Jen11, Kin14]. Blues [San17]. Blyth [Fer13]. Bob [Bar16b]. Bobbye [Bar14a]. Bodies [Ihd13, Moo16b, Ore12]. Bodily [Jac11b]. Body [Gue13, Lae16, Rab13a]. Boeing [Her16a]. Boel [Åbe15]. Bogaert [Cam19]. Bogdanov [Jos13]. Bold [BS13]. Bolting [Ald16a]. Bomb [Bre10b]. Bombay [Pat15]. Bombing [Mey19, Dor10]. Bombs [Mey16, Zim10]. Bonaccorso [Ans17]. Bonnell [Ran15]. Bonneuil [Har13]. Bonvillian [Gal17]. Book [Bor10, Dot12, Low10, Sin15]. Books [Amm12, How15a]. Boom [Cut11, Mor14a, Whi10]. Bootleg [Git10]. Borden [Har16]. Border [Mar19]. Border-crossing [Mar19]. Borderlands [Gro14, Les19, vdS19, Wu19]. Borders [vdV14]. Borgman [Ste16b]. Bosch [Gor11]. Boschi [Kat14]. Boston [Fei12, Tom14, Abb11]. Botanical [Sol11]. Bottom [For11]. Bouk [Car18]. Boulton [Eva15]. Boundaries [Dre10, Hoc10, Wes10b, Set16]. Boundary [Buc12b]. Bourgfontaine [BM12]. Bow [Fer11b]. Box [Eis17]. Boxed [Pra19]. Boyd [Man18]. Boyer [Dow15]. Boys [Cer11]. Bozak [Jac13]. BP [Col12c]. Brain [Ona18, Hui16]. Brainwashing [Lie18]. Brakes [Gan11]. Brasov [BBŁT15]. Braun [Big16]. Bray [Kim14]. Brazil [GKMV18, Cru19]. Brazos [Sum17]. Breakthrough [OF18]. Breast [App17]. Breastfeeding [App17]. Breathing [Big16]. Breeder [Cam18, Fja15, Mad19]. Breeding [Fin13, Cha12]. Brenna [Sch13c]. bresciani [Bal11a]. Breton [Bai12]. Brett [Fis15]. Brew [ZT18]. Brewing [Ung15]. Brezhnev [Smi10a]. Brian [Bus19a, Gar18, Hir14, Hoo19, Hus19, Jør13, Mun19, Wil17b]. Bridge [Bro12b, Bro14a]. Bridger [Sla16]. Bridges [Dre10]. Brief [Hor14, Nye15]. Bright [Swa13b]. Bringing [Ler10]. Brita [Sch13c]. Britain [Blu12, Car13b, DiM15, For10, For17, Fre15, Hep16, Hic12, Jaf11, Joh12, Luc18a, Rol17, Ros13, Sla14, Tin19, Coo11c, Edw12b, EW12, Gif15b, Jon13, Kap19, Bjö13, Joh15a, Nah13, Sch11c]. britisch [Lin13]. British [Arc12, Bar12, Han14, Jon15, Ket16, Sla10, Ste10a, Zim10, Bru19, Hir10, Sat10a, Sim18, Bhi13, Coo10b, Dic10, Dob16]. Britton [Joh16].

Broadcasting [Bal10, Bir19, Bru19, Oka17a, Vai13, Wol10, Arc12, Jon15]. Brock [Cer16, Hu14]. Broich [Tom17]. Broken [Mck13a, Pfa13]. Brookline [Abb15]. Brøvig [Kra17]. Brown [vdH17, Aid19, Car14a, Gua13]. Bruce [Ole17a]. Bruno [Fis13]. Brunton [Coo14]. Bryan [Mac17]. Buchanan [Blo13]. Buckley [Cru19]. Bucky [Seg10]. Build [Dav17]. Building [DeB11, De 16, DiM10, Die10, Dre10, Dus10, Ert12, Gat18, Jos16a, Lan18b, LeC11, Les10, Mil13b, Mor18, Sch14c, Smi10a, Tom11, Yan19, Moo16b, Gol14b]. Buildings [Pet13]. Built [Bro17b, Fir12]. Bulbous [Fer11b]. Bulldozer [Amm12, Van18]. Bulliet [Nye17]. Bunker [Zia17]. Bunn [Gol14a]. Burdens [Pra19]. Bureau [Lag17, Ald16a]. Bureaucracy [Riz10]. Bureaucratic [Wel12]. Bureaus [Swa18]. Burke [Aga16, Car13a]. Burleson [Kru13a]. Busch [Kno14a]. Buse [Pul16]. Business [Coc18, Hei15a, Hir12, Hug15, Jep18, O’R17, Sin15, SV19, Yos16, Ree18, Hin10]. Buss [De 19]. Bust [Cut11]. Busting [Hur15]. Butte [Mun19]. Butter [Boy16]. Button [Bet14, Plo12]. Buying [Fra18]. Byers [Uzi19]. Byrd [Cro16].

C [Amm14, Ang17, Ber18, Blo13, Bow15, Car15, Cer16, Cou17b, Joh17, Joh13, Kum16, Lan15, Lau13, New18, Pet13, Ric13, Rob13, Sla14, Sti18, Str13, Tay17, Whi13b, Wu19, Zar17, Hea12, Muk19]. C. [Hok17]. c.1750 [Tay14]. ca [Mae13, Par18]. ca.1630 [Rat12]. Cable [Bri10, Mül16a]. Cables [Joh16]. Cabling [Mül16a]. Cain [Bed18]. Calamitous [Aik19]. Calculating [Der17]. California [Bar14a, Man15, Mar13, Kin14, Low13]. Call [Jos19]. Callahan [Neu14a]. Calling [Rus11b]. Camaros [Luc14]. Camels [Pat11]. Camera [Jac13, Pul16, Tuc16]. Cameralism [Vos10]. Camouflage [Mac13]. Campaign [Pre19]. Campaigning [Her17b]. Campbell [Cor16, Fro10, Str13]. Campbell-Kelly [Cor16]. Camprubí [And15]. Can [Gra16]. Canada [Boc16, The10a, AH18, Ham17, Ran15, Doe12]. Canal [Gra11, Sch10c, Car16]. Canals [Bie16, Chu15, Sau11]. Cancer [Neu14b]. Cancer-Causing [Neu14b]. Canoes [SP17]. Can’t [Nel16]. Canton [Kim12]. Capacity [See15b]. Capital [Arc13a, Sen17, Qui17]. Capitalism [Coc18, Fro10, Ole17a, Pop13a, Ros17, Str12, dV12, Hoc10]. Capitalist [Lim19]. Capitalists [Rob10b]. Capitals [Shu15]. Captivating [Ava15]. Car [Kre14b, Lad14, Luc18b, Mom14, Rie16, Ros12a, Vol11, Zar13, Nye13]. Cara [Edw17]. Carbide [Gif15b]. Carbon [Bar16b]. Card [Ens11]. Cardiac [Mck13a]. Cards [Aga16]. Care [Mck13a]. Career [Big16]. Caren [Ste19]. Carl [Tom14, Kur19b]. Carlson [Bow14, Moh16]. Carnage [Sim17a]. Carolina [Sau11]. Carolyn [App14]. Carpets [Reg18]. Carrera [Cos17]. Carroll [Wac17]. Cars [Fra13, Mor14a, Neu14b, Rus11b, SP17, Fra18]. Carse [Car16]. Cartels [vM16]. Carter [Joh17]. Cartography [Ham18a, Lig11, Mor10b]. Casale [Whi13a]. Case [CC18, De 10, Ens10, Had10, Lae15, Les19, Plo12, Whi12]. Casey [Hjo16]. Casid [Ban16]. Casper [Bhi13]. Cassandra [Olw10].

Castaing [Bat14]. Castaing-Taylor [Bat14]. Catalog [Jon13, Var18]. Catalogue [Fre15]. Catalyst [Smi11a]. Catastrophic [Jen15]. Categories [God16a, Ler10]. Catherine [Dob16, Mul16b]. Causing [Neu14b]. Cavallo [Gue13]. Cedeño [Col19b]. Cell [Sin15, Tim18]. Cells [Hög14c, Nor15a]. Center [Kru15]. Central [Ric13, Sch10a, Shu19, Wil17a, vdS19]. Centric [Wol11a]. Centrifuge [Kem12]. Centrism [Pra19]. Centuries [Wro10, Łot15]. Century [App14, Ber18, Ber16, Ber14, Blu12, Bra13, Bro11, Cru19, Eds10, Fer19, Fra18, Gal15, Gor12, Hin10, Hir12, Hou10, Hur15, Kas17, Kin14, Kre14c, Lac16, Lov11, MF17, McC15, Mor18, Neu11, Ole17a, Pan15, Per12, Ram14, Ros10, Rus11a, Smi17, Whi13b, WH11, Ans17, Ber13, CR19, Der17, EW12, FP16, HPR13, Jep18, Mac19, Mas17, Nor15b, OL17, Pre10a, Rob17, Van16, Wer12, Wol11b, Zan13, Bed18, Coc18, Fer16, Gam15, Hag15b, Ham18a, Hea13b, Lau13, Mil13b, SH15, Die10, For11, Fro10, Gor10, Ran11, Sla12b, Sta10a]. Ceramic [Bhi10, Mat17b]. CESBIO [CC18]. Chad [Cas16a]. Chairman [Hu14]. Chakanetsa [Edw18, Hoa15]. Challenge [Hin10, Sim10, Swa15, Hir16, Wol15]. Challenger [Pom10]. Challenges [BSA16, Dis15, Kai11, Sch11d]. Chang [McL17]. Change [Bhi10, Clo11, God16a, Gul19, Oat16, Oka17b, Pre10a, Pre10b, Reu14, Vin15, Wes11, Wro10, Sim10]. Changed [Gor12, Kel10, Kem12, Moh16, Pul16, Ryd11]. Changes [Jac11b, Kur19a, Mul12b, Sin15, Dio12]. Changing [Fre11, Hay14, Lee13, Mey19, Sut19]. Channeling [Tuc12]. Chanute [Bug13]. Chapel [Sta14b]. Chapman [Luc12]. Chapoutier [Sch19]. Characteristically [Gru16b]. Charis [Tim18]. Charles [Abb11, Gal17, Wes15, GR10, Leg11, Rob13, Phi14]. Charterhouse [BM12]. Chasing [Gop15]. Chaussées [Gui16]. Cheap [Fro10, Sin15]. Cheat [Rat12]. Check [Gai19]. Checkered [Coo11b]. Checkpoint [Hri16]. Chemical [Ham18b, Hay18]. Chemicals [Ham16a]. Chemiker [Tod11]. chemin [Pan18b]. Chemistry [Kle16, Kni13, Smi11a]. Chen [Mun14, HV15]. Chen-Morris [HV15]. Chen-Pang [Mun14]. Cheng [Tel13]. Chernobyl [Kru16]. Chesapeake [Coo11a]. Chicago [Cut10, Hou10, SH19b, Mac15, Tom14]. Chikako [Hog14b]. Chilcotin [Wal10]. Child [Hep16]. Childhood [Rud18]. Children [Amm12, Olw13]. Chile [Med18]. chimie [Lev11]. chimistes [Lev11]. China [Kim14, Sch14a, Ste14, Whi11b, Wu19, Yan19, Bra19, Mul18a, Wan10, Hai15, Ham18d, Hu14, Ree16, Son17, Bra10, Per12]. Chinese [Bus19b, Wan13, BM19, Bus19b, Mul12a, Mul18a]. Choice [Ber12, Mac18]. Choices [Wis17]. Choreography [Tim18]. Chotkowski [Sch13d]. Christensen [Sla11b]. Christer [Tob13]. Christina [Sah13a]. Christine [Ste16b]. Christo [Bix18]. Christoph [Nye14]. Christophe [Har13]. Christopher [Bro15, Cas14, Cou17b, Jos19, Lag17, Maa16, Mar18, Nye13, Pri19, Spu13, Whi13b]. Chromatic [Jac12b]. Chrysler [War10]. Chunjuan [Hu14]. Church [Gru16a]. Churella [Sch14c]. CIA [Day11, Her16b].

Cifuentes [PT13]. Cinema [Bro17a, Cas16b, Ham16b, Var11, Jac12b, Sch12c]. Cinematic [Gar18, Jac13, Jac15]. circulation [Lev17]. Cities [Car14a, Jac16, Ore13, Rom10, Shu19, Wro10, Hig12]. Citizen [Ona18]. Citizenship [Goo19a, Kim15]. City [Bea19a, Bro12b, Dou19, Har16, Lae15, Les10, Mac16b, Mun19, Pik15, Tar15, Tom14, Tom15, Arc13a, Bea19a, Mul16b, Nem14, Tom17, KO12, Tom14]. Civil [Ang15, Ang17, Col16, Edw17, Iri12, Kni18, Mar11b]. Civilization [Bla18, Her11]. Civilizations [Sch14a]. Claiming [MF17]. Claims [Sid10]. Claire [Fre15, Wal18]. Clapperton [Edw18, Hoa15]. Clarified [Ber19]. Clark [O’R19]. Clash [Sch14b]. Class [Cam19, Cas16b, Kel11, Pat19]. Classes [Hir10]. Classification [Pet16]. Classroom [Acl17]. Classrooms [Rud12]. Clay [Jac19]. Clean [Jon10, Swa10]. Clearance [Amm12, Van18]. Cleveland [Nie12, Hor11, Soc12]. Cleveland-Cliffs [Nie12]. Clifford [Bro13]. Cliffs [Nie12]. Climate [Bar12, Clo11, Coo11b, Eli12, Oat16, Pat11, Wea18, Cah19, Whi11a]. Clinch [Cam18]. Clive [Ada17]. Clocks [Fra13]. Close [Lig14]. Closer [Sla12b]. Cloth [Got12]. Clothing [Gro19]. Club [Smi11a]. Clunkers [Luc14]. Co [Sop14]. Co-Production [Sop14]. Coal [Ald16a, Bar10, Cas10, Sov16, Ziz19, Sov16, Fer19]. Coast [SDJ19]. Coastal [Miz11]. Coats [Sch12c]. Cobre [Spu13]. Coca [vdHP13]. Cocoa [Mac16a]. Code [Abb18, Jac19]. Codebreaker [Swa13a]. Codes [Col12a, Pan12]. Coding [Nem14]. Coen [Cah19, Jen13]. Cognition [ZT18]. Cohen [Eri15]. Cohen-Cole [Eri15]. Cohn [Fre18]. Colbey [Rae16]. Cold [Ala12, Car13a, DK18, Day11, Eng12, Eri15, Fir12, Fja12, Iri12, Kar17, Kli19, Kru14, Lag17, Las18, Rub18, Sla12a, Sla16, Wea16, vdV12, Cam15, DiM10, Gro19, Mat19, Rot16, Tin18, God15, Jør10, Kat11]. Cole [Eri15]. Coleman [Arn18]. Colette [Kre14c]. Colin [Aga16, Ber16, Kni13, MF16, Luc12]. Collaboration [Neu14a, Zia18, Sis12]. Collapse [Bro12b]. Collection [Low16b]. Collections [Mar18, Pan18b]. Collider [Hag14, Wes16]. Colliery [Hau11]. Colligan [Kre14c]. Collins [Jos19, Low14, Hag15b, Sla19]. Colonial [Coo11a, Gua13, Hel13, Hos13, Ket16, Las18, Lee13, Mav10, McL17, Pra19, Sah13a, Sah14, vdHP13]. Colonialism [Yan19]. Colonialists [Vai19]. Colonies [Sah13a, Tho15]. Color [Kal14, Kin11, Pop19]. Column [Hen18b]. Com [Cor19]. Comamala [PT13]. Combat [Gai19, Gro19]. Comfort [Mom14, Zar17]. Coming [Sch15a]. Command [Fel14]. Commemoration [Pre13]. Comment [Cou17a]. Commerce [Pop19]. Commercial [Cor19, Mei13, Van14]. Commercialization [Bon16, SK16]. Commission [Wel12]. Committee [BBŁT15, BBB^{+}19, Ber12, BGW^{+}17, ERST13, GKMV18, KLŁ14, KNSZ16, Wil11]. Commodification [Tei11]. Commodities [Ber19, Gud18]. Commodity [Swa13b]. Commodore [Nov14]. Common [Lau11, Rat12]. Common-Weale [Rat12]. Commons [Cum11b, Dow17, vdV14]. Communicating [Hau19]. Communication [Aga14, Arc13a, Eis10, Hug15, Mar19, Pat19, Pop11, Tuc12, Wen17, Wes10b, Sat13b, Ste11]. Communications [Aga11, Ste14, Bri10]. Communism [Fre12]. Communist [Sid12]. Communities [CC18, Kai17, Mal17, Man19, MV15, Nye15]. Community [Joh13, Kra14, Kum19a, LeC11, Mal17, Spu13, Wu11, Mat17b, Abb15, Bhi10]. Compact [Dow10]. Company [Hoo15, Wil17a]. Comparative [Wij14, Dow16]. Compass [San11]. Compete [Gra16]. Competing [Sid10, Olw14]. Competition [Ber16, Der17, Rob10a]. Complex [Zim11]. Complexity [Car14b, Sov16]. Compound [Smi14a]. Computation [Lig11, McD10]. Computer [Abb18, Cer11, Cer15, Cor16, Eli12, Ens10, Hic12, Med18, NPA14, Nof15, Paj16, Sla11a, Tay17, Tin18, Wea18, ZT16, Mar13, De 17]. Computer-based [Wea18]. Computers [Alp16, Cer11, Ens12, Nov11, Tin18, Wes15, Fra13]. Computing [CK15, Col14, Cor13b, Eis17, Ens18, Geo15, Mar13, Mul18a, Mul18b, Sch13a, Hay14, Nof18, Tin19, Col12a]. Comunicaciones [Hea15]. Con [Rob10b]. Concave [Edw16]. Concept [God16b, Mar10, Nik15]. Conception [NPA14]. Concise [Mor12b]. Concrete [Lag17, Sla15]. Condensed [Fro10]. Condition [Hay12]. Conference [Wor10]. Conflicting [Web15]. Conflicts [Deu17]. Confounding [Smi11a]. Confrontation [Can17]. Confronting [Sch16, Bar12]. Congress [Cam18]. Congreve [Wer12]. Conlogue [Gor15, Ziz19]. Connecticut [Gor11]. Connecting [Hea13a]. Connectivity [Sim17a]. Connects [Wil14b]. Conners [Bow16]. Conquest [Hög15, Pop13a, Mor10a]. Conquistadors [Hoo10]. Consalvo [Jin17]. Consequences [Aik19, Man18]. Conservation [Bru19, Str11]. Conservatoire [Pan18b]. Considerations [DGS16]. Considering [Buh17]. Consolidation [Bru19]. consommation [Yat16]. Conspiracy [Pua17]. Constitution [Jep18]. Constraints [Kor18]. Constructing [Bor14, Leg11, Wu19, Bro14b]. Construction [Bal11c, Cho17, Ens12, Hau10, Hen18b, Jeo10, Kor18, Mom14, Gui16, Pet13]. Consulting [Cas10]. Consumer [Ber12, Boy16, Fra13, Mei10]. Consumerism [Low10]. Consumers [App14, Gag15]. Consumption [Wes10a, Edw12a]. Contact [Pop11]. Contagion [Deu17]. Contagions [Leu14]. Container [Sam13]. Contamination [Cas19, Ona18]. Contemporary [BGW^{+}17, Cho18, Sla11a, Wak11]. Contentious [Bou19, Smi11a, vdH17]. Contestable [Joh15a]. Contested [FP16]. Context [Hah17, Sca11, Sim10, Sla11b, Sla11b]. Contexts [KNSZ16, Jin17]. Continental [Sla12a]. Continents [Vee10]. continuità [Col17b]. Continuity [Pre10a]. Contraception [Pre19]. Control [Ber18, Fel14, GR10, Hog18a, Kat10, Pre19, Sat10b, Wen17, Whi11a, Deu17, Abn19, Coo11b, Jac10b, Jac11a, Low10, Swa10]. Controversies [Aik19]. Conventional [Sim17a]. Conversion [Hah15]. Conway [Mes16]. Cook [Ver15]. Cooking [Moh12]. Cookson [Hah19]. Cool [Ree19]. Cooperate [Rub18]. Cooperation [Mat19]. Coopersmith [Pop16]. Copeland [Joh15b]. Copenhagen [Ano13b, Hey10]. Copper [Kor18, LeC11, Spu13]. Copyright [Día16, Got12, Cum19, Git10]. Corbusier [Har12]. Córdoba [PT13]. Corinna [Åbe15]. Corn [Fra13]. Cornell [Pen18]. Cornering [Ree18]. Corporate [Kou10, dV12]. Corporation [Spu11, War10]. Corps [McC14a]. Corpses [De 10]. Correo [Hea15]. Corridors [Mei13]. Corruption [Ole17a]. Cortada [McC13]. Cosmography [Bui10]. Cosmonauts [Jos16b, Kil15]. Cosmopolitan [vdV14]. Cosmopolitical [Dis15]. Cosmos [Ger11, Hen18a, Smi13, Zak15]. Cost [Hos13]. costruzione [Mor18]. Cotton [Coc16, Dav13, Mar11b, Ole17a, Pat11]. Count [Col14, Pfa13]. Counterculture [Her17a]. Counterfactual [Bro14a]. Counterfeit [Sea11]. Counterfeiters [Rob10b]. Countries [God18]. Country [Ham10b, Kle10, Nye13, Ziz19, Lan18b]. Countryside [Fin11]. Course [Moh16]. Courtney [Ham18c]. Crabgrass [Whi13b]. Craft [Edw12a, Lan17, Lon10]. Crafting [Per12, Rab13b]. Craftspeople [Bar14a]. Craig [Ang15, Bye13]. Crash [Gan13, Var11]. Cravens [God15]. Create [Lub18]. Creating [Sis12, Swa13b, CC18, App14, Ken14]. Creation [Bek18, Ham10b, Hir14]. Creative [Cum12, ZT18]. Creators [Mer17a, Hjo16]. crédit [Yat16]. Cressy [Buc14]. Crim [Bus19a]. Crime [Gro18, Jac10b, Pua17]. Crimes [Med18]. Crises [Joh16]. Crisis [Eis17, Hea12, SH17, Wel12, Har16]. Crispin [De 17]. Critical [Ste11, Dug10a]. Cronon [Cut10]. Crooked [Lon15]. Crooning [Wij14]. Crops [Sah13b, Sto13a]. Crosbie [Fer19]. Cross [BM19, Fra16, Hah15, Luc18b]. Cross-Cultural [BM19, Hah15]. Crossing [Set16, Mar19]. Crossroads [MF11]. Crucible [Whi13b]. Cryptocurrency [Zim17]. Crystal [Kat10]. CT [Dus11]. CTRL [Mul18b]. Cuba [Hea15]. Cuisine [Coo15]. Cult [Fer12, Fis13]. Cultivating [Sah13a]. Cults [Mun16]. Cultural [Abn19, BM19, Bek18, Ber17, Boo17, Cha11b, Cor13a, Cum12, Fir17, Hac15, Hah15, Ham16b, Hu14, KNSZ16, Mai12, McD10, Mul18b, Nik15, Pur10, Rob10a, Sca14, Wea18, Wel10, Zel11, dS17, MF16]. Cultural-Historical [Zel11]. Culture [Amm12, BH17, Cor13a, Cro14b, ERST13, Fer12, Gol15, Gop15, Kai12, Kas17, Kim12, Low10, Mac19, Mac18, Mom14, OL17, Pik15, Pos10, Pra12, Rud18, Sla15, Ste16a, Swa10, Tho16, Ver12, Ver15, Ves19, Wei13, Arc19, Bar16b, Gal15, Kar18, Pop13a, Smi13, Vol17, Cas16b, Coo12a, Cor12a, Dug10b, Rus11a, Sch11a]. Cultures [Lac16, Leu14, Moh15, Mun16, Paj16, Sch11d, Zim13, Ver13, Wil14b, Suo12, Wan13, Wea18]. Cupers [Par17]. Curcio [Cas15]. Curing [Zan13]. Curtis [Dry19, Bro17a, Gro15]. Cushing [Abb15]. Cushman [Whi14]. Cutting [Gif15b, Joh18a, Dus11]. Cyber [Yos16]. Cybernetic [Mir12, Pet18, Pic14]. Cyclades [RS14]. Cycle [Han14, Kru13a, Swa15]. Cyrus [Joh13]. Czechoslovak [Fre12].

D [Bea19b, Cor16, Cro14b, Dér18, Hay18, Kra14, Mey16, O’R19, Oka18, Pop13a, Rae16, Rol10c]. D.C [Seg14]. D.C. [Rol10b, Wil13]. Dagmar [Wan13]. Daguerre [Cla13, Cla13]. Daily [Dus18, Sin15]. Dale [Acl17]. Dallas [Mil13b]. Dam [Sti18, Jac10a]. Dams [Amm14, Lag17]. Dan [Car18, Gar13]. Danbom [Hur15]. Dancing [Ruu11]. Dangerous [Fai16, Ham11, Odd15]. Dangers [Rub18]. Daniel [Bor17, Cor16, Her17b, Kar17, Kri19a, Lac16, Lae16]. Danielle [Low18]. Danielsen [Kra17]. dans [But12]. Danube [Gat18]. Daqing [Tsu12]. Darryl [Hu14]. Darwin [Jen15]. Daryl [Pan15]. Dasgupta [Nof15, Sha17]. Data [BT19, CC18, Eli12, Hag15a, Her17b, Jac19, SW15, Ste16b]. Data-Driven [Hag15a]. Database [Var18]. Davenport [See15a]. David [Arc13b, Buc14, Car13a, Cer16, Cha14, Hac15, Hay18, Her17a, Hur15, Ken19, Kra14, LeC17, Leu14, Mac16b, Mck13a, Nah13, O’R19, Pos15, Sim17a, Smi10b, Son17, SP17, Van14, Wei15, Her16b, Mis11b]. Davis [Man13, ZT18]. Dawn [Bye13, Col17a, Ham13, Hic12, Ros11]. Dawson [How15b]. Days [Bow10, Car18, Kru12]. DDT [Gor12]. Deadliest [Bar10]. Deal [Kle10, Nee17]. Dean [Mat17b]. Dearborn [Ano15b]. Dearen [Eva18]. Death [Gra15, Han14, Mor15, Sla11c, Dor10]. Debate [Cam18]. Debates [Gul19]. Debating [Fin17]. Deborah [Cah19, Jen13, Sch13b]. Decades [Cam16]. December [Mis11b, Rhe14, Mod17]. Decentered [Med18]. Deception [Bre10a]. Decimals [Aga16]. Decision [Wis11]. Declassified [Mey16]. Decline [Gam15, Sim17a]. Decoration [Edw12b]. Deep [Hun16, Sch16]. Defending [Mai10]. Defense [Col14, Gru12, Iri12, Kri19a, Sla12a, FR18]. dégâts [MS16]. DeGolyer [MM11]. Degradation [For11]. DeHaven [Gan13]. Del [Hea15, PT13, Spu13]. Delahousse [Sla11b]. Delegating [Wet15]. Delete [Cer15]. Delhi [Arn18]. Delia [Pur17]. Delivery [Lan11]. della [CG16]. Delta [Ert12]. Demise [Gor11, Kle10]. Democracies [Col13]. Democracy [God15, Her17b, Jas12, Kri19a]. Democratic [Geo15]. Demoscenes [Paj16]. Demotion [Van18]. Demystifying [Mul12b]. DeNardis [Mue15]. Denis [Sch19]. Denise [McA14]. Denmark [Mar19, Søn12]. Dennis [Rae16]. Department [Arc10]. dépasser [Has16]. Dependence [Lag13]. Depersonalizing [Kle14]. Depression [Mir14, Edw17]. depuis [Lev17, MS16]. Deregulation [Oka17b]. Derek [Kim19]. Describing [Tin15]. Desert [Cam18, Cas16a, Cit12, Pra13]. Design [Abb15, Bar14a, Bjö13, Bre18, Cam16, Cer15, Con15, Cor19, Cro14a, Epp15, Gan11, Lee14, Low13, Low18, Par16, Pen18, Ten14, Ten15, Tho10a, UHW16, Hah15, Mat17c, Kat11, Rae16]. Designers [Bar14a, Mal17, UHW16]. Designing [Ber11, De 11a, Hei15b, Iri12, Mae13, Vin15, Bre18, Hes11, Koe18, Mal17, Nye11, O’R19]. Designs [Bre16]. Desire [Fra16]. Desmond [Boy16]. Despina [Ham16b]. Destruction [Shu11]. Detachment [Wen17]. Detectives [Pua17]. Detector [Gol14a]. Détente [Rot16]. Determinism [Ham14b, Pla16]. Deterministic [Rol10d]. Detroit [Dou19, Sch10a, She14a]. deutsch [Lin13]. deutsch-britisch-amerikanischen [Lin13]. deutscher [Kre14a, Tod11]. Deutsches [JJ16, Mor12a]. Deutschland [Die12]. Devastating [New18]. Developed [Mul12a, Wel15]. Developer [Hjo16]. Developers [Mal17]. Development [Abb15, Ala12, Ans17, Bal11b, Bes18, Cho18, Cro14a, Cru19, Gan13, God16b, Hah15, Lan11, Mat17a, McK13b, Mun14, Oka17b, Reu14, Sch12d, Sut12, Tya13, Kim19, Maa16].

Developments [SK16]. Devil [Woo19]. Devin [Gar13]. devint [Cou17a]. Dewalt [Mac17]. Dewey [Aga16, Ada17]. Diagnosis [Dus11]. Diagnostic [Hog13a, Kre14b]. Dial [Bat11, Kre14b]. Dialogue [Tar15]. Diameter [Can17]. Diane [Wis17]. Diatonic [Ava15]. Dick [Rol10c, Bim18, Eva15]. Dickens [Phi14]. Dictatorship [Cur12]. Dictionnaire [Pan18c]. Dienel [Pop13b]. Dienesch [Dav17]. Dierikx [Bye19]. Diesel [Lag11, Neu14b]. Dieselautos [Die12]. Dieselization [Sch12e]. Dietmar [Kai17]. Differ [Wis17]. Difference [Ler10]. Diffusion [McC13]. Dig [Jac11a]. Digital [All11, Alp16, Boo17, Bow12, BL12, Cur12, Dow15, Dug10a, Ens12, Gal16, McC13, OF18, OS16, Pan12, Ret16, Ves19, Wij12, Ihd16, CK15, Dug10b, Kar18, Kra17, Suo12, Sza15, Tho16]. Digitale [Spe11]. digitali [Col17b]. Digitality [Abn19]. Digitization [Kra17]. Digitize [Bor10]. Digitized [McC14b]. Digitizing [Sch13a, Sch19]. Dilemma [Hjo16, Jon10, Whi10]. Dilemmas [Tya12]. DiMoia [Moo16b]. Dinesh [Kum16]. Diogo [Vai19]. Diplomatic [Whi11b]. Directions [Moh15, ZT18]. Dirk [Vai19]. Dirt [Jac19]. Dirty [Jon10]. Disability [Alp16]. Disaster [Bro17b, Col12c, Jen13, Kno14b, Luc13, Nan12, Oka18, Cro14a, Sti18, Pom10]. Disasters [OS16, Car14a]. Disc [Dow10]. Discarded [Tin19]. Discipline [Lin11]. Disco [vdV14]. Discord [Sch13e]. Discourse [Mal17, MV15, Ret16]. Discovering [Wan10]. Discovery [Hag15b, Hun16]. Discriminating [Lig11]. Discussion [Buh17, TAF^{+}18]. Disease [Dec18, Deu17, Sta10a]. Disko [Woo19]. Disordered [Vos10]. Display [Bed18, Gil11a, TAF^{+}18]. Displaying [Cho18]. Disrupting [Car15]. Disruptive [Bix18]. Distance [Lig14]. Distillery [Coo11a]. Distribution [Bhi10]. Ditch [Car16]. Divall [Ber16, MF16]. Divergence [Til19]. Divided [Cas16b]. Divining [Ham12b]. Division [Łot16]. Dixie [Col16]. DIY [Dou19]. Do [Dou19, Dou10, Edw18]. Dobraszczyk [For17]. Doctored [Bra13]. Doctors [Hai15]. Documentary [Rio11, Rus12, SM17]. Documents [Bir19]. DoD [Her16b]. Does [Boy16, Coo10a, Pat19, Pul16, Bil15]. Doesn’t [Got12]. Dog [Hag15b]. Doing [Dus18, Rud13, Smi10c]. Dolly [Smi14b]. Domain [Cum11b]. Domenic [Zyl15]. Domestic [Ore11, Par16]. Domesticating [Gol12]. Dominic [Dow15, Jin19, SH19b]. Dominique [Gro18, MS16]. Dommann [Cum19]. Dompter [Pan18a]. Don [Wol14]. Donald [Amm14, Sti18, Ten15]. Donna [Car15, Pet16]. Donoghue [Kou10]. Don’t [Kat14]. Donzé [Has16]. Doogab [Low16a]. Doomed [Rub18]. Door [Ore13]. Dorn [McC12]. Doron [Sin15]. Dorothy [Ham18d, Jos16a]. Dot [Cor19]. Dot-Com [Cor19]. Doubt [Rob12]. Douglas [Pfa13]. Dow [Lee14]. Down [Ver13]. Downey [Cro14b]. Dragnets [Rus11b]. Dragons [Van16]. Drama [Kim13]. Drawing [Com12]. Dreadnought [Fer13]. Dreadnoughts [Zim10]. Dream [Big12, Tob13, Uzi19]. Dreams [Bar16a, Bre16, Cer10, Est10, Hen19, Kim15, Rei12, Var18, Wri12a]. Dreamscapes [Wil19]. Dress [Mas17]. Dribble [Car13a]. Drive [BBB^{+}19]. Driven [Hag15a]. Drivers [Wel10]. Driving [Fra18, Die10, Sjö11, Zel11]. Drones [Gra18, Mey19]. Drought [Cru19]. Droughts [Van16]. Drucker [Pet16]. Drugs [SH17]. Duarte [Lan18b]. Dufaux [Pan18b]. Dufresne [Gro18]. Dumont [SS13]. Dunn [Wol14]. Duong [Pur17]. DuPont [Cer13, Joh11b]. during [And17, Ben17, Boc16, Cam16, Col16, Cos17, Mar19, Mul12a, Par16, Rot16]. Dust [Cla10]. Dustheap [Zim13]. Dutch [vdHP13]. Dutchman [Bye19]. Dying [Whi10]. Dynamic [Gul19, Rud12]. Dynamics [Vee10].

Each [Kit12]. Eads [Bro14a]. Ear [Cah14]. Early [Arc19, Bet14, Bre10a, Bro17a, Chu15, Cla10, Coc13, De 16, Ens11, FR18, Gol15, Gue13, Haf11, Ham12b, Ham18d, Han13, KO13, Lim19, Mau11, Pat11, SK16, Sim18, dS17, DK18, Hau19, Rab13b, Rob17, Wer12, Kni13, Lov11]. Earth [Ham18c, Ham12b, JJ16, Jos13, Rus11a, Ver13, Scr12, Tur10]. Earthquake [Jen13, Wil17c, Mur12]. Earthrise [Tur10]. East [Bro14b, Gol15, Maa16]. Eastern [Gil15, Zar13, De 12, Shu19, Łot15]. Easton [Mey16]. Easy [Olw18]. Eat [Cla10]. Eating [Moh12, Wu11]. Eccentricity [Wri12b]. Ecclesiastical [Bon16]. Eclipse [Lau11]. Eco [Neu14b]. Eco-Friendly [Neu14b]. Ecological [Hos13, Mul18b, Rom10, ZT18]. Ecology [Bea19a, Car16, Coo11c, Ore11]. Economic [Bac15, Coc13, Dob10, Jaf11, Ken14, Las18, Mir14, Ros10, Sin12, Tin18, Vee10, Wak11]. Economic-Planning [Tin18]. Economics [App14, Lif18, Ros12b, Zim10]. Economy [Gif17b, Jaf11, Mad19, Pla16, Kle10]. Écrire [Sch19, Gui16]. Ed [Nof18, Åbe15, Anb18, Ang17, Ber16, Bim18, Blo13, Car15, Coc18, Col17a, Col15, Cro14b, DB13, Dis15, Dob16, Dow18, Dow16, Dup15, Edw18, Fer13, Gal15, Gar13, God15, Gor17, Gra15, Gro18, Hac15, Hah15, Hea13a, Her17a, Hic14, Hle19, Hu14, Hug15, Kai17, Kil15, Kin14, Kre14c, Lac16, Lau13, Lau17, LeC17, Leu14, Lev17, Łot15, Low13, Mac17, Mai16, MF16, Mar17a, Mey19, Mil13a, Moh15, MS16, Oak19, Oka18, Paj16, Pan15, Pan18a, Pla17, Pop13b, Pur17, Rae16, Ran15, Rub18, Rud13, Sah13a, Sca14, Sch13c, Set16, Shu15, Smi13, Sny17, Sta17, SP17, Sta14a, Sto13b, Tar15, Ver15, Ver13, Wad16, Wan13, Wea18, Wel15, Wil14b, Wil19, Wol14, Wor16a, Zar13, Zim13, vdV14]. Eda [vdV14]. Eden [Abb11]. Edgar [Acl17]. Edge [Joh18a, Pan16, Tin19]. Edgerton [Nah13]. Edison [Gop15, Rhe14, Hir14, Hin11]. Edisons [Lee13]. Edited [Rhe14, Eva15, Kat14, Low14, Sch14b, Smi14b, Gol15, Swa10]. Editor [Ali14, Ano19b, Haf17, Mar16, McD12, Mis11c, She14b, Uye11, Yan13]. Edo [Fru14]. Édouard [Fea19]. eds [Rol10c]. eds. [Sla11b]. Education [Goo19a, Hir10, Mil13b, Sch15a, SGB16]. Educational [Gar13]. Edward [Hoo15, Kli19, Oka18, Sla12b, Vit17, Man13]. Edwardian [Fer13]. Edwards [Mar18, Tuc16]. Edwin [Ham14a, Kra10]. Effect [McN14, Str12]. Effects [Kru16]. Efficiency [Web15]. Efforts [SV19]. Effosse [Yat16]. Egyedi [Car15]. Egyptian [Gru16b]. Eiffel [SS13]. Eighteenth [Gor10, Pan15, WH11, Ber13, EW12, HPR13, OL17, Van16, Zan13]. Eighteenth-Century [Pan15, WH11, Ber13, EW12, HPR13, OL17, Van16, Zan13]. Eighth [Ber12]. Einführung [Die11]. Einstein [Can15, Bey12, Joh11a]. Eisenhower [Dav17, Day11, Mey17, Sla12a, Neu13]. Eisler [Hög14c]. Elder [Smi14a]. Electric [Bow14, Cur13, Hin11, Hir17, Hug15, Hun16, Lif18, Oka17b, Plo12, Wil17a]. Electrical [Joh15a, KNSZ16]. Electricity [Boc16, Gol12, Goo19b, Hau10, Hau15, Hir14, Kap19, Zan13]. Electrification [Arn18]. Electrifying [Hau10]. électrique [But12, Lag15]. Electronic [Arc10, Cor12b, Gru16a, Mir19, Cor12b]. Electrotherapy [Ham17]. Elektrotechniker [Kai10]. Element [Ham11]. Elemental [Kim19]. Elena [Han13, Kat14, Lan18b]. Eleventh [Eds10]. Eleventh-Century [Eds10]. Eli [Gru14, Wes10a]. Eliot [Kar18]. Elite [Kum19a, Tho15]. elites [Pan18c]. Elizabeth [Abb15, Alp16, Ken14, Mar18, Ret16, Tuc16, Zim13]. Elkind [LeC17]. Ellen [Ore13]. Elliott [Hic12, Wil17a]. Elliott-Automation [Hic12]. Elman [De 12]. Els [Hic14]. Embargo [Can17]. Embedding [Aga14]. Embellishment [dSS19]. Embodied [Dow17, ZT18]. Embodiment [Lae16]. Embody [Hah17]. Embrace [Wis17]. Embracing [BF19]. Embryo [Col13]. Emergence [Col15, Gar18, Low16a, Mar10, Nee18, Rie16, Set16]. Emergencies [Tay14]. Emiliano [Dow18]. Emily [Hea13a]. Emmerson [Rae16]. Emory [Swa19]. Empire [Ada17, And17, Bro14b, Cah19, Coc16, Cos17, Dal15, Dec18, Eng12, Fer19, Fir12, Hea12, Hos13, Jon15, McD15, Sat10a, Sch14c, Sch10c, Sti11, Sul19, Tsu12, Hac15, Sov16, Cor14, Mac16b, Ram10, Ram11, Sch12a, Sto19, Tsu12]. Empires [Buc12a]. Empirical [Ver15]. Empowerment [Abb18]. Empty [Sch11b]. Enchantment [For17]. Enclosing [Cum11b]. Encounters [Ste14, Sch14a]. End [And17, Bow10, Dal15, Kem12, Ran11]. Endangered [KL19]. Endangerment [KL19]. Endocrinology [Tob13]. Enduring [Pla10, Jac12a]. Enemy [Kru13b]. énergie [Lag15]. Energy [Bla18, Bro15, Col12c, DiM10, Gor17, HS13, Hir16, Jon10, Kar17, Lag13, Mit18, Sov16, Wel12, Wil14b, Wol15, vMS15, vdH17, Hau15, Doe12, Hau15, Hir16, Shu15, Wol15]. Engaging [Sta17]. Engelke [Jør17]. Engine [Smi14a, Ken14]. Engineer [Joh11b, Ric13, Coo12b]. Engineered [Ore11]. Engineering [And17, Bhi13, Bra19, Bro11, Con15, Fer19, Gan11, Gat18, Gra11, Hah19, Kin11, Kir12, Lan15, Low16a, Mes16, Nan12, Pet10, SC17, Sch15a, Sla11b, Vos16, Wel12, vM12b, See15a, Kni18, Moo15, Sut12, Zyl15]. Engineers [Ald16a, Fri16, Gro14, Hir10, Jos16b, Koc16, Koe18, Mey18, Olw13, Reu14, And15]. Engines [Lag11, Rol17]. England [Bra15a, San11, Zel11, Bon16]. English [Bil15, Dup15, Rab13a, Rat12]. ENIAC [De 17]. Enigma [Pla10]. Enigmatic [Fea19]. Enlightened [Ber13, Jaf11]. Enlightenment [Aso12, Ban16, Cas13, EW12, Eva15, Kle16, Ben10, Smi11b]. Enlightenments [Sul19]. enregistrer [Sch19]. Entangled [Eng12]. Enterprise [BS13, Cla13, Fer19, Rol16, Wil14a, Fea16, Ros11].

Enterprising [Eva15]. Entertainment [Jin19]. Enthusiasm [Coo10b]. Enthusiast [Luc14]. Enthusiasts [Mey18]. entre [Lag15]. Environment [Buc12b, Coc18, Fin10, Gud13, Hir12, Oat16, Spr19, dSS19, Sta14a]. Environmental [Aik19, Bar12, Bea19b, Dér18, Ens18, Gor12, Jør17, Maa16, Mac16b, Man18, Man19, Mar11a, Mit18, Mul16b, Nye13, RAF^{+}11, Sah13a, Smi14b, Sti11, Str10, Sum17, Sut12, Tar14, Wak11, War14, Zim11, Fja12]. Environmentalism [Whi13b, vMS15, Ham13, Jen15, Nee18]. Environments [Bor14, Bro17b]. Envirotech [Fin10, SDJ19]. Envirotechnical [Par18]. Epic [Moh16]. Epistemic [Wea18]. Equal [Ava15]. Equal-Tempered [Ava15]. Equality [Lev13a]. Equipment [Kre14b]. Equity [Maa16]. Era [Aik19, Bro14b, DeB11, Gua13, Kil15, Lan18a, Li16, Sat13b, Sla12a, Ore11, Pla16, Ben17, Car13a, Col16, Dow15, Dow16]. Erastus [Reg18]. Erfahrungs [Tod11]. Erfahrungs-geschichte [Tod11]. Erhard [Epp15]. Eric [Gal16, Lan15, Mey16]. Erickson [Nov11]. Erie [Sch10c]. Erik [Gif17b, Mes16]. Erkki [McA14]. Ernst [Sza15]. Espagne [Lag15]. Espionage [Ber13, Day11]. esposizioni [Bal11a]. ESS [Wor16a]. Essay [Dis10, Jas12, Mül16a]. Essays [Ber11, Blo13, Cro14b, Dio12, Ihd11, Pur17]. Essen [Hau11]. Essential [Val10]. Establish [Lin11, Wor16a]. etal [Bra15b, Dis15, Shu15]. Ethical [Mey19, Mod11, Tim18]. Ethics [Col13, Hir16, Pet10, Sla16, Wol15, Nan12, Obe14]. Ethnographic [Hoo19]. Étienne [BBB^{+}19]. Étudier [Gil10]. Eureka [Mer17a]. Euro [Pra19]. Euro/West [Pra19]. Euro/West-Centrism [Pra19]. Europe [DB13, Die10, Haf11, Hau10, McC13, Paj16, Pan10, Set16, Sha10, TS12, Wel15, Whi11b, Zia18, vM16, Hau19, HPR13, Tay14, DB13, Gag15, Hui16, Lac16, Mor14b, Sid18, Sul19, Bre10a, Eds10, Fin13, Gif17b]. European [Ben10, Dio11, Jør10, Lag13, Lau13, Łot15, Rei12, Wor16a, Lon10, Shu19]. Europeans [Vai19]. Evaluating [Ona17]. Evangelista [Mey19]. Evans [Jos19]. Eve [Cru19, Gan17b]. Evenden [Boc16]. Everette [MM11]. Every [Coo11a]. Everyday [Col17a, Hoa15, Kat14, Low10, Mer17a, Ros17, Ten15, Cha14]. Everything [Ryd11]. Evidence [Åbe15]. Evolution [Bhi10, Cur13, Mat17b]. Evolutionary [Scr12]. Evolving [Bri10, Tho16]. Example [Pet19]. Excess [For17]. Exchange [Cor12b, Kri19b, Mac15, Rob10a, Whi11b, BM14]. Excitable [Cre17]. Exemplary [vdHP13]. Exercise [Gul19]. Exhibit [Cho18]. Exhibiting [Zak15]. Exhibition [CG16, Hey10, TAF^{+}18]. Exile [Kri19a]. Expanding [Mun19]. Expanses [Har19]. Expansion [Hel13]. Expect [Kit12]. Experience [Goo19a, Ihd11, Jac15, Lin11, Zyl11a, Haf18, vR12]. Experiences [SSH18]. Experiment [Gru14, Hag14, Pet12, Fle19]. Experimental [Cah14, Kur19b]. Experimentation [Cam16]. Experimenting [Wet15]. Experiments [Cah14, Hun16, Rud13, Rud18]. Expertise [Abb18, Gol12, Haf11, Joh11b, Sis12, Zia17, Kru13b, Cer11, Sut12]. Experts [BSA16, Luc13, Nah13, vM16]. Explanation [Edw16]. Explicit [Bij11].

Exploding [Eis17]. Exploration [Bro14a, Cro16, Las18, Sid10, Smi13, Whi13a, Mes16]. Explorations [Gro18, vdS19]. Explorer [Cro16]. Exploring [Sie14]. Explosion [Ens11]. Explosives [Kru14]. Expositions [Gil11a]. Exposure [Ona17]. Extraordinary [Sut19]. Extremely [Smi10c]. Eye [For10]. Eyeing [Dav17]. Eyes [Day11].

F [Bro15, Kni18, Mey16, Mis11b, Olw13, Pla17]. Faber [Cou17a]. Fabric [Ert16, Mar11b]. Fabrication [Wil19]. Face [Col12b]. Faces [Edw16, Fru15]. Fact [vdH17]. Factish [Fis13]. Factors [Gif15b]. Factory [Les19, Ros13]. Failed [McK13b]. Failure [Aso12, Kli19, Ten14]. Fair [Hou10]. Fairchild [Rio11]. Fairs [Gan17b, Nye11]. Faith [Ham18b]. Fall [Fin13, Hul16, Pop16, Ros13, Vol11, Wes16]. Fallan [Bre18]. Fallon [Pur17]. Fallout [Iri12]. Fame [Hor11]. Families [Car14a, Gue13, Hur15]. Family [Ham14b, Zyl15, Mom14]. Famine [Kim12]. Fandom [Boo17]. Fang [Hai15]. Fantasies [Fer12]. Far [Fin10]. Farm [Ham14b, Swa18, Dry19]. Farmer [Haf18]. Farmers [Swa18, Swa18]. Farms [Hur15, Moo16a]. Fascism [Coc17a]. Fascist [Coc17a]. Fashion [Bet14, Wos17]. Fast [Fja15, Mad19]. FastLane [Sla18]. Fatal [New18]. Fate [Cam15, Jac16, Kou10, Rud12]. Fatigue [Gor11]. Fats [Sch12b]. Faulkner [Hic14]. Fault [Til19]. Fax [Pop16]. FAXED [Pop16]. FDA [Boy16]. Fears [Fer12]. Features [LM19]. Federal [Hir12, Jac16, Mar11b, Lig11]. Federation [vM12b]. Feed [Lan11]. Feenberg [Ihd18]. Feminine [Tuc12]. Feminism [Bou19]. Feminist [Pre19]. fer [Pan18b]. Ferdinand [Vol10]. Ferguson [Mey15]. Fermi [Wes10b]. Fermilab [Mod10]. Ferriss [Vol17]. Ferry [Sim10]. Fertilizers [Cam15, Sah13b, Whi10]. Fetish [Dug11]. Fetus [Hog14a, Hog14a]. Fickers [DB13]. Fiction [Fer16, Kin10, Sch15b, Tel11, Tel13, vdH17, Sla10]. Field [Kui19, Mir14, Ben17]. Fields [Sat13a]. Fifteenth [Ans17]. fifth [BBB^{+}19]. Fifty [Neu14a, Pre13]. Fight [Mai10, Wol11a]. Fighter [Gai19]. Figure [Wel17]. File [Rob17]. Film [Gar13, Hok17, Tel11, Wei13, Wos10b, Fro14]. Films [She14a]. Fin [Hui16]. Fin-de [Hui16]. Final [DeB11]. Finance [Hoc10, Mir19]. Financial [Bra15a, FR18]. Financing [Reg18]. Finding [Jør13]. FinFET [OF18]. Finland [Wil11, MF11]. Finlay [And14]. Finn [Coo14, Sam13, Smi14b, Nof18]. Finnegan [Edw17]. Finnish [Mat19]. Finnström [Gra15]. Fino [Gai19]. Fire [Goo11, Wel15]. Firearms [Hac15, TAF^{+}18]. Fires [Sca11, SH15]. Firm [Gif15a]. First [Bac15, Dav17, Gai13, Gal15, Hag15b, Łot15, Mar13, McA14, Mey15, Mor12b, Sch14a, Tom11, Tur10, BBŁT15]. Fish [Fin17, Hun16]. Fishing [Lau17]. Fit [Ree13]. Five [Jac19, Kel10]. Fix [Joh18b, Jos19]. Fixation [Bix18]. Fixed [PR16]. Fixing [Fra18, Coo11b]. Flagships [Mat19]. Flanders [DH11]. Flash [Kor18]. Flashy [Mat19]. Flawed [Ada17]. Fleet [Wol11b]. Fleming [Hog14a]. Flexible [Gai15]. Flickering [Nye14]. Flier [Jeo15]. Flight [Bye13, Ded18, Her16a, Jeo15, McC15, Mey18]. Flood [McC13, San17, Hob14]. Floor [Jaf10]. Floracrats [Aso12]. Florentine [Ans17]. Florman [Pet13]. Flow [Bro17b]. Flown [MV15]. Flusser [vM12a]. Fluted [Hen18b]. Fly [Jac11a]. Flying [Bye19, Mey19, O’D12, Uzi19, vR12]. Flyswatter [SR19]. FM [Lov11]. Foetus [Hog14a]. Fogel [Bye13]. Fokker [Bye19]. Folke [Sah13a]. Follett [Gud18]. Folly [Smi10a]. Food [Boy16, Mer17b, Moh12, Ona18, Sah13b, Wil13, Gre11]. Foote [Zim13]. Footprint [Jac13, Mar11a]. Force [Dou16, Gai19, Gai13]. Forces [Rub18]. Ford [Mac18]. Ford. [Cas15]. Forecasters [Bac15]. Foreign [Ala12]. Forensic [Bou19, Med18, Gan17a]. Forest [Fro14, Koc16]. Forestry [Koc16, Joh17]. Forests [Dér18]. Forever [Dow10]. Forged [Cit12, Kra14]. Forgery [Lim19]. Forgive [Ten14]. Forgotten [Arc18]. Form [Abb15, Che19]. Format [Cum13]. Formation [Sim18]. Formats [Col19b]. Forms [Sid18]. Fors [Kle16]. Forschungsergebnisse [Die11]. Fortieth [KLŁ14]. Fortin [Ran15]. Fortresses [Mey19]. Fortune [Bij13, Hög13b, Bac15]. Forty [BBŁT15, BGW^{+}17, BBB^{+}19, GKMV18, KNSZ16, Sla15]. Forty-fifth [BBB^{+}19]. Forty-first [BBŁT15]. Forty-fourth [GKMV18]. Forty-second [KNSZ16]. Forty-third [BGW^{+}17]. Forward [Cow10, Mir14, Roh19]. Fossil [Bar16b, Qui17]. Foster [McC14a]. Foundational [Kli13]. Foundations [BL12]. Fountains [KO12]. Four [Nor15b, Wro10]. Fourth [KNSZ16, GKMV18]. Fragile [Mar11b]. Fragments [Hey10]. Frames [Fis13]. Framing [Hoc10]. France [BBB^{+}19, Fox11, Lag15, Yat16, Zar17, Ber13, CC18, OL17, Pan15, Par17, WH11]. Francesca [Kim14, Van18]. Francesco [Coo17]. Francis [Hok17, Sti18]. Francois [Cas19, Cou17a, Hel13, Pan18a]. Francoist [And15]. Frank [Fai16, Kre14a, Gal11, Hal16, Kou10]. Franklin [Abn19, Nor15a]. Franz [Gif17a]. Fraud [Gro14]. Frazier [Mey16]. Fred [Geo15]. Frédéric [Har13]. Free [Kir12, Pri19]. Freedom [Zyl11b]. Freeze [OS12]. Frehner [Jør13]. Freiheit [Kre10]. French [Dou16, Hel13, Lev13a, Vai13]. Frequency [Kat10, Mac15]. Fresh [Mor10b, Ham10a]. Freudian [Bow12]. Fridenson [Hei15a]. Friedman [Bac15]. Friendly [Fin13, Neu14b]. Front [Kat11]. Frontier [Cas16a, Kru12, Mod10, Wes10b, Wu19, Vet19]. Froth [Sch10b]. Fuel [Hög14c, Kru13a]. Fuels [Bar16b]. Fukushima [Oka18, Ona18]. Fuller [Seg10]. Fullfillment [Bar18]. Fullilove [Ham18c]. Fulton [CF10]. Fun [Pri19]. Fundamental [Cer13]. Funeral [Sta14b]. Furniture [Edw12b]. Further [MSPLC16]. Future [BF19, BBB^{+}19, BSA16, Bil15, Bow12, Buh17, Col12c, Der17, Kim15, Kno14b, Lon10, Nor15a, Nov14, Ste11, Tho15, Ber11, Dus10, Kra19]. Futures [Pic14, Dis15]. Futurism [Dry19, Hög14c, Joh18a].

G [Abb15, Dow18, Hal16, Hok17, Kur19a, Luc18a, Mey15, See15a, Smi17, dS17]. Gabriel [Luc13]. Gabriele [Col17b, Wea18]. Gabrielle [Abr13]. Gagnon [Lak13]. Gal [HV15]. Gale [Dou16]. Galea [Cam16]. galerie [Pan18b]. Galileo [Ver12]. Gallaga [Dow18]. Gallamore [Gam15]. Galloway [Ihd16].

Gambling [Lee14, Gro15]. Game [And17, O’D12]. Games [Bes13, Fis13]. Gaming [Sch12a]. Gandy [Ert16]. Ganson [McC15]. Gap [Mir12, Pre19]. Garage [DH11, De 16]. García [Kat14, Cor16]. Garcia-Swartz [Cor16]. Garden [Pat15, Str11]. Garments [Ret16]. Gary [Bye13, Fra16, Luc18b, Sat13b, Hig12]. Gas [Ham18b, Kem12, Lag11, Lag13, Tar14, Tom11]. Gaslight [Cas13]. Gate [Bor17, BL12]. Gates [Kra11]. Gatling [Ryd11]. Gauge [Kre14b, Vee10]. GDR [Wes10a]. Gearheads [Mor17]. Gelehrte [Gif17a]. Gender [Coo11a, De 12, Gol12, Hay14, Hic14, Kim14, Ler10, Mai12, Ona18, Col12a, De 11b, Sch11c]. Gendered [Haf18]. Gene [Fin17, Hog18b, Fea16]. Genealogy [Hog18b, McL17]. Genentech [Ona14]. General [Cur13]. Generation [Bey12, Kap19]. Generic [Dae19]. Genes [Fin17, Har13]. Genesis [Nof15, Vol10]. Genetic [Had10, Low16a]. Genetics [Big16]. Geneva [CR19]. Genius [Ryd11, Vol10]. Genome [Com12, Wad16]. Genomics [Hog18a, Ste15]. Genres [Sch12d]. Gents [De 11b]. Geoff [Gol14a]. Geographies [Bie16, Eng12, PR16]. Geography [Ran14, Sop11, Van16, Zel11]. Geology [Jør13, MM11, Ree16]. Geophysical [Clo12]. Geopolitics [Cam15, Joh16, Lag17]. George [Gam13]. Georges [Sch19]. Georgia [Lak13]. Geppert [Lau13]. Gerard [Paj16]. Germ [Kir12]. Germ-Free [Kir12]. Germain [Mit19]. German [Gif15a, Sjö11, Sla11a, Vos10, Kre14b, Les19, Vos16, Col16]. Germans [Bus19a]. Germany [De 15, Neu11, Rol17, Sjö11, Str10, Bir19, Gif15b, Hau11, Neu14b, Pre10a, Tka19, Vos16, Bro17a, Ham12b, Hes11]. Germuska [Łot16]. Gerolamo [Cam16]. Gerovitch [Fir17, Jos16b]. geschichte [Tod11]. Gesellschaft [Gif17a, Kre14a]. geste [Ora15]. Get [SH19a]. Getting [Seg10]. Ghana [Got12]. Ghee [Ber19]. Ghiberti [Ans17]. Ghost [Gal15, Hag15b, Ruu11]. Giacomo [Hac15]. Giancarlo [Whi13a]. Giffard [Rol17]. Giguere [Gru16b]. Gijs [Rie16]. Gil [Mit19]. Gilded [Mou10]. Gillian [Hah19]. Gilliom [Mac14]. Gillis [Cou17b]. Gillispie [Rob13]. Giovan [Coo17]. Girl [Cla11]. Girls [Sch15a]. GIS [Ran15]. Glaciers [Clo11]. Glasgow [Ber12]. Glass [Sea11, Whi11b]. Glassmaking [Ber17, Sea11]. Glean [Tho16]. Gleick [Hob14]. Glenn [JI16]. Global [Abr13, Bek18, Bil15, Cam15, Eli12, Eng12, EW15, Fru17, GKMV18, Gol15, Gor12, Gud18, Ham18c, Hir16, Hog14b, Hor18, Hri16, Jin17, Lad14, Mai10, Mar11a, Mar11b, Mor12b, Mue15, Mul18b, O’D12, Rob12, Shu15, Sid10, Sla19, Ste10a, Sto13a, Sut12, Til19, Wol15, Wor16b, Zia17, vdH17, Hug15]. Globalization [Gre11, Lag11, Mat17a, Bre18]. Globalized [Cam19]. Globalizing [Mac17, Clo12, Vai19]. Globe [Kim13]. glorious [Kra11]. Go [Jac11a, Pin10]. Godfrey [Hok17]. Gods [Fis13, Hen19]. Goes [Ber17]. Goffman [Pin10]. Going [Spu11]. Gold [Dér18, Spu11]. Golden [Bat11, Cos17]. Goldfields [Pet19]. Goldsmith [Bra15a, Sch13e]. Goldstein [App14]. González [Lev17]. Good [Bij13, Lau11, Pet13, Tim18]. Gooday [Joh15a]. Goods [Par16]. Goodyear [GR10, Low14]. Gordon [Cer16, Coc17b]. Gorman [Swa15]. Gorsky [Sta14a]. Gotham [Koe18].

Gould [JI16]. Gourmets [Kim12]. gouvernement [MS16]. Gouverner [MS16]. Governance [Sha17, Moh15, Mue15, Pla17]. Government [Dec18, Hul16, Jac16]. GPS [Mey16]. Grace [Ree16, Vin12]. Grade [Man13]. Graeme [Joh15a]. Graham [Gra16]. Gramelsberger [Wea18]. Grand [Ric13]. Grandfather [Ger11]. Grantedness [Aga14]. Graphics [Mei13]. Grasslands [Hay18]. Gravity [Hag15b]. Great [Car14a, Cut10, Dou16, Edw17, Jør17, Kim14, Lan11, Mir14, Seg14, Sin15, Sla14, Til19, Edw16, Gil15]. Greatest [Hag14, MV15, Rub18]. Greed [Har16]. Green [Big12, Fin13, Ham13, Sam13, Coc18]. Greene [Sop15, Dae19]. Greening [Sop14]. Greenwich [Ste10b]. Greg [Gan17a]. Gregory [Cro14b, Whi14]. Grid [Fre18]. Grocers [Ree18]. Groovy [Her17a]. Grossman [Phi14]. Ground [Fin11, LeC11, Sti18]. Groundbreaking [Shu19]. Group [Tho15]. Growing [And10]. Growth [Bea19a, Cho18, Ros11, Mir14]. Grusin [Sny17]. Gryphon [O’D12]. Guano [Whi14]. Guardianship [Mad19]. Guatemala [Mat17a]. Gudmestad [Dav13]. Guide [The10b]. Guldi [Car13b]. Gulf [Col12c]. Gun [Ryd11]. Gunmaker [Ans17]. Gunners [Cos17]. Gunpowder [Buc12a, Buc14]. Guns [Sto19, Hle19, Mav10]. Gupta [Oak19]. Guru [Fri16]. Gustafson [Hög13b]. Gusto [Ber14]. Gutenberg [Kra11]. Guts [Car11]. Guys [Pet13]. Gyeongbu [Jeo10]. Gyung [Sul19].

H [Ban16, Cas16a, Cas14, Dus18, Gan17b, Ger13, Li16, Phi14, Ver15, Zar13]. Haaften [Pan18d]. Hård [Gag15]. Habits [Fru14]. Hacker [Ang17]. Hacking [Paj16]. Hadland [Epp15]. Haffner [Jac14]. Hafter [Pan15]. Hahn [Mai16, Ole17a, Swa13b]. Haigh [De 17]. Halcyon [Gru14]. Hall [Cho18, Her16a, Hor11, Seg14, Gud13]. Hallam [Cha18a, Hag15a, Wad16]. Hallenbeck [MF17]. Hallonsten [Wor16a]. Hamblin [Jen15]. Hamilton [God15, Kle10]. Hamlet [Lau17]. Hampf [Hug15]. Hand [Hau19, Pac18]. Hand-pump [Pac18]. Handbook [Bar14a, Luc11, Mil13a]. Handful [Rob12]. Handicraft [Wu11]. Handler [SP17]. Handloom [MB18]. Hands [JJ16]. Handy [Kre10]. Hanford [Kru12]. Hanna [Mac13]. Hannah [Oak19]. Hans [Bra15a, Epp15, Kri19a, Mai16, Pop13b]. Hans-Erhard [Epp15]. Hans-Joachim [Bra15a]. Hans-Liudger [Pop13b]. Hans-Ulrich [Pop13b]. Hanssen [Kra17]. Happen [Dou10]. Happy [Hus10]. Hard [Gif15b]. Harding [Sto13b]. Harhoff [Kai17]. Harold [Gif15a, McC12, Ver15]. Harper [Ber18]. Harpers [Sim10]. Harriet [Ham13]. Harris [Gol14b]. Harry [Hag15b, Jos19]. Hart [Sch14a]. Hartmut [Coc18]. Harvesting [Coc17b]. Harwood [Bye13, Fin13]. Hatch [Mer17b]. Hawai’i [Meh10]. Hawking [Ihd13]. Hayao [Bre16]. Hayles [Dow16]. Hazardous [Mar10]. Hazlett [Aga18]. HDTV [Mei10]. Healing [SH17, Rob10a]. Health [Bra19, Gor12, Hay18, Hep16, Kru16, Moo16b, Ore12, Pre19, Ros13, Sta14a, Tya13, Gil11a]. Healthcare [Tim17]. Healthy [Ros13, Sch12b]. Hearing [Luc18a]. Heart [Col16]. Hearth [Mei10]. Hearts [McK13b, Mck13a]. Heat [New18]. Heather [Hog14b]. Heavens [Dor10, Rus11a, Sie14]. Heavy [Sti18]. Hecht [Abr13]. Hecker [Rub18]. Hedonizing [Har11]. Heering [Rud13]. Heidegger [Str13, Fee11]. Height [Mul12a]. Heikkerö [Obe14]. Heinlein [Dis15]. Helen [Har19, Ree19]. Hélène [Ihd13]. Helping [Swa18]. Henry [Łot15, Mey19, Ten14, Cas15]. Herbert [Aga18]. Here [Got12, Nov14, Gor15]. Heredity [Had10]. Heritage [Mei19, Sca14]. Hermione [Rol17]. Heroic [Cro16, Kil15]. Heroics [Gal11]. Hersch [Ger13]. Heyck [Bes18]. Heymann [Wea18]. Hicks [Tin19]. Hidden [Gai18, Wol10]. Hide [Mac13]. Hidetoshi [Oka17a]. Higgins [Jos19]. High [Col15, KNSZ16, Mac15, Sin12, Vos16]. High-Frequency [Mac15]. High-Industrial [Vos16]. High-Speed [Sin12]. High-Technologies [KNSZ16]. Highway [Fei12, Jeo10]. Hilaire [Lev17, Ora15]. Hilaire-Pérez [Ora15]. Hilgartner [Hog18a]. Hilgers [Bes13]. Hill [Odd15]. Hillel [Mor14c]. Hills [Shu11]. Hinterland [Bra19]. Hirata [Ona18]. Hirsh [McD17]. Hirt [Hir17]. Hiss [Aga16]. Histoire [Has16, Lag15, Cas19]. Historian [Moh16, SW15, Tuc16]. Historians [Hir11, Kai11]. Historic [Sau11]. Historical [BGW^{+}17, Fja13, Gul19, Jaf10, OM17, Ona17, SK16, Sca11, Sch16, SV19, Sov16, Tuc16, Zel11, Zim17, Bim18, Ran15]. Historicizing [Anb18]. Histories [Bou19, Bre18, DiM15, Fja12, Geo15, Git10, Goo11, MF16, OS16, Til19, Ham12a, Zim13]. Historiographical [DGS16]. Historiography [Edg10]. History [Arc12, BM19, BBŁT15, Bea19b, Bea10, Bed18, BS13, BBB^{+}19, Ber12, BGW^{+}17, Ber14, Bes13, Big12, Bla18, Bra12, Bra15b, Bro14a, Bus19b, Cer15, Cha12, Cho17, Coc13, Coc17a, Col19b, Coo11b, Coo15, Coo14, Cor16, Cou17b, Cro18, Cro14b, Cum19, Dio12, Div10, Dor10, Dru12, Edg10, Edw17, ERST13, Ens12, Ens18, EW15, Fei12, Fer11b, Fle19, For10, Gai18, GKMV18, Gif15a, Gol14a, Gua13, Hac15, Hag15a, Ham11, Ham14a, Har19, Hir11, Hir17, Hob14, HNG16, Hog13a, Hoo10, Jaf11, JJ16, Jør17, Jos16a, KLŁ14, KNSZ16, Kim14, Kin11, Kno14b, Kui19, Kur19a, Lac16, Lad14, Lag11, Ler10, Lon10, Lub18, Luc10, Mac16b, Mar13, Mat17c, Mck13a, Med18, Mei13, Mil11, Mis11a, Moh16, New10, Nik17, Nye14, Nye15, OF18, OS16, Osl10, Pac18, Pos10]. History [Pri19, Rio11, Rob13, Rol10c, Rol10b, Rud13, RAF^{+}11, Sch13b, Sch15a, Scr12, Sha10, Sid10, Smi14b, Spi15, Spr19, Spu11, Sta17, Sti11, Str11, Sum17, SSH18, Swa15, Tin15, TAF^{+}18, Vee10, Vol17, War14, Wel10, Wil10, Wil11, Wil13, Wil14a, Wri12b, Wu11, dS17, dV12, dlP10, vdS19, Cho18, Bor17, Epp15, Hei15a, Joh17, Kan13, Nye13, Pet18, Sch13d, Sla15, Sta14a, Wan13, Bau11, Buc12b, Ham10a, Kru10, Mar11a, McC11, Mor12b, Pat11, Pet11, Vol11]. Hitting [Gan11]. HIV [Hog18b]. HIV/AIDS [Hog18b]. Hjalmar [Kle16]. Hobbies [Har11]. Hochfelder [Arc13b]. Hochman [Hoo19]. Hocking [Col15]. Hoddeson [Wes16]. Högselius [Gif17b, Lag13]. Hollowed [LeC11]. Hollywood [Sch12c, She14a]. Holtan [Hic14]. Holy [Ram10, Ram11, Wak11]. Home [App14, Blu12, Coo11a, Gol14b, Kat11, Ore11, Pat19, Swa18, Zar17, Pat19, SH15]. Home-Based [Pat19]. Homily [Sta14b]. Hommels [Moh15]. Homo [Cou17a]. Honour [Blo13]. Hoover [Aga18]. Hope [Abb15, Had10]. Hopper [Vin12]. horlogère [Has16]. Hormones [Tob13]. Horning [Gop15]. Horse [Ruu11]. Horses [Sop15]. Hostile [Kli19]. Hotel [Wri12a]. House [Bar14b, Kar17, Pat15]. Houses [Moh12]. Housing [Bar14b, De 12, Par17, Pen18, Zar17]. Howard [Nye15, War14]. Hudson [Nee18, Kni13]. Huggard [Spu13]. Hugh [Swa15, Gan13]. Hughes [Car14b, Gre14, Kai14, Ona14, Sta14b]. Hugo [Uzi19]. Huhtamo [McA14]. Hui [Cah14]. Human [Com12, Dal15, DL18, Dus18, Eri15, God15, Hay12, Hos13, Mar11a, Med18, Lie18]. Humanities [BT19, Dow16, Kou10, Pan12, Gal16]. Humanity [Lan18a, Var18]. Humans [Kli17, Whi11a]. Humble [Spu13]. Humboldts [Ole17b]. Hundley [Sti18]. Hungary [Łot16]. Hunter [Bes18]. Hunting [Ham12b]. Hurricane [Miz11]. Hwee [McL17]. Hybrid [Cha12, Wes15]. Hydro [Boc16]. Hydro-Electricity [Boc16]. Hydroelectric [Cas14]. Hydrogen [Zim10]. Hygiene [Leu14]. Hyldgaard [Sla11b]. Hyman [Dup15]. Hyun [Wil19].

ICBMs [Gai18]. Ice [Wil14a, SH19a]. Icebreaker [Mat19]. Iceland [Kai12, Ros10]. ICHSTM [KLŁ14]. ICOHTEC [BBB^{+}19, KLŁ14]. Icons [Fis13]. Idea [Spr19]. Ideal [Jac12a, Kir12]. Idealism [Vos16, Bix18]. Ideas [Fai16, Gra16, Tom14]. Identification [Bre10a, LM19, Med18]. Identities [Gue13, Joh15a]. Identity [Boy16, Gru16b, Jeo15, Mor14a, Fir17, Olw13]. Ideology [Bro14b, Cor12a]. Igniting [Mau11]. ihre [Die11]. II [Ham14b, Ore11, Rol17, BF19, Mar19]. III [Hel13, Fai16]. Illuminations [Ken19]. Illusion [Joh12]. Illusions [McA14]. Illusory [Buc12b]. Illustrated [Epp15]. im [Lin13, Ole17b, Tod11]. Image [Kre14c, Rab13a, vM12a]. Images [Col19b, Fir17, Goo19b, Kil15, Lan17]. Imaginaries [Gra15, Wil19]. Imaginary [Coo12a]. Imagination [Fis15, Ket16, Nof18, Sch16, Tuc16, Ert16, San17, SP17, Dis10, Rei12]. Imagined [Kai12, Sul19, O’H10, Sch14a]. Imaging [Hog14a]. Imagining [Hog14a, Kap19, Tel13, Lau13]. Imhotep [Kli19]. Imitation [Edw12b, Sea11]. Immunization [Bra19]. Impact [Bim18, Kra17, Lag11, Rol10c, Shu15, SV19, Tar14]. Imperial [Kim14, DiM15, Ket16, Leu14]. Imperialism [Hea12, Ven12]. Implications [Sie14, Bri10]. Importance [Hec14]. Imported [Cho17]. Impossible [Gra11]. Impressionism [Pop19]. Improper [Str13]. Improve [Pan18c]. Improvement [Gol14b]. Improving [Oka18, Edw12b]. Impulse [SM17]. In-Between [DH11]. Inclusion [Hic14]. Inclusive [God16b]. Incorporated [Ihd13]. Incremental [Lov11]. Indefensible [Mai10]. Independent [Ree18]. Index [Aga16, Ano12d, Ano15c]. India [Cha14, Kum16, Lau17, Pra12, Sat10b, Ven12, Ber19, Coc13, Kum19a, Tya12, Wer12, DiM15, Ram14, Sah14, Sha17, Kum12]. Indian [Hea12, Lan18b, Sin15, MB18, Sah13b]. Indias [Cos17]. Indigo [Sah14]. Individual [Car18]. Indonesia [Aso12]. Indus [Ada17]. Industrial [Ber13, Bow16, Cer13, Cha11b, Coo12a, Cur13, ERST13, Fru15, GR10, Hah19, Hau15, Hep16, Hor18, Lak13, Lev13a, Low18, Par16, Rab13b, Ros11, Ros13, SC17, Sch11c, Ste10a, Sto19, Tod10, Tya13, Vos16, Wil11, Zyl15, vdS19, De 10, Ros12b]. industriali [Bal11a]. Industrialism [vdHP13]. Industrialist [Eva15]. Industrialization [Mit18, Par18, Hor14]. industrie [But12, Has16]. industriel [Cas19]. industrielles [Rey12]. Industries [Haf18, Łot16, Ver13]. Industry [Bro11, Cas13, Cro18, Cut13, Gol14b, Gud13, Ham16a, Kas17, Mai10, O’D12, OF18, Oka17b, Par16, Rob13, Sin12, Sla14, Smi11b, Tar14, Tob13, CK15, Cor16, Han14, Hoo15, Ens10, Her11, Lai12, Lan11]. Inequality [Tin19]. Influence [Shu15]. Influencing [Gif15b]. Influential [Sla12b]. Informatic [Dow15]. Information [Coo15, Cur12, Ens11, FR18, Gla17, Hic14, Hob14, Jac10b, Joh16, Joh15b, Jon13, Ket16, McC13, Rob17, Sha17, Sla11a, SSH18, Vin12, Wor16b, Acl17, Dow17, Mar18, Aga16, Fel14]. Information-Age [Coo15]. Infrastructure [Car13b, Car16, FR18, Gif17b, Gul19, McD17, Muk12, Pla16, Pla17, Sam13, Tom14, Anb18, Oak19]. Infrastructures [Car15, DB13]. Ingenieure [Tod11]. Ingenious [Bow16]. Ingenium [Kel10]. Ingham [Wal18]. Inherent [Git10]. Initiating [Tya12]. Initiative [vdV12]. Injury [Gan13]. Inkstones [Ham18d]. Innocent [Rud18]. Innovation [BGW^{+}17, Bet14, Cha11b, Cur13, De 16, Dow10, Edg10, Edw18, EW12, Gal17, God16a, God16b, God18, Haf18, Her17a, Hoa15, Kai17, Lin10, MB18, Moh16, OF18, Pre10b, Ree18, Reg18, Rol16, RV18, Smi11a, Vol10, ZT16, But12, Fox11, Coo17, Kai17, Mac18, Wal18]. Innovative [Hin10]. Innovator [Luc12]. Innovators [Mer17a]. Innumerate [Ros17]. Inquiry [Kel11]. ins [Kai10]. Insights [Hin10]. Institute [SC17]. Institution [DiM10, God16a, Wil13, Her16a]. Institution-Building [DiM10]. Institutional [Kru13b]. Instrumental [Joh13]. Instruments [Jon13, Rud13]. Insurance [Lig11]. Intangible [Ran14]. Integrating [McC14a]. Integration [McD15]. Intellectual [Car14b, Coo10a, Cum12, Got12, Kra11, Mei19, Mer17a, Kri19a]. Intellectuals [Cow10]. Intelligent [Ber13]. Intensive [Her17b]. Interaction [Wac17]. Interactional [Sis12]. Interactions [Whi11b]. interconnexion [Lag15]. Intercontinental [Gai13]. Interdependency [Ber16]. Interest [Sid18, Rus12]. Interests [Kru13b, Web15]. Interface [Plo12]. Internacionales [Hea15]. International [BBB^{+}19, Clo12, Cor16, GKMV18, Gil11a, Joh16, Kai14, KNSZ16, Kra14, Kri19b, Kum12, Mor18, Neu11, Neu13, Neu14a, Pet19, Ste14, Tar15, Vai13, Wij14, Yos16, vM12b, vM16, BBŁT15, Ber12, BGW^{+}17, ERST13, KLŁ14, Wil11]. Internationalism [And17]. Internationalization [Bra17]. Internet [Fer12, Lan18b, Mue15, Pla17, Sla11a, Sla18, FR18, Pin10, RS14, Tho16, Coo14, Nik17, O’R19, Str12]. Interrelation [Bai18]. Intersections [Fin10, Sop11]. Interventions [Dec18]. Interwar [Ber19, Bir19, Bjö13, Cer13, Gif15b, Tel13, Tka19, Cro16, Pop11]. Intolerable [Kor18]. Intrigue [Aga16]. Introducing [Mor15]. Introduction [Aga11, Bor14, EW15, HS19, Kre14b, Kum19b, Mac14, Nan12, Oka17a, OM17, Rab13b, Roc17, She14b, Cha18a, Hor14]. Invent [Pan18c]. Invented [Bea10, Ryd11]. inventeurs [Fox11]. Inventing [Gal11, Bro12a, Est10, Ger13, Wol14]. Invention [Bal10, Gri12, Hoo15, How15b, Kim13, Lee13, TS12, Vin12, Sch19, Mar17a, Hin11, Mar17b]. Inventions [Moh16, Nye17]. Inventive [Bow16]. Inventor [Bow14, Joh15a]. Inventors [Lev13a, Rat12]. Invents [Car13b]. Inverse [Car15]. Investigating [Åbe15]. Invisibility [Kru16]. Invisible [HS13, Joh11b, Pin10]. Ippolito [Low16b]. Iran [Pat11]. Ireland [Goo19b]. Iridium [Sla19]. Iron [Bar16a, Cut13, Gor10, Man13, Nie12, Ruu11, For17, Lan11, Nie12]. Islam [BS13, Cur12]. Islamic [Pat11, Buc12a]. Island [Sim18]. Isolation [New18]. Israel [KNSZ16, Rhe14]. Issue [Ano10b, Ano10c, Ano10d, Ano11b, Ano12b, Ano12c, SS10]. Issues [Mod11, Rob12]. Istanbul [Kar18, BS13]. Italian [Bal10, Par16]. Italy [Eva16, Gue13, Bet14, Par18, Ver12]. Itself [Mun19]. IUD [Hog14b]. Iurii [Coo12b]. Ivey [Gru14]. IVF [Nor15a].

J [Bim18, Bow16, Bye13, CK15, Chu15, Cla13, Cro14b, Fer13, Fra13, Fra18, Hoo15, Hor18, Hun16, Jon15, Jør17, Kil15, Kin14, Lak13, Lub18, Mat17a, Mir14, Olw14, Pet16, Pri19, Ric13, Rol10c, Sch15b, Sch14c, Shu11, Sla18, Spu13, Ver15, Yos16, ZT16]. Jack [Joh15b]. Jackson [Amm14, Hog18b, Sti18]. Jacob [Bie16, Jen15]. Jacobson [Gar18]. Jacolin [Łot15]. Jahnavi [Ram14]. Jahr [Kai10]. Jahrhundert [Pop13b, Tod11]. Jakle [De 16]. James [Hah15, Hel13, Her16b, Hob14, Joh11a, Luc18a, McC13, Mun19, Smi13, Ung15, Ver13, Vet19, Gif15a, Nee17]. Jamie [Eri15, Roh19]. Jan [De 15]. Jane [Kru13a]. Janeiro [GKMV18]. Janet [Hay14, Lan17]. Janis [Tsu12]. Jansen [Mil13b]. January [HHH12, Seg14, Ano10g, Ano11e, Ano12e, Ano13d, Ano14d, Ano15d, Ano16d, Ano17d, Ano18d, Ano19e]. Japan [Bro14b, Gru12, Jin17, Moo15, Oka17a, CR19, Fru14, Ona17]. Japanese [Oka17a, Oka17b, Tsu12]. japonaise [Has16]. Jared [De 19]. Jarrige [Cas19, Pan18a]. Jasanoff [Wil19]. Jason [Sim17b]. Java [vdHP13]. Javanese [vdHP13]. Jazz [Est10]. Jean [BBB^{+}19, Sch13b]. Jeanne [Jac14]. Jeff [Arc18]. Jeffrey [Bar16a, Sla18, Sin15]. Jenifer [Cor14]. Jenkins [Hok17]. Jennifer [Ran15]. Jenny [Moo16a]. Jens [Sop14]. Jeremy [Ben17, Dae19, Rol18]. Jervis [Cor10]. Jesse [Bar18]. Jessica [App17, Dow16, Mer17a, Moh15]. Jet [Mes16, Rol17]. Jiat [McL17]. Jiat-Hwee [McL17]. Jill [Ban16]. Jimenez [Col15]. Jimmy [Nov14]. Jo [Car13b, Sta17]. Joachim [Bra15a]. Joan [How15b]. Joanot [PT13]. Jobs [Fis15]. Jockeys [Fea16]. Johan [vM16]. John [Bye13, Caf18, Cro14a, De 16, Gam15, Hog14a, Kni13, Lif18, Mac14, Moo16b, Neu14a, Shi17, Tel13, Tom17, Zia18, Cow10, Olw10, Smi14a, Low18, Mou10]. Johns [Kra11]. Johnson [OS12, Bar16b, Mod17]. Johnston [Man15, Olw13]. Joined [Rub18]. Joining [Hey10, Smi14b]. Jolivette [Dob16]. Jon [Koe18, Low16b]. Jonathan [Cum13, Fin13, Hle19, Hor14, Phi14, Pop16, SH19a, Wel15, Wil14a]. Jones [Hac15, Pfa13, Bro15, Cer16, Fre15, Hen18a, Kli19, Mck13a]. Jones-Imhotep [Kli19]. Jørgensen [Smi14b, Sam13]. José [Hea15]. Joseph [Fra13, Sch13a, Shu15, Wea16, Wil17a, BM19, Mei19]. Josephson [Hög15]. Josh [Sch14b]. Joshua [Fer16]. Journey [Car14b, Lau17, Zyl11a]. Journeys [KL19]. jours [Has16]. Joy [Gru16b, Ung14, Jac11b]. Jr [Cor10, Sti18]. Jr. [Kra10, San17]. Juan [Mur12]. Judd [Wu19]. Judging [Sla12b]. Judgments [Bea10]. Jules [Bai12]. Jules-Louis [Bai12]. Julia [Pan18d]. Julian [Gal11]. Julie [Fre18, Hog14a]. July [BBŁT15, BBB^{+}19, BGW^{+}17, ERST13, GKMV18, KLŁ14, Mis11b, Seg14, Ano10i, Ano11g, Ano12g, Ano13f, Ano14f, Ano15f, Ano16f, Ano17f, Ano18f, Ano19h]. Junctures [Ste11]. June [Ano17b, Swa13a]. Jungle [Bea11]. Junkers [Uzi19]. Junkyards [Mor17]. Jurek [Wei15]. Justice [Hir16, Mer17b, Mit18, Wol15, Man19, Zim13]. Jütte [Bor17].

Kaijser [Gif17b]. Kaiser [Her17a, vM16]. Kaiserfeld [God16a]. Kakoudaki [Ham16b]. Kalba [Pop19]. Kalic [Bur13]. Kaplan [Low13, Ste19]. Kapsch [Chu15]. Kara [Kan13, Sch14b]. Karen [Bed18, Hac15, OS12]. Kargon [Gan17b]. Karim [Kai17]. Karin [Bra15b, Mil13a, Sca14, Wil14b]. Karush [Cas16b]. Karvellas [Mar17a]. Kassabian [Kat14]. Kastner [Mai16]. Kate [Car14a]. Katherine [Dow16, Fra18]. Kathryn [Ham16a]. Kathy [Col15]. Katz [Mat17c]. Kaufman [Kru14]. Keeling [Sch14b]. Keely [Mac19]. Keiko [Gol15]. Keith [Abb15, De 16]. Keller [New18]. Kellison [Joh17]. Kelly [Cor16, Cut13]. Kemp [Swa19]. Kendall [Gua13]. Kendra [War14]. Kenna [Sum17]. Kennedy [Kra19]. Kenneth [Eva15]. Kenny [Par17]. Kent [Gro15]. Kente [Got12]. Kenya [LL15]. Kenzo [Oka17a]. Kept [SH15]. Key [Tim17]. Keyboard [Ava15]. Keywords [Lon12b]. Killer [Ald16a]. Killing [Gra15, Bar10]. Kim [Gol15, Sul19, Wil19]. Kimberley [Dou19]. Kimberly [De 12]. Kimura [Ona18]. Kinchy [Sto13a]. Kind [Ala12]. Kinder [Dou19]. Kinds [Sis12]. King [Fei12, Fro14, Kra11, Mou10, Nee18, Coc16]. Kingdom [Cit12, Gud18, Dav13]. Kings [Ole17a]. Kinney [Rol18]. Kinsey [Pet16]. Kinship [Nor15a]. Kinzley [Wu19]. Kirby [Hep16]. Kirschenbaum [Kur19a]. Kisses [Bij13]. Kitchen [Jør10, Pen18, Ree19]. Kjetil [Bre18]. Klangmaschinen [Pet12]. Klaus [Tar15]. Klein [Har17, Ole17b]. Knights [Mor15]. Know [O’R17]. Know-How [O’R17]. Knowing [Dus18, Jac11b, Ver15]. Knowledge [De 10, God15, Hah15, Hau19, Hea13b, Her17a, Hog18a, Hor18, Kri19b, Kru16, Kui19, Lan17, Mal17, PR16, Per12, Set16, Wan13, Zia18, Pet16, Ver15, Åbe15, Bij11, Gal16, Gan11]. Knowles [Cut13, Luc13]. Knut [Hic14]. Ko [Ham18d]. Koenraad [Cam19]. Koerner [Han14]. Kofferradio [Kre10]. Kolb [Wes16]. Koloski [Eva16]. Koloski-Ostrow [Eva16]. König [Gif17a]. Konsum [Die12]. Konzepte [Die11]. Korea [Moo16b, Cho17, Jeo10, Lee13, Sch12a]. Korean [Cho18, DiM10, Lee13]. Kosnik [Boo17]. Krajewski [Wor16b]. Kranakis [vdV14]. Kreiss [Her17b]. Krementsov [Jos13]. Krige [Neu14a, Zia18]. Kristin [Mai16]. Kristine [Ber18]. Krupp [Gif15a]. Kruse [Åbe15]. Krysko [Ste14]. Kuchinskaya [Kru16]. Kuhlberg [Hul16]. Kultur [Kre10]. Kultur- [Kre10]. Kumar [Sah14]. Kun [Sch14b]. Kurgan [Lig14]. Kurt [Ric13]. Kushner [Pan15].

L [Ang15, Bil15, Blo13, Cla13, Cro14b, De 15, Deu17, Fai16, Fre15, Gra15, Gud13, Lif18, Roh19, Ste16b, Swa19]. Laak [Vai19]. Lab [Cha11a, Sch12c]. Labor [Bar10, Kar18, Łot16, SDJ19]. Laboratoire [Rey12, But12]. Laboratories [Bea11]. Laboratory [Cur13, Kir12, Mes16, Wes15, Haf11, Wes10b]. Labs [Tay17]. Lachmund [Sop14]. Ladies [De 11b]. Lady [Ste10b]. L’Aérostation [Gil11b]. LaFollette [Sch13d]. Laguës [Sch19]. Lake [Bea19a, LeC11, Sut19, McN14]. Lakes [Lan11, McN14]. Lakhani [Kai17]. L’amphithéâtre [Pan18b]. Land [Kim12, Ziz19]. Landings [Con15]. Landmarks [Sim17a]. Lands [Hay18]. Landscape [Dio12, Jac12a, Man15, Sjö11, Amm14, LeC17, Van18, Blo13]. Landscapes [Gor15, Sch13c, dV12]. Landschaften [Spe11]. Laney [Col16]. Lang [Sum17]. Langdon [Cor10]. Langston [Sut19]. Language [Bil15, Edw12a, Eis10, Fis13, NPA14, Tka19]. Lanzara [Coo17]. Lara [Tim17]. Large [Can17, Lag17, McN14]. Large-Diameter [Can17]. L’Arte [Ber17]. Laruelle [Ihd16]. Laser [Sla11c]. Last [Bow10, Cam16, Kou10]. Late [Low10, New10, Smi10a, Tya13, Ver12, Ans17, Fru14, Wol11b]. Late-Nineteenth-Century [Wol11b]. Lathers [Wei13]. Latin [Gua13, Hoo10]. Latour [Fis13]. Lattice [Dre10]. Launius [Rol10c]. Laupahoehoe [Meh10]. Laura [Lig14, Mue15, Pop19]. Laureate [Sid12]. Lauterbur [How15b]. Law [Bou19, Cer16, Día16, Kru13a, O’R17, Yos16]. Lawns [Whi10]. Lawrence [Kno14a, Olw10]. Laws [Mac19]. Lawyers [Bea10]. Layering [Gro19]. Layout [dSS19]. Layton [Kra10]. Leach [Hah15]. LeAnn [Nov11]. Leap [Mir14]. Learn [vMS15]. Learned [Wol11a]. Learning [BM14, Kno14b, MVÁ16, Rob17, Gar13, Jaf10, Oka18, Rud13]. Leaving [Col12a]. Leben [Kai10]. LeCain [Lub18, Shu11]. LeCavalier [Bar18]. Lectures [Ber11]. Lee [Hah15, MM11, Fro14]. Leech [Mun19]. Legacies [Sah13a]. Legacy [Gal17, Smi17, Sol11, Swa13a, Tar14, Zim11, Bro13, Kni13]. Legal [Cum12, Mey19]. Legendary [Gif15a]. Leggett [Wol14]. Leigh [Moo16a]. Leisure [Tho10b, Har11]. Lem [Pic14]. Lemov [Var18]. Lenin [Sid12]. Lennon [Hus19]. Lens [Hal16]. Leo [Arn18, Mer12]. Léon [Fea19]. Leonardi [Wis17]. Leonardo [CG16]. LERTS [CC18]. Leslie [Cas13, Cro18]. Less [Kit12, Ros12a]. Lessing [Epp15]. Lessons [Cor13b, McD15, Nan12]. Letter [Uye11]. Letterpress [Ste16a]. Leviathan [Bat14, Osl10]. Levittown [Mil12b]. Lewis [Zar13, Smi10b]. Ley [De 19]. Libby [Spr19]. Liberal [Geo15, God15]. Liberalization [Oka17b]. Liberation [Aga18, Dru12]. Liberty [Lau11, Lev13a]. Library [Bru19]. Lie [Gol14a]. Lies [Pom10]. Life [Ans17, Ben17, Cer16, Com12, Cre17, Dow15, Fea16, Ham18d, Hog18a, Hor14, Ihd18, Kir12, Kra10, Kru15, Mai16, New10, Pac18, Pan18d, Pet13, Ros12b, Sch13a, Scr12, Sid12, Sid18, Str13, Swa13a, Tob13, Tom14, Col17a, Low18, Sin15, Bed18, Cha11a, Gan10, Hag15a]. Life-Worlds [Mai16]. Lifset [Gor17, Nee18]. Light [Dow18, For10, Hin11, Joh11a, Nye14, Sla11c]. Lighting [Ken19]. Lightning [Goo11]. Lights [Gar13, Jac13, Mis11a]. like [Lan17]. Liliane [Lev17, Ora15]. Lillard [Caf18]. Lillian [Wes16]. Lim [Pua17]. Limestone [San12]. Limitless [Tho15]. Limits [Jac11b, Kle16]. Linda [Rol16]. Linden [Cha11a]. Line [FP16, For11, Pos15, Kin11]. Linear [Ona17]. Lines [Til19]. Ling [Aga14]. Linguistic [NPA14]. Linguistics [KL19]. Link [Jeo15]. Linker [Car13c]. Linley [Kre14c]. Lino [And15]. Linthead [Ham10b]. L’invention [Sch19]. Lionel [Pan18b]. Lipset [SP17]. Lisa [Ver13]. Lisbon [Jen13, SC17, dSS19]. Listener [Han13]. Listening [BT19, Bra15b, Kur19b, Sch14b]. Literary [Ket16, Kur19a]. Literature [Coo12a, Ham16b, Kre14c, Dup15, Cry11, Sla11a]. Lithography [Ste16a]. Little [Ste16b]. Liudger [Pop13b]. Live [De 10, Fre11]. Lived [O’H10]. Lives [Bow16, Coo12a, Hle19]. Living [Low13, Mir19, Pat15, Rei12, Whi10, Dus18, Dob10]. Llave [PT13]. Lluís [PT13]. Local [GKMV18, Hir12, HS13, Ros10, Shu15]. Localities [PR16]. Lock [Tim17]. Lockport [Sch10c]. Locksmith [Hau19]. Locomotive [Bug13]. Logel [Koe18]. Logemann [De 15]. Logic [Gri12, McD10, Abn19]. Logistics [Bar18, McC14a]. Logsdon [Shi17]. Lomonosov [Coo12b]. London [Cro18, Rab13a, Ros12b, Hag14, Swa13a, Tom15, Van16, Tom17]. Londres [Ora15]. Lone [McC15]. Lonely [Gra16]. Long [Las18, Neu14a, Ben14, Ros13]. Look [Sla12b]. Looking [Cow10, Kra10, Roh19]. Lorance [Seg10]. Lords [Sat13a]. Lordship [Bon16]. Loren [Gra16]. Loretta [Seg10]. Losses [Rhe14]. Lost [Fei12, Fir12, Pet14, Tin19, Var18, Arc18]. Lot [BM14, Dug11]. Louder [WH11]. Louis [Ano19c, RS14, Bai12]. Louisiana [Miz11]. Love [Bra10, vMS15]. Loved [Bra10]. Lovely [Sta12]. Low [Man13]. Low-Grade [Man13]. Lowell [Bro11]. Loyalties [Rhe14]. LP [Gop15]. Lucas [Bon16]. Lucien [Bat14]. Lucsko [Mor17]. Lukens [Eis10]. Lumières [Gil11b]. Lunar [Wei15]. Lund [Wor16a, Coo11c]. Lundy [Big16]. Luxury [Wri12a]. Lynn [Rab13a]. Lynne [De 12].

M [App14, Bed18, Bek18, Caf18, Car15, Cla13, Coc17b, Dav17, Gru16b, Har19, How15b, Hug15, Joh13, Leu14, Mat17c, Mes16, Pan15, Rae16, San17, Shi17, Spu13, Wis17, Gro19]. M-1943 [Gro19]. M1 [Zel11]. MacDonald [Las18]. Machina [Wos17]. Machine [Big16, BH17, Cro16, Día16, Eli12, Gal15, Gam11, Gol14a, Hel13, Kim13, Kum19a, Lae16, Lee14, LM19, Luc10, Nah13, Pop11, Sam13, Sim10, Sla11a, SSH18, Vai19, Pop16]. Machine-Age [Pop11]. Machinery [Hah19, HPR13]. Machines [BM19, Cha14, Coo12a, DL18, Fra13, Gal16, Gri12, Hal12, Ham17, Hen19, JI16, Kel10, Kli17, Pet18, Pra19, Rei12, Zan13, Car11, Luc18b]. Machinists [Bow16]. Mackay [Mou10]. Macola [Hac15]. Made [Fei12, Hur15, SH19b]. Madhavan [Fri16]. Magaudda [Col17b]. Magic [Kan16]. Magical [Gra15]. Magnificent [Gam11]. Maharaj [Neu14a]. Maher [Col19a, Nov14]. Mahony [Wea18]. Mainframes [Cor16]. Maintenance [FR18, Pac18, RV18]. Make [DL18, Sen17, Ber18, Mat17c]. Maker [Tur18]. Makers [SV19, Rio11]. Makes [Sen17]. Making [Abb11, And15, Bar12, Bar16b, Bea11, BL12, Cas16b, Cha14, Col16, De 17, Dou19, Fir17, Fle19, Har16, Hög14c, Joh11b, Les10, Man15, Mat17a, Mit12, Muk19, Ore11, PR16, Pra13, Rob10b, Sah13b, Sla11c, Sla19, Smi14a, Sto19, Tom17, Ver15, War10, Dow10, Alp16, Arc19, Cas16a, Cha11a, Cum12, Edw17, Mor14c, Rol17, Sam13, Swa13b]. Malaria [Sta10a]. Malcolm [Dow17, Eva15, Hog14a]. Malin [Man19]. Mall [Mat17a]. Malleability [JI16]. Malm [Qui17]. Man [Bra10, Cha18b, Fre11, Tur10, Uzi19]. Management [BM12, Gat18, HPR13, Nee17, Smi17]. Manager [Gif17a]. Managers [Mal17]. Managing [Ber19, Sla18, Yos16]. Manchester [Ham13, KLŁ14]. Manhattan [Mul16b, Kem12, Ric13]. Manoranjan [Lau17]. Mansell [Luc18a]. Manual [PT13]. Manuel [Bar16a]. Manufactured [Tar14, Dow18]. Manufacturers [Bar14a]. Manufacturing [Bet14, BL12, De 12, Ros10, Tin18, Wes10a]. Manus [Wos17]. Manuscripts [Lon11]. Many [Día16]. Mao [Hu14]. Maoism [Mul12a]. Map [Com12, Ham18a, Lin10, Lon12a]. Mapping [Jac10b, Lig14, MM11]. Maps [Jac11a, Lan15, Mor14b, Hor17]. Marc [Bye19, Cha18b, Dow18]. Marcel [Ger16, Ran15, Sch13d]. March [OS12, Oka17b, OM17]. marchands [Ora15]. Marco [Hor17]. Marconi [Sat10a, Cha18b]. Marez [Dry19]. Margaret [Kre14c, Low14]. Maria [Cam16, Vai19]. Marie [Tin19, Wei13]. Marilyn [vdH17]. Marine [Leg11, Smi14a]. Marisa [Lan18b]. Maritime [Wil10, Wol14, Rol10b]. Marítimo [Hea15]. Mark [De 17, Gan19, God15, God18, Hul16, McN14, And14]. Market [Gol14b, Gru12, Gud13, Haf11, Ken14, Lau11, Ree18, Tya13]. Marketing [Fra16, Mae13, Wei15]. Marketplace [Hog13a]. Markets [Hoc10, Sid18]. Marklund [Anb18]. Marks [Tim17]. Markup [Lon12b]. Markus [Wor16b]. Marquetry [Edw12b]. Mars [Con15, Lan17, Mes16]. Marsha [Gar13]. Mart [Kle10]. Marta [Kat14]. Martian [Jos13]. Martin [Cor16, Gro18, Mei13, Olw14, Sla11b, Sla19, Sta14a, Wea18]. Martini [Ham14a]. Martinville [Fea19]. Martucci [App17]. Marvel [Ryd11]. Mary [Sta17, Mun16]. Marzec [Oat16]. Masco [Wea16]. Masculinities [Gue13]. Masculinity [Cro16, McC17, Ste16a]. Mask [Ham18b]. Mason [Joh17]. Mass [Mar19, Sat13b, Wes10a, Zar17, Shu11]. Massachusetts [Abb15]. Mastering [Luc13, Cut13]. Material [Dob10, Sen17, Sla15, Ste16a, Ver15, Vol17, Wen17, Wol11b, Zim13]. Materializing [DB13]. Materials [Fle19, Kui19, Rud12, Haf11]. Mathematical [San11, Ver12]. Matheson [Gal15]. Matter [Che19, Edw17, Kle16, Lub18, Cre17]. Mattern [Jac19]. Matters [DGS16, Ves19]. Matthew [Bea19a, Boc16, Cas16b, Ert16, Ger13, Hög14c, Kur19a, Mac18, Mey19, Reu14, Eva15]. Matthias [Wea18]. Mavhunga [Edw18, Hoa15]. Max [McD17]. Maxwell [War10]. May [Hag14, HHH12, Mod17]. Maya [Bhi10, Mat17b]. Mayan [Mat17a]. Mayor [Hen19]. Mays [Gal15]. Mazur [Hau15]. Mazzolini [Zim13]. McClellan [Hel13]. McCormack [Kim19]. McCray [Her17a, Tho15]. McCullough [Dow17]. McDermott [Cow10, Cow12]. McElwee [Dér18]. McLaren [Bjö13]. MCM [Bro12a]. MCM/70 [Bro12a]. McNeill [Jør17]. McNeur [Mul16b]. Mean [Coo10a, Edw18]. Meaning [Cum13, KO13]. Means [Dru15]. Measures [Ziz19]. Mechanical [Cur13, Hau19, Pre10b]. Mechanics [Kre14b]. Mechanisation [Lae16]. Mechanism [Hen18a, Kan16]. Mechanization [Edw12b, Kir12]. Mechanizing [Mac15]. Media [Arc10, Boo17, Bor10, Bow12, Can15, Col19b, Cum19, Dow16, Fir12, Fis15, Gan17a, Hoo19, Ket16, Low16b, Mac13, Mul18b, OS16, Pop11, Wri12b, Col17b, Jac19, Aga11, Arc13a, Cum11a, Kre14c, Rob10c, Ste11, Wij12]. Mediated [Sca14]. Mediation [Rae16, Ven19]. Mediators [Hog13a]. Medical [Car11, Fre15, Ham17, Hog13a, Jon13, Mae13, Zan13, Hey10]. Medicina [PT13]. Medicine [Bra13, GKMV18, Gil11a, Hai15, Leu14, MSPLC16, Rab13b, Rob10a, Tim17, Dae19, Bar12, The10a]. Medientechnik [Pet12]. Medieval [BM12, Bon16, Fig10, Haf18, Lon11, San12, Wal18, Kan16]. Mediterranean [Hen18b]. Medium [Tuc12, Spi15]. Medosch [Gla17]. Meerman [Wei15]. Meet [Mun16]. Meeting [Ano10e, Ano11c, Ano13b, Ano14b, Ano15b, Ano16a, Ano17b, Ano18b, Ano19c, KNSZ16, Soc12]. Meets [Dob10, SW15]. Megalopolis [Pla10]. Megan [Cor19]. Meganck [Sla11b]. Megaprojects [Cam19]. Megascience [Mod10]. Mehos [Car15]. Meikle [Kas17]. Mekong [Ert12]. Melosi [Olw14]. Melting [Clo11]. Melts [Kou10]. mémoire [Sch19]. Memorial [Gru16b]. Memories [Fir17, Meh10]. Memory [Boo17, Jos16a, Pre13, Sza15, Sch19, Low16b]. Men [Rob10b, Woo19]. Menagerie [Tya13]. Mendel [Har13]. Mentoring [Hec14]. Merc [Mac15]. Mercantile [Mac15, Whi11b]. Merchants [Rob12]. Mercury [Hos13, Man15]. Merit [Wel17]. Merritt [Sim10]. Merry [Pin10]. Merry-Go-Round [Pin10]. Mesman [Moh15]. Mesoamerican [Dow18]. Messinger [Sat13b]. Messy [Car14b]. Metal [Gro15, Gor11]. Metalworking [Gor11]. Metaphors [Sen17, SP17]. Méthodes [Lev11]. Methods [Mor10b, Tho16]. métiers [Pan18b]. Metropolis [Cut10, Est10, Kin14, Wos10a, Wos10b, Zyl15]. Metropolitan [Wos17]. Mexican [Bea11]. Mexico [Bea19a, Gro14, Bea19b, Spu13, Vit17]. Meyer [Gal16, Gam15, McC17]. Meyers [Ver15]. Mi [Sul19]. Mialet [Ihd13]. Michael [Dis15, Hir14, Hui16, Kil15, Lak13, Mey17, Spi15, Ste14, Wes16, dS17]. Michaela [Hug15]. Michel [Sch19]. Michelle [Dru15, Pan16]. Michigan [Sch10a]. Micro [Wen17]. Micro-Study [Wen17]. Microchip [Rio11]. Microcosm [Bow15]. Microlessons [Coo15]. Microphone [Olw18].

Microscopy [Joh13]. Microworld [Cer13]. Midcentury [Kat11]. Middle [Jac12a, Mor10b, Pat19]. Middle-Class [Pat19]. Midi [Gra11]. Midnight [Ros11]. Midwest [BH17, Sim17b, DeB11]. Mieczkowski [Neu13]. Migration [Haf18, O’H10]. Mikael [Bea19b, Gag15]. Mike [Fro14, Sch13e]. Milano [CG16]. Miles [Tar15, ZT18]. Milestones [Her16a]. Militarization [Bur13]. Militarizing [Oat16]. Military [Gro19, Gru12, Lin10, Łot16, Mac18, Sim17a]. Milk [War14]. Mill [Wel15]. Millenarian [Tur18]. Miller [Col15]. Milling [Bon16]. Miltown [Hus10]. Mimesis [Kan13]. Mimura [Tsu12]. Mind [Eri15, Mit19, Sat13b, Cum11b]. Minding [Mir12]. Minds [Fri16]. Mines [Ald16a]. Minimum [Har12]. Mining [Ald16a, Bar16a, Gro14, Gro15, Gua13, Hos13, Kle16, Lan15, LeC11, Man15, Mun16, Nie12, SV19, Spu11, Spu13, Gol15]. Minneapolis [Bro12b]. Minnesota [Bar16a, CK15]. Mirko [Cor13a]. Mirror [Kra10]. Mirrors [Bij13, Dow18]. Misa [Sla18, CK15, Kon11]. Misfits [Luc10]. Misreads [Mit19]. Missile [Col14, Gai13]. Mission [SK16]. Missions [MV15]. Mississippi [Car13a, Jos16a, San17]. Misunderstood [Ryd11]. Mithen [KO14]. Mixed [OL17]. Miyazaki [Bre16]. Mizelle [San17]. Mobile [Aga14, Fel14, Gai18, Mil19, Ore11, Aga11, Sin12]. mobiler [Kre10]. Mobilités [Lev17]. Mobilities [MF16]. Mobility [McD17, Mom14, Sop11, UHW16, dSS19, Bie16, Pra12]. Mobilizing [Boc16, Div10, Hog14a]. Model [CR19, Ona17]. Modelers [Cre17]. Modeling [Tin18]. Modelling [Wea18]. Models [Cre17, De 10, Eli12, Gul19, Lan15, Mae13, Tin18]. Modern [Arc19, Bes18, Bet14, Bra12, Bre10a, Bro15, Cas16a, Cho17, Cit12, Coc13, Dae19, De 17, Est10, Fis13, Fru15, Gol15, Gue13, Haf11, Ham13, Ham12b, Hau19, Hes11, Hoc10, Hoo19, Hu14, Kar17, Kim12, Koe18, KO13, Lac16, Les10, Low13, Low18, Luc13, Mat17a, Mau11, Mit12, Moo15, Muk19, Nee18, New10, Pik15, Pua17, Ree16, Ric13, Roh19, Rol18, Sah13b, Set16, Son17, Tei11, Tho10a, Tom17, Vit17, War14, Wu19, Zar17]. moderne [Pop13b]. Modernism [Hui16, Luc11]. modernismo [Mor18]. Modernità [Bal11a]. modernité [Lan10]. Modernity [Amm11, Cha11b, De 12, Dis15, Ert16, For17, Gan17b, Pra12, Wil19, vdS19, Cha14, Gan17a, Luc18a, Ihd11]. Modernization [Ala12, Jeo10]. Modernize [Cut13]. Mody [Joh13]. Mogens [Anb18]. Mohun [Kno16]. Moldavia [Les19]. Molecular [Cre17]. Molella [Mar17a]. Möllers [Mac17, Wil14b]. Mom [Rie16]. Moment [Neu13, Pat11]. Momentous [Mul12b]. Moments [Wis11]. Moms [Ona18]. Mona [Col15]. Monahan [Mac14]. monde [Cas19]. mondes [Lev17]. Money [Sen17]. Monies [Gol15]. Monika [Cum19]. Monique [Col16]. Monitoring [Rot16]. Monk [Kru15]. Monmonier [McN14]. Monnet [BBB^{+}19]. Monoclonal [Tim17]. Monographs [Ros12b]. Monopoly [Pla16]. Montana [Mun19, Col19b]. Montel [Gui16]. Montgomery [Bye13, Bil15]. Monumental [Amm14]. Monuments [San12]. Moon [Wei15]. Moore [Cer16, Bro14b, Cer16]. Moral [Arn18, New10, SP17]. Morality [Hei15b]. Moralizing [Hei15b]. Morat [Lac16]. Morgan [Abb15].

Morning [Hog14b]. Morocco [Wil17c, Cam19]. Morris [Dal15, HV15]. Mortgage [Lig11]. Morton [Mar18]. Moscow [Fir12]. Most [Fle19, Gam11, Ham11, Hig12, Sla12b]. Mother [Buc14, Jen15]. Motherhood [App17]. Motion [Cah19, Mac17, Mai16, McA14]. Motor [Han14, Mac19, Mor15, War10]. Motorcycles [Vol17]. Motorcycling [Woo19, Tho10b]. Motorists [Cla10]. Motorola [Sla19]. Motorrad [Kre14a]. Motorway [Zel11]. Mountain [Bal11b, Mar13, Nee18, WM11]. Mountains [MM11, Gud13]. Mouth [Gat18]. Move [Cow10]. Moveable [Mul12a]. Movement [Tur18, Wil17b]. Movements [Pik15, Wen17]. Movers [Lag11]. Movie [Jac15]. Moving [Col19b, PR16, Hic12]. Moyen [Lev17]. MP3 [Cum13]. Mr. [Man13, Hu14, Ryd11]. MRI [DiM15, How15b]. Mueller [Olw10]. Mughals [Buc12a]. Mules [Bie16]. Mulholland [Sti18]. Mullaney [Bus19b]. Müller [Hug15, Bek18]. Munich [JJ16]. Munitions [Sla14]. Munns [Kra14]. Munro [Hog14b]. Murphy [Dru15]. Museion [Hey10]. Museo [CG16]. Museum [BS13, Bow15, Cho18, Hag14, Hau11, Her16a, Hor11, JJ16, Mar13, Meh10, Mor12a, Rol10b, Seg14, Swa13a, Wil13, Wos17, AH18]. Museums [Bed18, Mar18, Søn12]. Musiani [Pla17]. Music [Ham10b, Kra17, Pop13a, ZT16, ZT18, Van14]. Musical [Cah14]. Musics [Kat14]. Muskingum [Swa19]. Must [Hah17, Pre19]. Mutio [Ver12]. My [BH17, Cla10, Pac18]. Myers [Cre17]. Myles [Hog18b]. Mysteries [Hor17]. Myth [For17, Mer17a]. MythBusters [Ung14]. Mythologies [Fir17]. Myths [Hen19].

N [Abb15, Bur13, Dow16, Fra16, Hög14c, Mor17, Swa15, Wes15]. Nadia [Jac13]. Nafus [Col17a]. Nagase [Gol15]. Nagase-Reimer [Gol15]. Naked [Fin17]. Nancy [Hu14, Sut19, Swa18]. Nanny [Gol15]. Nanoethics [Mod11]. Nanotechnologies [Tho15]. Nanotechnology [Joh13]. Narrative [Sla10]. Narratives [BGW^{+}17, Kum19a, Mai12, Sid10]. NASA [Her16b, Kru10, McC14a, Mey15, Neu14a]. Natasha [Cre17, Lee14]. Nathalie [Gui16]. Nation [Bar16b, De 11a, Ert12, Goo19b, Jos16a, Kli19, Kum12, MVÁ16, Mil13b, Moo16b, Nik17, Ree16, Rob10b, Wil14a, Yan19, vM12b]. Nation-Building [Ert12, Jos16a, Yan19, Moo16b]. National [And10, Bea11, Bre18, Bus19a, Cas14, Cho18, Fai16, Gru16b, Her16a, Kri19b, Rol10b, Rol16, Sid10, Wea16, Wes10b, Wes15, Wil13]. nationalen [Tod11]. Nationalism [Ree16]. Nationalsozialismus [Lin13]. Nations [Arc12, Hau15]. NATO’s [Can17]. Natural [App17, Bed18, Jac13, Pla16, Rob13, Str11, Wu19]. Nature [De 10, Ert12, Gro15, Jen15, Jør13, Kai12, Kan16, Kli19, McL17, Pet10, Rom10, RAF^{+}11, SV19, Sti11, Sut12, Vet19, Whi13b, Eri15, Gif17b, God15, Sop14, Cut10, Ore13]. Natures [Día16, Smi14b]. Naval [Wol11b]. Navies [Wol11a]. Navigating [Yan19]. Navigation [Ham18a]. Navy [Caf18, Fel14]. Nazi [Col16]. Nazionale [CG16]. Neal [Lan18a]. Necromedia [Ger16].

Nederland [Dis11]. Need [Bil15, Bor10]. Needham [BM19, Che19, Mei19, NW19]. Neel [Dec18]. Negotiating [Kno16, McK13b, Tya12]. Neil [Col19a, Gal15, Gra15]. Nelles [Cas14]. Nelson [ZT16]. Neoliberal [Sha17]. Neon [Nye14]. Neri [Ber17]. Net [Str12]. Network [Bri10, FR18, Nik17, RS14, Tho10a, Tom11, Hug16, Arc12, Gro18, Lan18b, Wol11a]. Network-Centric [Wol11a]. Networked [Cha18b, Ste16b, Tom15]. Networking [Ste15]. Networks [Car15, Coc17b, Ket16, Kum12, Mac16a, McL17, OF18, Phi14, Bek18, Bat11, Hau10]. Neufeld [Kil15, Lif18]. Neutron [Olw13]. Never [MV15]. Newman [Spi15]. Newmont [Spu11]. News [Hug15]. Newsmaking [Dow15]. Newsprint [Hul16]. Next [Ore13]. Nice [Nel16]. Nicholas [Hos13, Mey17]. Nicolas [Fea16]. Nicole [Hug16, Zar17]. Nicolson [Hog14a]. Nikhil [Oak19, Pat15]. Nikolai [Jos13]. Nil [vdV14]. Nina [Mac17, Pan15, Wil14b]. Nineteenth [Bra13, Bro11, Fer19, Gal15, Hea13b, Kas17, Kre14c, Łot15, MF17, Mil13b, Rus11a, Wer12, Wol11b, FP16, Jep18, Mac19, Mas17]. Nineteenth- [Gal15]. Nineteenth-Century [Bra13, Bro11, Fer19, Kas17, MF17, Rus11a, Mac19, Mas17]. Nintendo [Jin19]. Ninth [ERST13]. Nishiyama [Moo15]. Nitrogen [Swa15]. Nitrogenous [Whi10]. Nixon [Shi17]. No [Ald16a, Jac11a, Ona17, Pat15, Ste16b, Lag15, Jac11a]. No-Threshold [Ona17]. Noble [Mis11b]. Nocturne [SC17]. Noir [Ran11, Sch12d]. Noise [Luc18a, Mor14c, Kra19, Sch13e, Mir19]. Non [Fru15, Kui19, Pan18c, Whi12]. Non-Asian [Kui19]. Non-elites [Pan18c]. Non-Industrial [Fru15]. Nonalignment [Tya12]. Nonhuman [Sny17]. Noninvasive [Dus11]. Nonpoint [Whi10]. Nonproliferation [Zia18]. Nordlund [Tob13]. Norman [Ten15]. Norms [Mey19]. Norris [Sti18]. North [Hay18, Kai12, Sha10, Vet19, Wro10, SDJ19, Tay14, UHW16]. Northeast [Ore13]. Northwest [Hir17]. Norton [Sop15]. nos [Has16]. Note [Ano19b]. Notebook [Eis17]. Notes [Sta14b, Dis10, Fin11]. Notice [Kat14]. Novecento [Mor18]. Novel [Sla10, Phi14]. November [Ano15b, Hag14, Soc12, Sch13a]. Nowhere [Kru13a]. Nu [Dis11]. Nuclear [Abr13, Car13a, Car14a, Dob16, Fja13, Hec10, Joh11b, Joh12, Kem12, Kin10, Kru13a, Kru13b, Kru14, Mad19, Man19, Mat19, Oka18, Olw10, Olw13, Olw14, Rol10d, Rub18, Sla12a, Wel17, vMS15, Joh12]. Nucleus [Kum12]. Number [Ano10h, Ano10g, Ano10i, Ano10j, Ano11f, Ano11e, Ano11g, Ano11h, Ano12f, Ano12e, Ano12g, Ano12h, Ano13e, Ano13d, Ano13f, Ano13g, Ano14e, Ano14d, Ano14f, Ano14g, Ano15e, Ano15d, Ano15f, Ano15g, Ano16e, Ano16d, Ano16f, Ano16g, Ano17e, Ano17d, Ano17f, Ano17g, Ano18e, Ano18d, Ano18f, Ano18g, Ano18h, Ano19f, Ano19e, Ano19h, Ano19i, Ano19g]. Numbered [Car18]. Numbers [Ros17]. numériser [Sch19]. Nützliches [Har17]. Nye [LeC17, Ken19, Pos15]. Nystrom [Lan15].

O [Ben14, Pom10]. O-Rings [Pom10]. Obach [Wil17b]. Obelisk [Pet11]. Obert [Hle19]. Objects [Ber14, Sta17, Mac17]. O’Brien [Goo19b].

Obscured [Rob12]. Observation [Ham12a]. Observatories [Dic10, Rus11a]. Observers [Jen13]. Obsession [Luc18b, Olw10]. Obsolescence [Fri13]. Occult [Tuc12]. Occupational [Kin11]. Occupations [Wis17]. Oceans [Har19]. Octave [Bug13]. October [And14, Ano10e, Ano11c, Ano13b, Ano14b, Ano16a, Ano18b, Ano19c, OS12, Swa13a, Ano10j, Ano11h, Ano12h, Ano13g, Ano14g, Ano15g, Ano16g, Ano17g, Ano18g, Ano18h, Ano19i]. Odd [BM14]. Odd-Lot [BM14]. Oddi [Ver12]. Oddities [Wei13]. O’Donnell [Hjo16]. Ofer [HV15]. Off [Gar13]. Offer [SV19]. Offset [Ste16a]. Ogle [Fru17]. OGM [Har13]. O’Gorman [Ger16]. Oil [Cas10, Cas16a, Cit12, Col12c, Hög13b, Jør13, The10b]. Okinawa [Fel14]. Oklahoma [Gru16a, MM11]. Okuda [Oka17a]. Old [Ald10, Hel13]. Oldenziel [Gag15, Paj16]. Oleksijczuk [McA14]. Olfactory [SSH18]. Olga [Eva16, Kru16]. Olmstead [Deu17]. Olmsted [Abb15]. Olof [Wor16a]. Olshin [Hor17]. Olson [Smi17]. Once [WM11]. One [Ald16a, Sla12b, Ros12a]. Ongoing [BSA16]. Onion [Rud18]. Online [Mal17, Sch12a, Dug11]. Ontario [Hul16]. Open [Bor10, Eri15, Kai17, Pri19]. Opening [Whi14]. Operations [Gri12, Wea16]. Opiate [Cow10]. Oppenheimer [Kru15]. Opportunity [BE13, Pet14]. Opposition [HS13]. Optical [Sla14]. Optics [Bau11]. Oral [Gru16a]. Orange [Ham14a]. Orchestrating [Mom14]. Order [Car14b, Cow10, Kle10, Vos16]. Ore [Gro15, Lan11, Man13, Nie12]. Organic [Ham16a, Wil17b]. organique [Lev11]. organisation [Fox11]. Organisms [Coc17a]. Organization [Mat17b, Pet16, See15b]. Organizational [See15b]. Organizations [vM16]. Orgeron [Gar13]. Oriented [Hir11]. Origins [Arc10, Bey12, Cas13, Cor12b, Cur12, Fri13, God16b, Hoo19, Lag13, Las18, Mar11b, NPA14, SR19, Ste15, Tay17, Wol10, Lev11]. Orleans [Ole17a]. Ornament [For17]. Orvell [Tar15]. Ostrow [Eva16]. Other [Mey17, Rob10c, SP17, Kit12]. Ottoman [Whi13a]. Ottomans [Buc12a]. Our [Car18, Dus10, JJ16, Wil14c, Bus19a, Dus18, Kru13b]. Outer [Lau13, Wei13]. Outrages [Fri13]. Outsourcer [Kum16]. Over/Through [O’D12]. Overpotential [Hög14c]. Overseas [Hel13]. Overwhelm [Mor10a]. Owens [Mal17]. Own [Kru13b]. Owns [Bow15]. Oxford [Mil13a, Tay14].

P [Kim19, Kra14, Lau17, Mar17a, Mey18, Moo16b, Nye15, Oat16, Sch15b]. Pace [Pik15]. Pacific [Whi14]. Packaged [Fra16]. Pacyga [SH19b]. Pál [Łot16]. Palazzos [Wil17a]. Pamela [Ben14, Dér18, Ver15]. Panacea [Hög14c]. Panama [Car16]. Pang [Mun14]. Panic [Hea12]. Panoramas [McA14]. Panoramic [McA14]. Pants [Mas17]. Paolo [Col17b]. Paper [Sen17, Bes13]. Paperclip [Bus19a]. Papermakers [Wu11]. Papermaking [Man18]. Papers [Rhe14]. Paradigms [Nor15b, Vai13]. Paradise [Ret16]. Paravel [Bat14]. Parenthood [Pre19]. Paris [Haf18, New18, Rey12, Zan13]. Park [Cas14, Bed10]. Parke [Car14b]. Parker [Cas16a]. Parkin [Fra18]. Parks [Ver13]. parlementaire [Bai12]. Parliament [Gul19]. Parr [Jac11b]. Parsons [Leg11]. Parthasarathy [Sid18]. Participation [Cor13a, Hay14, Maa16, Pet19]. Passage [Gat18]. Passenger [vR12]. Passes [SS13]. Passing [Sta10b]. Passwords [Hus19]. Past [BBB^{+}19, Bow15, Buh17, Col16, Lon10, Sca14, Wil11, Lub18, Mor17, Wil14b, Wri12b]. Pastoral [Jac12a, Amm14, dV12]. Pasts [Wal10]. Patent [Coo10a, Sid18]. Patentees [Rat12]. Patenting [Mer12]. Patently [Joh15a]. Patents [Bea10, GR10]. Path [Joh13, Lon15]. Paths [Har11, Vee10]. Patients [Hog18b]. Patricia [Wal18]. Patrick [Eva18, Hei15a, Her17a, SH15, Tho15]. Patterson [Tay17]. Paul [Deu17, For17, Gru14, Hei15b, Hir17, Hög15, Rhe14, Ros11, Spr19, Ves19, Wis17, Olw18, How15b]. Paula [Vai19]. Payment [Cor12b]. PC [Bro12a]. Peace [Moo15, Sat13b, vM12b, Ham18b, Yos16]. Peaceful [Kru14]. Peacetime [Kat10]. Peanut [Boy16]. Peasant [Fin13]. Peasant-Friendly [Fin13]. Peasants [Bea11]. Peavitt [Ree19]. Peckham [Leu14]. Pecos [Eva18]. Peddling [Lai12]. Pencil [All11]. Penner [Mol15]. Pennsylvania [Gor15, Sch14c, Chu15, Mil12b]. People [Ber15, Dér18, Gag15, Lad14, Sah13b, Ham16b]. Perception [Aid19, ZT18]. Perceptions [Neu14b]. Peréz [Lev17, Ora15]. Perfect [Sch12b, Dow10]. Performance [Vin15]. Peril [Dru12]. Perilous [Miz11]. Period [KO13]. Periodic [Bla10]. périodiques [Pan10]. Peripheral [Tay17]. Peripheries [BSA16]. Periphery [Dio11, HNG16, MVÁ16, MSPLC16]. Perishable [Ham10a]. Peronism [Mur12]. Perpetuity [Coo10a]. Persistence [Rie16]. Personal [Eis17, Fra13]. Perspective [Kui19, Ona17, Wij14, Cum12, Neu11, Sch11d, Ste10a]. Perspectives [EW15, Gol15, Jaf10, Mor10b, Sha10, Wad16, Wor16a, Zim17, Mai16, Bre10b, Fee11]. Persuasion [Bij12]. Pesticide [Gor12]. Pesticides [Hay18]. Pestre [MS16]. Peter [Bra15a, Hei15b, Hep16, Jør17, Kin14, Mai16, Pul16, Rud13, Coo11c, Sov16]. Peter-Paul [Hei15b]. Peters [Nik17]. Petrick [Alp16]. Petroleum [Jør13, Whi11a]. Petroski [Mit11, Ten14]. Petrostate [Jos11]. Peutinger [Lon12a]. Phalkey [Ram14]. Pharmaceutical [Tob13, Tya12]. pharmaciens [Lev11]. Pharmacology [Mer17b]. Phenomenon [BH17]. Philadelphia [Tom14, Wil17a, Zyl15, Ano18b, Zyl11b]. Philip [Hei15a, Pra13]. Philipp [Bes13]. Phillips [Nel16]. Philology [Hus19]. Philosophical [Obe14]. Philosophy [Vos16]. Phoenix [Pra13]. Phonautograms [Fea19]. Phone [Sin15]. Phones [Mil19]. Phonogramm [Kur19b]. Phonogramm-Archiv [Kur19b]. Photographers [Tuc16]. Photographic [Zyl11a]. Photographs [Mar18]. Photography [Bra13, Edw17, Lim19, Mac13, Pan16, Pan18d, Pul16]. Physicians [Mey18]. Physics [Col14, Joh11a, Mac19, Mod10, Wes11, Neu11]. Physiological [Hui16]. Picture [Kon10]. Pictures [Mor14b]. pièce [Ora15]. Piedmont [Ham10b]. Pierre [Has16]. Pierre-Yves [Has16]. Pietruska [Roh19]. Pietz [Son17]. Pigeons [Bie16]. Pigs [Coc17a]. Pilar [Lev17]. Pill [Dru12, Pre13, Bea11]. Pills [Hus10]. Pilot [Gai19, Jeo15]. Pilots [McC17, Mey18]. PIMBY [BH17]. Pinch [Mil13a]. Pink [Mai12]. Pinson [Cla13]. Pioneer [Hok17, Joh15b]. Pipe [Can17]. Piracy [Kra11]. Pirate [Kra11]. Pit [Mun19]. Pitfalls [Bix18]. Pittsburgh [Ano10e]. Place [Cor12a, Miz11, Wal10, Pur17]. Places [Nem14, Mar17a, Mar17b]. Plagues [Whi11a]. Plains [Hur15]. Planet [Lau17]. Planned [Pre19]. Planning [Leu14, Ric13, Tin18, Tka19, Tsu12]. Plans [Moo16a]. Plant [Cha12, Fin13]. Plantation [Big16, Gud18]. Plantations [Mac16a, Sah14]. Plants [And10]. Planung [Pop13b]. Plateau [Wal10]. Play [Wac17, Blu12]. Playback [Dow10]. Playgrounds [Wac17]. Playing [Caf18, Goo11]. PlayStation [Wac17]. Please [BH17]. Pleasure [Har11, Luc11]. Pleasures [Fra16, Sch11b]. Plows [Whi11a]. Plowshare [Kru14]. Plutonium [Car14a, Mad19, Ham11]. Plutopia [Car14a]. Podolsky [Li16]. Poe [Hle19]. Pohl [Hug15]. Point [Koe18, Sto12, Whi12]. Pointe [Lae15]. Points [Bat11]. Poisoned [Blu12]. Polar [Clo12, Cro16]. Polaroid [Pul16]. Police [Vos10]. Policing [Jac10b, Rus11b]. Policy [Ala12, Bal11b, Col13, Deu17, Gor17, Hir11, Hir12, Kru13a, Lif18, Mit11, Oka17a, Ros10, Sca11, SV19, Str10, Web15, vdH17, Gru12]. Policy-Oriented [Hir11]. Political [Arn18, Bea19a, Ber15, Coo11c, Cur12, For10, Mac16b, Pla16, Wij14, Wor16a, Aga18]. Politics [And10, Bor10, Boy16, Car16, Cer11, Col13, Cru19, Eli12, Eri15, Fin17, God18, Gor12, Hah15, Ham14a, Ham16a, Ham14b, Har16, Her17b, Hir12, Hug15, Kim12, Kru16, Maa16, Mar11b, McC14a, Neu11, Ona18, Pre13, Pur17, Ran14, Ros12a, Ros13, Sid18, Sin15, Sop14, Sto13a, Var11, Wil14b, Wil17c, Lig14]. Politik [Die12]. Pollution [Rot16, Whi10]. pollutions [Cas19, Rey12]. Polo [Hor17]. Pomeroy [Gai18]. Pomfret [Leu14]. Ponentina [Cam16]. Ponts [Gui16]. Poole [Gai15]. Popp [Ken14]. Popular [Kra17, Rud18, Sch12d, Wei13]. Popularization [MSPLC16]. Popularizing [Dio11]. Porsche [Vol10]. Portable [Hen18a, Col19b]. Porter [Arc18]. Portland [Ano14b]. Porto [BGW^{+}17]. Portugal [BGW^{+}17]. Posed [Lap12]. Post [BF19, Bow15, Fei12, Nan12, Pra19, Rub18, Ham14b, Ore11]. Post-Cold [Rub18]. Post-Disaster [Nan12]. Post-World [Ham14b]. Postcards [Amm14]. Postcolonial [Sto13b]. Postgenomics [Wad16]. Posthuman [Rae16]. Postindustrial [Eis17]. Postmodernism [Kin10]. Postphenomenological [Fee11]. Postprint [Dow16]. Postwar [Amm12, De 15, Hal16, Les19, McC17, Mor14a, O’R17, Oka17a, Oka17b, Par17, Pat19, Pen18, Rus12, Van18, Zar17]. Potosí [Zim17]. Potter [Jon15]. pouvoirs [Har13]. Pouzin [RS14]. Power [Boc16, Bon16, Bor17, Bro15, Car13b, Hau10, Hec10, Hir12, Hir17, Hög13b, Hög14c, Hul16, Jin19, Joh11a, Jon10, Jos11, Kap19, Kat11, KO14, Kum12, Ler10, Lif18, Man19, Mav10, Mor14b, Oka17a, Oka17b, Par18, Pur17, Qui17, RAF^{+}11, Smi14a, Wan10, Wil17a, Wol11b, Sim17a, Wil19, Hir14, Nee18]. Powered [Hun16, Sim18]. Powerful [Miz11]. Powering [Big12, Goo19b]. Practical [Hir14]. Practice [Bow15, Car11, Coo17, Cos17, Lon10, Nem14, Pur10, San11, Åbe15, Dis15, Sca11, Vos10]. Practices [Coo17, Hri16, Ste15]. Practitioners [OL17, Ben14]. Prairies [Sim17b]. Prakash [Sah14]. Prasad [DiM15]. pratiques [Lev11]. Pratt [Shu15]. Precise [Sla11c]. Preda [Mir19]. Prediction [Roh19, Wea18]. Predictive [Mul12a]. Premodern [Lon10]. Prenatal [Hog13a]. Preparations [Caf18]. Prescott [Hog14b]. Present [Pra19, Wil14b, Gua13, Lae16, Ala12]. Preservation [Hoo19]. Presidents [Bur13]. Press [Joh16]. Presse [Pan10]. Pressed [Tay10]. Pressing [Ves19]. Pressman [Dow16]. Prestige [Neu13]. Preußen [Ole17b]. prévoir [Gil10]. Price [Man19]. Pricing [Der17]. Priestley [De 17]. Prime [Fir12, Lag11]. Print [Cro14b, Low10, Ree13, Ung15, Cro14b]. Pritchard [Smi14b]. Privacy [Jep18]. Private [De 15, Fir17, McC17]. Priya [Sto19]. Pro [vMS15]. Pro-Nuclear [vMS15]. Probability [Wel17]. Probe [Joh13]. Probing [Mun14]. Problem [Ber15, Cah19, Mey18, Son17]. Proceedings [Wor10]. Process [Edw12b, Wor16a]. Processing [Col15, Kur19a]. Proctor [Fra16]. Producers [Hog14b]. Producing [Cos17]. Product [vdHP13]. Production [Bhi10, Cum12, Fro10, Gan11, God15, Mal17, Mat17b, Pur10, Ros10, Sop14, Tya12, Har13, Cor13a, Łot16]. Produktgeschichte [Lin13]. Profane [Mun16]. Profession [vM12b]. Professionalism [SH17]. Professions [Rom10]. Professors [Kou10]. Profit [Ham18c, Pri19]. Profiteers [Lev13a]. profits [Har13]. Program [Jos16b, Shi17, Wei15]. Programmed [Tin19]. Programmers [Cer11]. Programming [Arc19, Lie18, NPA14, Gri12]. Programs [Luc14]. progrès [MS16]. Progress [Ald10, Cry11, Gil15, Moo16a, Vit17, Hou10]. Progressive [Aik19, Pla16, Cas13]. Progressive-Era [Pla16]. Project [Bus19a, Car13a, DB13, Par17, Pet10, De 11a, DiM10, Mat19, Com12, Kru14]. Projection [Ban16]. Projections [Gil15]. Projectors [Rat12]. Projectory [MV15]. Projects [Bea11, Kum19a, Tin18, Wor16b]. Proletarian [Jos13]. Proliferation [Kru13b, Mad19]. Prométhée [Pan18a]. Prometheus [Bra15a]. Promise [Big12, Dru12, Oak19, Wor16a, Pet10]. Promoting [Esb15, Ham17]. Promotion [Ald16a, Joh18b, Man13]. Prone [Luc10]. Proof [OL17]. Proofs [Fea19]. Propeller [Rol18]. Propen [Rae16]. Properly [Kum19a]. Property [Coo10a, Cum12, Got12, Kra11, Mer17a]. Prophet [De 19]. Prophets [Gil15]. Propriété [Fox11]. Propulsion [Mes16]. Prose [Kin10]. Prosperity [De 15]. Protection [Coo10a, Miz11]. Prototype [Her17b]. Proximal [UHW16]. Prozac [Hus10]. Psychology [Luc10]. Psychophysical [Cah14]. Psychotechnics [Tka19]. PTW [Tho10b]. Public [Ber17, Ber15, Cum11b, De 11b, De 15, Fir17, Kru16, Phi14, Pla16, Rus12, Sid18, TAF^{+}18, Wol10, Han13, Cor12a]. Publics [Fin17]. publier [Gui16]. publique [Har13]. Pulse [Hui16]. pump [Pac18]. Punch [Ens11]. Punch-Card [Ens11]. Pure [War14]. Puritan [Tur18]. Purpose [Lan18a]. Pursell [Wac17]. Pursued [Tho15]. Pursuing [Ald10, Joh11a]. Pursuit [Coo10a, Fro10, Hir11, Li16, Wor16a, vM12b]. Push [Plo12]. Putin [Jos11]. Putting [Pet13]. Pyramid [Edw16]. Pyrénées [Lag15]. Pyrotechnics [Mau11].

Qing [Ham18d]. Quagmire [Ert12]. Quality [Bor14]. Quantified [Col17a].

Quantifying [Wel17]. Queen [Ded18]. Quest [Car13a, DL18, Hea13b, Kem12, Mes16, Mey17, Var18, Bye13]. Question [Dou10]. Questions [Bre16, vdH17]. Quickenden [Eva15]. Quiet [Cer16]. Quinlan [Bou19]. Quiñones [Kat14]. QWERTY [Mul18a].

R [Alp16, And14, Blo13, Cah19, Gam15, Gar18, Hög15, Ihd16, Jen13, Joh17, Jør17, Kai17, Kan16, Rol18, Sla18, Ver15]. Raboy [Cha18b]. Race [Big16, Ler10, Man15, Mav10, Mey17, Mit18, Neu13, Rol10d, dlP10, Aid19, Hog18b, Kin11]. Rachel [Cer16]. Racial [Mer17b, Zyl11b]. Racing [Vol10]. Rader [Bed18]. Radiance [Gru14]. Radiation [Kru16, Ona17, Ona18]. Radical [Mul18a, Pre10a]. Radio [Arc18, Arc19, Bal10, Bat11, Bir19, Cas16b, Fro14, Han13, Kra14, Lov11, Mun14, Rus11b, Ste14, Wij14, Rus12, Wol10]. Radioactive [Bal11b]. Radionavigation [Ran14]. Ragnhild [Kra17]. Rags [Sen17]. Rail [Ber16, Pan18b]. Railroad [Ben17, DeB11, Gam11, Gan19, Sch14c, Ran11]. Railroader [Cor10]. Railroads [Chu15, Gam15, Sch12e, Ald10, Zyl11a]. Rails [DeB11]. Railway [Coo10b, De 11a, Kni13, McD15, Pra12, Vee10, Gam13]. Railways [Łot15, Coo12b]. Rain [Ber18]. Rainfall [Van16]. Ralf [Ber16, Łot15]. Ralph [Gal16]. Randal [Gud13]. Range [Bar16a]. Ranieri [Wel15]. Rankin [Ham18a]. Rao [Lau17, Pat15]. Rape [Bou19]. Raphael [Ver12]. Rasmussen [Fea16]. Ratio [Eds10]. Rational [Li16]. Rationality [Jac10b]. Rattraper [Has16]. Raulerson [Fer16]. Ray [Kru15]. Rays [Sla11c]. Raz [HV15]. Razlogova [Han13]. Re [Dow10, Low16b, Wol14]. Re-Collection [Low16b]. Re-inventing [Wol14]. Re-making [Dow10]. Reach [DK18]. Reactor [Cam18, Mad19, Wel17]. Reactors [Fja15]. Reader [Ros12b, Sto13b, Dug10a]. Readers [All11]. Reading [Bij12, Gor15, How15a, NW19]. Readings [Leg11, Val10]. Real [Cer10, Hir11]. Realities [Moo16a]. Reality [Fja15, Joh12, Kno14a, Mit19]. Realizations [Gil15]. Realm [Dow18]. Reaper [Coc17b]. Reason [Ihd11, Ihd18, Bra12]. Reasons [HS13]. Reassessing [Gai13]. Rebecca [Col14, Hri16, Rud18, Var18]. Rebels [Gag15]. Recapitalizing [Wol11b]. Recasting [Ban16]. Recherche [But12, Har13]. Recipes [Kno14a]. Reciprocal [Lae16]. Reckless [Mad19]. Reclamation [Lag17]. Recoding [Hay14]. Recognition [Joh11b, LM19]. Recombinant [Low16a]. Reconfiguring [Rob17]. Reconnaissance [Mac13]. Reconsidered [Kim14, Lon12a]. Reconsidering [Clo12]. Reconstructing [Gan17a, Ona17, Moo16b]. Reconstruction [Wil17c]. Recorded [Bij12]. Recording [Bru19, Gop15, Kar18, ZT18, Sch19]. Recordings [Kur19b]. Records [Ves19]. Recreational [Tho10b]. Recurrent [McN14]. Recycling [Jac12a, Sam13]. Red [Dav17, Lau17, Joh18a, Ger11, Lag13]. Redefined [Mey18]. Redefining [GR10, Lan18a]. Reduced [SV19]. Reed [Abb15]. Reenvisioned [Wil10]. Rees [Hor14, SH19a, Wil14a]. Referees [Jos19]. Reflections [Ava15, Cho17, Coo17, Gre14, Hir11, KO13, Kra10, Mei19, Pac18]. Reform [Bea19b, Bix18, Li16, Mas17]. Refrigeration [Wil14a]. Refrigerator [SH19a, Ree19]. Regime [And15, Hel13]. Regimentation [Lae15]. Regimes [Mac16a, Har13]. Regina [Kal14]. réglementation [Yat16]. Regourd [Hel13]. Regulating [Kri19b]. Regulation [DK18]. Rehabilitation [Car13c]. Reid [Rae16]. Reilly [Kan13]. Reimagining [Hei15a]. Reimer [Gol15]. Reinhardt [Yan19]. Reinventing [Rol18]. Reinventions [Nye17]. Relations [Ber15, Cam19, Yos16]. Relationships [Her16b]. Relatives [Nor15a]. Relativity [Bey12, Can15]. Religion [Gru14, Wak11]. Religious [Fer12, Fis15, Gil15, Sie14]. Remaining [Hin10]. Remaking [De 17, Mur12]. Remedy [Tob13]. Remembering [Buh17, Wor10]. Removing [Les19]. Renaissance [Dup15, Gam15, Lae16, Mor14b, Ver12, Ans17]. Renan [Lag15]. Renata [Ber14]. Rendering [Cre17]. Reordering [Hog18a]. Reorganization [Oka17b]. Replacement [Sch12b]. Reproducibility [Rob10c]. Reproduction [Dru15, Bjö13]. Reproductive [Sat10b]. Republic [Woo19, Wel10]. Republican [Ree16]. République [Fox11]. Research [BM19, BSA16, Cer13, Cur13, DK18, Joh18a, Kat10, Kno14b, Kri19b, Moh15, Neu11, Nik15, Pan18c, Pen18, Ran15, Sah13b, Sta10a, Tei11, Tka19, Bil15, Sla16, Tim18]. Reservoir [Cas14]. Resistance [Li16, dlP10, Dry19]. Resonance [Bra12]. Resource [Cam15]. Resources [Nah13, Wu19, vdV14, Jac13]. Response [Ald16b, Cow12, Gai14, Gai15, Kon11]. Responses [Ded18]. Rest [Pul16]. Restraints [Sat10b, Wet15]. Restrict [Jac11a]. Rethinking [Bai18, Coc13, Ens12, Tar15]. Retirement [Luc14]. Reuleaux [Gif17a]. Reusing [Wil11]. Revealing [Hen18a]. Revere [Ros11]. Reverse [Bro17b]. Review [Mül16a, Åbe15, Abb11, Abb15, Abn19, Abr13, Aga11, Aga14, Aga16, Aga18, Aid19, Ala12, Ald10, All11, Alp16, Amm11, Amm14, Anb18, And15, And10, Ang15, Ang17, App14, App17, Arc12, Arc13a, Arc13b, Arc18, Arc19, Arn18, Aso12, Bac15, Bai12, Bal11a, Bal11b, Ban16, Bar14a, Bar18, Bar12, Bar10, Bar16a, Bar16b, Bat11, Bau11, Bea11, Bea19a, Bea19b, Bed18, Bed10, Bek18, Ben10, Ben14, Ben17, Ber11, Ber18, Ber16, Ber14, Bes13, Bes18, Bey12, Bhi10, Bhi13, Bie16, Big12, Big16, Bij11, Bij12, Bil15, Bim18, Bix18, Bjö13, Bla10, Blo13, Blu12, Boc16, Bon16, Boo17, Bor10, Bor17, Bou19, Bow10, Bow14, Bow15, Bow16, Bow12, Bra15a, Bra13, Bra12, Bra15b, Bre10a, Bre18, Bre10b, Bri10, Bro12a, Bro17a, Bro11, Bro12b, Bro13]. review [Bro14b, Bro15, Buc12a, Buc14, Buc12b, Bug13, Bui10, Bur13, Bus19a, But12, Bye13, Bye19, Caf18, Cah14, Cah19, CK15, Cam19, Car13a, Car14a, Car13b, Car16, Car15, Car18, Car11, Car13c, Cas15, Cas10, Cas13, Cas16a, Cas14, Cas19, Cas16b, Cer11, Cer15, Cer16, Cha14, Cha11a, Cha18a, Cha18b, Cha11b, Cha12, Chu15, Cit12, Cla13, Cla10, Cla11, Clo11, Clo12, Coc17b, Coc17a, Coc18, Col12a, Col13, Col17a, Col14, Col16, Col19a, Col15, Col19b, Col17b, Col12c, Com12, Coo10b, Coo11a, Coo11b, Coo12a, Coo12b, Coo14, Coo17, Cor13a, Cor14, Cor10, Cor12a, Cor19, Cor12b, Cor16, Cou17a, Cou17b, Cre17, Cro18, Cro14a, Cro14b, Cru19, Cry11, Cum11b, Cum11a, Cum12, Cum13, Cum19, Cur12, Cut11, Cut13, Dae19, Dal15, Dav17, Dav13]. review [Day11, DeB11, DB13, De 17, De 19, De 11b, De 15, De 16, Dec18, Dér18, Deu17, DiM15, Dic10, Die10, Die11, Die12, Dio11, Dio12, Dis10, Dis11, Dis15, Dob16, Dob10, Doe12, Dor10, Dou19, Dou16, Dow18, Dow15, Dow16, Dow17, Dru12, Dru15, Dry19, Dug10a, Dug10b, Dug11, Dup15, Dus18, Dus10, Dus11, Eds10, Edw18, Edw12a, Edw17, Eis10, Eli12, Eng12, Ens10, Ens11, Epp15, Eri15, Ert12, Ert16, Est10, Eva15, Eva16, Eva18, Fai16, Fea16, Fee11, Fei12, Fel14, Fer11a, Fer13, Fer19, Fer12, Fer16, Fig10, Fin11, Fin13, Fir12, Fir17, Fis15, Fja12, Fle19, For10, For17, For11, Fox11, Fra13, Fra16, Fra18, Fre11, Fre15, Fre18, Fri16, Fro10, Fro14, Fru17, Gag15, Gai15, Gai18, Gai19, Gal17, Gal16, Gal15, Gam11, Gam13, Gam15]. review [Gan10, Gan11, Gan17a, Gan19, Gan17b, Gar13, Gar18, Ger16, Ger11, Ger13, Gif15a, Gif17b, Gif17a, Gil11a, Gil10, Gil11b, Git10, Gla17, God15, God16a, God18, Gol14a, Gol15, Gol12, Gol14b, Goo11, Goo19b, Gop15, Gor10, Gor11, Gor15, Gor12, Gor17, Got12, Gra11, Gra16, Gra15, Gre11, Gri12, Gro18, Gro15, Gru14, Gru16b, Gru12, Gua13, Gud13, Gud18, Gue13, Gui16, Hac15, Had10, Haf11, Haf18, Hag15a, Hag15b, Hah15, Hah19, Hai15, Ham11, Ham18a, Ham12a, Ham14a, Ham16a, Ham18b, Ham10a, Ham10b, Ham18c, Ham13, Ham12b, Ham18d, Ham16b, Han13, Han14, Har16, Har19, Har11, Har12, Har13, Has16, Hau10, Hau15, Hay18, Hay14, Hay12, Hea11, Hea12, Hea13b, Hea13a, Hea15, Hei15a, Hei15b, Hel13, Hen18a, Hen19, HV15, Hep16, Her16b]. review [Her17a, Her17b, Her11, Hes11, Hic12, Hic14, Hig12, Hin10, Hin11, Hir12, Hir14, Hir17, Hir16, Hjo16, Hle19, Hoa15, Hob14, Hoc10, Hog18b, Hog18a, Hög13b, Hög14c, Hög15, Hok17, Hoo10, Hoo15, Hoo19, Hor14, Hor18, Hor17, Hos13, Hou10, How15b, Hri16, Hu14, Hug15, Hug16, Hui16, Hul16, Hun16, Hur15, Hus19, Hus10, Ihd11, Ihd16, Ihd18, Iri12, Jac10a, Jac10b, Jac11a, Jac12a, Jac14, Jac16, Jac19, Jac12b, Jac13, Jac15, Jaf10, Jaf11, Jas12, Jen11, Jen13, Jen15, Jin17, Jin19, Joh16, Joh17, Joh11a, Joh12, Joh13, Joh15a, Joh15b, Jon10, Jon15, Jør10, Jør13, Jør17, Jos11, Jos13, Jos16a, Jos16b, Jos19, Kai12, Kai10, Kai17, Kal14, Kan13, Kan16, Kar18, Kar17, Kas17, Kat11, Kat14, Kel10, Kel11, Ken14, Ken19, Ket16, Kil15]. review [Kim12, Kim14, Kim19, Kin10, Kin11, Kin14, Kit12, Kle16, Kli19, Kni13, Kni18, Kno14a, Kno16, Koe18, KO12, KO14, Kra14, Kra17, Kra19, Kre10, Kre14a, Kre14c, Kri19a, Kru10, Kru12, Kru13a, Kru13b, Kru14, Kru15, Kru16, Kum12, Kum16, Kur19a, Lac16, Lad14, Lae16, Lag11, Lag13, Lag15, Lag17, Lai12, Lak13, Lan17, Lan18a, Lan10, Lan11, Lan15, Lan18b, Lap12, Las18, Lau11, Lau13, Lau17, LeC11, LeC17, Lee14, Les10, Leu14, Lev11, Lev17, Li16, Lie18, Lif18, Lig14, Lin13, Lon11, Lon12a, Lon12b, Łot15, Łot16, Lov11, Low16a, Low10, Low13, Low14, Low18, Low16b, Lub18, Luc10, Luc13, Luc18a, Luc11, Luc12, Luc18b, Maa16, Mac17, Mac16b, Mac18, Mac13, Mac14, Mai10, Mai12, Mai16, Mal17, Man18, MF16, MF17, Man15]. review [Man19, Mar18, Mar11a, Mar11b, Mar17a, Mar17b, Mat17a, Mat17b, Mat17c, Mav10, McC14a, McC15, McC17, McC13, McC11, McD10, McD17, McL17, McN14, Mck13a, Mei10, Mei13, Mer17a, Mer17b, Mes16, Mey12, Mey15, Mey16, Mey17, Mey18, Mey19, Mil12a, Mil13a, Mil12b, Mil13b, Mir14, Mir19, Mis11a, Mit19, Mit12, Miz11, Mod10, Mod11, Moh12, Moh15, Mol15, Moo16b, Moo16a, Moo15, Mor18, Mor12a, Mor10a, Mor12b, Mor10b, Mor14b, Mor14c, Mor17, Mou10, Mue15, Mul16b, Mun19, Mun14, Mur12, MS16, Nah13, Nan12, Nee18, Nel16, Nem14, Neu11, Neu13, Neu14a, New10, New18, Nie12, Nik17, Nof15, Nof18, Nor15a, Nov14, Nye11, Nye13, Nye14, Nye15, Nye17, O’H10, O’R19, Oak19, Oat16, Obe14, Oka17a, Oka18, Ole17a, Ole17b, Olw13, Ona14, Ona18, Ora15, Ore12, Ore13, Osl10]. review [Paj16, Pan12, Pan10, Pan15, Pan18a, Pan18b, Pan16, Pan18d, PT13, Par17, Pat15, Pat11, Per12, Pet11, Pet12, Pet18, Pet10, Pet13, Pet16, Pfa13, Phi14, Pik15, Pla17, Pom10, Pop19, Pop13a, Pop16, Pop13b, Pos15, Pra12, Pra13, Pre10b, Pri19, Pua17, Pul16, Pur17, Rab13a, Rae16, Ram14, Ram10, Ram11, Ran11, Ran15, Ree16, Ree18, Ree19, Rei12, Ret16, Rey12, Rhe14, Ric13, Rie16, Rio11, Riz10, Rob10a, Rob12, Rob10b, Rob10c, Roh19, Rol10a, Rol16, Rol17, Rol18, Rom10, Ros12a, Ros11, Ros12b, Ros13, Ros10, Rub18, Rud13, Rud18, Rus11a, Rus11b, Rus12, Ryd11, Sah13a, Sah14, Sam13, San11, San12, San17, Sat13b, Sat10b, Sau11, Sca14, Sca11, Sch14a, Sch19, Sch12a, Sch13a, Sch11a, Sch11b, Sch13b, Sch15a, Sch14b, Sch13c]. review [Sch11c, Sch11d, Sch12c, Sch12d, Sch13d, Sch15b, Sch14c, Sch10b, Sch10c, Sch12e, Sch13e, Scr12, Sea11, See15a, Set16, Sha17, Sha10, Shi12, Shi17, Shu15, Sid12, Sid18, Sim17b, Sim17a, Sin12, Sin15, Sis12, Sjö11, Sla10, Sla11a, Sla15, Sla12a, Sla12b, Sla14, Sla16, Sla18, Sla19, SH15, SH19a, SH19b, Smi10a, Smi11b, Smi13, Smi14b, Smi17, Sny17, Sol11, Son17, Sop14, Sop15, Spe11, Spi15, Spr19, Spu11, Spu13, Sta17, SP17, Sta10a, Sta14a, Ste10a, Ste19, Ste10b, Ste11, Ste14, Ste16b, Sti11, Sti18, Sto13a, Sto13b, Sto19, Str10, Str11, Str12, Str13, Sul19, Sum17, Suo12, Sut12, Sut19, Swa10, Swa13b, Swa15, Swa18, Swa19, Sza15, Tar15, Tay17, Tei11, Tel11, Tel13, Ten14, Ten15, TS12, The10a, The10b]. review [Tho10a, Tho15, Tho10b, Tim17, Tim18, Tin19, Tob13, Tod10, Tod11, Tom14, Tom17, Tuc16, Tuc12, Tur10, Tya13, Ung15, Uzi19, Val10, Van14, Van18, Var11, Var18, Vee10, Ven12, Ver12, Ver15, Ver13, Ves19, Vet19, Vin12, Vit17, Vol10, Vol11, Vol17, Vos10, Wac17, Wad16, Wak11, Wal10, Wal18, Wan13, War10, War14, Wea16, Wea18, Wei13, Wei15, Wel10, Wel15, Wes10b, Wes11, Wes15, Wes16, Whi11a, Whi13a, Whi14, Whi11b, Whi13b, WH11, Wij12, Wil17a, Wil10, Wil14b, Wil14a, Wil17b, Wil19, Wis11, Wis17, Wol10, Wol15, Wol11a, Wol14, Woo19, Wor10, Wor16a, Wor16b, Wos17, Wri12a, Wri12b, Wro10, Wu11, Wu19, Yan19, Yat16, Yos16, ZT16, Zar13, Zar17, Zel11, Zia18, Zim10, Zim13, Ziz19, Zyl11a, Zyl15, dS17, dV12, vM16]. review [vR12, vdH17, vdV12, vdV14]. Revill [Gam13]. Revival [Oka17a, Gru16b]. Reviving [Sop15]. Revolt [Bie16]. Revolting [Reu14]. Revolution [Ala12, Bau11, Bea19b, Bow10, Bra15a, Fin13, Gla17, Hah19, Hor18, Hu14, Lag17, Lau11, Lev13a, Mar13, Pra19, Sch11c, Ste10a, Ste11, Van14, Wes11, dS17, De 10, Bow16, Bug13, Hog18a, Kal14, Kum16, Pri19, Sto19, ZT16, All11, Bey12, Cha11b, Cor10, Tod10]. Revolutionary [Coo12b, Sul19, Cer16]. Revolutionized [Fra16]. Revolutionizing [Bed18, Kai17]. Revolutions [Spi15]. revue [Gui16]. Rhetoric [MF17, Wes10b, Wij14]. Rhode [Deu17]. Ribbat [Nye14]. Ribbon [Wel15]. Ricardo [Col19b, PT13]. Rice [Kim12, Wu11]. Rich [Aga14]. Richard [Gol14b, Gud18, Kru13a, Low16b, Mey16, Mor14b, New18, Nye17, San17, Shi17, Smi17, Sny17, SP17, Uzi19, Wei15, Wol14, Cro16, Tho16]. Richardson [Wad16]. Richest [Shu11]. Richter [Jen13]. Rid [Pet18]. Ride [Ros11]. Rider [Tho10b]. Riedler [Gif17a]. Rieger [Lad14]. Right [Ala12, Seg10, Zar17]. Rights [Alp16, Col16, Med18]. Rigor [Kin11]. Rima [Cro14b]. Rinehart [Low16b]. Rings [Pom10]. Rio [GKMV18]. Riordan [Wes16]. Rise [Arc19, Ben14, Bir19, Car18, Cer13, Cor19, Dav13, Fea16, Fin13, Gol14b, Hul16, Jon13, Kle10, Koe18, Kri19a, Lai12, Mac15, Pop16, Ros13, Swa18, Tho10a, Vol11, Wes16, Whi13b, Zia17, Pet18]. Rises [Bre16]. Risk [Buh17, Car18, Luc13, McK13b, Wil17c, Kno16]. Risks [Lap12, vdV14]. Rita [Spu13]. Ritual [Arn18, Dug11]. Ritvo [Ham13]. River [Bro12b, Cam18, Gat18, Gor11, Son17, Eva18, Sum17]. Rivers [Lan13, Jos16a, Jos16a, Sha10]. Rivoluzioni [Col17b]. Roach [Hoo15]. Road [Bal11b, Ber16, EW15, Esb15, Fei12, Gai18, Jeo10, Kat10, Kle10, LL15, Vol10, Web15]. Road-Safety [Web15]. Roads [Die10, Fre11, Sch13c, Car13b]. Robert [Bow15, Chu15, Dav17, Fer13, Fra16, Gam15, Gan17b, Gor17, Gro18, Hui16, Joh17, Jos19, Leu14, Mey15, Nee18, Oat16, CF10, Kru15]. Roberts [Gru16a, Hog14a, See15a]. Robeson [Olw18]. Robin [Sin15, Spr19]. Robins [Hos13]. Robinson [Fis15]. Robot [Bow12, Ham16b]. Robots [Bjö13, DL18, Hen19, Kan16, Lan17]. Rock [Hor11]. Rocket [Wer12]. Rocketeers [Col16]. Rocketry [Ger11]. Rockets [dS17]. Rod [Goo11]. Roe [Sim10]. Roebling [Bro13]. Roger [Abb15, Jac16, Rol10c, Sch14a, HHH12]. Rogers [Jac15, Tho16]. Rogue [Boo17]. Roland [Rud13]. Role [Bro17b, Hog13a]. Roll [Hor11]. Rolt [Blo13]. Roman [Eva16, Ram10, Ram11]. Romance [Lae15]. Romania [BBŁT15]. Romantic [SC17]. Romanticism [Str12]. romantique [Pan18a]. Rome [KO12, Coc18, Ber14, Hea11, Lon12a]. Rommes [Hic14]. Roof [Ald16a]. Roots [Nee17, OM17, Tom15, Tur18, Wu11, Cut11]. Rope [De 17]. Rosalind [Dal15]. Rose [Mac13]. Rosenberg [Hea13a]. Rosies [Nov11]. Rossi [Mor18]. Roth [Ber16, Łot15]. Round [Pin10]. Roundtable [TAF^{+}18]. Route [Fei12]. Routes [Fre11, Bro15, Sch13c]. Roux [Cas19]. Rover [Lan17]. Rozwadowski [Har19]. Rubber [And10]. Rubble [Nan12]. Rubin [Wes10a]. Rüdiger [Anb18]. Rudolph [Zar17]. Ruggero [Wel15]. Ruhr [Hau11]. Ruins [Amm11, Mur12]. Rule [Dér18, San11, Bar18]. Rules [vM16]. Run [Lan13]. Ruptures [HPR13]. Rural [BH17, DH11, MF11, Wu11]. Russell [Kni13]. Russelo [Van18]. Russia [Coo12b, Gra16, Lag13, De 12, Hög13b, Jos11]. Russian [Hög15, Rub18, Nik15, UHW16]. Rust [Mor17]. Ruth [Gag15, Paj16, Boy16, Ste10b]. Ryan [Mer17b, Ret16].

S [Bus19b, Bye13, De 19, Fel14, Fra16, Gai15, Hea13a, Luc18b, Mck13a, Sat13b, Swa15, Tay17, Wad16, Whi12]. S-C [Tay17]. Sabers [Sla11c]. Sabine [Yat16]. Saccharin [Sch11b]. Sacred [Mun16]. Safavids [Buc12a]. Safe [Bra15b]. Safety [Ber15, EW15, Esb15, Gan13, Gan19, Kno16, LL15, Nor15b, Oka18, Tay14, Tin15, Web15]. Sagan [Oka18]. Sahara [Cam15]. Saint [BBB^{+}19, Ano19c]. Saint-Étienne [BBB^{+}19]. Sale [DiM10]. Sally [Eva15, Ona14]. Salmon [Fin17]. Saltpeter [Buc14]. Salvaging [Mor17]. Sambaluk [Mey17]. Sambrook [Sla14]. Samson [Pua17]. Samuel [Pet13]. Samuels [Pur17]. San [Mur12]. Sanatorium [Ven19]. Sandra [Gue13, Sto13b]. Sang [Wil19]. Sang-Hyun [Wil19]. Sanitation [Eva16]. Santa [Spu13]. Santos [SS13]. Santos-Dumont [SS13]. Sapnar [Cor19]. Sara [Smi14b]. Sarah [LeC17, MF17, Nor15a, Sla16, Wad16]. Saraiva [Coc17a]. Sarat [Hle19]. Sarkar [Hea13b]. Sas [Gal15]. Sasha [Woo19]. Satélite [Hea15]. Satellite [Ver13, Dav17]. Satellites [CC18]. Satia [Sto19]. Satire [Rat12]. Saturated [Sch12b]. Saudi [Cas16a, Cit12]. Savage [Hoo19]. Savagery [Her11]. savant [Bai12, Gil10]. savoir [Ora15]. savoirs [Gui16, Har13, Lev17]. savoirs-mondes [Lev17]. Saw [Tur10]. Scale [Cah19]. Scenes [Ban16]. Schäfer [Cor13a, Wan13]. Scheme [Goo19b]. Schiedt [Pop13b]. Schiffer [Hir14]. Schlichting [Ric13]. Schmidt [Gop15]. Scholar [Kai14]. Scholars [Ham18d, OL17]. Scholarship [Ste16b]. School [Bix18, Cos17, Goo19a, Ros12b]. Schot [vM16]. Schottky [Kai10]. Schroeder [Gal16]. Schuh [Lin13]. Schüll [Lee14]. Schulman [Kas17]. Schultz [Mey18]. Schwartz [Mor14c]. Schwoch [Ver13, Vet19]. Science [AH18, Aso12, BM19, Bed18, BS13, Ben17, Bil15, Bro17a, Bru19, Bui10, Cah19, Cha12, Clo12, Cro14b, Deu17, Dio11, Dio12, Dob10, Edw18, Fea16, Fer16, Fig10, Fin11, Fin17, Fle19, Gan17b, GKMV18, God15, God18, Gri12, Gro18, Hag14, Ham14a, Ham17, Ham12b, Hel13, Her17a, Hey10, Hu14, Jac14, Jen13, Ken14, Kra14, Kru10, Lee13, Lev13b, Mai12, Mau11, MSPLC16, Mil13b, Pet14, Ram14, Rud13, Rud18, Rus11a, Sah14, San12, Sch12c, Sch10b, Set16, Sla18, Smi14b, Sop14, Sta17, Sto13a, Sto13b, Swa13a, Tei11, Tel11, Tel13, Til19, Tim18, Tsu12, Ung15, Vos10, Wea18, Wes15, Wol14, Bes18, Bou19, Gru14, HV15, Jos13, Mun14, Nof15, Ben10, Cro14b, Sch13d, Sch15b, Sid12, Ven12, Wes11, vdV12]. Sciences [Ben14, BT19, Eri15, Gal16, Har17]. Scientific [Dus18, Gil15, Hag15b, Ham12a, Hen18a, Hir14, Hun16, New10, Pra19, Pre13, Smi17, Til19, Sch11d]. Scientism [Smi17]. Scientists [Kum12, Ona18, Rob12, Rub18, Sch12c, Sla12b, Tho15, Cas10, Sla16]. Scienza [CG16]. Scott [Bil15, Bro17a, Fea19, Joh17, Koe18, Kru14, Li16, Lie18, Luc13, Man15, Oka18, Wei15, Yos16, vdH17]. Scranton [Hei15a]. Screening [Had10]. Script [dSS19]. Sculle [De 16]. Sea [Bat14, Fel14, Gat18, SDJ19, Ang15]. Sean [Bur13, Olw13, SH15]. Search [Vit17, Yos16]. Searching [OM17]. Seas [Bri10, Fer19]. Seawalls [SDJ19]. Seb [Abn19]. Second [Boc16, Cha11a, Cha11b, McC11, Nah13, Sch16, Ven12, KNSZ16, Mil12b]. Secret [Her16b, Hjo16, Nov11, Bui10]. Secrets [Ber13]. Sectors [Gal17].

Security [Bus19a, Hus19, Kri19b, Lap12, Oat16, Oka18, Rol16, Sov16, Wea16, Fai16, Mey17, And10]. Seduction [Gra18]. Seeds [Ham18c, Sto13a]. Seeing [Goo19a, Koc16, WH11, Lan17, Lan15]. Seek [Mac13]. Seen [Kap19]. Segal [Nye15]. Seigneurial [Bon16]. Sein [Kai10]. Seismic [Wil17c]. Seismological [Buh17]. Seizing [Dru15]. Selbst [Tod11]. Selectivity [Kin11]. Self [HPR13, JI16]. Self-Machinery [HPR13]. Selisker [Lie18]. Sell [Gif15b]. Sellers [Zyl15, Whi13b]. Selling [Wei15, Lif18, Van14]. Semi [Yan19]. Semi-Colonialism [Yan19]. Semiconductor [OF18, Rio11]. Sense [Col16]. Senses [Kre14b]. Sensing [Jac11b]. September [And14, Ano11c, Tay14]. Sequence [Hag15a, Ste15]. Sequence-Sharing [Ste15]. Serenissima [Cam16]. Service [Ham18b]. Services [Dou19]. ses [MS16, Pan18b]. Set [Hog13a]. Seventeenth [Ber14, Per12, CR19, Der17]. Seventeenth-Century [Ber14, Per12, CR19]. Seventh [Wil11]. Seville [Cos17]. Sewers [Eva16]. Sewing [Fra13]. Sex [Bjö13, Pet16]. Shackelford [Yos16]. Shackled [Bra15a]. Shadow [Clo11, Coo14, RS14, Sol11]. Shane [Kle10]. Shannon [Goo19b]. Shape [Bro17a, Fai16, Hir12]. Shaped [Bye19, Gag15, Sim17b]. Shaping [Bro17b, Olw13, Sch10c, Zia18]. Shapiro [Low18]. Sharing [Ste15, vdV14, Zia18]. Sharma [Kum16]. Sharon [Haf18, Kru13b]. Shawn [Arc19, Pan16]. Sheehan [Bra13]. Sheila [Wil19]. Shell [Bie16, Mac13]. Shelter [Iri12]. Shen [Ree16]. Shephard [Maa16]. Sherry [Col15]. Shibata [Oka17a]. Shift [Mul18b]. Shifting [Boy16, Coo17]. Shifts [Wea18]. Ship [Cam16, Wil10, Wol14, Fer11a]. Shipboard [Fel14]. Shipbuilding [FP16]. Shipping [Lon11, Yan19]. Ships [Fer19, Lon11, Mor14b]. Shobita [Sid18]. Shocking [Hun16]. Shop [Edw12a, Jaf10, Kel11]. Short [Bug13]. SHOT [Bra17, See15b]. Should [vMS15]. Shredder [Zim11]. Shue [Mey19]. Shulman [Sov16]. Shuttle [Lan18a, Pom10]. Shuttles [Her16b]. Siam [Lim19, Pua17]. Sichuan [Wu11]. Siddiqi [Smi13]. Siècle [Hui16, Gui16, Ora15, Pan18b, Gil10]. siècles [But12]. Siegel [Gan17a]. Siegelbaum [Zar13]. Siegfried [Rub18]. Sigaut [Cou17a]. Sight [Goo19a, Pan16]. Sight-Seeing [Goo19a]. Sights [Kas17]. Siglo [PT13]. Sign [Dug11]. Signatures [Kra17]. Significance [Bla10]. Signs [Mei13]. Silbey [Mer17a]. Silicon [BL12, Cer16, Mat17c, Ore12]. Silicon-Gate [BL12]. Silk [Ber13, Haf18]. Silver [Hos13, Mou10, Ded18, Zim17]. Silverware [Jin19]. Simanti [Sha17]. Simine [Bug13]. Simmonds [Coo11c]. Simon [Fle19, Jon15, Bra10]. Simone [Bek18, Hug15]. Simons [Pfa13]. Sims [Bix18]. Simulation [Con15, Jeo15, Wea18]. Since [Fin13, Hay18, Jør17, Moo16b, Rol10a, War14]. Singapore [Ano17b]. Singing [Olw18]. Single [Kra14]. Singularities [Fer16]. Siobhan [See15a]. Skies [Ded18]. Skills [Cos17, De 10, Kui19]. Sky [Coo11b, Day11, Hoo10, Kin14, Kra14, McC14b, Mun14]. Skye [Har16]. Skygirls [Cla11]. Skyscraper [Aid19, Les10]. Slain [Man18]. Slaughterhouse [SH19b]. Slava [Fir17, Jos16b]. Slave [Ald10]. Slavery [Sol11]. Slayton [Col14]. Slums [Cam19]. Small [Mac18, Mod11, Ree18]. Smart [Mey16]. Smarter [Wol11a]. Smartphone [Aga18]. Smartphones [Cor16, Mey16]. Smelling [SSH18]. Smelter [Aik19]. Smelting [Kor18]. Smita [Tya13]. Smith [Fai16, Ona14, Ver15, Fer19, Moo16a, Pan16, Sim10, Tom14, War14, dS17]. Smith-Howard [War14]. Smithsonian [Bow15, Her16a, Seg14, Sta17, Wil13]. Smoke [Aik19, Rob12, Str10]. Sneddon [Lag17]. Snows [McN14]. Soaring [O’D12]. Sociabilities [MF16]. Social [Bal11c, Bek18, Bes18, Cho17, Cor19, Fer11b, Fer12, God15, God16a, Gol14a, Hag15b, Ham18d, Ihd18, Jac14, Kim13, Low16b, OS16, Par17, Ros12b, Shi12, Sum17, Swa15, Vos16, Wu11, Zim13, Bhi10]. Socialism [Jen11, Wes10a, Smi10a]. socialismes [Pan18a]. Socialist [De 12, Zar13]. Societal [Bim18, Rol10c]. Societies [Gol15, Mor10a, Wil14b]. Society [Bre10b, Cha18a, Dis10, Dus10, Ens12, Fin11, KO13, Mei10, Ros10, Aga14, Hic14, Kno16, Mac14, Ald10, Clo11]. Socio [KNSZ16, Sin12]. Socio-Cultural [KNSZ16]. Socio-Economic [Sin12]. Sociology [Dob10, Hau15, Pin10]. Sociopolitical [Mul18b]. Sociotechnical [Bor14, Dus10, Wil19]. Sod [Hur15]. Soft [Kat11]. Software [Día16, Lin10, Mal17, Nem14, Pri19]. Soil [Fin11, Sah13b]. Solar [Bar14b, Kar17]. Soldiers [Jos16b]. Solid [Kou10]. Soll [Mac16b]. Solovey [God15]. Solution [Sch12b]. Some [Dou10, God18, Mei19, Mul12b, Rub18]. Something [SV19]. Sonderweg [Kre14a]. Sonic [Bij12]. Sorcha [Goo19b]. Sørensen [Hic14]. Sorge [Rab13a]. Sorge-English [Rab13a]. Sörlin [Spr19]. Sorts [Smi10c]. Soulcraft [Kel11]. Sound [Arc18, Arc19, Bij12, BT19, Bir19, Dow10, Gop15, KL19, Kre14c, Mil13a, Mor14a, Sca14, Vai13, Wij14, ZT16, Kra17, Bra15b, Sch14b]. Sounds [Bru19, Cah14, Kat14, Kra19, Kur19b, Pop13a, Van14, Lac16]. Soundscapes [Sca14]. Soup [Fro10]. Source [Pri19, Whi10, Wor16a]. South [Ald10, Gud18, Jeo10, Joh17, Lak13, Mav10, Moo16b, Nem14, Rob10a, Sau11, DiM10, LL15, Gud13, Man18, Ham10b]. Southern [Ded18, Kin14]. Sovacool [vdH17, Hir16, Wol15]. Sovereignty [Lan18b, Yan19]. Soviet [Car14a, Fir12, Ger11, Jos16b, Łot16, Mat19, Moo16a, Nik17, Smi10a, Smi13, De 12, Koc16, Les19, Fir17]. Soviets [Olw14]. sozialistischen [Tod11]. Space [Bur13, CC18, Cor12a, De 19, Ded18, Ert16, Fir17, Gru12, Her16a, Jos16b, Las18, Lau13, Lau17, Mey17, Neu13, Pom10, Rol10c, Shi17, Sid10, Sie14, Smi13, Tho15, Wei13, Wor10, Nee17, Gar18, Jac14, Jos16b, Lan18a, Neu14a, Wei13]. Space-based [CC18]. Spacecraft [Con15]. Spacefarers [Kil15]. Spaceflight [Bim18, Kil15, Rol10c, dS17, Lan18a]. Spaces [MF16, Zel11]. Spain [ERST13, Pre10a]. Spallation [Wor16a]. SPAM [Coo14]. Spanish [Bui10]. Spark [Hun16]. Spatial [Par18]. Spatializing [Pur17]. Special [Hey10]. Spéciale [Dou16]. Specialty [Rol18]. Species [Dec18]. Spectacle [Ben10, Leg11]. Spectatorship [Bro17a]. Spectrum [Aga18]. Speculating [Cla13]. Speculative [Dry19]. Speech [LM19]. Speed [Luc11, Sin12, Vai13, Var11]. Speier [Kri19a]. Spellbinding [Wij14]. Spellman [Ree18]. Sphere [CR19]. Spies [Her16b]. Spill [Col12c]. Spiral [Hen18b]. Spirit [Dus18, Str11]. Spiritualism [Gal15]. Spirometer [Big16]. Splenda [Sch11b]. Splice [Hey10]. Split [Hey10]. Sponsored [Aso12]. Spoony [Jin19]. Spoorwegen [Dis11]. Spot [Wer12]. Sprague [Gal11]. Spraying [Hay18]. Spread [Cor13b, Sut12]. Sputnik [Neu13, Bal11c, Rud12]. Spy [Dav17]. Srinivas [Tya13]. St [Sti18]. Stage [Kan13]. Staged [Sca14]. Stairway [Sch10c]. Stalter [Pik15]. Stalter-Pace [Pik15]. Stamper [Pet19]. Standard [Mac16a]. Standardization [Vee10, Kir12]. Standards [Boy16, Vin15, Kno14a]. Stanford [Low16a, ZT16]. Stanislaw [Pic14]. Star [McC15]. Starosielski [Hug16]. Start [Pos10]. Stasis [Var11]. State [Aso12, Ber18, CK15, Hor18, Kri19b, Ram14, Rol10c, Sat10a, See15b, Tsu12, Vos10, vdV12, Bus19a, Car13b, McC15, Oat16, Rol16, Sti11]. State-Sponsored [Aso12]. States [DiM15, Gil11a, Gro15, Joh12, Kin10, Pre10b, Sah13a, Sid18, Str10, DK18, Gif15b, Ham14b, Neu14b, Nor15b, Pop11, Tar14, De 15, Gar13, Rol17, Rud18, Sch13a, Tya13, Wil17b, Bal11b, Jon10, Kin10, Rob10b, Wol10]. Stathis [Joh15a]. Station [Sch10a]. Stationary [Bow10]. Stations [Wil17a]. Statistical [Car18]. Status [Sin12]. Staymaking [Rab13a]. Stays [Rab13a]. Steam [Fer19, Qui17, Sch12e, Wan10, Bow10]. Steamboats [Ada17, Dav13]. Steamship [Gil15]. Steel [Cut11, Eis10, EW12, HPR13]. Steen [Sla11b, Ham16a]. Steinbeck [Kre14a]. Stem [Nor15a, Tim18]. Step [Hag14, MSPLC16, DGS16]. Stephanie [Man19, Zim13]. Stephen [Bro14b, Cla13, Cro14b, Hog18a, Kra19, Sla14, Ihd13]. Sterility [Tob13]. Sterne [Cum13]. Stethoscope [Muk19]. Steve [Fis15, Han14]. Steven [Bim18, Gai18, Gai19, KO14, Vol17, Rol10c]. Stevens [Cha18a, Hag15a, Wad16]. Stevenson [Dal15]. Stewart [Kru13a, Kru13a]. stigmatisation [Yat16]. Still [Cou17b]. STM [HNG16]. Stock [BM14, SH19b]. Stomp [Ham10b]. Stop [Boy16, vMS15]. Storage [Cas14, Col19b]. Store [Arc10]. Storefronts [Mei13]. Storia [Col17b]. Stories [Kli13, Rol10b, Tay10, Cla11]. Storm [Dav17, Nee18]. Storms [Miz11]. Story [Bac15, Bla10, CK15, Kum16, Ree19, Sch11b, Sch13e, Swa15, Bro12a]. Storytelling [Kai12, Arc18]. Strait [Bor17]. Stranded [Jos13]. Strange [Han14]. Strasse [Pop13b]. Strategic [And10, Dor10, Sin12]. Strategies [Web15]. Strategy [Gai18, Zim10]. Streetcars [Zyl11b]. Streets [Mei13, dSS19]. Streible [Gar13]. Strip [Wel15]. Strongly [Hun16]. Stroud [Ore13]. Structure [Mal17, Sin12]. Structures [Pre10b]. Struggle [Bar16a, Cut13, Mer12, Sto13a, Wil17b, Zyl11b]. Struggles [Eva18, Fra13, Hir14]. Students [MM11]. Studies [Bim18, Bir19, Dug10a, Gan13, Gro18, Mil19, Sch14b, Smi14b, SV19, Sto13b, Mil13a, Dob10]. Studio [Gop15, Kar18]. Studios [Gar18]. Study [Ens10, Les19, Wen17, Obe14]. Stuff [Sch13b]. Stumpf [Kur19b]. Style [Fru14, Mor18]. Subject [Ban16]. Sublime [TS12, Rei12]. Subrata [Nof15]. Suburb [Mil12b]. Suburban [DH11, Whi10, Whi13b, dV12]. Suburbs [Pat15]. Subversion [Hah15].

Subversive [Bie16]. Subway [Pik15]. Sucralose [Sch13b]. Sudrow [Lin13]. Sue [Sch15a]. Sugar [Les19, Mer17b, Sch13b]. Suisman [Van14]. suisse [Has16]. Suite [Dus11]. Sultans [Pet14]. Summa [Pic14]. Sumner [Ung15]. Sun [Kar17]. Sunny [Pik15]. Super [Jin19, Wes16, Jin19]. Superconducting [Wes16]. Superior [LeC11, Sut19]. Superpower [Zia17]. Supervision [Mac14]. Supplement [Ano18h, Ano19g]. Supply [Shu19, Tom15, Mac16b]. Support [Neu11]. Surgeon [Hau19]. Surgically [Sla11c]. Surprising [Big16]. Surround [Geo15]. Surveillance [Bre10a, Mac14, Mar19, Lap12]. Survival [Mai10]. Surviving [Ruu11]. Susan [Gop15, Ree18, Ret16]. Suspense [Kin10]. Sustain [Bar16a]. Sustainability [BGW^{+}17, Swa15]. Sustainable [Wil17b]. Sustaining [Sut19]. Suvobrata [Hea13b]. Suzanne [Vol17]. Sven [Coc16]. Sverker [Gra15, Spr19]. Swartz [Cor16]. Sweden [Fja15, Wor16a, vdV12]. Swedish [Lin11]. Sweet [Meh10, Sch13b]. Sweeteners [Sch11b, Sch13b]. Swindlers [Cas10]. Swinfield [Cro14a]. Switch [Abb18]. Switching [Pan12]. Switzerland [Zia17]. Symbols [Ham17]. Symonds [Ang15]. Sympathetic [Mac19, Tuc12]. Symposium [BBŁT15, BBB^{+}19, Ber12, BGW^{+}17, ERST13, GKMV18, KLŁ14, KNSZ16, Wil11]. Synchrotron [Wes11]. Syndrome [Ore11]. Synergies [Pre10a, Til19]. Syntheses [Olw14]. Synthesize [Gul19]. Synthesizing [RAF^{+}11]. Synthetic [Ham16a, Wes10a]. System [Bes18, Gar18, Gro19, Joh16, Lan11, Jin19, Cor12b]. Systems [Eis17, Ens11, FR18, Koc16, Par18, Shu19, Van16, Eva16].

T [Bar16a, Blo13, Fis15, Gal16, Kra10, Lau13, Mat17a, Smi13]. Tōhoku’s [Buh17]. Table [Bla10, Wil13]. Tacit [Bij11]. Tacoma [Ano11c]. Taconite [Man13, Bar16a]. Tactical [Cum11a]. Tailfins [De 15]. Taiwan [Tin18]. Takashi [Moo15]. Take [Cer11]. Taken [Aga14]. Takes [McC15]. Takeshita [Hog14b]. Takhteyev [Nem14]. Taking [Ber15]. Tale [Olw10]. Tales [McN14]. Talking [Edw12a]. Taming [Cer13, Mul16b, Swa19]. Tampere [Wil11]. Tangled [Mck13a]. Tank [Dou16]. Tanya [Bra13]. Targets [Hic12]. Taylor [Bat14, God18, Pop13a]. Teaching [Rud12, Rud13]. Teams [Lan17]. Tech [Sch15a, Sch12d]. Tech-Noir [Sch12d]. Technical [Cam16, Cer11, Dus18, Eis10, Hau19, Hea13b, Jas12]. Technics [Ona17]. Technik [Gif17a, Ole17b]. Technikgeschichte [Kre10, Die11, Die11]. technique [Ora15]. techniques [Pan10]. Techno [Hay12, Tsu12, Vai13, Bix18]. Techno-Aesthetic [Vai13]. Techno-Human [Hay12]. Techno-Idealism [Bix18]. Techno-Science [Tsu12]. Technocracy [Nee17, Smi17]. Technocrats [Cru19]. Technoculture [Fer16]. Technologiae [Pic14]. Technological [Bea19b, BBB^{+}19, Cro16, Dio12, God16b, Gra18, Haf18, Ham14b, Jen11, Joh18b, JI16, Kli19, Kor18, Les19, MV15, Mit19, Moh15, Pet19, Pla16, Pre10b, Rob10c, SC17, Sch11d, Ste16a, Vin15, Whi11b, Zia18, Wil14c, Ens10, Gal17]. Technologie [Die12]. Technologies [Åbe15, Ava15, Bra19, Buh17, Cam15, Can15, Cho17, Col17a, Cor13b, Der17, Dry19, Fee11, Fra13, Gra15, Har11, Hoa15, Joh15a, KNSZ16, Ket16, Mae13, Mit12, Mit18, Olw18, Pan18a, Pin10, Pur10, Sid10, Tuc12, Ver13, Wil14b, Lap12, Jac15, Hic14, Wri12b]. Technologique [Pan18c]. Technologist [SW15]. Technologist-Historian [SW15]. Technologists [Tin19]. Technology [Aga18, AH18, Ang17, Arn18, BM19, BBŁT15, Bat14, BS13, BBB^{+}19, Ber12, BGW^{+}17, Blo13, BM12, BM14, Bre16, BL12, Bro14b, Bro14a, Buc12b, Cha14, Cho17, Coc17b, Coo11a, Coo17, Cow10, Cry11, Cur12, Dio11, Dis10, Div10, Dob10, Edg10, Edw18, ERST13, Ens12, Fig10, Fin10, Fis15, Fra16, Fre12, Gai18, Gal15, Gan17b, Gar18, GKMV18, God16a, Goo19a, Gop15, Gra11, Gro19, Hau15, Hei15b, Her17b, Hir11, HS13, Hog13a, Hu14, Ihd11, Jac10b, Jac11b, Joh18a, Joh11a, Jør10, JJ16, Jos19, Kai11, KLŁ14, KNSZ16, Kim15, Kit12, Kni18, Kno14b, Kre14c, Lee13, Ler10, Lev13a, Lev13b, Lig14, Lon10, Lov11, Luc10, MVÁ16, MF11, MF17, Mar10, Mas17, McC17, McC13, MSPLC16, Mit19, Moh16, Mom14, Mun14, Nik15, NPA14, OF18, Obe14, Oka17a]. Technology [OS16, Pan12, Pat19, Per12, Pop19, Pos10, RAF^{+}11, Rus11b, Ruu11, SK16, San12, Sha17, Sie14, Sim10, Sla10, Smi14b, Søn12, Sop11, Sop15, Ste14, Sto13b, Str13, SSH18, Swa10, The10a, Ung14, Ung15, Vos16, Wac17, Wan13, Wel15, Wer12, Wil11, Wol14, Wri12a, ZT18, Zyl11b, dlP10, vdS19, vdV12, Alp16, Fre18, God18, Gro18, Hen19, Hoo19, Mac17, Wis17, Wos17, Big12, Car11, Dio12, Dus10, Kim14, Mai12, Mod11, Rol10a, Sch11a, Shi12, Tsu12, Val10, Vit17, Wis17]. Technology-Intensive [Her17b]. Technopolitics [Eng12]. Technoscience [DiM15, McL17]. technosciences [MS16]. Technoscientific [Bou19, Coc17a, Mac16a, Pre10a]. Technosystem [Ihd18]. Tecnologia [CG16]. Tel [KNSZ16]. Telecommunications [Can15, Sin12]. Telefonico [Bal10]. Telegraph [Arc13b, Bek18, Jep18, Mül16a, Vet19, Wen17]. Telegraphic [Hea12]. Telegraphy [Hug15]. Telephone [Bea10, Sla19]. Television [Col15, Fro14, Hal16, Sch13d, Tho10a, Hok17, Tel11, Wij12]. Teller [Sla12b]. Tellers [Bac15]. Telotte [Sch15b]. Temin [Bra15a]. tempête [Gil10]. Tempered [Ava15]. Temporality [Kra19]. temps [Gil10, Gil11b]. Tendencies [Gla17]. Tennessee [Jac12a]. Tenth [Eds10, McC11]. Tenth- [Eds10]. Terrence [Spu13]. Terrible [Ryd11]. Territoriality [Gra11]. Territory [Ham18a, Mor14a]. Terror [Gra15, Hri16, Wea16]. Tesla [Bow14]. Test [Bjö13, Vin15]. Test-Tube [Bjö13]. Testing [Car13a, Gro19, OL17]. Texas [McC15]. Text [Día16, Mul12a, vM12a]. Textbooks [SGB16]. Texts [Sen17]. Textual [Dow16]. Thackray [Cer16]. Thailand [Pua17]. Thane [Hög13b]. Theater [Fre12, Wea16]. Theatre [Kan13]. Their [Hir10, KNSZ16, Sah13a, Sat13a, Sch14a, TAF^{+}18, Wis17]. Theme [Bed10]. Theories [Vos16]. Theory [Ber15, Cos17, FP16, Gro18, Hob14, Sch12d, Shi12]. Therapeutics [Li16]. There [Gor15]. Thermodynamics [Mac19]. Thing [Got12, Kra11]. Things [Ber14, Dou10, Hec10, Kim19, Lub18, Per12, Smi10c, Hei15b, Nel16, Ten15]. Think [Fri16]. Thinking [Eis17, Pan12, Mit19]. third [BGW^{+}17]. Thirlmere [Ham13]. Thirst [KO14]. Thirsty [Har16]. Thirty [Ber12, ERST13, Wil11]. Thirty-Eighth [Ber12]. Thirty-Ninth [ERST13]. Thirty-Seventh [Wil11]. Thomas [Aga18, CK15, Cas19, Col13, De 17, God16a, Ham18b, Kni18, Pet18, Rhe14, Sla18, Bus19b, Car14b, Har13]. Thompson [Tim18]. Thought [Dis15, Ros10]. Thoughts [Car14b, Dou10]. Thousand [Jac19]. Three [Ihd13, Wol11a]. Threshold [Gla17, Ona17]. Thresholds [Bor17]. Thrifty [Fle19]. Tiago [Coc17a]. Tiger [Gai19]. Tigerman [Bar14a]. Tim [Dou16]. Time [Cry11, Dug11, Fir12, Fja13, Fru14, Fru17, Kra10, Ste10b]. Timekeeping [Fru14]. Times [BBŁT15, Fru15]. Timor [Maa16]. Timothy [Fel14, Lub18, Mey18, Pop13a, Str13, Shu11]. Tineke [Car15]. Tinkerers [Gag15]. Tinkering [Tur18]. Tintorería [PT13]. Tobacco [Rob12, Swa13b]. Tobias [Cor13a]. Today [Kri19b, Fre11]. Together [Kit12]. Toilets [De 11b, Eva16]. Tolerating [Kor18]. Tom [Gre14, Kai14, Sta14b]. Tomorrow [Bar14b, Nye11]. Tomory [Cas13, Cro18]. Tony [Epp15]. Took [Jos16b, Whi11a]. Tools [Col15, Gif15b, Mai12, OF18, Ree13, Smi10c, Ada17]. Tooth [Jen11]. Top [Nov11]. Topi [Obe14]. Topic [Gul19]. Torin [Mac14]. Toronto [SK16]. Torpedoes [CF10]. tot [Dis11]. Touch [Kre14c]. Touring [Mom14]. Tower [SS13]. Town [Cut10, Cut11]. Toxic [Tar14]. Toys [Mai12]. Tozzi [Pri19]. Traces [Tka19]. Track [Gan19, Kur19a]. Tracking [Mit12, Pra12]. Tracks [Vee10]. Tractors [Fin10]. Trade [Fre15, Jon13, Mae13, Rab13a, Abr13, Hor18]. Trades [EW12]. Trading [BM14, Mac15, Mir19, Sis12]. Tradition [Bet14, Kar18, MF11, Nik15, Ruu11]. Traditions [Rob10a]. Traffic [Ber15, Nor15b, Tin15]. Trail [Kai11]. Train [Meh10]. Trainer [Jeo15]. Training [Hir10]. Trams [De 15]. Trans [Ded18, Sch12b]. Trans-African [Ded18]. Transatlantic [Mer12, Sch11d, vM12b]. Transfer [Les19]. Transformation [Fru17, Ham18a, Hor14, Tim17]. Transformations [Gal16, Kim14]. Transforming [Dow16, Wil13]. Transforms [Cor13a]. Transfusions [Jos13]. Transgenic [Sto13a]. Transhumanism [Fer16]. Transient [Hoa15]. Transition [BBŁT15, Gif17b, Łot15, Sch16, Lak13]. Transitions [Ste16a]. Translating [Fru14]. Translation [Bra17]. Transmutation [Sea11]. Transnational [Arc12, DB13, DiM15, Mac16a]. Transport [Ber16, Lan11, Wen17]. Transportation [Bug13, Cor10, Phi14, Pop11, Vol17, Bie16]. Trapped [BE13]. Travel [Ber13, Fig10]. Travels [Cut10]. traversée [Lag15]. Trawling [Mar19]. Treasure [Ham12b]. Treaty [Car13a]. Trees [Bjö13, Dér18, Koc16, Ore13]. Treu [Mei13]. Trevor [Mal17, Mil13a]. Trial [Joh15a]. Trials [Wer12]. Tribute [OS12]. Triumph [Dal15, Rol18]. Tropical [McL17]. Trotsky [Jen11]. Trucking [Kle10]. Truitt [Kan16]. Trusses [HPR13]. Truth [Gol14a, Rob12, Pom10]. Tsiolkovskii [Ger11]. Tube [Bjö13, Dre10]. Tulsa [Gru16a]. Tumultuous [Aga18]. Tungsten [Gif15b]. Tunnel [Wes16]. Turbine [Leg11]. Turbines [BH17, HS13, Lag11]. Turf [Ens10]. Turin [Mor15]. Turing [Swa13a, Joh15b]. Turkel [Hun16]. Turn [Hen18b, Ole17a, Pla17, RV18, Sny17]. Turn-of-the-Century [Ole17a].

Turner [Geo15]. Turning [Cou17b]. TV [Oka17a, Sch15b, Tho10a]. Twentieth [App14, Bed18, Ber18, Coc18, Cru19, Die10, Fro10, Gam15, Ham18a, Hin10, Hir12, Lau13, Lov11, Mor18, Neu11, Ram14, Ran11, Rob17, Sla12b, Smi17, Sta10a, Whi13b, Nor15b, Pre10a]. Twentieth-Century [App14, Ber18, Cru19, Hin10, Hir12, Lov11, Neu11, Ram14, Whi13b, Nor15b]. Twenty [Gal15, Hag15b, Łot15]. Twenty-First [Hag15b, Łot15]. Twenty-First-Century [Gal15]. Twilight [DeB11]. Two [Ang17, Bow16, Hir10, Olw10, Vol17]. Two-Wheeled [Vol17]. Typewriter [Bus19b, Mul12a, Bus19b]. Typists [Mul12a].

U.S [Ala12, Bed18, Caf18, Fel14, Gai19, Ham18b, Joh17, Kin11, Kru13a, Lag17, McC13, Mir14, Oka17a, Wol11b, Zia18]. U.S. [Cam18, Gro14, Jep18, Vai13, Vin15]. Ubiquitous [Kat14]. UK [Ber12, KLŁ14, Tay14]. Ulrich [Pop13b]. Umpires [Jos19]. Unarticulated [HS13]. Unbranding [Dae19]. Uncanny [DH11]. Uncertain [Jas12]. Uncertainty [Gol12, Roh19, Wel12, Ham14a]. Unconscious [Bow12]. Underground [Dis10, Lan15, Zia17, Pik15]. Undermined [Ziz19]. Undersea [Hug16]. Understand [Scr12]. Understanding [Bes18, Hei15b, Lee13, Ten14, Tho10b, Moh16]. Understandings [Lae16]. Underwater [Kim15]. Unearthed [Shu11]. Unearthing [Amm12, Wal10, Ree16]. Uneasy [Nik17]. UNESCO [And17]. Unfreedom [Lie18]. Unfriendly [Fra13]. Unger [Mor14b]. Unglued [Col15]. Unification [Jeo10]. Unified [Łot16]. Union [Fir12, Mar11b, SH19b]. United [Bal11b, DK18, De 15, DiM15, Gar13, Gil11a, Gro15, Ham14b, Joh12, Jon10, Kin10, Nor15b, Pop11, Pre10b, Rob10b, Rol17, Rud18, Sch13a, Sid18, Str10, Wil17b, Wol10, Gif15b, Neu14b, Tar14]. Uniting [Scr12]. Universal [Sid10]. University [BBB^{+}19, Dic10, Gru16a, Ken14, Kou10, Kri19b, Low16a, SK16, SK16, Sta14b, Tei11]. Unmaking [Cum12]. Unmasked [Maa16]. Unreliable [Kli19]. Unruly [Sum17]. Unseen [Pan16]. Untaken [Gai18]. Untilled [Str11]. Untimely [Amm11]. Update [Wos10b]. upon [Lin10]. Uranium [Abr13, Man19]. Urban [Amm11, Bea19a, Cam19, Ert16, HNG16, Jac16, Jac19, Ken19, Rom10, Sca14, Sop14, Wri12a, Wro10, dSS19, Cha11b, Mai16]. Urbanism [McD17]. Urbanizing [Tom14]. Ursula [Ole17b, Har17]. USA [Die12]. Usable [Cor19]. Use [KO13, Kru14, ZT18]. Used [SH19a]. Useful [Har17]. User [CC18, Cor13a, Lin10, Fra13]. Users [Hog13a, Hog14b, Kai17, UHW16, Jør10]. Uses [Sla11b]. Using [Gul19, ZT18]. USSR [Jos16b]. Utilities [Lif18]. Utility [Pla16]. Utilization [Coo11c]. Utopia [Rus12]. Utopian [Gru14]. Utopianism [Jen11]. Utopias [Nye15].

V [Cas14, Coo12b, Lau17, Olw14, Tim17, Wil17a, vdH17]. Vail [Hay18]. Valencia [PT13]. Valentine [vdH17]. Valerie [Lan18a]. Valero [PT13]. Valley [Cer16, Mat17c, Par18, Gor11, Jac12a, Ore12]. Value [Cor12b, Kel11].

Values [Val10]. VanCour [Arc19]. Vanderburg [Dus18]. VanderMeer [Pra13]. Vanessa [Fru17, Kas17]. Vanishing [Bie16]. Vapourware [Cer15]. Vassos [Low18]. Vast [Eli12, Har19]. Vauban [Lan10]. Vaughn [Cro14b]. Vegas [Lee14]. Vehicle [Luc14]. Vehicles [Mor15, SP17]. Velocity [Luc11]. Velvet [Wes11]. Verbeek [Hei15b]. Véréna [Bat14]. Vergleich [Lin13]. Verhältnis [Die12]. Verification [Car13a]. Verkehr [Pop13b]. Vernacular [MF11]. Verne [Dal15]. Version [Swa10]. Verso [Bal11a]. Versprechen [Kre10]. versus [Kre14b, Neu14b, OL17]. Vertesi [Lan17]. Veterinary [SH17]. Vetraria [Ber17]. Vetter [Ben17]. Vice [Git10]. Vicky [Ros13]. Victoria [Bed18]. Victorian [Blu12, Coo12a, For10, For17, Coo11c, Ham17, Kap19]. Victory [Kni18]. Video [Col15, Spi15]. Videogame [Hjo16]. Videogames [Jin17]. Videotape [Git10]. Vietnam [Dér18]. View [Jac14, Mar13]. Viewer [Edw17]. Views [Olw10]. Viguié [Lag15]. Vincent [Cas15]. Vinci [CG16]. Vinyl [Ves19]. Violent [Sto19]. Virginia [Sta14a, Sta14b]. Virtual [Cha11a, Goo19a, Jeo15, Nye15, Gra15]. Virtuous [Lev13a]. VISA [Cor12b]. Visible [Joh11b]. Vision [Fja15, For10, Sim17b, Tay17]. Visionary [Cor10]. Visioneers [Tho15]. Visions [Abb18, Est10, Kim15, McA14, Pra13, Rom10, Wes16]. Visual [Goo19a, Hec14, Hri16, Kas17, Wri12b]. Visualised [Hog14a]. Visualization [SW15]. Visualizing [Pua17]. Vital [Ven19]. Vitiello [Zyl15]. Vitz [Bea19a]. Vleck [Cor14]. Vleuten [Gif17b]. Vocal [LM19]. Vocation [New10]. Voice [Han13, LM19, Olw18, Wol10]. Voices [Jos16b]. Voiture [Har12]. Vol [Ros12b]. Volga [Jos16a]. Volkswagen [Lad14]. vols [Col15]. Volume [Ano10f, Ano10h, Ano10g, Ano10i, Ano10j, Ano11d, Ano11f, Ano11e, Ano11g, Ano11h, Ano12d, Ano12f, Ano12e, Ano12g, Ano12h, Ano13c, Ano13e, Ano13d, Ano13f, Ano13g, Ano14c, Ano14e, Ano14d, Ano14f, Ano14g, Ano15c, Ano15e, Ano15d, Ano15f, Ano15g, Ano16c, Ano16e, Ano16d, Ano16f, Ano16g, Ano17c, Ano17e, Ano17d, Ano17f, Ano17g, Ano18c, Ano18e, Ano18d, Ano18f, Ano18g, Ano18h, Ano19d, Ano19f, Ano19e, Ano19h, Ano19i, Ano19g, Sch14c]. Volunteer [See15b]. Volunteer-based [See15b]. vom [Pop13b, Tod11]. Votes [Pfa13]. Voth [Bra15a]. Vralsted [Jos16a]. vs [Wak11]. Vulcanization [GR10]. Vulnerability [Moh15].

W [Bow14, Bro13, Deu17, Gua13, Her17a, Hir17, Hog18b, Man13, McC13, Nye13, Nye17, Pfa13, Pua17, Tho15, Ver15, Wes16]. W. [Moh16, Smi10b]. Wages [Sch11c]. Wal [Kle10]. Wal-Mart [Kle10]. Walkman [Kre10]. Wallace [Arc13a]. Wallinger [Joh17]. Walmart [Bar18]. Walter [Bac15, Gai15, Kai10]. Wanamaker [Arc10]. Wandering [Cam18]. Want [Nof18]. War [Ala12, Ang15, BF19, Bes13, Caf18, Car13a, Car13c, DK18, Day11, Dou16, Edw17, Eri15, Gan17b, Gif17b, God15, Gra15, Gra18, Ham16a, Ham18b, Ham14b, Hri16, Jør10, Kat11, Kat10, Kru14, Lag17, Mar19, Moo15, Mor12b, Mue15, Nah13, Ore11, Rol10a, Rol17, Rub18, Sat10a, Sat13b, Sla16, Tin18, Ven12, Wea16, Wor16b, vdV12, Cam15, DiM10, Gro19, Lin10, Mat19, Rot16, Ang17, Boc16, Kar17, Kli19, Kni18, Las18, Nah13, Bar10, Bes13, Car13c, Eng12, Fir12, Fja12, Iri12, Mar11b, Sla12a, Sta10a, Ven12]. Warde [Spr19]. Warfare [Dou16, Ham18b, Sat10a, Wol11a, vdV12]. Wargaming [Caf18]. Warm [SH15]. Warming [Eli12, Rob12]. Warner [Sch13b]. Wars [Ens10, Kra11, Wak11, Wol11a]. Wartime [Bra19, Bro14b, Ste19]. Was [Blu12, Nov14, Rol10d]. Washington [Rol10b, Seg14, Wil13]. Waste [Bal11b, Car13c, Coo11c, Fja13, Mit18, Zim13, Kru13a]. Water [BM12, Cit12, Cro18, Ert16, Eva16, Har16, KO13, KO14, Mac16b, Par18, Rol10b, Sha17, Shu19, Son17, Sut12, Tom14, Tom15, Tom17, Van16]. Watering [Bea19b]. Waterlines [FP16]. Waterpower [Cas14, Bro11]. Waters [Eva18, KO12, Sum17, Whi10]. Waterways [Sau11, Sha10]. Watt [Joh11a]. Wave [FP16, New18]. Wave-Line [FP16]. Waves [Mun14]. Way [Fre11, Mey19, Rol10a, Wil10, Mat17a, Low13]. Ways [Jac11b, WH11, Ver15]. Weale [Rat12]. Weapon [Mac18]. Weapons [Car13a, Fai16, Joh12, Kem12, Kru13b, Nah13, Sla11c, Sla16]. Wear [Jen11]. Wearable [Ret16]. Wears [Mas17]. Weather [Coo11b]. Weaving [MB18]. Web [Cor19, Mal17]. Webb [Nee17]. Weekend [McC17]. Weems [Sim17b]. Wei [Hu14]. Weimar [Woo19]. Weinberg [Joh18b]. Weiner [Kru13b]. Weiss [Rae16, Gal17, Rol16]. Weitekamp [Low14]. Welch [Hac15]. Welcome [JJ16]. Welcoming [BH17, Sta12]. Welfare [Tay14, vdV12]. Well [Ore12, Tay10]. Well-Being [Ore12]. Well-Pressed [Tay10]. Wells [Nye13]. Wendel [Pur17]. Wendy [Dup15, Hic14, Low13]. Went [Mis11a]. Were [Aik19, Nov11]. Werk [Kai10]. Werrett [Fle19]. West [Ben17, De 15, Koe18, Ran11, Sch14a, Wro10, Fre12, Mac16a, Neu14b, Bye13, Cut10, Kru12]. West-Centrism [Pra19]. Western [Bor17, Hai15, Koc16, Wro10, Cam15, Fru14]. Western-Style [Fru14]. Westwick [Kin14]. Whaling [Osl10]. Wheel [Fra18, Nye17, Bra15b, Hög13b]. Wheeled [Vol17]. Wheels [Woo19]. Where [Mun16, Pat19]. White [Ded18, Par16]. White-Goods [Par16]. Whitehead [Gra15]. Whiteness [Zyl11b, dlP10]. Whitney [Nel16]. Who [Bea10, Bow15, Bra10, Bye19, Cha18b, Gag15, Jos16b, Ryd11, Kap19, Bre10a]. Widescreen [Mil12a]. Wiebe [Moh15]. Wiederaufbau [Mor12a]. Wilderness [Mit12, Cas14]. Wildlife [Bru19, Mit12]. Wilgus [Ric13]. Will [Nie12, Pfa13]. Willem [Dus18]. William [Gal17, Ham18a, Hor18, Hun16, Man18, Ric13, Sti18, Cut10]. Williams [Dal15]. Willy [De 19]. Wilson [Hah15]. Winchester [Bra10]. Wind [Bre16, BH17, HS13, Kap19, See15a, See15a]. Winder [Coc17b]. Windmills [Cou17b]. Winds [McN14]. Wing [McC15]. Winslow [Aga18]. Winter [Olw10]. Winters [Ves19]. Wired [Hir17, Mit12, Vet19]. Wireless [Aga18, Mun14, Sat10a]. Wires [Mar19]. Wiretapping [Lap12]. Wiring [Mül16a, Bek18]. Wiseguys [Pet13]. Wisnioski [Reu14]. Wissen [Har17]. Wissenschaft [Gif17a, Ole17b]. within [Pat19]. without [Dou19, Lee13]. Witness [Bou19]. Witnessing [Leg11, Smi14a]. Wittje [Rud13]. Wizard [See15a]. Wolfe [Olw14, Bea19b]. Wolfgang [Gif17a, Sza15]. Wolfram [vM16]. Wolters [Fel14]. Women [Cla10, Col12a, Hay14, MF17, McC14a, Nov11, Tin19, Wei13, WH11, Sch15a, Fra18, Pan15]. Won [Kni18, Wro10]. Wonder [Hen18a, Tel13]. Wonders [Seg14]. Wood [Man18, Mit18]. Woods [Ore11]. Wool [Ros10]. Woolworth [Les10]. Word [Kur19a]. Words [WH11]. Work [Blu12, Cla13, Dus11, Eis10, For11, Got12, Ham16b, Kra10, Pan15, Pat19, Rob10c, Rob17, Ros12b, Sat13a, Sch11c, SDJ19, Kas17, Kel11, Sop15]. Worker [Dry19]. Workers [Hep16, WH11]. Working [Gor15, Kum19a, Ore12, Pac18, Ziz19, San12]. Works [Moo16a]. Workspaces [Hoa15]. World [BF19, Bek18, Boc16, Car13c, Dob10, Dot12, Fre11, Goo19a, Gor10, Hag14, Ham11, Hea13a, Hen18b, Hir11, Hjo16, Hou10, Hug15, Jas12, Jep18, Jon15, Lon12a, Moh16, Mor12b, Mül16a, Nah13, Neu14a, Nye11, Ore11, Pat11, Reg18, Rol17, Sin12, Sla19, SH19b, Sta12, Ven12, Wol11a, Wol14, Wor16b, vM12b, Ham14b, Mar19, Cha18b, Dal15, Hen18a, KO14, Sla18, Ste16b, Sut19, Whi14, Bui10, Gan17b, Gor12, Kel10, Sol11, Wor16b]. Worlds [Ang17, Bre18, Cha11a, Mai16]. Worldwide [Mil12a]. Worrying [vMS15]. Worst [Kru13b, Vol11]. Worth [Ket16]. Would [Jen11]. Wright [Hoo15]. Write [HNG16]. Writers [All11]. Writing [Sla10, Sch19, vM16]. Writings [Nye15, Rob10c]. Wrong [Kon10]. Wunsch [Wil17a]. WWII [Caf18].

x [Wos17]. Xiaoping [Hai15]. XIX [But12, Gil10, Pan18b]. XIXe [Gui16]. XV [PT13]. XVIII [Ora15]. XX [But12].

Yanek [Neu13]. Yang [Tsu12]. Yankee [O’H10]. Yard [SH19b]. Yeang [Mun14]. year [DL18]. Years [BBB^{+}19, Jac19, Mar13, Neu14a, Tin15, Clo12]. Yellow [Son17]. Yen [Ree16]. Yi [Low16a]. Yood [Wes15]. York [Arc13a, BM14, Koe18, Lae15, Mac16b, Ole17a, Pik15, Sti11, Les10]. Yorkshireman [O’H10]. Yost [Sla18]. Youth [Luc18b]. Yucca [Bal11b]. Yugo [Vol11]. Yuri [Nem14]. Yves [Has16].

Z [Spi15]. Zabet [Tay17]. Zachary [God18]. Zachmann [Wil14b]. Zarecor [De 12]. Zealand [Pet19]. Zeisler [Jos16a]. Zeisler-Vralsted [Jos16a]. Zelda [Jin17]. Zhores [Sid12]. Zimbabwe [Hoa15]. Zink [Bro13]. Zollverein [Hau11]. Zone [Wes15]. Zones [Sis12]. Zoom [Hal16]. Zoomar [Hal16]. Zum [Die12, Pop13b, Tod11]. Zur [Kre10, Tod11]. zwischen [Pet12].

References

Abbott:2011:ECM

[Abb11] Carl Abbott. Eden on the Charles: The Making of Boston (review). Technology and Culture, 52(3):644–645, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447801.

Abbott:2015:CDO

[Abb15] Carl Abbott. Community by Design: The Olmsted Form and the Development of Brookline, Massachusetts by Keith N. Morgan, Elizabeth Hope Cushing, and Roger G. Reed (review). Technology and Culture, 56(2):544–545, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/582329.

Abbate:2018:CSA

[Abb18] Janet Abbate. Code switch: Alternative visions of computer expertise as empowerment from the 1960s to the 2010s. Technology and Culture, 59(4S):S134–S159, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/712116.

Aaberg:2015:KEI

[Åbe15] Anna Åberg. Knowledge and Evidence: Investigating Technologies in Practice ed. by Boel Berner, Corinna Kruse (review). Technology and Culture, 56(3):749–750, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/593072.

Abnet:2019:CDC

[Abn19] Dustin Abnet. Control: Digitality as Cultural Logic by Seb Franklin (review). Technology and Culture, 60(4):1128–1129, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/741399.

Abraham:2013:BNA

[Abr13] Itty Abraham. Being Nuclear: Africans and the Global Uranium Trade by Gabrielle Hecht (review). Technology and Culture, 54 (4):996–997, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530537.

Acland:2017:AAE

[Acl17] Charles R. Acland. American AV: Edgar Dale and the Information Age classroom. Technology and Culture, 58(2):392–421, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662982.

Adas:2017:FTE

[Ada17] Michael Adas. The flawed “Tools of empire”: Clive Dewey, steamboats on the Indus. Technology and Culture, 58(2):578–580, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662991.

Agar:2011:MCI

[Aga11] Jon Agar. Mobile Communications: An Introduction to New Media (review). Technology and Culture, 52(3):664–665, July 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/447811.

Agar:2014:TGE

[Aga14] Jon Agar. Taken for Grantedness: The Embedding of Mobile Communication into Society by Rich Ling (review). Technology and Culture, 55(4):1035–1037, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562867.

Agar:2016:III

[Aga16] Jon Agar. Information and Intrigue: From Index Cards to Dewey Decimals to Alger Hiss by Colin B. Burke (review). Technology and Culture, 57(4):1019–1020, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/637946.

Agar:2018:PST

[Aga18] Jon Agar. Political Spectrum: The Tumultuous Liberation of Wireless Technology, from Herbert Hoover to the Smartphone by Thomas Winslow Hazlett (review). Technology and Culture, 59 (3):806–808, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/703159.

Anderson:2018:CST

[AH18] Katherine Anderson and Jan Hadlaw. The Canada Science and Technology Museum. Technology and Culture, 59(3):781–786, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/703147.

Aidoo:2019:BSA

[Aid19] Fallon Samuels Aidoo. The Black Skyscraper: Architecture and the Perception of Race by Adrienne Brown (review). Technology and Culture, 60(1):324–326, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719959.

Aiken:2019:ECL

[Aik19] Katherine G. Aiken. “The environmental consequences … were calamitous”: Smelter smoke controversies in Progressive Era America, 1899–1918. Technology and Culture, 60(1):132–164, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/719948.

Alacevich:2012:RKR

[Ala12] Michele Alacevich. The Right Kind of Revolution: Modernization, Development, and U.S. Foreign Policy from the Cold War to the Present (review). Technology and Culture, 53(3):727–729, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/481931.

Aldrich:2010:ROS

[Ald10] Mark Aldrich. Railroads in the Old South: Pursuing Progress in a Slave Society (review). Technology and Culture, 51(2):510–511, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381265.

Aldrich:2016:EAN

[Ald16a] Mark Aldrich. Engineers attack the “No. One Killer” in coal mining: The Bureau of Mines and the promotion of roof bolting, 1947– 1969. Technology and Culture, 57(1):80–118, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611801.

Aldrich:2016:RM

[Ald16b] Mark Aldrich. Response to mark. Technology and Culture, 57 (3):717–719, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630962.

Alic:2014:E

[Ali14] John A. Alic. To the Editor. Technology and Culture, 55(1): 282–285, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538939.

Allen:2011:BPR

[All11] Frederick E. Allen. A Better Pencil: Readers, Writers, and the Digital Revolution (review). Technology and Culture, 52(2):418– 420, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431386.

Alper:2016:MCA

[Alp16] Meryl Alper. Making Computers Accessible: Disability Rights and Digital Technology by Elizabeth R. Petrick (review). Technology and Culture, 57(4):1043–1044, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/637960.

Ammon:2011:URA

[Amm11] Francesca Russello Ammon. Untimely Ruins: An Archaeology of American Urban Modernity, 1819–1919 (review). Technology and Culture, 52(4):828–829, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/454551.

Ammon:2012:UBB

[Amm12] Francesca Russello Ammon. Unearthing Benny the Bulldozer: The culture of clearance in postwar children’s books. Technology and Culture, 53(2):306–336, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/476807.

Ammon:2014:PMD

[Amm14] Francesca Russello Ammon. Pastoral and Monumental: Dams, Postcards, and the American Landscape by Donald C. Jackson (review). Technology and Culture, 55(4):1012–1014, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/562854.

Anbinder:2018:HIE

[Anb18] Jacob Anbinder. Historicizing Infrastructure ed. by Andreas Marklund and Mogens Rüdiger (review). Technology and Culture, 59 (3):791–792, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/703150.

Anderson:2010:GAR

[And10] J. L. Anderson. Growing American Rubber: Strategic Plants and the Politics of National Security (review). Technology and Culture, 51(4):1033–1034, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/403290.

Anderson:2014:MRF

[And14] J. L. Anderson. Mark R. Finlay, 15 September 1960–6 October 2013. Technology and Culture, 55(1):223–226, January 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/538916.

Andersen:2015:EMF

[And15] Casper Andersen. Engineers and the Making of the Francoist Regime by Lino Camprubí (review). Technology and Culture, 56 (3):767–768, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/593082.

Andersen:2017:IEU

[And17] Casper Andersen. Internationalism and engineering in UNESCO during the end game of empire, 1943–68. Technology and Culture, 58(3):650–677, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669464.

Angevine:2015:CWS

[Ang15] Robert G. Angevine. The Civil War at Sea by Craig L. Symonds (review). Technology and Culture, 56(2):538–540, April 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/582326.

Angevine:2017:ATW

[Ang17] Robert G. Angevine. Astride Two Worlds: Technology and the American Civil War ed. by Barton C. Hacker (review). Technology and Culture, 58(4):1081–1083, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/679830.

Anonymous:2010:A

[Ano10a] Anonymous. Awards. Technology and Culture, 51(1):164–180, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
375779.

Anonymous:2010:Ia

[Ano10b] Anonymous. In this issue. Technology and Culture, 51(2):a–c, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381252.

Anonymous:2010:Ib

[Ano10c] Anonymous. In this issue. Technology and Culture, 51(3):a–c, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/390835.

Anonymous:2010:Ic

[Ano10d] Anonymous. In this issue. Technology and Culture, 51(4):a–d, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
408073.

Anonymous:2010:PMO

[Ano10e] Anonymous. The Pittsburgh Meeting, 15–18 October 2009. Technology and Culture, 51(1):181–198, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/375778.

Anonymous:2010:V

[Ano10f] Anonymous. Volume 51 (2010). Technology and Culture, 51(4): 1063–1088, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403280/pdf.

Anonymous:2010:VNJa

[Ano10g] Anonymous. Volume 51, number 1, January 2010. Technology and Culture, 51(1):??, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2010:VNA

[Ano10h] Anonymous. Volume 51, number 2, April 2010. Technology and Culture, 51(2):??, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2010:VNJb

[Ano10i] Anonymous. Volume 51, number 3, July 2010. Technology and Culture, 51(3):??, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2010:VNO

[Ano10j] Anonymous. Volume 51, number 4, October 2010. Technology and Culture, 51(4):??, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2011:A

[Ano11a] Anonymous. Awards. Technology and Culture, 52(1):143–158, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
416411.

Anonymous:2011:I

[Ano11b] Anonymous. In this issue. Technology and Culture, 52(3):i, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447776.

Anonymous:2011:TMS

[Ano11c] Anonymous. The Tacoma Meeting, 30 September–3 October 2010. Technology and Culture, 52(1):159–170, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/416412.

Anonymous:2011:V

[Ano11d] Anonymous. Volume 52 (2011). Technology and Culture, 52(4): 866–890, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454536.

Anonymous:2011:VNJa

[Ano11e] Anonymous. Volume 52, number 1, January 2011. Technology and Culture, 52(1):??, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2011:VNA

[Ano11f] Anonymous. Volume 52, number 2, April 2011. Technology and Culture, 52(2):??, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2011:VNJb

[Ano11g] Anonymous. Volume 52, number 3, July 2011. Technology and Culture, 52(3):??, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2011:VNO

[Ano11h] Anonymous. Volume 52, number 4, October 2011. Technology and Culture, 52(4):??, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2012:A

[Ano12a] Anonymous. Awards. Technology and Culture, 53(1):146–160, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466775.

Anonymous:2012:Ia

[Ano12b] Anonymous. In this issue. Technology and Culture, 53(3):a–b, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481904.

Anonymous:2012:Ib

[Ano12c] Anonymous. In this issue. Technology and Culture, 53(4):viii, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
490726.

Anonymous:2012:IV

[Ano12d] Anonymous. Index: Volume 53 (2012). Technology and Culture, 53 (4):965–989, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490761.

Anonymous:2012:VNJa

[Ano12e] Anonymous. Volume 53, number 1, January 2012. Technology and Culture, 53(1):??, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2012:VNA

[Ano12f] Anonymous. Volume 53, number 2, April 2012. Technology and Culture, 53(2):??, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2012:VNJb

[Ano12g] Anonymous. Volume 53, number 3, July 2012. Technology and Culture, 53(3):??, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2012:VNO

[Ano12h] Anonymous. Volume 53, number 4, October 2012. Technology and Culture, 53(4):??, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2013:A

[Ano13a] Anonymous. Awards. Technology and Culture, 54(1):130–143, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500791.

Anonymous:2013:CMO

[Ano13b] Anonymous. The Copenhagen Meeting, 4-7 October 2012. Technology and Culture, 54(1):144–157, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500792.

Anonymous:2013:V

[Ano13c] Anonymous. Volume 54 (2013). Technology and Culture, 54(4): 1011–1034, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530545.

Anonymous:2013:VNJa

[Ano13d] Anonymous. Volume 54, number 1, January 2013. Technology and Culture, 54(1):??, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2013:VNA

[Ano13e] Anonymous. Volume 54, number 2, April 2013. Technology and Culture, 54(2):??, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2013:VNJb

[Ano13f] Anonymous. Volume 54, number 3, July 2013. Technology and Culture, 54(3):??, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2013:VNO

[Ano13g] Anonymous. Volume 54, number 4, October 2013. Technology and Culture, 54(4):??, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2014:A

[Ano14a] Anonymous. Awards. Technology and Culture, 55(1):186–197, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
538912.

Anonymous:2014:PMO

[Ano14b] Anonymous. The Portland Meeting, 10–13 October 2013. Technology and Culture, 55(1):198–210, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/538913.

Anonymous:2014:V

[Ano14c] Anonymous. Volume 55 (2014). Technology and Culture, 55(4): 1038–1057, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562868.

Anonymous:2014:VNJa

[Ano14d] Anonymous. Volume 55, number 1, January 2014. Technology and Culture, 55(1):??, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2014:VNA

[Ano14e] Anonymous. Volume 55, number 2, April 2014. Technology and Culture, 55(2):??, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2014:VNJb

[Ano14f] Anonymous. Volume 55, number 3, July 2014. Technology and Culture, 55(3):??, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2014:VNO

[Ano14g] Anonymous. Volume 55, number 4, October 2014. Technology and Culture, 55(4):??, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2015:A

[Ano15a] Anonymous. Awards. Technology and Culture, 56(1):209–223, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572117.

Anonymous:2015:DMN

[Ano15b] Anonymous. The Dearborn Meeting, 6–9 November 2014. Technology and Culture, 56(1):224–240, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/572118.

Anonymous:2015:IV

[Ano15c] Anonymous. Index: Volume 56 (2015). Technology and Culture, 56 (4):1016–1035, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601503.

Anonymous:2015:VNJa

[Ano15d] Anonymous. Volume 56, number 1, January 2015. Technology and Culture, 56(1):??, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2015:VNA

[Ano15e] Anonymous. Volume 56, number 2, April 2015. Technology and Culture, 56(2):??, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2015:VNJb

[Ano15f] Anonymous. Volume 56, number 3, July 2015. Technology and Culture, 56(3):??, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2015:VNO

[Ano15g] Anonymous. Volume 56, number 4, October 2015. Technology and Culture, 56(4):??, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2016:AMO

[Ano16a] Anonymous. The Albuquerque Meeting, 8–11 October 2015. Technology and Culture, 57(1):196–215, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611805.

Anonymous:2016:A

[Ano16b] Anonymous. Awards. Technology and Culture, 57(1):183–195, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
611804.

Anonymous:2016:V

[Ano16c] Anonymous. Volume 57 (2016). Technology and Culture, 57(4): 1047–1074, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637962.

Anonymous:2016:VNJa

[Ano16d] Anonymous. Volume 57, number 1, January 2016. Technology and Culture, 57(1):??, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2016:VNA

[Ano16e] Anonymous. Volume 57, number 2, April 2016. Technology and Culture, 57(2):??, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2016:VNJb

[Ano16f] Anonymous. Volume 57, number 3, July 2016. Technology and Culture, 57(3):??, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2016:VNO

[Ano16g] Anonymous. Volume 57, number 4, October 2016. Technology and Culture, 57(4):??, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2017:A

[Ano17a] Anonymous. Awards. Technology and Culture, 58(1):206–217, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
648258.

Anonymous:2017:SMJ

[Ano17b] Anonymous. The Singapore Meeting, 22–26 June 2016. Technology and Culture, 58(1):218–249, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648259.

Anonymous:2017:V

[Ano17c] Anonymous. Volume 58 (2017). Technology and Culture, 58(4): 1115–1138, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679849.

Anonymous:2017:VNJa

[Ano17d] Anonymous. Volume 58, number 1, January 2017. Technology and Culture, 58(1):??, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2017:VNA

[Ano17e] Anonymous. Volume 58, number 2, April 2017. Technology and Culture, 58(2):??, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2017:VNJb

[Ano17f] Anonymous. Volume 58, number 3, July 2017. Technology and Culture, 58(3):??, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2017:VNO

[Ano17g] Anonymous. Volume 58, number 4, October 2017. Technology and Culture, 58(4):??, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2018:A

[Ano18a] Anonymous. Awards. Technology and Culture, 59(2):439–453, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/697238.

Anonymous:2018:PMO

[Ano18b] Anonymous. The Philadelphia Meeting, 26–29 October 2017. Technology and Culture, 59(2):410–438, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/697237.

Anonymous:2018:V

[Ano18c] Anonymous. Volume 59 (2018). Technology and Culture, 59(4): 1003–1025, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711792.

Anonymous:2018:VNJa

[Ano18d] Anonymous. Volume 59, number 1, January 2018. Technology and Culture, 59(1):??, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2018:VNA

[Ano18e] Anonymous. Volume 59, number 2, April 2018. Technology and Culture, 59(2):??, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2018:VNJb

[Ano18f] Anonymous. Volume 59, number 3, July 2018. Technology and Culture, 59(3):??, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2018:VNO

[Ano18g] Anonymous. Volume 59, number 4, October 2018. Technology and Culture, 59(4):??, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2018:VNS

[Ano18h] Anonymous. Volume 59, number 4 supplement, October 2018. Technology and Culture, 59(4S):??, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

/journal/194.

Anonymous:2019:A

[Ano19a] Anonymous. Awards. Technology and Culture, 60(1):278–292, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
719953.

Anonymous:2019:EN

[Ano19b] Anonymous. Editor’s note. Technology and Culture, 60(2):553, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
726950.

Anonymous:2019:SLM

[Ano19c] Anonymous. The Saint Louis Meeting, 11–14 October 2018. Technology and Culture, 60(1):293–316, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/719954.

Anonymous:2019:V

[Ano19d] Anonymous. Volume 60 (2019). Technology and Culture, 60(4): 1132–1154, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741388.

Anonymous:2019:VNJa

[Ano19e] Anonymous. Volume 60, number 1, January 2019. Technology and Culture, 60(1):??, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2019:VNA

[Ano19f] Anonymous. Volume 60, number 2, April 2019. Technology and Culture, 60(2):??, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2019:VNS

[Ano19g] Anonymous. Volume 60, number 2 supplement, April 2019. Technology and Culture, 60(2S):??, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/

194.

Anonymous:2019:VNJb

[Ano19h] Anonymous. Volume 60, number 3, July 2019. Technology and Culture, 60(3):??, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2019:VNO

[Ano19i] Anonymous. Volume 60, number 4, October 2019. Technology and Culture, 60(4):??, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Ansani:2017:LRG

[Ans17] Fabrizio Ansani. The life of a Renaissance gunmaker: Bonaccorso Ghiberti and the development of Florentine artillery in the Late Fifteenth Century. Technology and Culture, 58(3):749–789, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/669467.

Apple:2014:CCH

[App14] Rima D. Apple. Creating Consumers: Home Economics in Twentieth-Century America by Carolyn M. Goldstein (review).

Technology and Culture, 55(1):259–261, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/538927.

Apple:2017:BBN

[App17] Rima D. Apple. Back to the Breast: Natural Motherhood and Breastfeeding in America by Jessica Martucci (review). Technology and Culture, 58(2):595–596, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/663000.

Arceneaux:2010:WDS

[Arc10] Noah Arceneaux. Wanamaker’s department store and the origins of electronic media, 1910–1922. Technology and Culture, 51 (4):809–828, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403266.

Arceneaux:2012:NNT

[Arc12] Noah Arceneaux. Network Nations: A Transnational History of British and American Broadcasting (review). Technology and Culture, 53(4):951–952, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/490753.

Arceneaux:2013:MCA

[Arc13a] Noah Arceneaux. Media Capital: Architecture and Communication in New York City by Aurora Wallace (review). Technology and Culture, 54(4):983–985, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530530.

Arceneaux:2013:TAD

[Arc13b] Noah Arceneaux. The Telegraph in America, 1832–1920 by David Hochfelder (review). Technology and Culture, 54(3):665–666, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520293.

Arceneaux:2018:LSF

[Arc18] Noah Arceneaux. Lost Sound: The Forgotten Art of Radio Storytelling by Jeff Porter (review). Technology and Culture, 59(1): 199–200, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/692163.

Arceneaux:2019:MRE

[Arc19] Noah Arceneaux. Making Radio: Early Radio Programming and the Rise of Modern Sound Culture by Shawn VanCour (review). Technology and Culture, 60(1):335–337, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/719965.

Arnold:2018:MTE

[Arn18] David Arnold. A Moral Technology: Electrification as Political Ritual in New Delhi by Leo Coleman (review). Technology and Culture, 59(2):478–480, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/697244.

Aso:2012:FSS

[Aso12] Mitch Aso. The Floracrats: State-Sponsored Science and the Failure of the Enlightenment in Indonesia (review). Technology and Culture, 53(1):223–225, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/466794.

Avanti:2015:CTR

[Ava15] Peter Avanti. Captivating technologies: Reflections on the equaltempered diatonic keyboard. Technology and Culture, 56(1): 150–183, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572114.

Backhouse:2015:FTS

[Bac15] Roger E. Backhouse. Fortune Tellers: The Story of America’s First Economic Forecasters by Walter A. Friedman (review). Technology and Culture, 56(2):547–549, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/582331.

Baillargeon:2012:JLB

[Bai12] Denyse Baillargeon. Jules-Louis Breton 1878–1940: Un savant parlementaire (review). Technology and Culture, 53(2):495–496, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476815.

Bailey:2018:RAI

[Bai18] Robert Bailey. Rethinking art as interrelation. Technology and Culture, 59(2):464–472, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/697240.

Balbi:2010:RBR

[Bal10] Gabriele Balbi. Radio before radio: Araldo telefonico and the invention of Italian broadcasting. Technology and Culture, 51(4): 786–808, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403265.

Balbi:2011:VMB

[Bal11a] Gabriele Balbi. Verso la Modernità: I bresciani e le esposizioni industriali 1800–1915 (review). Technology and Culture, 52(4): 831–833, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454553.

Balogh:2011:RYM

[Bal11b] Brian Balogh. The Road to Yucca Mountain: The Development of Radioactive Waste Policy in the United States (review). Technology and Culture, 52(2):417–418, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/431385.

Balzer:2011:SCS

[Bal11c] Harley Balzer. The social construction of Sputnik. Technology and Culture, 52(3):614–617, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/447786.

Bantjes:2016:SPR

[Ban16] Rod Bantjes. Scenes of Projection: Recasting the Enlightenment Subject by Jill H. Casid (review). Technology and Culture, 57(4): 1010–1012, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637941.

Barnhill:2010:KCA

[Bar10] John Barnhill. Killing for Coal: America’s Deadliest Labor War (review). Technology and Culture, 51(1):257–258, January 2010.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375816.

Barnes:2012:CCB

[Bar12] David S. Barnes. Confronting the Climate: British Airs and the Making of Environmental Medicine (review). Technology and Culture, 53(4):927–928, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/490739.

Barber:2014:HCD

[Bar14a] Daniel A. Barber. A Handbook of California Design, 1930–1965: Craftspeople, Designers, Manufacturers by Bobbye Tigerman (review). Technology and Culture, 55(1):263–265, January 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/538929.

Barber:2014:THS

[Bar14b] Daniel A. Barber. Tomorrow’s house: Solar housing in 1940s America. Technology and Culture, 55(1):1–39, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/538907.

Barnhill:2016:TDS

[Bar16a] John H. Barnhill. Taconite Dreams: The Struggle to Sustain Mining on Minnesota’s Iron Range, 1915–2000 by Jeffrey T. Manuel (review). Technology and Culture, 57(4):1015–1017, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637944.

Barrett:2016:CNF

[Bar16b] Ross Barrett. Carbon Nation: Fossil Fuels in the Making of American Culture by Bob Johnson (review). Technology and Culture, 57 (1):258–259, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611819.

Barber:2018:RLW

[Bar18] Daniel A. Barber. Rule of Logistics: Walmart and the Architecture of Fullfillment by Jesse LeCavalier (review). Technology and Culture, 59(4):995–997, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711788.

Battles:2011:PDG

[Bat11] Kathleen Battles. Points on the Dial: Golden Age Radio beyond the Networks (review). Technology and Culture, 52(3):654–655, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447806.

Battles:2014:TSL

[Bat14] Matthew Battles. Technology at sea: Lucien Castaing-Taylor and Véréna Paravel, Leviathan. Technology and Culture, 55(2):479–481, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546874.

Baucom:2011:AOR

[Bau11] Donald Baucom. The Adaptive Optics Revolution: A History (review). Technology and Culture, 52(2):412–413, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/431382.

Belot:2019:FFS

[BBB^{+}19] Robert Belot, Yoel Bergman, Hans-Joachim Braun, Jan Hadlaw, Stefan Poser, and Magdalena Zdrodowska. Forty-fifth Symposium of the International Committee for the History of Technology: “Technological Drive from Past to Future? 50 Years of ICOHTEC,” Jean Monnet University, Saint-Étienne, France, 17–21 July 2018. Technology and Culture, 60(4):1083–1092, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741384.

Bartha:2015:FFS

[BBŁT15] Biborka Bartha, Yoel Bergman, Sławomir Łotysz, and Anna Turza. Forty-first symposium of the International Committee for the History of Technology: “Technology in Times of Transition,” Brasov, Romania, 29 July–2 August 2014. Technology and Culture, 56 (1):241–247, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572119.

Bix:2013:TO

[BE13] Amy Sue Bix and Taner Edis. Trapped by opportunity. Technology and Culture, 54(1):170–174, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500799.

Beauchamp:2010:WIT

[Bea10] Christopher Beauchamp. Who invented the telephone?: Lawyers, patents, and the judgments of history. Technology and Culture, 51 (4):854–878, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403268.

Beatty:2011:JLM

[Bea11] Edward Beatty. Jungle Laboratories: Mexican Peasants, National Projects, and the Making of the Pill (review). Technology and Culture, 52(4):840–842, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/454557.

Beatty:2019:CLU

[Bea19a] Edward Beatty. A City on a Lake: Urban Political Ecology and the Growth of Mexico City by Matthew Vitz (review). Technology and Culture, 60(3):904–905, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/731464.

Beatty:2019:WRE

[Bea19b] Edward Beatty. Watering the Revolution: An Environmental and Technological History of Agrarian Reform in Mexico by Mikael D. Wolfe (review). Technology and Culture, 60(1):317–319, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/719955.

Bednarek:2010:TPR

[Bed10] Janet R. Daly Bednarek. Theme Park (review). Technology and Culture, 51(1):270–272, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375815.

Bedi:2018:LDR

[Bed18] Joyce Bedi. Life on Display: Revolutionizing U.S. Museums of Science and Natural History in the Twentieth Century by Karen A. Rader, Victoria E. M. Cain (review). Technology and Culture, 59(1):168–169, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/692172.

Bektas:2018:WWS

[Bek18] Yakup Bektas. Wiring the World: The Social and Cultural Creation of Global Telegraph Networks by Simone M. Müller (review). Technology and Culture, 59(3):794–796, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/703152.

Benhamou:2010:SSE

[Ben10] Reed Benhamou. Science and Spectacle in the European Enlightenment (review). Technology and Culture, 51(2):493–495, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381260.

Bennett:2014:APR

[Ben14] Jim Bennett. Artisan/Practitioners and the Rise of the New Sciences, 1400–1600 by Pamela O. Long (review). Technology and Culture, 55(1):254–256, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/538924.

Benson:2017:FLS

[Ben17] Etienne Benson. Field Life: Science in the American West during the Railroad Era by Jeremy Vetter (review). Technology and Culture, 58(4):1083–1084, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679831.

Berger:2011:BLE

[Ber11] Molly Berger. The Banham Lectures: Essays on Designing the Future (review). Technology and Culture, 52(4):860–862, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454569.

Bergman:2012:TES

[Ber12] Yoel Bergman. Thirty-eighth symposium of the International Committee for the History of Technology: “Consumer Choice and Technology,” Glasgow, UK, 2–7 August 2011. Technology and Culture, 53(1):177–184, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466773.

Bertucci:2013:ESS

[Ber13] Paola Bertucci. Enlightened secrets: Silk, intelligent travel, and industrial espionage in Eighteenth-Century France. Technology and Culture, 54(4):820–852, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530513.

Bertucci:2014:GTH

[Ber14] Paola Bertucci. Gusto for Things: A History of Objects in Seventeenth-Century Rome by Renata Ago (review). Technology and Culture, 55(3):742–743, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551620.

Bernardin:2015:TPP

[Ber15] Stève Bernardin. “Taking the problem to the people”: Traffic safety from public relations to political theory, 1937–1954. Technology and Culture, 56(2):420–439, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/582315.

Bernardin:2016:RRB

[Ber16] Stève Bernardin. From Rail to Road and Back Again? A Century of Transport Competition and Interdependency ed. by Ralf Roth, Colin Divall (review). Technology and Culture, 57(4):1014–1015, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
637943.

Beretta:2017:GGP

[Ber17] Marco Beretta. Glassmaking goes public: The cultural background to Antonio Neri’s L’Arte Vetraria (1612). Technology and Culture, 58(4):1046–1070, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/679823.

Bergman:2018:MIR

[Ber18] James Bergman. Make It Rain: State Control of the Atmosphere in Twentieth-Century America by Kristine C. Harper (review). Technology and Culture, 59(4):988–990, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711784.

Berger:2019:CCM

[Ber19] Rachel Berger. Clarified commodities: Managing ghee in interwar India. Technology and Culture, 60(4):1004–1026, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/741381.

Bessner:2013:WGH

[Bes13] Daniel Bessner. War Games: A History of War on Paper by Philipp von Hilgers (review). Technology and Culture, 54(2):395– 396, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506080.

Bessner:2018:ASU

[Bes18] Daniel Bessner. Age of System: Understanding the Development of Modern Social Science by Hunter Heyck (review). Technology and Culture, 59(2):494–496, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/697253.

Bettoni:2014:FTI

[Bet14] Barbara Bettoni. Fashion, tradition, and innovation in button manufacturing in early modern Italy. Technology and Culture, 55(3):675–710, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551609.

Beyler:2012:EGO

[Bey12] Richard H. Beyler. Einstein’s Generation: The Origins of the Relativity Revolution (review). Technology and Culture, 53(2):492–493, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476816.

Bayer:2019:EAF

[BF19] Florian Bayer and Ulrike Felt. Embracing the “Atomic future” in post–world war II Austria. Technology and Culture, 60(1):165–191, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
719949.

Bergman:2017:FTS

[BGW^{+}17] Yoel Bergman, Francesco Gerali, Joanna Walewska, Dick van Lente, and Liliia Zemnukhova. Forty-third symposium of the International Committee for the History of Technology: “Technology, Innovation, and Sustainability: Historical and Contemporary Narratives,” Porto, Portugal, 26–30 July 2016. Technology and Culture, 58(1):250–256, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/648260.

Brinkman:2017:WWT

[BH17] Joshua T. Brinkman and Richard F. Hirsh. Welcoming wind turbines and the PIMBY (“Please in My Backyard”) phenomenon: The culture of the machine in the rural American Midwest. Technology and Culture, 58(2):335–367, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/662980.

Bhimull:2010:SCE

[Bhi10] Chandra D. Bhimull. Social Change and the Evolution of Ceramic Production and Distribution in a Maya Community (review). Technology and Culture, 51(3):770–771, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/390866.

Bhimull:2013:BEA

[Bhi13] Chandra D. Bhimull. British Engineering and Africa, 1875–1914 by Casper Andersen (review). Technology and Culture, 54(1): 193–194, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500807.

Biehler:2016:TRP

[Bie16] Dawn Biehler. Transportation and Revolt: Pigeons, Mules, Canals, and the Vanishing Geographies of Subversive Mobility by Jacob Shell (review). Technology and Culture, 57(3):678–680, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/630941.

Biggs:2012:PDH

[Big12] Lindy Biggs. Powering the Dream: The History and Promise of Green Technology (review). Technology and Culture, 53(3):735– 736, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481951.

Biggs:2016:BRM

[Big16] Adam Biggs. Breathing Race into the Machine: The Surprising Career of the Spirometer from Plantation to Genetics by Lundy Braun (review). Technology and Culture, 57(1):252–254, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611816.

Bijker:2011:TEK

[Bij11] Wiebe E. Bijker. Tacit and Explicit Knowledge (review). Technology and Culture, 52(4):809–810, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/454541.

Bijsterveld:2012:SPR

[Bij12] Karin Bijsterveld. Sonic Persuasion: Reading Sound in the Recorded Age (review). Technology and Culture, 53(3):713–715, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481923.

Bijker:2013:GFM

[Bij13] Wiebe E. Bijker. Good fortune, mirrors, and kisses. Technology and Culture, 54(3):600–618, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/520283.

Billings:2015:DSN

[Bil15] Linda Billings. Does Science Need a Global Language? English and the Future of Research by Scott L. Montgomery (review). Technology and Culture, 56(1):261–263, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/572122.

Bimm:2018:HSS

[Bim18] Jordan Bimm. Historical Studies in the Societal Impact of Spaceflight ed. by Steven J. Dick (review). Technology and Culture, 59 (1):183–185, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/692181.

Birdsall:2019:RDB

[Bir19] Carolyn Birdsall. Radio documents: Broadcasting, sound archiving, and the rise of radio studies in interwar Germany. Technology and Culture, 60(2S):S96–S128, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/727300.

Bix:2018:DFS

[Bix18] Amy Sue Bix. Disruptive Fixation: School Reform and the Pitfalls of Techno-Idealism by Christo Sims (review). Technology and Culture, 59(3):818–820, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/703166.

Bjorkman:2013:RDS

[Bjö13] Maria Björkman. Reproduction by Design: Sex, Robots, Trees, and Test-Tube Babies in Interwar Britain by Angus McLaren (review). Technology and Culture, 54(2):415–417, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/506092.

Brock:2012:DFM

[BL12] David C. Brock and Christophe Lécuyer. Digital foundations: The making of silicon-gate manufacturing technology. Technology and Culture, 53(3):561–597, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481906.

Black:2010:PTS

[Bla10] David V. Black. The Periodic Table: Its Story and Its Significance (review). Technology and Culture, 51(1):287–289, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375814.

Black:2018:ECH

[Bla18] Brian Black. Energy and civilization: A history. Technology and Culture, 59(3):787–788, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/703148.

Bloomfield:2013:LTE

[Blo13] Gerald T. Bloomfield. Landscape with Technology: Essays in Honour of L. T. C. Rolt ed. by R. Angus Buchanan (review). Technology and Culture, 54(1):188–189, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500804.

Blum:2012:ACH

[Blu12] Deborah Blum. The Arsenic Century: How Victorian Britain Was Poisoned at Home, Work, and Play (review). Technology and Culture, 53(4):929–930, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/490740.

Bonde:2012:TMW

[BM12] Sheila Bonde and Clark Maines. The technology of medieval water management at the charterhouse of bourgfontaine. Technology and Culture, 53(3):625–670, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481908.

Bradford:2014:TLA

[BM14] Phillip G. Bradford and Paul J. Miranti. Technology and learning: Automating odd-lot trading at the New York Stock Exchange, 1958–1976. Technology and Culture, 55(4):850–879, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562834.

Baichun:2019:JNR

[BM19] Zhang Baichun and Tian Miao. Joseph Needham’s research on Chinese machines in the cross-cultural history of science and technology. Technology and Culture, 60(2):616–624, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/726943.

Bocking:2016:APM

[Boc16] Stephen Bocking. Allied Power: Mobilizing Hydro-Electricity during Canada’s Second World War by Matthew Evenden (review). Technology and Culture, 57(3):686–687, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/630945.

Bonde:2016:ELS

[Bon16] Sheila Bonde. Ecclesiastical Lordship, Seigneurial Power, and the Commercialization of Milling in Medieval England by Adam Lucas (review). Technology and Culture, 57(4):1005–1006, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637938.

Booth:2017:RAD

[Boo17] Paul Booth. Rogue Archives: Digital Cultural Memory and Media Fandom by Abigail De Kosnik (review). Technology and Culture, 58 (4):1111–1113, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679847.

Borgman:2010:DBP

[Bor10] Christine L. Borgman. Digitize This Book! The Politics of New Media, or Why We Need Open Access Now (review). Technology and Culture, 51(3):768–770, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/390865.

Borg:2014:ICS

[Bor14] Kevin L. Borg. Introduction: Constructing sociotechnical environments — aurality, air quality, and automobiles. Technology and Culture, 55(2):287–298, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/546866.

Borys:2017:SGT

[Bor17] Ann Marie Borys. The Strait Gate: Thresholds and Power in Western History by Daniel Jütte (review). Technology and Culture, 58 (4):1072–1074, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679825.

Bourke:2019:TWR

[Bou19] Joanna Bourke. The Technoscientific Witness of Rape: Contentious Histories of Law, Feminism, and Forensic Science by Andrea Quinlan (review). Technology and Culture, 60(1):342–343, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
719969.

Bowditch:2010:ERL

[Bow10] John Bowditch. The End of a Revolution: The Last Days of Stationary Steam (review). Technology and Culture, 51(1):259–260, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
375813.

Bowker:2012:FRD

[Bow12] Geoff Bowker. The Freudian Robot: Digital Media and the Future of the Unconscious (review). Technology and Culture, 53(3):745– 747, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481957.

Bowditch:2014:TIE

[Bow14] John Bowditch. Tesla: Inventor of the Electric Age by W. Bernard Carlson (review). Technology and Culture, 55(4):1009–1010, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562852.

Bowditch:2015:SMA

[Bow15] John Bowditch. The Smithsonian as a Microcosm of American Museum Practice: Robert C. Post, Who Owns America’s Past? (review). Technology and Culture, 56(1):248–251, January 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/572120.

Bowditch:2016:IMT

[Bow16] John Bowditch. Ingenious Machinists: Two Inventive Lives from the American Industrial Revolution by Anthony J. Conners (review). Technology and Culture, 57(2):473–474, April 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/619066.

Boyce:2016:WDI

[Boy16] Angie M. Boyce. “When does it stop being peanut butter?”: FDA food standards of identity, Ruth Desmond, and the shifting politics of consumer activism, 1960s–1970s. Technology and Culture, 57 (1):54–79, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611800.

Bray:2010:HBL

[Bra10] Francesca Bray. How blind is love?: Simon Winchester’s The Man Who Loved China. Technology and Culture, 51(3):578–588, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/390837.

Braun:2012:RRH

[Bra12] Hans-Joachim Braun. Reason and Resonance: A History of Modern Aurality (review). Technology and Culture, 53(1):207–209, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466785.

Brandt:2013:DMP

[Bra13] Marisa Brandt. Doctored: The Medicine of Photography in Nineteenth-Century America by Tanya Sheehan (review). Technology and Culture, 54(3):666–669, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520294.

Bradford:2015:PSG

[Bra15a] Phillip G. Bradford. Prometheus Shackled: Goldsmith Banks and England’s Financial Revolution after 1700 by Peter Temin and Hans-Joachim Voth (review). Technology and Culture, 56 (1):269–270, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572126.

Braun:2015:SSH

[Bra15b] Hans-Joachim Braun. Sound and Safe: A History of Listening behind the Wheel by Karin Bijsterveld, etal. (review). Technology and Culture, 56(2):555–556, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/582335.

Bray:2017:SIA

[Bra17] Francesca Bray. SHOT: Internationalization and the art of translation. Technology and Culture, 58(3):815–834, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669469.

Brazelton:2019:EHT

[Bra19] Mary Augusta Brazelton. Engineering health: Technologies of immunization in China’s wartime hinterland, 1937–45. Technology and Culture, 60(2):409–437, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726932.

Breckenridge:2010:WYI

[Bre10a] Keith Breckenridge. Who Are You? Identification, Deception, and Surveillance in Early Modern Europe (review). Technology and Culture, 51(2):490–491, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381255.

Breyman:2010:ABA

[Bre10b] Steve Breyman. The Atomic Bomb and American Society: New Perspectives (review). Technology and Culture, 51(4):1036–1038, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403292.

Breen:2016:DDQ

[Bre16] Deborah Breen. Designs and dreams: Questions of technology in Hayao Miyazaki’s The Wind Rises. Technology and Culture, 57 (2):457–459, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/619059.

Breen:2018:DWN

[Bre18] Deborah Breen. Designing Worlds: National Design Histories in an Age of Globalization by Kjetil Fallan (review). Technology and Culture, 59(4):999–1000, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711790.

Britton:2010:CUS

[Bri10] John A. Britton. Communications under the Seas: The Evolving Cable Network and Its Implications (review). Technology and Culture, 51(4):1045–1046, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/403297.

Brown:2011:WLE

[Bro11] John K. Brown. Waterpower in Lowell: Engineering and Industry in Nineteenth-Century America (review). Technology and Culture, 52(1):199–200, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416425.

Brock:2012:IPM

[Bro12a] David C. Brock. Inventing the PC: The MCM/70 Story (review). Technology and Culture, 53(2):518–519, April 2012. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/476817.

Brown:2012:CRB

[Bro12b] John K. Brown. The City, The River, The Bridge: Before and After the Minneapolis Bridge Collapse (review). Technology and Culture, 53(3):737–738, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481952.

Brown:2013:RLC

[Bro13] John K. Brown. The Roebling Legacy by Clifford W. Zink (review). Technology and Culture, 54(1):200–201, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500811.

Brown:2014:EBE

[Bro14a] John K. Brown. Not the Eads bridge: An exploration of counterfactual history of technology. Technology and Culture, 55(3):521–559, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
551605.

Brown:2014:CEA

[Bro14b] Philip C. Brown. Constructing East Asia: Technology, Ideology, and Empire in Japan’s Wartime Era, 1931–1945 by Aaron Stephen Moore (review). Technology and Culture, 55(4):1019–1021, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/562858.

Brown:2015:RPE

[Bro15] John K. Brown. Routes of Power: Energy and Modern America by Christopher F. Jones (review). Technology and Culture, 56(2):567– 568, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582342.

Brockmann:2017:SSA

[Bro17a] Stephen Brockmann. The Shape of Spectatorship, Art, Science, and Early Cinema in Germany by Scott Curtis (review). Technology and Culture, 58(2):583–584, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/662993.

Brown:2017:RFR

[Bro17b] Philip C. Brown. Reverse flow: The role of built environments in shaping disaster. Technology and Culture, 58(1):170–181, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/648255.

Bruyninckx:2019:SBC

[Bru19] Joeri Bruyninckx. For science, broadcasting, and conservation: Wildlife recording, the BBC, and the consolidation of a British library of wildlife sounds. Technology and Culture, 60 (2S):S188–S215, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/727303.

Bektas:2013:BNE

[BS13] Yakup Bektas and Roger Sherman. A bold new enterprise: The Istanbul Museum of the History of Science and Technology in Islam. Technology and Culture, 54(3):619–639, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/520284.

Bertomeu-Sanchez:2016:EPO

[BSA16] José Ramón Bertomeu-Sánchez and Stathis Arapostathis. Experts and peripheries: Ongoing research and future challenges. Technology and Culture, 57(4):951–965, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/637933.

Birdsall:2019:LAS

[BT19] Carolyn Birdsall and Viktoria Tkaczyk. Listening to the archive: Sound data in the humanities and sciences. Technology and Culture, 60(2S):S1–S13, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/727296.

Buchanan:2012:IGE

[Buc12a] Brenda J. Buchanan. Islamic Gunpowder Empires: Ottomans, Safavids, and Mughals (review). Technology and Culture, 53(4): 923–925, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490737.

Buckley:2012:IBE

[Buc12b] Eve Buckley. The Illusory Boundary: Environment and Technology in History (review). Technology and Culture, 53(2):470–471, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476818.

Buchanan:2014:SMG

[Buc14] Brenda J. Buchanan. Saltpeter: The Mother of Gunpowder by David Cressy (review). Technology and Culture, 55(4):1000–1002, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562847.

Bugos:2013:LAO

[Bug13] Glenn Bugos. Locomotive to Aeromotive: Octave Chanute and the Transportation Revolution by Simine Short (review). Technology and Culture, 54(2):414–415, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506091.

Buhrman:2017:RFR

[Buh17] Kristina Buhrman. Remembering future risk: Considering technologies of the archive for discussion of Tōhoku’s seismological past after 2011. Technology and Culture, 58(1):159–169, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/648254.

Buisseret:2010:SSS

[Bui10] David Buisseret. Secret Science: Spanish Cosmography and the New World (review). Technology and Culture, 51(4):1021–1023, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403283.

Burke:2013:UPM

[Bur13] David Allen Burke. US Presidents and the Militarization of Space, 1946–1967 by Sean N. Kalic (review). Technology and Culture, 54(3):685–686, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520304.

Buss:2019:OGP

[Bus19a] Jared S. Buss. Our Germans: Project Paperclip and the National Security State by Brian E. Crim (review). Technology and Culture, 60(3):917–919, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/731451.

Bussotti:2019:AHC

[Bus19b] Michela Bussotti. The amazing history of the Chinese typewriter: Thomas S. Mullaney, The Chinese Typewriter. Technology and Culture, 60(3):896–899, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/731448.

Butrica:2012:AAL

[But12] Andrew Butrica. De l’atelier au laboratoire: Recherche et innovation dans l’industrie électrique, XIX^{e}–XX^{e} siècles (review). Technology and Culture, 53(4):933–935, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/490743.

Byers:2013:QFJ

[Bye13] Richard Byers. Quest for Flight: John J. Montgomery and the Dawn of Aviation in the West by Craig S. Harwood and Gary B. Fogel (review). Technology and Culture, 54(3):669–670, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520295.

Byers:2019:AFF

[Bye19] Richard Byers. Anthony Fokker: The Flying Dutchman Who Shaped American Aviation by Marc Dierikx (review). Technology and Culture, 60(3):912–913, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/731452.

Caffrey:2018:PWW

[Caf18] Matthew Caffrey. Playing War: Wargaming and U.S. Navy Preparations for WWII by John M. Lillard (review). Technology and Culture, 59(2):482–483, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/697246.

Cahan:2014:PEM

[Cah14] David Cahan. The Psychophysical Ear: Musical Experiments, Experimental Sounds, 1840–1910 by Alexandra Hui (review). Technology and Culture, 55(1):256–258, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/538925.

Cahan:2019:CMS

[Cah19] David Cahan. Climate in Motion: Science, Empire, and the Problem of Scale by Deborah R. Coen (review). Technology and Culture, 60(4):1103–1105, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/741396.

Camprubi:2015:RGC

[Cam15] Lino Camprubí. Resource geopolitics: Cold War technologies, global fertilizers, and the fate of Western Sahara. Technology and Culture, 56(3):676–703, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/593068.

Campana:2016:TES

[Cam16] Lilia Campana. Technical experimentation in ship design during the last decades of the Serenissima: Gerolamo Maria Balbi’s Galea alla Ponentina. Technology and Culture, 57(1):144–182, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611803.

Camp:2018:WDC

[Cam18] Michael Camp. “Wandering in the desert”: The Clinch River Breeder Reactor Debate in the U.S. Congress, 1972–1983. Technology and Culture, 59(1):26–47, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/692166.

Camprubi:2019:GAM

[Cam19] Lino Camprubí. Globalized Authoritarianism: Megaprojects, Slums, and Class Relations in Urban Morocco by Koenraad Bogaert (review). Technology and Culture, 60(1):326–328, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/719960.

Canales:2015:MRE

[Can15] Jimena Canales. The media of relativity: Einstein and telecommunications technologies. Technology and Culture, 56(3):610–645, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593066.

Cantoni:2017:WPN

[Can17] Roberto Cantoni. What’s in a pipe?: NATO’s confrontation on the 1962 large-diameter pipe embargo. Technology and Culture, 58(1):67–96, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648250.

Caruso:2011:TMP

[Car11] David Caruso. Technology and Medical Practice: Blood, Guts and Machines (review). Technology and Culture, 52(4):862–863, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454571.

Carlisle:2013:ATM

[Car13a] Rodney Carlisle. Atomic Testing in Mississippi: Project Dribble and the Quest for Nuclear Weapons Treaty Verification in the Cold War Era by David Allen Burke (review). Technology and Culture, 54(3):691–692, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520307.

Carroll:2013:RPB

[Car13b] Patrick Carroll. Roads to Power: Britain Invents the Infrastructure State by Jo Guldi (review). Technology and Culture, 54(2):405–407, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506086.

Caruso:2013:WWR

[Car13c] David Caruso. War’s Waste: Rehabilitation in World War I America by Beth Linker (review). Technology and Culture, 54 (1):198–200, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500810.

Carlisle:2014:PNF

[Car14a] Rodney Carlisle. Plutopia: Nuclear Families, Atomic Cities, and the Great Soviet and American Plutonium Disasters by Kate Brown (review). Technology and Culture, 55(1):269–270, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538932.

Carlson:2014:OMC

[Car14b] W. Bernard Carlson. From order to messy complexity: Thoughts on the intellectual journey of Thomas Parke Hughes. Technology and Culture, 55(4):945–952, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/562837.

Carse:2015:IID

[Car15] Ashley Carse. Inverse Infrastructures: Disrupting Networks from Below ed. by Tineke M. Egyedi and Donna C. Mehos (review). Technology and Culture, 56(2):570–572, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/582344.

Carroll:2016:BBD

[Car16] Patrick Carroll. Beyond the Big Ditch: Politics, Ecology, and Infrastructure at the Panama Canal by Ashley Carse (review). Technology and Culture, 57(2):467–469, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/619063.

Carson:2018:HOD

[Car18] John Carson. How Our Days Became Numbered: Risk and the Rise of the Statistical Individual by Dan Bouk (review). Technology and Culture, 59(2):496–497, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/697254.

Castaneda:2010:SSC

[Cas10] Christopher J. Castaneda. Scientists and Swindlers: Consulting on Coal and Oil in America, 1820–1890 (review). Technology and Culture, 51(3):743–745, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390850.

Castaneda:2013:PEO

[Cas13] Chris Castaneda. Progressive Enlightenment: The Origins of the Gaslight Industry, 1780–1820 by Leslie Tomory (review). Technology and Culture, 54(3):661–663, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520291.

Castonguay:2014:WWH

[Cas14] Stéphane Castonguay. Wilderness and Waterpower: How Banff National Park Became a Hydroelectric Storage Reservoir by Christopher Armstrong and H. V. Nelles (review). Technology and Culture, 55(4):1011–1012, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562853.

Casey:2015:HFV

[Cas15] Robert Casey. Henry Ford. by Vincent Curcio (review). Technology and Culture, 56(1):274–276, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/572129.

Castaneda:2016:MDM

[Cas16a] Christopher J. Castaneda. Making the Desert Modern: Americans, Arabs, and Oil on the Saudi Frontier, 1933–1973 by Chad H. Parker (review). Technology and Culture, 57(3):687–689, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/630946.

Castro:2016:CCR

[Cas16b] J. Justin Castro. Culture of Class: Radio and Cinema in the Making of a Divided Argentina by Matthew B. Karush (review). Technology and Culture, 57(4):1022–1023, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/637948.

Castonguay:2019:CDM

[Cas19] Stéphane Castonguay. La Contamination du monde: Une histoire des pollutions à l’âge industriel by François Jarrige and Thomas Le Roux (review). Technology and Culture, 60(3):902–904, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/731455.

Cirac-Claveras:2018:SWC

[CC18] Gemma Cirac-Claveras. Satellites for what? Creating user communities for space-based data in France: The case from LERTS to CESBIO. Technology and Culture, 59(2):203–225, April 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/697230.

Ceruzzi:2010:WDB

[Cer10] Paul E. Ceruzzi. When dreams become real. Technology and Culture, 51(3):675–679, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390841.

Ceruzzi:2011:CBT

[Cer11] Paul Ceruzzi. The Computer Boys Take Over: Computers, Programmers, and the Politics of Technical Expertise (review). Technology and Culture, 52(3):662–664, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/447810.

Cerveaux:2013:TMD

[Cer13] Augustin Cerveaux. Taming the microworld: DuPont and the interwar rise of fundamental industrial research. Technology and Culture, 54(2):262–288, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/506070.

Ceruzzi:2015:DDH

[Cer15] Paul Ceruzzi. Delete: A Design History of Computer Vapourware by Paul Atkinson (review). Technology and Culture, 56(4): 1013–1015, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601502.

Ceruzzi:2016:MLL

[Cer16] Paul E. Ceruzzi. Moore’s Law: The Life of Gordon Moore, Silicon Valley’s Quiet Revolutionary by Arnold Thackray, David C. Brock, and Rachel Jones (review). Technology and Culture, 57(2):499–500, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
619081.

Cutcliffe:2010:RFT

[CF10] Stephen H. Cutcliffe and Kimberly Fabbri. Robert Fulton’s torpedoes. Technology and Culture, 51(4):879–888, October 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/403269.

Chakraborty:2016:FEM

[CG16] Anwesha Chakraborty and Anna Guagnini. The #FoodPeople Exhibition at the Museo Nazionale della Scienza e della Tecnologia Leonardo da Vinci, Milano. Technology and Culture, 57(3):649– 656, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
630932.

Chan:2011:MVW

[Cha11a] Anita Say Chan. Making Virtual Worlds: Linden Lab and Second Life (review). Technology and Culture, 52(1):226–228, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416442.

Chant:2011:UMC

[Cha11b] Colin Chant. Urban Modernity: Cultural Innovation in the Second Industrial Revolution (review). Technology and Culture, 52 (4):833–835, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454552.

Charnley:2012:HHS

[Cha12] Berris Charnley. Hybrid: The History and Science of Plant Breeding (review). Technology and Culture, 53(2):474–475, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476819.

Chakrabarti:2014:ETM

[Cha14] Pratik Chakrabarti. Everyday Technology: Machines and the Making of India’s Modernity by David Arnold (review). Technology and Culture, 55(4):1017–1019, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562857.

Channell:2018:BSI

[Cha18a] David F. Channell. Biotechnology and Society: An Introduction by Hallam Stevens (review). Technology and Culture, 59(2):499–501, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
697256.

Channell:2018:MMW

[Cha18b] David F. Channell. Marconi: The Man Who Networked the World by Marc Raboy (review). Technology and Culture, 59(4):973–975, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
711775.

Chen:2019:NMF

[Che19] Buyun Chen. Needham, matter, form, and us. Technology and Culture, 60(2):574–582, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/726939.

Choi:2017:SCI

[Cho17] Hyungsub Choi. The social construction of imported technologies: Reflections on the social history of technology in modern Korea. Technology and Culture, 58(4):905–920, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679818.

Choi:2018:DGD

[Cho18] Hyungsub Choi. Displaying “Growth and Development”: Exhibit Hall 3, National Museum of Korean Contemporary History. Technology and Culture, 59(2):454–463, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/697239.

Churella:2015:AEC

[Chu15] Albert J. Churella. Over the Alleghenies: Early Canals and Railroads of Pennsylvania by Robert J. Kapsch (review). Technology and Culture, 56(2):540–542, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/582327.

Citino:2012:DKH

[Cit12] Nathan J. Citino. Desert Kingdom: How Oil and Water Forged Modern Saudi Arabia (review). Technology and Culture, 53(1): 234–236, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466800.

Campbell-Kelly:2015:DSS

[CK15] Martin Campbell-Kelly. Digital State: The Story of Minnesota’s Computing Industry by Thomas J. Misa (review). Technology and Culture, 56(1):296–298, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/572142.

Clarke:2010:EMD

[Cla10] Deborah Clarke. Eat My Dust: Early Women Motorists (review). Technology and Culture, 51(1):265–266, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/375812.

Clarke:2011:BSA

[Cla11] Deborah Clarke. From Birdwomen to Skygirls: American Girls’ Aviation Stories (review). Technology and Culture, 52(1):218–220, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
416437.

Clark:2013:SDA

[Cla13] Catherine E. Clark. Speculating Daguerre: Art and Enterprise in the Work of L. J. M. Daguerre by Stephen Pinson (review). Technology and Culture, 54(4):977–978, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530526.

Cloud:2011:SMG

[Clo11] John Cloud. In the Shadow of Melting Glaciers: Climate Change and Andean Society (review). Technology and Culture, 52(4): 842–843, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454560.

Cloud:2012:GPS

[Clo12] John Cloud. Globalizing Polar Science: Reconsidering the International Polar and Geophysical Years (review). Technology and Culture, 53(2):493–495, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/476820.

Coclanis:2013:REH

[Coc13] Peter A. Coclanis. Rethinking the economic history of early modern India. Technology and Culture, 54(2):389–392, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/506078.

Coclanis:2016:KCS

[Coc16] Peter A. Coclanis. King cotton: Sven beckert, empire of cotton. Technology and Culture, 57(3):661–667, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/630934.

Coclanis:2017:FPT

[Coc17a] Peter A. Coclanis. Fascist Pigs: Technoscientific Organisms and the History of Fascism by Tiago Saraiva (review). Technology and Culture, 58(4):1088–1090, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679834.

Coclanis:2017:ARH

[Coc17b] Peter A. Coclanis. The American Reaper: Harvesting Networks and Technology, 1830–1910 by Gordon M. Winder (review). Technology and Culture, 58(1):267–269, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648264.

Coclanis:2018:GCB

[Coc18] Peter A. Coclanis. Green Capitalism: Business and the Environment in the Twentieth Century ed. by Hartmut Berghoff and Adam Rome (review). Technology and Culture, 59(3):803–804, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/703157.

Colatrella:2012:GCW

[Col12a] Carol Colatrella. Gender Codes: Why Women Are Leaving Computing (review). Technology and Culture, 53(4):958–959, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490757.

Cole:2012:FB

[Col12b] Simon A. Cole. The face of biometrics. Technology and Culture, 53 (1):200–203, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466768.

Colten:2012:BGB

[Col12c] Craig E. Colten. Blowout in the Gulf: The BP Oil Spill Disaster and the Future of Energy in America (review). Technology and Culture, 53(1):236–237, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/466801.

Cole:2013:EPE

[Col13] Simon A. Cole. Embryo Politics: Ethics and Policy in Atlantic Democracies by Thomas Banchoff (review). Technology and Culture, 54(2):430–431, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/506100.

Collins:2014:ACP

[Col14] Martin Collins. Arguments that Count: Physics, Computing, and Missile Defense, 1949–2012 by Rebecca Slayton (review). Technology and Culture, 55(4):1028–1030, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/562863.

Collopy:2015:EVP

[Col15] Peter Sachs Collopy. The Emergence of Video Processing Tools: Television Becoming Unglued. 2 vols ed. by Kathy High, Sherry Miller Hocking, and Mona Jimenez (review). Technology and Culture, 56(4):1008–1010, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/601499.

Collins:2016:GRH

[Col16] Martin Collins. German Rocketeers in the Heart of Dixie: Making Sense of the Nazi Past during the Civil Rights Era by Monique Laney (review). Technology and Culture, 57(3):691–692, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/630948.

Cole:2017:QBT

[Col17a] Simon A. Cole. Quantified: Biosensing Technologies in Everyday Life ed. by Dawn Nafus (review). Technology and Culture, 58(1):293–295, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648279.

Colombo:2017:SMD

[Col17b] Fausto Colombo. Storia dei media digitali: Rivoluzioni e continuità by Gabriele Balbi, Paolo Magaudda (review). Technology and Culture, 58(3):897–898, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/669491.

Collins:2019:AAA

[Col19a] Martin J. Collins. Apollo in the Age of Aquarius by Neil Maher (review). Technology and Culture, 60(2):640–642, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726954.

Collopy:2019:PMI

[Col19b] Peter Sachs Collopy. Portable Moving Images: A Media History of Storage Formats by Ricardo Cedeño Montaña (review). Technology and Culture, 60(1):337–338, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/719966.

Comfort:2012:DML

[Com12] Nathaniel Comfort. Drawing the Map of Life inside the Human Genome Project (review). Technology and Culture, 53(1):240–241, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466803.

Conway:2015:SSD

[Con15] Erik M. Conway. Simulation and spacecraft design: Engineering Mars landings. Technology and Culture, 56(4):812–838, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/601469.

Cooper:2010:DIP

[Coo10a] Carolyn C. Cooper. Does intellectual property protection mean pursuit of patent perpetuity? Technology and Culture, 51(2):486– 489, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381256.

Cooper:2010:BRE

[Coo10b] Martin Cooper. British Railway Enthusiasm (review). Technology and Culture, 51(1):260–262, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375811.

Cooper:2011:EHD

[Coo11a] Carolyn Cooper. Every Home a Distillery: Alcohol, Gender, and Technology in the Colonial Chesapeake (review). Technology and Culture, 52(1):187–188, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/416418.

Cooper:2011:FSC

[Coo11b] Gail Cooper. Fixing the Sky: The Checkered History of Weather and Climate Control (review). Technology and Culture, 52(4): 844–845, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454559.

Cooper:2011:PLS

[Coo11c] Timothy Cooper. Peter Lund simmonds and the political ecology of waste utilization in Victorian Britain. Technology and Culture, 52(1):21–44, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416405.

Cooper:2012:LMI

[Coo12a] Timothy Cooper. The Lives of Machines: The Industrial Imaginary in Victorian Literature and Culture (review). Technology and Culture, 53(4):930–932, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/490741.

Coopersmith:2012:ERR

[Coo12b] Jonathan Coopersmith. Engineer of Revolutionary Russia: Iurii V. Lomonosov (1876–1952) and the Railways (review). Technology and Culture, 53(3):707–708, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/481919.

Coopersmith:2014:SSH

[Coo14] Jonathan Coopersmith. SPAM: A Shadow History of the Internet by Finn Brunton (review). Technology and Culture, 55(4): 1034–1035, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562866.

Cooper:2015:MTH

[Coo15] Ken Cooper. Microlessons: Toward a history of information-age cuisine. Technology and Culture, 56(3):579–609, July 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/593065.

Coopersmith:2017:SPR

[Coo17] Jonathan Coopersmith. Shifting Practices: Reflections on Technology, Practice, and Innovation by Giovan Francesco Lanzara (review). Technology and Culture, 58(4):1071–1072, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/679824.

Cornillie:2010:VRJ

[Cor10] Thomas Cornillie. Visionary Railroader: Jervis Langdon Jr. and the Transportation Revolution (review). Technology and Culture, 51(1):262–263, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375810.

Cornillie:2012:PSI

[Cor12a] Thomas C. Cornillie. Public Space and the Ideology of Place in American Culture (review). Technology and Culture, 53(3):709– 710, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481920.

Cortada:2012:EVE

[Cor12b] James W. Cortada. Electronic Value Exchange: Origins of the VISA Electronic Payment System (review). Technology and Culture, 53(1):247–248, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/466807.

Corn:2013:BCH

[Cor13a] Joseph J. Corn. Bastard Culture! How User Participation Transforms Cultural Production by Mirko Tobias Schäfer (review). Technology and Culture, 54(1):223–225, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500824.

Cortada:2013:HNT

[Cor13b] James W. Cortada. How new technologies spread: Lessons from computing technologies. Technology and Culture, 54(2):229–261, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506069.

Corn:2014:EAA

[Cor14] Joseph J. Corn. Empire of the Air: Aviation and the American Ascendency by Jenifer Van Vleck (review). Technology and Culture, 55(4):1025–1026, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/562861.

Cortada:2016:MSH

[Cor16] James W. Cortada. From Mainframes to Smartphones: A History of the International Computer Industry by Martin Campbell-Kelly and Daniel D. Garcia-Swartz (review). Technology and Culture, 57(3):703–704, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630955.

Corry:2019:DCD

[Cor19] Frances Corry. Dot-Com Design: The Rise of a Usable, Social, Commercial Web by Megan Sapnar Ankerson (review). Technology and Culture, 60(1):339–340, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/719967.

Cossart:2017:PSE

[Cos17] Brice Cossart. Producing skills for an empire: Theory and practice in the Seville School of Gunners during the Golden Age of the Carrera de Indias. Technology and Culture, 58(2):459–486, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662984.

Coupaye:2017:CHD

[Cou17a] Ludovic Coupaye. Comment Homo devint Faber by François Sigaut (review). Technology and Culture, 58(4):1076–1077, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/679827.

Courtwright:2017:STH

[Cou17b] Julie Courtwright. Still Turning: A History of Aermotor Windmills by Christopher C. Gillis (review). Technology and Culture, 58(4): 1092–1093, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679836.

Cowan:2010:LBO

[Cow10] Ruth Schwartz Cowan. Looking back in order to move forward: John McDermott, “Technology: The Opiate of the Intellectuals”. Technology and Culture, 51(1):199–215, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/375777.

Cowan:2012:RM

[Cow12] Ruth Schwartz Cowan. Response to McDermott. Technology and Culture, 53(2):530–531, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/476847#sub01. See [McD12].

Cullen:2019:GSA

[CR19] Christopher Cullen and Hiraoka Ryuji. The Geneva Sphere: An astronomical model from Seventeenth-Century Japan. Technology and Culture, 60(1):219–251, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/719951.

Creager:2017:RLM

[Cre17] Angela N. H. Creager. Rendering Life Molecular: Models, Modelers, and Excitable Matter by Natasha Myers (review). Technology and Culture, 58(4):1103–1104, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/679842.

Crouch:2014:ADD

[Cro14a] Tom D. Crouch. Airship Design, Development, and Disaster by John Swinfield (review). Technology and Culture, 55(2):501–502, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546885.

Crowther:2014:SPE

[Cro14b] Kathleen Crowther. Science in Print: Essays on the History of Science and the Culture of Print ed. by Rima D. Apple, Gregory J. Downey, and Stephen L. Vaughn (review). Technology and Culture, 55(2):491–492, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/546879.

Cronin:2016:RBT

[Cro16] Marionne Cronin. Richard Byrd, technological explorer: Polar exploration, the machine, and heroic masculinity in Interwar America. Technology and Culture, 57(2):322–352, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619054.

Crook:2018:HLW

[Cro18] Tom Crook. The History of the London Water Industry, 1580–1820 by Leslie Tomory (review). Technology and Culture, 59(3):792–794, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703151.

Cruz:2019:TPD

[Cru19] Felipe Fernandes Cruz. Technocrats and the Politics of Drought and Development in Twentieth-Century Brazil by Eve Buckley (review). Technology and Culture, 60(1):319–320, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/719956.

Crystal:2011:APT

[Cry11] Lisa Crystal. The Assault on Progress: Technology and Time in American Literature (review). Technology and Culture, 52(3):669– 671, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
447814.

Cummings:2011:TMR

[Cum11a] Alex Cummings. Tactical Media (review). Technology and Culture, 52(3):667–669, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447813.

Cummings:2011:PDE

[Cum11b] Alex Cummings. The Public Domain: Enclosing the Commons of the Mind (review). Technology and Culture, 52(4):857–858, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454568.

Cummings:2012:MUI

[Cum12] Alex Sayf Cummings. Making and Unmaking Intellectual Property: Creative Production in Legal and Cultural Perspective (review). Technology and Culture, 53(4):922–923, October 2012. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/490736.

Cummings:2013:MMF

[Cum13] Alex Sayf Cummings. MP3: The Meaning of a Format by Jonathan Sterne (review). Technology and Culture, 54(4):1007–1008, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/530543.

Cummings:2019:AAM

[Cum19] Alex Sayf Cummings. Authors and Apparatus: A Media History of Copyright by Monika Dommann (review). Technology and Culture, 60(4):1123–1124, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/741398.

Curran:2012:DOD

[Cur12] James Curran. The Digital Origins of Dictatorship: Information Technology and Political Islam (review). Technology and Culture, 53(3):743–745, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481956.

Curry:2013:IEM

[Cur13] Helen Anne Curry. Industrial evolution: Mechanical and biological innovation at the general electric research laboratory. Technology and Culture, 54(4):746–781, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/530510.

Cutcliffe:2010:TTW

[Cut10] Stephen H. Cutcliffe. Travels in and out of town: William Cronon’s Nature’s Metropolis: Chicago and the Great West. Technology and Culture, 51(3):728–737, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390846.

Cutcliffe:2011:RSB

[Cut11] Stephen Cutcliffe. Roots of Steel: Boom and Bust in an American Steel Town (review). Technology and Culture, 52(1):204–205, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
416428.

Cutcliffe:2013:MIS

[Cut13] Stephen H. Cutcliffe. Mastering Iron: The Struggle to Modernize an American Industry, 1800–1868 by Anne Kelly Knowles (review). Technology and Culture, 54(4):975–976, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530525.

Daemmrich:2019:GUM

[Dae19] Arthur Daemmrich. Generic: The Unbranding of Modern Medicine by Jeremy A. Greene (review). Technology and Culture, 60(4): 1129–1131, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741400.

Daly:2015:THE

[Dal15] Nicholas Daly. The Triumph of Human Empire: Verne, Morris, and Stevenson at the End of the World by Rosalind Williams (review). Technology and Culture, 56(3):760–761, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/593078.

Davis:2013:SRC

[Dav13] Robert Scott Davis. Steamboats and the Rise of the Cotton Kingdom by Robert Gudmestad (review). Technology and Culture, 54 (1):190–191, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500805.

David:2017:ERS

[Dav17] James David. Eyeing the Red Storm: Eisenhower and the First Attempt to Build a Spy Satellite by Robert M. Dienesch (review). Technology and Culture, 58(3):879–881, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669481.

Day:2011:ESE

[Day11] Dwayne Day. Eyes in the Sky: Eisenhower, the CIA, and Cold War Aerial Espionage (review). Technology and Culture, 52(2):413–415, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431383.

DeBlock:2013:MET

[DB13] Greet De Block and Matthias Blondia. Materializing Europe: Transnational Infrastructures and the Project of Europe ed. by Alexander Badenoch and Andreas Fickers (review). Technology and Culture, 54(1):211–212, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/500817.

DeMunck:2010:CLM

[De 10] Bert De Munck. Corpses, live models, and nature: Assessing skills and knowledge before the Industrial Revolution (case: Antwerp). Technology and Culture, 51(2):332–356, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/381273.

DeBlock:2011:DNB

[De 11a] Greet De Block. Designing the nation: The Belgian Railway Project, 1830–1837. Technology and Culture, 52(4):703–732, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454530.

DeVos:2011:LGP

[De 11b] Els De Vos. Ladies and Gents: Public Toilets and Gender (review). Technology and Culture, 52(1):220–221, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/416438.

DeVos:2012:SHE

[De 12] Els De Vos. Socialist housing in the Eastern Bloc: Lynne attwood, gender and housing in Soviet Russia kimberly Elman zarecor, manufacturing a socialist modernity. Technology and Culture, 53 (2):465–469, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476814.

DeVos:2015:TTP

[De 15] Els De Vos. Trams or Tailfins? Public and Private Prosperity in Postwar West Germany and the United States by Jan L. Logemann (review). Technology and Culture, 56(1):283–284, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/572134.

DeVos:2016:GAB

[De 16] Els De Vos. The Garage: Automobility and Building Innovation in America’s Early Auto Age by John A. Jakle and Keith A. Sculle (review). Technology and Culture, 57(1):264–265, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611822.

DeMol:2017:EAM

[De 17] Liesbeth De Mol. ENIAC in Action: Making and Remaking the Modern Computer by Thomas Haigh, Mark Priestley, and Crispin Rope (review). Technology and Culture, 58(2):602–604, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/663004.

DeSyon:2019:WLP

[De 19] Guillaume De Syon. Willy Ley: Prophet of the Space Age by Jared S. Buss (review). Technology and Culture, 60(2):639–640, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/726946.

DeBlasio:2011:TRF

[DeB11] Donna DeBlasio. Twilight Rails: The Final Era of Railroad Building in the Midwest (review). Technology and Culture, 52(3):635– 636, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
447796.

Decoteau:2018:BDI

[Dec18] Claire Laurier Decoteau. Bioinsecurities: Disease Interventions, Empire and the Government of Species by Neel Ahuja (review). Technology and Culture, 59(2):480–482, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/697245.

Dedering:2018:BSW

[Ded18] Tilman Dedering. Blue skies into white space: Southern African responses to the trans-African flight of the Silver Queen, 1920. Technology and Culture, 59(2):289–312, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/697233.

Deringer:2017:PFS

[Der17] William Deringer. Pricing the future in the Seventeenth Century: Calculating technologies in competition. Technology and Culture, 58(2):506–528, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/662986.

Dery:2018:FGT

[Dér18] Steve Déry. Forests Are Gold: Trees, People, and Environmental Rule in Vietnam by Pamela D. McElwee (review). Technology and Culture, 59(1):181–183, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/692180.

Deutsch:2017:ACS

[Deu17] Chris Deutsch. Arresting Contagion: Science, Policy, and Conflicts over Animal Disease Control by Alan L. Olmstead and Paul W. Rhode (review). Technology and Culture, 58(1):280–281, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/648271.

Diogo:2016:SMH

[DGS16] Maria Paula Diogo, Kostas Gavroglu, and Ana Simões. STEP matters: Historiographical considerations. Technology and Culture, 57(4):926–929, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637930.

DeVos:2011:UBG

[DH11] Els De Vos and Hilde Heynen. Uncanny and in-between: The garage in rural and suburban Belgian Flanders. Technology and Culture, 52(4):757–787, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/454532.

Diaz:2016:TMA

[Día16] Gerardo Con Díaz. The text in the machine: American copyright law and the many natures of software, 1974–1978. Technology and Culture, 57(4):753–779, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/637924.

Dick:2010:BUO

[Dic10] Steven J. Dick. British University Observatories, 1772–1939 (review). Technology and Culture, 51(1):249–251, January 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/375809.

Dienel:2010:DEB

[Die10] Hans-Liudger Dienel. Driving Europe: Building Europe on Roads in the Twentieth Century (review). Technology and Culture, 51 (2):500–502, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381262.

Dienel:2011:TBT

[Die11]  Hans-Liudger Dienel.  Technikgeschichte: Basistexte, and:

Technikgeschichte: Eine Einführung in ihre Konzepte und Forschungsergebnisse (review). Technology and Culture, 52(4): 864–865, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454570.

Dienel:2012:DDU

[Die12] Hans-Liudger Dienel. Dieselautos in Deutschland und den USA: Zum Verhältnis von Technologie, Konsum und Politik, 1949–2005 (review). Technology and Culture, 53(4):937–939, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490745.

DiMoia:2010:ASC

[DiM10] John DiMoia. Atoms for sale?: Cold War institution-building and the South Korean Atomic Energy Project, 1945–1965. Technology and Culture, 51(3):589–618, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/390838.

DiMoia:2015:ITT

[DiM15] John P. DiMoia. Imperial Technoscience: Transnational Histories of MRI in the United States, Britain, and India by Amit Prasad (review). Technology and Culture, 56(2):562–563, April 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/582339.

Diogo:2011:PST

[Dio11] Maria Paula Diogo. Popularizing Science and Technology in the European Periphery, 1800–2000 (review). Technology and Culture, 52(2):406–408, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431379.

Diogo:2012:CTL

[Dio12] Maria Paula Diogo. Changes in the Technological Landscape. Essays in the History of Science and Technology (review). Technology and Culture, 53(2):475–477, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/476821.

Disco:2010:NUE

[Dis10] Cornelis Disco. Notes on the Underground: An Essay on Technology, Society, and the Imagination (review). Technology and Culture, 51(3):761–763, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390861.

Disco:2011:SNV

[Dis11] Cornelis Disco. Spoorwegen in Nederland van 1834 tot Nu (review). Technology and Culture, 52(3):633–635, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447795.

Disco:2015:FMC

[Dis15] Nil Disco. Futures of Modernity: Challenges for Cosmopolitical Thought and Practice ed. by Michael Heinlein etal. (review). Technology and Culture, 56(4):973–975, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601478.

Divall:2010:MHT

[Div10] Colin Divall. Mobilizing the history of technology. Technology and Culture, 51(4):938–960, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/403273.

Daniels:2018:BRR

[DK18] Mario Daniels and John Krige. Beyond the reach of regulation? “Basic” and “Applied” research in the Early Cold War United States. Technology and Culture, 59(2):226–250, April 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/697231.

Dittmann:2018:RYQ

[DL18] Frank Dittmann and Nicolas Lange. Robots: The 500-year quest to make machines human. Technology and Culture, 59(1):159–167, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692171.

delaPena:2010:HTR

[dlP10] Carolyn de la Peña. The history of technology, the resistance of archives, and the whiteness of race. Technology and Culture, 51 (4):919–937, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403272.

Dobres:2010:LMW

[Dob10] Marcia-Anne Dobres. Living in a Material World: Economic Sociology Meets Science and Technology Studies (review). Technology and Culture, 51(1):284–286, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375808.

Dobraszczyk:2016:BAN

[Dob16] Paul Dobraszczyk. British Art in the Nuclear Age ed. by Catherine Jolivette (review). Technology and Culture, 57(2):488–490, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/619075.

Doern:2012:ECR

[Doe12] Bruce Doern. Energy in Canada (review). Technology and Culture, 53(3):733–735, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481961.

Dorn:2010:DHH

[Dor10] Harold Dorn. Death from the Heavens: A History of Strategic Bombing (review). Technology and Culture, 51(4):1035–1036, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403291.

Dotson:2012:WB

[Dot12] John Dotson. A world in a book. Technology and Culture, 53 (3):688–691, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481911.

Douglas:2010:STQ

[Dou10] Susan J. Douglas. Some thoughts on the question “How do new things happen?”. Technology and Culture, 51(2):293–304, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381270.

Doughty:2016:FAT

[Dou16] Robert Doughty. The French Army’s Tank Force and Armoured Warfare in the Great War: The Artillerie Spéciale by Tim Gale (review). Technology and Culture, 57(2):479–481, April 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/619070.

Doucet:2019:DDM

[Dou19] Brian Doucet. DIY Detroit: Making Do in a City without Services by Kimberley Kinder (review). Technology and Culture, 60(2):645– 647, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
726952.

Downes:2010:PSF

[Dow10] Kieran Downes. “Perfect sound forever”: Innovation, aesthetics, and the re-making of compact disc playback. Technology and Culture, 51(2):305–331, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381297.

Downey:2015:LIN

[Dow15] Greg Downey. The Life Informatic: Newsmaking in the Digital Era by Dominic Boyer (review). Technology and Culture, 56(1):298–299, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572143.

Downey:2016:CTM

[Dow16] Gregory J. Downey. Comparative Textual Media: Transforming the Humanities in the Postprint Era ed. by N. Katherine Hayles and Jessica Pressman (review). Technology and Culture, 57(1):242–243, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
611810.

Downey:2017:ACA

[Dow17] Greg Downey. Ambient Commons: Attention in the Age of Embodied Information by Malcolm McCullough (review). Technology and Culture, 58(2):606–607, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/663006.

Dowd:2018:MLM

[Dow18] Anne S. Dowd. Manufactured Light: Mirrors in the Mesoamerican Realm ed. by Emiliano Gallaga and Marc G. Blainey (review). Technology and Culture, 59(2):473–475, April 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/697241.

Dreicer:2010:BBB

[Dre10] Gregory K. Dreicer. Building bridges and boundaries: The lattice and the tube, 1820–1860. Technology and Culture, 51(1):126–163, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
375780.

Drucker:2012:APH

[Dru12] Donna J. Drucker. America and the Pill: A History of Promise, Peril, and Liberation (review). Technology and Culture, 53(2):515– 516, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476822.

Drucker:2015:SMR

[Dru15] Donna J. Drucker. Seizing the Means of Reproduction by Michelle Murphy (review). Technology and Culture, 56(3):782–784, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593091.

Dryer:2019:FWF

[Dry19] Theodora Dryer. Farm Worker Futurism: Speculative Technologies of Resistance by Curtis Marez (review). Technology and Culture, 60(3):924–926, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/731460.

deSyon:2017:RRC

[dS17] Guillaume de Syon. Rockets and Revolution: A Cultural History of Early Spaceflight by Michael G. Smith (review). Technology and Culture, 58(1):276–278, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/648269.

daSilva:2019:SNU

[dSS19] Álvaro Ferreira da Silva and M. Luísa Sousa. The “Script” of a new urban layout: Mobility, environment, and embellishment in Lisbon’s streets (1850–1910). Technology and Culture, 60(1):65–97, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
719946.

Dugdale:2010:CDS

[Dug10a] Timothy Dugdale. Critical Digital Studies: A Reader (review). Technology and Culture, 51(3):766–768, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390864.

Dugdale:2010:DCR

[Dug10b] Timothy Dugdale. Digital Culture (review). Technology and Culture, 51(1):277–279, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/375807.

Dugdale:2011:OLT

[Dug11] Timothy Dugdale. Online a Lot of the Time: Ritual, Fetish, Sign (review). Technology and Culture, 52(2):426–427, April 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/431390.

Dupre:2015:AER

[Dup15] Sven Dupré. The Automaton in English Renaissance Literature ed. by Wendy Beth Hyman (review). Technology and Culture, 56(3):751–752, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/593073.

Dussauge:2010:TSB

[Dus10] Isabelle Dussauge. Technology and Society: Building Our Sociotechnical Future (review). Technology and Culture, 51(3):778– 780, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
390871.

Dussauge:2011:CSW

[Dus11] Isabelle Dussauge. CT Suite: The Work of Diagnosis in the Age of Noninvasive Cutting (review). Technology and Culture, 52(1): 225–226, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416441.

Dusek:2018:OBH

[Dus18] Val Dusek. Our Battle for the Human Spirit: Scientific Knowing, Technical Doing, and Daily Living by Willem H. Vanderburg (review). Technology and Culture, 59(3):820–822, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/703167.

deVos:2012:PCH

[dV12] Els de Vos. Pastoral Capitalism: A History of Suburban Corporate Landscapes (review). Technology and Culture, 53(4):942–943, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
490748.

Edgerton:2010:ITH

[Edg10] David Edgerton. Innovation, technology, or history: What is the historiography of technology about? Technology and Culture, 51(3):680–697, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390842.

Edson:2010:AAA

[Eds10] Evelyn Edson. Aspects of the Astrolabe: “Architectonica Ratio” in Tenth- and Eleventh-Century Europe (review). Technology and Culture, 51(3):776–777, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390869.

Edwards:2012:TSL

[Edw12a] Clive Edwards. Talking Shop: The Language of Craft in an Age of Consumption (review). Technology and Culture, 53(4):932–933, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
490742.

Edwards:2012:IDF

[Edw12b] Clive Edwards. “Improving” the decoration of furniture: Imitation and mechanization in the marquetry process in Britain and America, 1850–1900. Technology and Culture, 53(2):401–434, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476810.

Edwards:2016:CFG

[Edw16] James Frederick Edwards. The concave faces of the Great Pyramid: An explanation. Technology and Culture, 57(4):909–925, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637929.

Edwards:2017:MPM

[Edw17] Elizabeth Edwards. Making Photography Matter: A Viewer’s History from the Civil War to the Great Depression by Cara A. Finnegan (review). Technology and Culture, 58(2):581–583, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662992.

Edward:2018:WDS

[Edw18] Frank Edward. What Do Science, Technology, and Innovation Mean from Africa? ed. by Clapperton Chakanetsa Mavhunga (review). Technology and Culture, 59(3):815–816, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/703164.

Eisler:2010:LWT

[Eis10] Matthew N. Eisler. The Language of Work: Technical Communication at Lukens Steel, 1810 to 1925 (review). Technology and Culture, 51(1):252–254, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375806.

Eisler:2017:EBB

[Eis17] Matthew N. Eisler. Exploding the black Box: Personal computing, the notebook battery crisis, and postindustrial systems thinking. Technology and Culture, 58(2):368–391, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/662981.

Elichirigoity:2012:VMC

[Eli12] Fernando Elichirigoity. A Vast Machine: Computer Models, Climate Data, and the Politics of Global Warming (review). Technology and Culture, 53(1):243–245, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/466805.

Engerman:2012:EGE

[Eng12] David C. Engerman. Entangled Geographies: Empire and Technopolitics in the Global Cold War (review). Technology and Culture, 53(3):729–731, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481932.

Ensmenger:2010:TTW

[Ens10] Nathan Ensmenger. Technological Turf Wars: A Case Study of the Computer Antivirus Industry (review). Technology and Culture, 51 (4):1054–1056, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403303.

Ensmenger:2011:PCS

[Ens11] Nathan Ensmenger. Punch-Card Systems and the Early Information Explosion, 1880–1945 (review). Technology and Culture, 52 (2):420–422, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431387.

Ensmenger:2012:DCT

[Ens12] Nathan Ensmenger. The digital construction of technology: Rethinking the history of computers in society. Technology and Culture, 53(4):753–776, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/490727.

Ensmenger:2018:EHC

[Ens18] Nathan Ensmenger. The environmental history of computing. Technology and Culture, 59(4S):S7–S33, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/712112.

Epperson:2015:BDI

[Epp15] Bruce Epperson. Bicycle Design: An Illustrated History by Tony Hadland and Hans-Erhard Lessing (review). Technology and Culture, 56(4):983–984, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/601484.

Erickson:2015:OMC

[Eri15] Paul Erickson. The Open Mind: Cold War Politics and the Sciences of Human Nature by Jamie Cohen-Cole (review). Technology and Culture, 56(3):775–777, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/593087.

Emgin:2013:TNS

[ERST13] Bahar Emgin, Georgina Ruff, Matteo Serafini, and Jytte Thorndahl. Thirty-ninth symposium of the International Committee for the History of Technology: “Technology, the Arts, and Industrial Culture,” Barcelona, Spain, 10-14 July 2012. Technology and Culture, 54(1):158–166, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/500793.

Ertsen:2012:QNB

[Ert12] Maurits Ertsen. Quagmire: Nation-Building and Nature in the Mekong Delta (review). Technology and Culture, 53(4):944–945, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
490749.

Ertsen:2016:FSW

[Ert16] Maurits Ertsen. The Fabric of Space: Water, Modernity, and the Urban Imagination by Matthew Gandy (review). Technology and Culture, 57(3):684–685, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630944.

Esbester:2015:PRS

[Esb15] Mike Esbester. Promoting road safety. Technology and Culture, 56(2):493–497, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582319.

Estrada:2010:IAD

[Est10] Gilbert Estrada. Inventing Autopia: Dreams and Visions of the Modern Metropolis in Jazz Age Los Angeles (review). Technology and Culture, 51(4):1042–1043, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403295.

Evans:2015:MBE

[Eva15] Chris Evans. Matthew Boulton: Enterprising Industrialist of the Enlightenment, edited by Kenneth Quickenden, Sally Baggott, and Malcolm Dick (review). Technology and Culture, 56(1):271–272, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572127.

Evans:2016:ASR

[Eva16] Harry B. Evans. The Archaeology of Sanitation in Roman Italy: Toilets, Sewers, and Water Systems by Ann Olga Koloski-Ostrow (review). Technology and Culture, 57(2):469–470, April 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/619064.

Evans:2018:BWS

[Eva18] Sterling Evans. Bitter Waters: The Struggles of the Pecos River by Patrick Dearen (review). Technology and Culture, 59(1):175–176, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692176.

Evans:2012:ESI

[EW12] Chris Evans and Alun Withey. An enlightenment in steel?: Innovation in the steel trades of Eighteenth-Century Britain. Technology and Culture, 53(3):533–560, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/481905.

Esbester:2015:IGP

[EW15] Mike Esbester and Jameson M. Wetmore. Introduction: Global perspectives on road safety history. Technology and Culture, 56 (2):307–318, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582310.

Faith:2016:ABH

[Fai16] Thomas I. Faith. American Biodefense: How Dangerous Ideas about Biological Weapons Shape National Security by Frank L. Smith III (review). Technology and Culture, 57(2):495–497, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/619079.

Fearnley:2016:GJL

[Fea16] Lyle Fearnley. Gene Jockeys: Life Science and the Rise of Biotech Enterprise by Nicolas Rasmussen (review). Technology and Culture, 57(3):700–701, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630953.

Feaster:2019:EPA

[Fea19] Patrick Feaster. Enigmatic proofs: The archiving of Édouard–Léon Scott de Martinville’s phonautograms. Technology and Culture, 60 (2S):S14–S38, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/727297.

Feenberg:2011:HTP

[Fee11] Andrew Feenberg. Heidegger’s Technologies: Postphenomenological Perspectives (review). Technology and Culture, 52(4):811–812, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454540.

Fein:2012:KBH

[Fei12] Michael R. Fein. The King’s Best Highway: The Lost History of the Boston Post Road, the Route that Made America (review). Technology and Culture, 53(2):484–486, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/476823.

Felker:2014:ISS

[Fel14] Craig C. Felker. Information at Sea: Shipboard Command and Control in the U.S. Navy, from Mobile Bay to Okinawa by Timothy S. Wolters (review). Technology and Culture, 55(4):1027–1028, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562862.

Ferreiro:2011:SR

[Fer11a] Larrie D. Ferreiro. Ship (review). Technology and Culture, 52 (4):818–819, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454545.

Ferreiro:2011:SHB

[Fer11b] Larrie D. Ferreiro. The social history of the bulbous bow. Technology and Culture, 52(2):335–359, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/431364.

Ferro:2012:CII

[Fer12] David L. Ferro. The Culture of the Internet and the Internet as Cult: Social Fears and Religious Fantasies (review). Technology and Culture, 53(3):747–749, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/481958.

Ferreiro:2013:DEA

[Fer13] Larrie D. Ferreiro. The Dreadnought and the Edwardian Age ed. by Robert J. Blyth et al. (review). Technology and Culture, 54 (1):195–196, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500808.

Ferro:2016:STT

[Fer16] David L. Ferro. Singularities: Technoculture, Transhumanism, and Science Fiction in the 21st Century by Joshua Raulerson (review). Technology and Culture, 57(1):285–286, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611834.

Ferreiro:2019:CSS

[Fer19] Larrie D. Ferreiro. Coal, Steam and Ships: Engineering, Enterprise and Empire on the Nineteenth-Century Seas by Crosbie Smith (review). Technology and Culture, 60(4):1105–1107, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/741411.

Figg:2010:AST

[Fig10] Kristen M. Figg. The Art, Science, and Technology of Medieval Travel (review). Technology and Culture, 51(2):499–500, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381293.

Finlay:2010:FBT

[Fin10] Mark Finlay. Far beyond tractors: Envirotech and the intersections of technology, agriculture, and the environment. Technology and Culture, 51(2):480–485, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381257.

Finlay:2011:NGS

[Fin11] Mark Finlay. Notes from the Ground: Science, Soil, and Society in the American Countryside (review). Technology and Culture, 52 (1):188–190, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416419.

Finlay:2013:EGR

[Fin13] Mark Finlay. Europe’s Green Revolution and Others Since: The Rise and Fall of Peasant-Friendly Plant Breeding by Jonathan Harwood (review). Technology and Culture, 54(4):988–990, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/530533.

Finstad:2017:NGS

[Fin17] Terje Finstad. Naked gene Salmon: Debating fish, genes, and the politics of science in the “Age of publics”. Technology and Culture, 58(1):97–120, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648251.

First:2012:MPT

[Fir12] Joshua First. Moscow Prime Time: How the Soviet Union Built the Media Empire that Lost the Cold War (review). Technology and Culture, 53(3):731–733, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/481960.

First:2017:SSM

[Fir17] Joshua First. Soviet Space Mythologies: Public Images, Private Memories and the Making of a Cultural Identity by Slava Gerovitch (review). Technology and Culture, 58(2):600–602, April 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/663003.

Fischer:2013:IFL

[Fis13] Michael M. J. Fischer. Icons, frames, and language games: Bruno Latour, On the Modern Cult of the Factish Gods. Technology and Culture, 54(4):963–967, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530520.

Fischer:2015:AMT

[Fis15] Michael M. J. Fischer. Appletopia: Media Technology and the Religious Imagination of Steve Jobs by Brett T. Robinson (review). Technology and Culture, 56(1):291–293, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/572139.

Fjaestad:2012:EHC

[Fja12] Maja Fjaestad. Environmental Histories of the Cold War (review). Technology and Culture, 53(1):233–234, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466799.

Fjaestad:2013:NWH

[Fja13] Maja Fjaestad. Nuclear waste and historical time. Technology and Culture, 54(2):371–372, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506074.

Fjaestad:2015:FBR

[Fja15] Maja Fjaestad. Fast breeder reactors in Sweden: Vision and reality. Technology and Culture, 56(1):86–114, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/572112.

Fleetwood:2019:TSM

[Fle19] Lachlan Fleetwood. Thrifty Science: Making the Most of Materials in the History of Experiment by Simon Werrett (review). Technology and Culture, 60(4):1102–1103, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741412.

Forgan:2010:VEP

[For10] Sophie Forgan. The Victorian Eye: A Political History of Light and Vision in Britain, 1800–1910 (review). Technology and Culture, 51 (1):246–247, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375805.

Forrant:2011:ABL

[For11] Robert Forrant. At the Altar of the Bottom Line: The Degradation of Work in the 21st Century (review). Technology and Culture, 52 (1):207–208, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416430.

Forgan:2017:IOA

[For17] Sophie Forgan. Iron, Ornament and Architecture in Victorian Britain: Myth and Modernity, Excess and Enchantment by Paul Dobraszczyk (review). Technology and Culture, 58(1):271–273, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
648266.

Fox:2011:RIP

[Fox11] Robert Fox. La République des inventeurs: Propriété et organisation de l’innovation en France (1791–1922) (review). Technology and Culture, 52(3):631–633, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/447794.

Ferreiro:2016:CWW

[FP16] Larrie D. Ferreiro and Alexander Pollara. Contested waterlines: The wave-line theory and shipbuilding in the Nineteenth Century. Technology and Culture, 57(2):414–444, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619057.

Fidler:2018:FAI

[FR18] Bradley Fidler and Andrew L. Russell. Financial and administrative infrastructure for the early Internet: Network maintenance at the Defense Information Systems Agency. Technology and Culture, 59(4):899–924, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711766.

Franz:2013:UUC

[Fra13] Kathleen Franz. User Unfriendly: Consumer Struggles with Personal Technologies, from Clocks and Sewing Machines to Cars and Computers by Joseph J. Corn (review). Technology and Culture, 54(2):434–435, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506102.

Franz:2016:PPH

[Fra16] Kathleen Franz. Packaged Pleasures: How Technology and Marketing Revolutionized Desire by Gary S. Cross, Robert N. Proctor (review). Technology and Culture, 57(4):1027–1028, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637951.

Franz:2018:WWC

[Fra18] Kathleen Franz. Women at the Wheel: A Century of Buying, Driving, and Fixing Cars by Katherine J. Parkin (review). Technology and Culture, 59(4):981–983, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/711780.

Freed:2011:RMH

[Fre11] Libbie Freed. The Routes of Man: How Roads are Changing the World and the Way We Live Today (review). Technology and Culture, 52(4):850–852, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/454563.

Freeze:2012:CTT

[Fre12] Karen J. Freeze. Czechoslovak theater technology under Communism: Ambassador to the West. Technology and Culture, 53(2):442– 460, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476812.

French:2015:MTC

[Fre15] Mike French. The Medical Trade Catalogue in Britain, 1870–1914 by Claire L. Jones (review). Technology and Culture, 56(4):988–989, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
601487.

French:2018:GBA

[Fre18] Dan French. The Grid: Biography of an American Technology by Julie A. Cohn (review). Technology and Culture, 59(4):975–976, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
711776.

Friedel:2013:OOO

[Fri13] Robert Friedel. Obsolescence: Origins and outrages. Technology and Culture, 54(1):167–169, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500794.

Friedel:2016:AMH

[Fri16] Robert Friedel. Applied Minds: How Engineers Think by Guru Madhavan (review). Technology and Culture, 57(2):464–465, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/619061.

Frohlich:2010:CCC

[Fro10] Xaq Frohlich. Condensed Capitalism: Campbell Soup and the Pursuit of Cheap Production in the Twentieth Century (review). Technology and Culture, 51(2):523–525, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/381285.

Frost:2014:LFK

[Fro14] Gary Frost. Lee de Forest: King of Radio, Television, and Film by Mike Adams (review). Technology and Culture, 55(3):751–752, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
551625.

Frumer:2014:TTH

[Fru14] Yulia Frumer. Translating time: Habits of Western-style timekeeping in Late Edo Japan. Technology and Culture, 55(4):785–820, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562832.

Frumer:2015:NIF

[Fru15] Yulia Frumer. Non-industrial faces of modern times. Technology and Culture, 56(4):965–969, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601475.

Frumer:2017:GTT

[Fru17] Yulia Frumer. The Global Transformation of Time: 1870–1950 by Vanessa Ogle (review). Technology and Culture, 58(1):275–276, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
648268.

Gaggio:2015:CTR

[Gag15] Dario Gaggio. Consumers, Tinkerers, Rebels: The People Who Shaped Europe by Ruth Oldenziel and Mikael Hård (review).

Technology and Culture, 56(4):993–995, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601490.

Gainor:2013:AAF

[Gai13] Christopher Gainor. The Atlas and the Air Force: Reassessing the beginnings of America’s first intercontinental ballistic missile. Technology and Culture, 54(2):346–370, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/506073.

Gainor:2014:RA

[Gai14] Christopher Gainor. Response to alic. Technology and Culture, 55(1):285, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538940.

Gainor:2015:AFR

[Gai15] Christopher Gainor. Adapting to Flexible Response: 1960–1968 by Walter S. Poole (review). Technology and Culture, 56(4):998–1000, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
601493.

Gainor:2018:URS

[Gai18] Christopher Gainor. An Untaken Road: Strategy, Technology and the Hidden History of America’s Mobile ICBMs by Steven A. Pomeroy (review). Technology and Culture, 59(2):485–487, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/697248.

Gainor:2019:TCA

[Gai19] Christopher Gainor. Tiger Check: Automating the U.S. Air Force Fighter Pilot in Air-to-Air Combat, 1950–1980 by Steven A. Fino (review). Technology and Culture, 60(3):914–915, July 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/731453.

Gallimore:2011:IHF

[Gal11] Mark Gallimore. Inventing heroics: Frank Julian Sprague. Technology and Culture, 52(2):385–388, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/431368.

Galvan:2015:MGT

[Gal15] Jill Galvan. The Machine and the Ghost: Technology and Spiritualism in Nineteenth- to Twenty-First-Century Art and Culture ed. by Sas Mays and Neil Matheson (review). Technology and Culture, 56(1):286–287, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572136.

Galloway:2016:KMD

[Gal16] Patricia Galloway. Knowledge Machines: Digital Transformations of the Sciences and Humanities by Eric T. Meyer, Ralph Schroeder (review). Technology and Culture, 57(4):1045–1046, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637961.

Galambos:2017:TIL

[Gal17] Louis Galambos. Technological Innovation in Legacy Sectors by William B. Bonvillian and Charles Weiss (review). Technology and Culture, 58(2):611–613, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/663009.

Gamst:2011:MMM

[Gam11] Frederick C. Gamst. A Most Magnificent Machine: America Adopts the Railroad, 1825–1862 (review). Technology and Culture, 52 (4):826–827, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454550.

Gamst:2013:RGR

[Gam13] Frederick Gamst. Railway by George Revill (review). Technology and Culture, 54(2):407–408, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506087.

Gamst:2015:ARD

[Gam15] Frederick C. Gamst. American Railroads: Decline and Renaissance in the Twentieth Century by Robert E. Gallamore and John R. Meyer (review). Technology and Culture, 56(2):551–553, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582333.

Gangloff:2010:AAL

[Gan10] Amy Gangloff. The Automobile and American Life (review). Technology and Culture, 51(2):517–518, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/381258.

Gangloff:2011:HBE

[Gan11] Amy Gangloff. Hitting the Brakes: Engineering Design and the Production of Knowledge (review). Technology and Culture, 52 (2):429–431, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431392.

Gangloff:2013:SAH

[Gan13] Amy Gangloff. Safety in accidents: Hugh DeHaven and the development of crash injury studies. Technology and Culture, 54 (1):40–61, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500796.

Gangloff:2017:FMR

[Gan17a] Amy Gangloff. Forensic Media: Reconstructing Accidents in Accelerated Modernity by Greg Siegel (review). Technology and Culture, 58(1):295–296, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648280.

Ganz:2017:WFE

[Gan17b] Cheryl R. Ganz. World’s Fairs on the Eve of War: Science, Technology, & Modernity, 1937–1942 by Robert H. Kargon et al. (review). Technology and Culture, 58(1):273–274, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648267.

Gangloff:2019:BTA

[Gan19] Amy Gangloff. Back on Track: American Railroad Accidents and Safety, 1965–2015 by Mark Aldrich (review). Technology and Culture, 60(4):1117–1119, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/741391.

Garner:2013:LLE

[Gar13] Kenneth Garner. Learning with the Lights Off: Educational Film in the United States ed. by Devin Orgeron, Marsha Orgeron, and Dan Streible (review). Technology and Culture, 54(3):681–682, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520302.

Garner:2018:SBS

[Gar18] Kenneth Garner. Studios Before the System: Architecture, Technology, and the Emergence of Cinematic Space by Brian R. Jacobson (review). Technology and Culture, 59(3):799–801, July 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/703155.

Gatejel:2018:BBP

[Gat18] Luminita Gatejel. Building a better passage to the sea: Engineering and river management at the mouth of the Danube, 1829–61. Technology and Culture, 59(4):925–953, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711767.

Geoghegan:2015:TLH

[Geo15] Bernard Dionysius Geoghegan. Toward liberal histories of computing: Fred Turner, the democratic surround. Technology and Culture, 56(3):745–748, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/593071.

Gerovitch:2011:RCK

[Ger11] Slava Gerovitch. Red Cosmos: K. E. Tsiolkovskii, Grandfather of Soviet Rocketry (review). Technology and Culture, 52(2):408–410, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431380.

Gerovitch:2013:IAA

[Ger13] Slava Gerovitch. Inventing the American Astronaut by Matthew H. Hersch (review). Technology and Culture, 54(3):687–689, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520305.

Gere:2016:NMO

[Ger16] Charlie Gere. Necromedia by Marcel O’Gorman (review). Technology and Culture, 57(3):711–713, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/630960.

Giffard:2015:KHL

[Gif15a] Hermione Giffard. Krupp: A History of the Legendary German Firm by Harold James (review). Technology and Culture, 56(4): 990–991, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601488.

Giffard:2015:HSF

[Gif15b] Hermione Giffard. A hard sell: Factors influencing the interwar adoption of tungsten carbide cutting tools in Germany, Britain, and the United States. Technology and Culture, 56(4):895–927, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
601472.

Giffard:2017:GMF

[Gif17a] Hermione Giffard. Der Gelehrte und der Manager: Franz Reuleaux (1829–1905) und Alois Riedler (1850–1936) in Technik, Wissenschaft und Gesellschaft by Wolfgang König (review). Technology and Culture, 58(1):266–267, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648263.

Giffard:2017:EIT

[Gif17b] Hermione Giffard. Europe’s Infrastructure Transition: Economy, War, Nature by Per Högselius, Arne Kaijser, Erik van der Vleuten (review). Technology and Culture, 58(3):881–883, July 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/669482.

Gillispie:2010:STE

[Gil10] Charles C. Gillispie. Le savant et la tempête: Étudier l’atmosphère et prévoir le temps au XIX^{e} siècle (review). Technology and Culture, 51(1):243–244, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/375804.

Gilbert:2011:HMD

[Gil11a] James Gilbert. Health and Medicine on Display: International Expositions in the United States, 1876–1904 (review). Technology and Culture, 52(2):396–397, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/431373.

Gillispie:2011:LAT

[Gil11b] Charles C. Gillispie. L’Aérostation au temps des Lumières (review). Technology and Culture, 52(2):394–396, April 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/431372.

Gillin:2015:PPA

[Gil15] Edward Gillin. Prophets of progress: Authority in the scientific projections and religious realizations of the Great Eastern Steamship. Technology and Culture, 56(4):928–956, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/601473.

Gitelman:2010:IVB

[Git10] Lisa Gitelman. Inherent Vice: Bootleg Histories of Videotape and Copyright (review). Technology and Culture, 51(4):1051–1053, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403301.

Gerali:2018:FFS

[GKMV18] Francesco Gerali, Marcin Krasnodȩbski, Saara Matala, and Dick Van Lente. Forty-fourth Symposium of the International Committee for the History of Technology: “Science, Technology, and Medicine between the Global and the Local,” Rio de Janeiro, Brazil, 24–29 July 2017. Technology and Culture, 59(1):152–158, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692170.

Glanc:2017:NTA

[Gla17] Tomáš Glanc. New Tendencies: Art at the Threshold of the Information Revolution (1961–1978) by Armin Medosch (review). Technology and Culture, 58(4):1107–1108, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679844.

Godin:2015:CWS

[God15] Benoît Godin. Cold War Social Science: Knowledge Production, Liberal Democracy, and Human Nature ed. by Mark Solovey and Hamilton Cravens (review). Technology and Culture, 56(1): 284–286, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572135.

Godin:2016:BIT

[God16a] Benoît Godin. Beyond Innovation: Technology, Institution, and Change as Categories for Social Analysis by Thomas Kaiserfeld (review). Technology and Culture, 57(1):244–245, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611811.

Godin:2016:TIO

[God16b] Benoît Godin. Technological innovation: On the origins and development of an inclusive concept. Technology and Culture, 57(3):527– 556, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
630928.

Godin:2018:PIW

[God18] Benoît Godin. The Politics of Innovation: Why Some Countries Are Better Than Others at Science and Technology by Mark Zachary Taylor (review). Technology and Culture, 59(2):489–490, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
697250.

Goldstein:2012:DEE

[Gol12] Carolyn M. Goldstein. Domesticating Electricity: Expertise, Uncertainty, and Gender, 1880–1914 (review). Technology and Culture, 53(4):935–937, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/490744.

Golan:2014:TMS

[Gol14a] Tal Golan. The Truth Machine: A Social History of the Lie Detector by Geoff Bunn (review). Technology and Culture, 55(3):752– 754, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
551626.

Goldstein:2014:BMR

[Gol14b] Carolyn M. Goldstein. Building a Market: The Rise of the Home Improvement Industry, 1914–1960 by Richard Harris (review). Technology and Culture, 55(3):756–757, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/551628.

Golas:2015:MMC

[Gol15] Peter J. Golas. Mining, Monies, and Culture in Early Modern Societies: East Asian and Global Perspectives. edited by Nanny Kim and Keiko Nagase-Reimer (review). Technology and Culture, 56(1):267–268, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572125.

Gooday:2011:PFH

[Goo11] Graeme Gooday. Playing with Fire: Histories of the Lightning Rod (review). Technology and Culture, 52(3):626–627, July 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/447791.

Good:2019:SSS

[Goo19a] Katie Day Good. Sight-seeing in school: Visual technology, virtual experience, and world citizenship in American education, 1900– 1930. Technology and Culture, 60(1):98–131, January 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/719947.

Gooday:2019:PNI

[Goo19b] Graeme Gooday. Powering the Nation: Images of the Shannon Scheme and Electricity in Ireland by Sorcha O’Brien (review). Technology and Culture, 60(3):908–909, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731461.

Gopinath:2015:CST

[Gop15] Sumanth Gopinath. Chasing Sound: Technology, Culture, and the Art of Studio Recording from Edison to the LP by Susan Schmidt Horning (review). Technology and Culture, 56(3):763–765, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593080.

Gordon:2010:BIA

[Gor10] Robert B. Gordon. Baltic Iron in the Atlantic World in the Eighteenth Century (review). Technology and Culture, 51(3):756–757, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
390858.

Gordon:2011:MFA

[Gor11] Robert B. Gordon. Metal Fatigue: American Bosch and the Demise of Metalworking in the Connecticut River Valley (review). Technology and Culture, 52(1):206–207, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416429.

Gorman:2012:DAC

[Gor12] Hugh Gorman. DDT and the American Century: Global Health, Environmental Politics, and the Pesticide that Changed the World (review). Technology and Culture, 53(4):945–947, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490750.

Gordon:2015:HTR

[Gor15] Robert Gordon. Here and There: Reading Pennsylvania’s Working Landscapes by Bill Conlogue (review). Technology and Culture, 56(2):542–543, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582328.

Gorman:2017:AEP

[Gor17] Hugh Gorman. American Energy Policy in the 1970s ed. by Robert Lifset (review). Technology and Culture, 58(3):890–891, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/669487.

Gott:2012:CTD

[Got12] Suzanne Gott. The Copyright Thing Doesn’t Work Here: Adinkra and Kente Cloth and Intellectual Property in Ghana (review). Technology and Culture, 53(3):749–751, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/481959.

Guise-Richardson:2010:RVC

[GR10]  Cai Guise-Richardson.  Redefining vulcanization: Charles

goodyear, patents, and industrial control, 1834–1865. Technology and Culture, 51(2):357–387, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/381290.

Graber:2011:IET

[Gra11] Frédéric Graber. Impossible Engineering: Technology and Territoriality on the Canal du Midi (review). Technology and Culture, 52(3):620–622, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447788.

Gray:2015:VWM

[Gra15] Chris Hables Gray. Virtual War and Magical Death: Technologies and Imaginaries for Terror and Killing ed. by Neil L. Whitehead and Sverker Finnström (review). Technology and Culture, 56(4): 1012–1013, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601501.

Grant:2016:LIC

[Gra16] Jonathan Grant. Lonely Ideas: Can Russia Compete? by Loren Graham (review). Technology and Culture, 57(1):262–264, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611821.

Gray:2018:DWT

[Gra18] Chris Hables Gray. Drones, war, and technological seduction. Technology and Culture, 59(4):954–962, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711768.

Grew:2011:GFR

[Gre11] Raymond Grew. The Globalization of Food (review). Technology and Culture, 52(1):210–211, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416432.

Greene:2014:RTH

[Gre14] Ann N. Greene. Reflections on Tom Hughes. Technology and Culture, 55(4):958–963, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/562839.

Grier:2012:SOM

[Gri12] David Alan Grier. A Science of Operations: Machines, Logic, and the Invention of Programming (review). Technology and Culture, 53(3):741–742, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481954.

Grossman:2014:MEF

[Gro14] Sarah E. M. Grossman. Mining engineers and fraud in the U.S.– Mexico borderlands, 1860–1910. Technology and Culture, 55(4): 821–849, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562833.

Grossman:2015:GON

[Gro15] Sarah Grossman. Gambling on Ore: The Nature of Metal Mining in the United States, 1860–1910 by Kent A. Curtis (review). Technology and Culture, 56(4):985–986, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601485.

Gromme:2018:ANT

[Gro18] Francisca Grommé. Actor-Network Theory and Crime Studies: Explorations in Science and Technology ed. by Dominique Robert, Martin Dufresne (review). Technology and Culture, 59(1):201–202, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692164.

Gross:2019:LCW

[Gro19] Rachel S. Gross. Layering for a Cold War: The M-1943 combat system, military testing, and clothing as technology. Technology and Culture, 60(2):378–408, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726931.

Grunden:2012:DJM

[Gru12] Walter E. Grunden. In Defense of Japan: From the Market to the Military in Space Policy (review). Technology and Culture, 53 (4):949–950, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490752.

Grubiak:2014:RHU

[Gru14] Margaret M. Grubiak. Radiance from Halcyon: A Utopian Experiment in Religion and Science by Paul Eli Ivey (review). Technology and Culture, 55(3):754–756, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551627.

Grubiak:2016:AEC

[Gru16a] Margaret M. Grubiak. An architecture for the electronic church: Oral Roberts University in Tulsa, Oklahoma. Technology and Culture, 57(2):380–413, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/619056.

Grubiak:2016:CAM

[Gru16b] Margaret M. Grubiak. Characteristically American: Memorial Architecture, National Identity, and the Egyptian Revival by Joy M. Giguere (review). Technology and Culture, 57(1):256–257, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611818.

Guajardo:2013:HML

[Gua13] Guillermo Guajardo. A History of Mining in Latin America: From the Colonial Era to the Present by Kendall W. Brown (review). Technology and Culture, 54(3):659–660, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/520290.

Gudmestad:2013:MMI

[Gud13] Robert Gudmestad. Mountains on the Market: Industry, the Environment, and the South by Randal L. Hall (review). Technology and Culture, 54(2):409–410, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506088.

Gudmestad:2018:PKA

[Gud18] Robert Gudmestad. Plantation Kingdom: The American South and its Global Commodities by Richard Follett, et al. (review). Technology and Culture, 59(1):176–177, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/692177.

Guerrini:2013:ABE

[Gue13] Anita Guerrini. Artisans of the Body in Early Modern Italy: Identities, Families and Masculinities by Sandra Cavallo (review). Technology and Culture, 54(2):396–398, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/506081.

Guillerme:2016:EPS

[Gui16] André Guillerme. Écrire et publier des savoirs au XIXe siècle: Une revue en construction: Les Annales des Ponts et Chaussées (1831– 1866) by Nathalie Montel (review). Technology and Culture, 57 (4):1012–1013, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637942.

Guldi:2019:PDA

[Gul19] Jo Guldi. Parliament’s debates about infrastructure: An exercise in using dynamic topic models to synthesize historical change. Technology and Culture, 60(1):1–33, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/719944.

Hacker:2015:CHF

[Hac15] Barton C. Hacker. A Cultural History of Firearms in the Age of Empire ed. by Karen Jones, Giacomo Macola, David Welch (review). Technology and Culture, 56(3):754–755, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/593075.

Haddow:2010:HHC

[Had10] Gill Haddow. Heredity and Hope: The Case for Genetic Screening (review). Technology and Culture, 51(2):532–533, April 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/381278.

Hafter:2011:MEE

[Haf11] Daryl Hafter. Materials and Expertise in Early Modern Europe: Between Market and Laboratory (review). Technology and Culture, 52(3):622–624, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447789.

Hafter:2017:E

[Haf17] Daryl M. Hafter. To the Editor. Technology and Culture, 58 (4):1114, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679848.

Hafter:2018:SIM

[Haf18] Daryl M. Hafter. The Silk Industries of Medieval Paris: Artisanal Migration, Technological Innovation, and Gendered Experience by Sharon Farmer (review). Technology and Culture, 59(2):475–476, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
697242.

Hagmann:2014:CSI

[Hag14] Johannes-Geert Hagmann. Collider: Step Inside the World’s Greatest Experiment: 13 November 2013–5 May 2014, Science Museum, London. Technology and Culture, 55(3):716–722, July 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/551611.

Hagen:2015:LSD

[Hag15a] Joel B. Hagen. Life Out of Sequence: A Data-Driven History of Bioinformatics by Hallam Stevens (review). Technology and Culture, 56(1):301–303, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/572145.

Hagmann:2015:GGB

[Hag15b] Johannes-Geert Hagmann. Gravity’s Ghost and Big Dog: Scientific Discovery and Social Analysis in the Twenty-First Century by Harry Collins (review). Technology and Culture, 56(2):576–577, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582347.

Hahn:2015:SCD

[Hah15] Hans Peter Hahn. Subversion, Conversion, Development. Cross-Cultural Knowledge and the Politics of Design ed. by James Leach and Lee Wilson (review). Technology and Culture, 56(2):573–574, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582345.

Hahn:2017:MWE

[Hah17] Barbara Hahn. Must we embody context? Technology and Culture, 58(2):556–569, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/662989.

Hahn:2019:AME

[Hah19] Barbara Hahn. The Age of Machinery: Engineering the Industrial Revolution, 1770–1850 by Gillian Cookson (review). Technology and Culture, 60(2):630–631, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726948.

Haiyan:2015:BDW

[Hai15] Yang Haiyan. Barefoot Doctors and Western Medicine in China by Xiaoping Fang (review). Technology and Culture, 56(3):780– 782, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
593090.

Halpern:2012:AM

[Hal12] Orit Halpern. Affective machines. Technology and Culture, 53 (1):189–191, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466771.

Hall:2016:ZFG

[Hal16] Nick Hall. Zoomar: Frank G. Back and the postwar television zoom lens. Technology and Culture, 57(2):353–379, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619055.

Hamilton:2010:FPH

[Ham10a] Shane Hamilton. Fresh: A Perishable History (review). Technology and Culture, 51(4):1031–1033, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403289.

Hamilton:2010:LSC

[Ham10b] Shane Hamilton. Linthead Stomp: The Creation of Country Music in the Piedmont South (review). Technology and Culture, 51 (1):267–268, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375803.

Hamblin:2011:PHW

[Ham11] Jacob Darwin Hamblin. Plutonium: A History of the World’s Most Dangerous Element (review). Technology and Culture, 52(2):415– 417, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431384.

Hamerla:2012:HSO

[Ham12a] Richard Hamerla. Histories of Scientific Observation (review). Technology and Culture, 53(1):206–207, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466784.

Hamm:2012:DST

[Ham12b] Ernst Hamm. Divining Science: Treasure Hunting and Earth Science in Early Modern Germany (review). Technology and Culture, 53(4):925–926, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490738.

Hamlin:2013:DGM

[Ham13] Christopher Hamlin. The Dawn of Green: Manchester, Thirlmere, and Modern Environmentalism by Harriet Ritvo (review). Technology and Culture, 54(4):980–981, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/530528.

Hamerla:2014:AOH

[Ham14a] Rich Hamerla. Agent Orange: History, Science, and the Politics of Uncertainty by Edwin A. Martini (review). Technology and Culture, 55(4):1030–1032, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/562864.

Hamilton:2014:AFF

[Ham14b] Shane Hamilton. Agribusiness, the family farm, and the politics of technological determinism in the Post-World War II United States. Technology and Culture, 55(3):560–590, July 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/551606.

Hamerla:2016:ASO

[Ham16a] Rich Hamerla. The American Synthetic Organic Chemicals Industry: War and Politics, 1910–1930 by Kathryn Steen (review). Technology and Culture, 57(2):481–483, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619071.

Hampton:2016:ARL

[Ham16b] Gregory Hampton. Anatomy of a Robot: Literature, Cinema, and the Cultural Work of Artificial People by Despina Kakoudaki (review). Technology and Culture, 57(4):1024–1025, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637949.

Hamilton:2017:MMS

[Ham17] Vivien Hamilton. Medical machines as symbols of science?: Promoting electrotherapy in Victorian Canada. Technology and Culture, 58(4):1017–1045, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679822.

Hamblin:2018:AMC

[Ham18a] Jacob Darwin Hamblin. After the Map: Cartography, Navigation, and the Transformation of Territory in the Twentieth Century by William Rankin (review). Technology and Culture, 59(3):805–806, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703158.

Hamerla:2018:BGM

[Ham18b] Rich Hamerla. Behind the Gas Mask: The U.S. Chemical Warfare Service in War and Peace by Thomas I. Faith (review). Technology and Culture, 59(1):185–186, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/692182.

Hamilton:2018:PEG

[Ham18c] Shane Hamilton. The Profit of the Earth: The Global Seeds of American Agriculture by Courtney Fullilove (review). Technology and Culture, 59(4):990–992, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/711785.

Hammers:2018:SLI

[Ham18d] Roslyn Lee Hammers. The Social Life of Inkstones: Artisans and Scholars in Early Qing China by Dorothy Ko (review). Technology and Culture, 59(2):477–478, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/697243.

Hangen:2013:LVE

[Han13] Tona Hangen. The Listener’s Voice: Early Radio and the American Public by Elena Razlogova (review). Technology and Culture, 54(3):677–679, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520300.

Hannah:2014:SDB

[Han14] Leslie Hannah. The Strange Death of the British Motor Cycle Industry by Steve Koerner (review). Technology and Culture, 55(4): 1021–1023, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562859.

Haring:2011:HTP

[Har11] Kristen Haring. Hedonizing Technologies: Paths to Pleasure in Hobbies and Leisure (review). Technology and Culture, 52(1): 192–193, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416421.

Harp:2012:VMC

[Har12] Stephen L. Harp. Voiture Minimum: Le Corbusier and the Automobile (review). Technology and Culture, 53(1):215–216, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466789.

Harwood:2013:GPP

[Har13] Jonathan Harwood. Gènes, pouvoirs et profits: Recherche publique et régimes de production des savoirs de Mendel aux OGM by Christophe Bonneuil and Frédéric Thomas (review). Technology and Culture, 54(1):207–208, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500815.

Hardy:2016:TCP

[Har16] Eric M. Hardy. Thirsty City: Politics, Greed, and the Making of Atlanta’s Water Crisis by Skye Borden (review). Technology and Culture, 57(1):267–269, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/611824.

Harwood:2017:USU

[Har17] Jonathan Harwood. “Useful sciences”?: Ursula Klein, Nützliches Wissen. Technology and Culture, 58(3):862–865, July 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/669473.

Hardy:2019:VEH

[Har19] Penelope Hardy. Vast Expanses: A History of the Oceans by Helen M. Rozwadowski (review). Technology and Culture, 60(4): 1098–1099, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741410.

Hashimoto:2016:RDS

[Has16] Takehiko Hashimoto. “Rattraper et dépasser la suisse”: Histoire de l’industrie horlogère japonaise de 1850 a nos jours by Pierre-Yves Donzé (review). Technology and Culture, 57(3):672–673, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/630937.

Hausman:2010:EEP

[Hau10] William J. Hausman. Electrifying Europe: The Power of Europe in the Construction of Electricity Networks (review). Technology and Culture, 51(3):753–754, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/390856.

Hauser:2011:RMZ

[Hau11] Walter Hauser. The Ruhr Museum at the Zollverein Colliery in Essen, Germany. Technology and Culture, 52(1):171–179, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416413.

Hausman:2015:EEI

[Hau15] William J. Hausman. Energy and Electricity in Industrial Nations: The Sociology and Technology of Energy by Allan Mazur (review). Technology and Culture, 56(2):565–566, April 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/582341.

Hausse:2019:LSM

[Hau19] Heidi Hausse. The locksmith, the surgeon, and the mechanical hand: Communicating technical knowledge in Early Modern Europe. Technology and Culture, 60(1):34–64, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/719945.

Hayles:2012:THC

[Hay12] N. Katherine Hayles. The Techno-Human Condition (review). Technology and Culture, 53(4):920–921, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/490735.

Hayes:2014:RGW

[Hay14] Caroline Clarke Hayes. Recoding Gender: Women’s Changing Participation in Computing by Janet Abbate (review). Technology and Culture, 55(1):280–281, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/538938.

Hay:2018:CLP

[Hay18] Amy M. Hay. Chemical Lands: Pesticides, Aerial Spraying, and Health in North America’s Grasslands since 1945 by David D. Vail (review). Technology and Culture, 59(4):992–993, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/711786.

Headrick:2011:WAN

[Hea11] Daniel Headrick. Why America Is Not a New Rome (review). Technology and Culture, 52(4):814–816, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/454703.

Headrick:2012:TIC

[Hea12] Daniel Headrick. Telegraphic Imperialism: Crisis and Panic in the Indian Empire, c. 1830 (review). Technology and Culture, 53(2):488–490, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476824.

Headrick:2013:WCE

[Hea13a] Daniel R. Headrick. A World Connecting: 1870–1945 ed. by Emily S. Rosenberg (review). Technology and Culture, 54(4):985–986, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
530531.

Headrick:2013:QTK

[Hea13b] Daniel R. Headrick. The Quest for Technical Knowledge: Bengal in the Nineteenth Century by Suvobrata Sarkar (review). Technology and Culture, 54(2):412–413, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506090.

Headrick:2015:CIC

[Hea15] Daniel Headrick. Las Comunicaciones Internacionales de Cuba: Del Correo Marítimo al Satélite by José Altshuler (review). Technology and Culture, 56(4):975–976, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601479.

Hecht:2010:PNT

[Hec10] Gabrielle Hecht. The power of nuclear things. Technology and Culture, 51(1):1–30, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/375785.

Hecht:2014:IVL

[Hec14] Gabrielle Hecht. On the importance of the visual … and of mentoring. Technology and Culture, 55(4):964–969, October 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/562840.

Heide:2015:RBH

[Hei15a] Lars Heide. Reimagining Business History by Philip Scranton and Patrick Fridenson (review). Technology and Culture, 56(1):263–265, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572123.

Heikkero:2015:MTU

[Hei15b] Topi Heikkerö. Moralizing Technology: Understanding and Designing the Morality of Things by Peter-Paul Verbeek (review). Technology and Culture, 56(1):265–267, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/572124.

Heller:2013:CMF

[Hel13] Henry Heller. The Colonial Machine: French Science and Overseas Expansion in the Old Regime by James E. McClellan III and François Regourd (review). Technology and Culture, 54(2):400–401, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506083.

Henderson:2018:PCR

[Hen18a] Georgina Henderson. A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World by Alexander Jones (review). Technology and Culture, 59(3):789–790, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703149.

Henderson:2018:TTT

[Hen18b] Georgina J. Henderson. Turn, turn, turn: The construction of the architectural spiral fluted column in the ancient Mediterranean world. Technology and Culture, 59(2):363–409, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/697236.

Henderson:2019:GRM

[Hen19] Georgina J. Henderson. Gods and Robots: Myths, Machines, and Ancient Dreams of Technology by Adrienne Mayor (review). Technology and Culture, 60(4):1100–1101, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741406.

Hepler:2016:CWI

[Hep16] Allison L. Hepler. Child Workers and Industrial Health in Britain 1780–1850 by Peter Kirby (review). Technology and Culture, 57 (1):248–250, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611814.

Herzig:2011:AAS

[Her11] Rebecca Herzig. American Abyss: Savagery and Civilization in the Age of Industry (review). Technology and Culture, 52(1):202–204, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
416427.

Hersch:2016:BMF

[Her16a] Matthew Hersch. The Boeing Milestones of Flight Hall, National Air and Space Museum, Smithsonian Institution. Technology and Culture, 57(4):998–1004, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/637937.

Hersch:2016:SSN

[Her16b] Matthew Hersch. Spies and Shuttles: NASA’s Secret Relationships with the DoD and CIA by James E. David (review). Technology and Culture, 57(1):275–276, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/611828.

Hersch:2017:GSK

[Her17a] Matthew Hersch. Groovy Science: Knowledge, Innovation, and American Counterculture ed. by David Kaiser, W. Patrick McCray (review). Technology and Culture, 58(3):888–889, July 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/669486.

Hersh:2017:PPT

[Her17b] Eitan D. Hersh. Prototype Politics: Technology-Intensive Campaigning and the Data of Democracy by Daniel Kreiss (review). Technology and Culture, 58(2):608–609, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/663007.

Hessler:2011:DMG

[Hes11] Martina Hessler. Designing Modern Germany (review). Technology and Culture, 52(1):223–224, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416440.

Heymann:2010:JAS

[Hey10] Matthias Heymann. Joining art and science: The special exhibition split and splice: Fragments from the Age of Biomedicine at the Medical Museion, Copenhagen. Technology and Culture, 51 (4):995–1001, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403275.

Hahn:2012:RHJ

[HHH12] Ellen Hahn, Elisabeth Hahn, and Sophie Hahn. Roger Hahn, 5 January 1932–30 May 2011. Technology and Culture, 53(1):254–257, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466781.

Hicks:2012:MTE

[Hic12] Marie Hicks. Moving Targets: Elliott-Automation and the Dawn of the Computer Age in Britain, 1947–1967 (review). Technology and Culture, 53(3):738–740, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/481953.

Hicks:2014:TIG

[Hic14] Marie Hicks. Technologies of Inclusion: Gender in the Information Society ed. by Knut Holtan Sørensen, Wendy Faulkner, and Els Rommes (review). Technology and Culture, 55(2):509–510, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/546890.

Highsmith:2012:GMA

[Hig12] Andrew Highsmith. Gary, the Most American of All American Cities (review). Technology and Culture, 53(3):710–712, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/481921.

Hintz:2010:CRI

[Hin10] Eric S. Hintz. The Challenge of Remaining Innovative: Insights from Twentieth-Century American Business (review). Technology and Culture, 51(2):520–521, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/381280.

Hintz:2011:EEL

[Hin11] Eric Hintz. Edison’s Electric Light: The Art of Invention (review). Technology and Culture, 52(4):829–831, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/454549.

Hirose:2010:TCB

[Hir10] Shin Hirose. Two classes of British engineers: An analysis of their education and training, 1880s-1930s. Technology and Culture, 51 (2):388–402, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381288.

Hirsh:2011:HTR

[Hir11] Richard F. Hirsh. Historians of technology in the real world: Reflections on the pursuit of policy-oriented history. Technology and Culture, 52(1):6–20, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/416404.

Hirsh:2012:HLP

[Hir12] Richard Hirsh. How Local Politics Shape Federal Policy: Business, Power, and the Environment in Twentieth-Century Los Angeles (review). Technology and Culture, 53(4):940–942, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490747.

Hirsh:2014:PSS

[Hir14] Richard Hirsh. Power Struggles: Scientific Authority and the Creation of Practical Electricity before Edison by Michael Brian Schiffer (review). Technology and Culture, 55(2):499–501, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/546884.

Hirt:2016:EEJ

[Hir16] Paul Hirt. Energy and Ethics: Justice and the Global Energy Challenge by Benjamin K. Sovacool (review). Technology and Culture, 57(4):1029–1030, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/637952.

Hirsh:2017:WNH

[Hir17] Richard Hirsh. The Wired Northwest: The History of Electric Power, 1870s–1970s by Paul W. Hirt (review). Technology and Culture, 58(4):1090–1092, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679835.

Hjorth:2016:DDS

[Hjo16] Larissa Hjorth. Developer’s Dilemma: The Secret World of Videogame Creators by Casey O’Donnell (review). Technology and Culture, 57(2):504–505, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/619084.

Hlebinsky:2019:LGE

[Hle19] Ashley Hlebinsky. The Lives of Guns ed. by Jonathan Obert, Andrew Poe, and Austin Sarat (review). Technology and Culture, 60 (4):1119–1121, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741407.

Hochadel:2016:HWU

[HNG16] Oliver Hochadel and Agustí Nieto-Galan. How to write an urban history of STM on the “Periphery”. Technology and Culture, 57 (4):978–988, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637935.

Hoag:2015:TWT

[Hoa15] Heather J. Hoag. Transient Workspaces: Technologies of Everyday Innovation in Zimbabwe by Clapperton Chakanetsa Mavhunga (review). Technology and Culture, 56(4):1002–1003, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/601495.

Hobart:2014:IHT

[Hob14] Michael E. Hobart. The Information: A History, a Theory, a Flood by James Gleick (review). Technology and Culture, 55(2):489–490, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546878.

Hochfelder:2010:FFB

[Hoc10] David Hochfelder. Framing Finance: The Boundaries of Markets and Modern Capitalism (review). Technology and Culture, 51(4): 1023–1024, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403284.

Hogan:2013:SAP

[Hog13a] Andrew J. Hogan. Set adrift in the prenatal diagnostic marketplace: Analyzing the role of users and mediators in the history of a medical technology. Technology and Culture, 54(1):62–89, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/500797.

Hogselius:2013:WFB

[Hög13b] Per Högselius. Wheel of Fortune: The Battle for Oil and Power in Russia by Thane Gustafson (review). Technology and Culture, 54(3):701–703, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520313.

Hogan:2014:MFJ

[Hog14a] Andrew J. Hogan. Mobilizing the fetus: Julie Roberts, The Visualised Foetus; Malcolm Nicolson and John E. E. Fleming, Imaging and Imagining the Fetus. Technology and Culture, 55(3):730–733, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
551614.

Hogan:2014:PUA

[Hog14b] Andrew J. Hogan. Producers, users, and the actors between: Chikako Takeshita, The Global Biopolitics of the IUD; Heather Munro Prescott, The Morning After. Technology and Culture, 55 (1):241–244, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538919.

Hogselius:2014:OFC

[Hög14c] Per Högselius. Overpotential: Fuel Cells, Futurism, and the Making of a Power Panacea by Matthew N. Eisler (review). Technology and Culture, 55(3):765–767, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551633.

Hogselius:2015:CRA

[Hög15] Per Högselius. The Conquest of the Russian Arctic by Paul R. Josephson (review). Technology and Culture, 56(3):761–763, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593079.

Hogan:2018:RLK

[Hog18a] Andrew J. Hogan. Reordering Life: Knowledge and Control in the Genomics Revolution by Stephen Hilgartner (review). Technology and Culture, 59(2):501–503, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/697257.

Hogan:2018:GGP

[Hog18b] Andrew J. Hogan. The Genealogy of a Gene: Patients, HIV/AIDS, and Race by Myles W. Jackson (review). Technology and Culture, 59(4):997–998, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711789.

Hoke:2017:CFJ

[Hok17] Donald R. Hoke. C. Francis Jenkins: Pioneer of Film and Television by Donald G. Godfrey (review). Technology and Culture, 58(3):876–877, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669479.

Hoogerwerf:2010:CSH

[Hoo10] James J. Hoogerwerf. Conquistadors of the Sky: A History of Aviation in Latin America (review). Technology and Culture, 51 (1):272–273, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375802.

Hoogerwerf:2015:WCI

[Hoo15] James Hoogerwerf. The Wright Company: From Invention to Industry by Edward J. Roach (review). Technology and Culture, 56(3):765–766, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/593081.

Hooks:2019:SPE

[Hoo19] Tisha Hooks. Savage Preservation: The Ethnographic Origins of Modern Media Technology by Brian Hochman (review). Technology and Culture, 60(1):328–330, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/719961.

Horning:2011:ARR

[Hor11] Susan Schmid Horning. The architects of rock and roll: The Rock and Roll Hall of Fame and Museum, Cleveland. Technology and Culture, 52(3):598–605, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/447783.

Horn:2014:ITA

[Hor14] Jeff Horn. Industrialization and the Transformation of American Life: A Brief Introduction by Jonathan Rees (review). Technology and Culture, 55(3):746–747, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551622.

Horodowich:2017:MMP

[Hor17] Elizabeth Horodowich. The Mysteries of the Marco Polo Maps by Benjamin B. Olshin (review). Technology and Culture, 58(4): 1078–1079, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679828.

Horn:2018:IRS

[Hor18] Jeff Horn. The Industrial Revolution: The State, Knowledge and Global Trade by William J Ashworth (review). Technology and Culture, 59(4):967–968, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711771.

Hosler:2013:MME

[Hos13] Dorothy Hosler. Mercury, Mining, and Empire: The Human and Ecological Cost of Colonial Silver Mining in the Andes by Nicholas A. Robins (review). Technology and Culture, 54(4):973–975, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
530524.

Houltz:2010:CWF

[Hou10] Anders Houltz. The 1933 Chicago World’s Fair: A Century of Progress (review). Technology and Culture, 51(1):268–270, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375801.

Howard:2015:BR

[How15a] Nicole Howard. Books and reading. Technology and Culture, 56 (4):957–964, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601474.

Howell:2015:PLI

[How15b] Joel Howell. Paul Lauterbur and the Invention of MRI by M. Joan Dawson (review). Technology and Culture, 56(1):300–301, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/572144.

Hilaire-Perez:2013:SMS

[HPR13] Liliane Hilaire-Pérez and Christelle Rabier. Self-machinery?: Steel trusses and the management of ruptures in Eighteenth-Century Europe. Technology and Culture, 54(3):460–502, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/520279.

Hristova:2016:BCV

[Hri16] Stefka Hristova. Beyond the Checkpoint: Visual Practices in America’s Global War on Terror by Rebecca A. Adelman (review). Technology and Culture, 57(1):282–283, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611832.

Hirsh:2013:WTI

[HS13] Richard F. Hirsh and Benjamin K. Sovacool. Wind turbines and invisible technology: Unarticulated reasons for local opposition to wind energy. Technology and Culture, 54(4):705–734, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/530508.

Hsia:2019:I

[HS19] Florence Hsia and Dagmar Schäfer. Introduction. Technology and Culture, 60(2):554–561, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/726937.

Hu:2014:MSC

[Hu14] Danian Hu. Mr. Science and Chairman Mao’s Cultural Revolution: Science and Technology in Modern China ed. by Chunjuan Nancy Wei and Darryl E. Brock (review). Technology and Culture, 55 (1):270–273, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538933.

Hugill:2015:GCE

[Hug15] Peter J. Hugill. Global Communication Electric: Business, News, and Politics in the World of Telegraphy ed. by M. Michaela Hampf, Simone Müller–Pohl (review). Technology and Culture, 56(3):756– 757, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
593076.

Hugill:2016:UNN

[Hug16] Peter Hugill. The Undersea Network by Nicole Starosielski (review). Technology and Culture, 57(4):1020–1022, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637947.

Hui:2016:PMP

[Hui16] Alexandra Hui. The Pulse of Modernism: Physiological Aesthetics in Fin-de–Siècle Europe by Robert Michael Brain (review). Technology and Culture, 57(3):677–678, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/630940.

Hull:2016:PGR

[Hul16] James Hull. In the Power of the Government: The Rise and Fall of Newsprint in Ontario, 1894–1932 by Mark Kuhlberg (review). Technology and Culture, 57(2):476–478, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619068.

Hunt:2016:SDH

[Hun16] Bruce J. Hunt. Spark from the Deep: How Shocking Experiments with Strongly Electric Fish Powered Scientific Discovery by William J. Turkel (review). Technology and Culture, 57(1):247–248, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611813.

Hurt:2015:SBH

[Hur15] R. Douglas Hurt. Sod Busting: How Families Made Farms on the 19th-Century Plains by David Danbom (review). Technology and Culture, 56(4):982–983, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/601483.

Hussain:2010:HPA

[Hus10] Fahmida Hussain. Happy Pills in America: From Miltown to Prozac (review). Technology and Culture, 51(2):528–530, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381266.

Huskey:2019:PPS

[Hus19] Samuel J. Huskey. Passwords: Philology, Security, Authentication by Brian Lennon (review). Technology and Culture, 60(4): 1126–1127, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741405.

Henninger-Voss:2015:BSO

[HV15] Mary Henninger-Voss. Baroque Science by Ofer Gal and Raz Chen-Morris (review). Technology and Culture, 56(2):533–535, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582323.

Ihde:2011:BRE

[Ihd11] Don Ihde. Between Reason and Experience: Essays in Technology and Modernity (review). Technology and Culture, 52(2):434–436, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431395.

Ihde:2013:SHT

[Ihd13] Don Ihde. Stephen Hawking’s three bodies: Hélène Mialet, Hawking Incorporated. Technology and Culture, 54(3):650–653, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520286.

Ihde:2016:LAD

[Ihd16] Don Ihde. Laruelle: Against the Digital by Alexander R. Galloway (review). Technology and Culture, 57(3):708–709, July 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/630958.

Ihde:2018:TSL

[Ihd18] Don Ihde. Technosystem: The Social Life of Reason by Andrew Feenberg (review). Technology and Culture, 59(2):506–508, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/697260.

Irish:2012:FSD

[Iri12] Sharon Irish. Fallout Shelter: Designing for Civil Defense in the Cold War (review). Technology and Culture, 53(3):726–727, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/481930.

Jackson:2010:DR

[Jac10a] Donald C. Jackson. Dam (review). Technology and Culture, 51 (1):286–287, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375817.

Jackson:2010:TPC

[Jac10b] Shannon Jackson. The Technology of Policing: Crime Mapping, Information Technology, and the Rationality of Crime Control (review). Technology and Culture, 51(2):492–493, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/381269.

Jackson:2011:NDN

[Jac11a] Shannon Jackson. No Dig, No Fly, No Go: How Maps Restrict and Control (review). Technology and Culture, 52(2):432–434, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/431394.

Jackson:2011:TBL

[Jac11b] Shannon Jackson. Technology, bodily limits, and ways of knowing: Joy Parr sensing changes. Technology and Culture, 52(4):792–795, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454534.

Jackson:2012:EPR

[Jac12a] Donald C. Jackson. Enduring Pastoral: Recycling the Middle Landscape Ideal in the Tennessee Valley (review). Technology and Culture, 53(2):502–503, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/476825.

Jacobson:2012:CCR

[Jac12b] Brian R. Jacobson. Chromatic Cinema (review). Technology and Culture, 53(3):715–717, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481924.

Jacobson:2013:CFL

[Jac13] Brian R. Jacobson. The Cinematic Footprint: Lights, Camera, Natural Resources by Nadia Bozak (review). Technology and Culture, 54(2):424–426, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506097.

Jackson:2014:VAS

[Jac14] Shannon Jackson. The View from Above: The Science of Social Space by Jeanne Haffner (review). Technology and Culture, 55 (1):267–268, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538931.

Jacobson:2015:CAE

[Jac15] Brian R. Jacobson. Cinematic Appeals: The Experience of New Movie Technologies by Ariel Rogers (review). Technology and Culture, 56(2):558–560, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582337.

Jackson:2016:FCU

[Jac16] Shannon Jackson. The Fate of Cities: Urban America and the Federal Government, 1945–2000 by Roger Biles (review). Technology and Culture, 57(1):269–271, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/611825.

Jackson:2019:CCL

[Jac19] Shannon Jackson. Code + Clay … Data + Dirt: Five Thousand Years of Urban Media by Shannon Mattern (review). Technology and Culture, 60(2):647–648, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726956.

Jaffe:2010:LSF

[Jaf10] James Jaffe. Learning on the Shop Floor: Historical Perspectives on Apprenticeship (review). Technology and Culture, 51(2):495–497, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381268.

Jaffe:2011:EEE

[Jaf11] James Jaffe. The Enlightened Economy: An Economic History of Britain 1700–1850 (review). Technology and Culture, 52(4): 821–822, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454548.

Jasanoff:2012:AUW

[Jas12] Sheila Jasanoff. Acting in an Uncertain World: An Essay on Technical Democracy (review). Technology and Culture, 53(1):204–206, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466783.

Jenks:2011:WTW

[Jen11] Andrew Jenks. Would Trotsky Wear a Blue Tooth? Technological Utopianism under Socialism, 1917–1989 (review). Technology and Culture, 52(2):410–411, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/431381.

Jenks:2013:EOD

[Jen13] Andrew Jenks. The Earthquake Observers: Disaster Science from Lisbon to Richter by Deborah R. Coen (review). Technology and Culture, 54(4):982–983, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530529.

Jenks:2015:AMN

[Jen15] Andrew Jenks. Arming Mother Nature: The Birth of Catastrophic Environmentalism by Jacob Darwin Hamblin (review). Technology and Culture, 56(1):293–295, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/572140.

Jeon:2010:RMU

[Jeo10] Chihyung Jeon. A road to modernization and unification: The construction of the Gyeongbu Highway in South Korea. Technology and Culture, 51(1):55–79, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375783.

Jeon:2015:VFL

[Jeo15] Chihyung Jeon. The virtual flier: The link trainer, flight simulation, and pilot identity. Technology and Culture, 56(1):28–53, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572110.

Jepsen:2018:NBW

[Jep18] Thomas Jepsen. “A new business in the world”: The telegraph, privacy, and the U.S. Constitution in the Nineteenth Century. Technology and Culture, 59(1):95–125, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/692168.

Jones-Imhotep:2016:MMG

[JI16] Edward Jones-Imhotep. Malleability and machines: Glenn Gould and the technological self. Technology and Culture, 57(2):287–321, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
619053.

Jin:2017:AZJ

[Jin17] Dal Yong Jin. Atari to Zelda: Japan’s Videogames in Global Contexts by Mia Consalvo (review). Technology and Culture, 58(2):604– 606, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
663005.

Jin:2019:SPS

[Jin19] Dal Yong Jin. Super Power, Spoony Bards, and Silverware: The Super Nintendo Entertainment System by Dominic Arsenault (review). Technology and Culture, 60(2):652–654, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/726945.

Jorgensen:2016:AHT

[JJ16] Finn Arne Jørgensen and Dolly Jørgensen. The Anthropocene as a history of technology: Welcome to the Anthropocene: The Earth in Our Hands, Deutsches Museum, Munich. Technology and Culture, 57(1):231–237, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611808.

Johnston:2011:PPL

[Joh11a] Sean F. Johnston. Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein (review). Technology and Culture, 52(2):403–404, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/431377.

Johnston:2011:MIE

[Joh11b] Sean F. Johnston. Making the invisible engineer visible: DuPont and the recognition of nuclear expertise. Technology and Culture, 52(3):548–573, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447781.

Johnston:2012:NIN

[Joh12] Sean F. Johnston. Nuclear Illusion, Nuclear Reality: Britain, the United States, and Nuclear Weapons, 1958–1964 (review). Technology and Culture, 53(3):722–724, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/481928.

Johnston:2013:ICP

[Joh13] Sean Johnston. Instrumental Community: Probe Microscopy and the Path to Nanotechnology by Cyrus C. M. Mody (review). Technology and Culture, 54(1):221–223, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500823.

Johnston:2015:PCE

[Joh15a] Sean Johnston. Patently Contestable: Electrical Technologies and Inventor Identities on Trial in Britain by Stathis Arapostathis and Graeme Gooday (review). Technology and Culture, 56(1):276–277, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572130.

Johnstone:2015:TPI

[Joh15b] Adrian Johnstone. Turing: Pioneer of the Information Age by B. Jack Copeland (review). Technology and Culture, 56(3):772–773, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
593085.

John:2016:CCP

[Joh16] Richard R. John. Cables, Crises, and the Press: The Geopolitics of the New International Information System in the Americas, 1866–1903 by John A. Britton (review). Technology and Culture, 57(2):474–476, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/619067.

Johnson:2017:FUS

[Joh17] Christopher Johnson. Forestry in the U.S. South: A History by Mason C. Carter, Robert C. Kellison, and R. Scott Wallinger (review). Technology and Culture, 58(2):620–621, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/663014.

Johnson:2018:TCE

[Joh18a] Ian Johnson. Technology’s cutting edge — futurism and research in the Red Army, 1917–1937. Technology and Culture, 59(3):689–718, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703144.

Johnston:2018:AWP

[Joh18b] Sean F. Johnston. Alvin Weinberg and the promotion of the technological fix. Technology and Culture, 59(3):620–651, July 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/703142.

Jones:2010:DED

[Jon10] Christopher Jones. The Dirty Energy Dilemma: What’s Blocking Clean Power in the United States (review). Technology and Culture, 51(3):746–748, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390852.

Jones:2013:IMI

[Jon13] Claire L. Jones. Instruments of medical information: The rise of the medical trade catalog in Britain, 1750–1914. Technology and Culture, 54(3):563–599, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/520282.

Jones:2015:BEB

[Jon15] Allan Jones. Broadcasting Empire: The BBC and the British World, 1922–1970 by Simon J. Potter (review). Technology and Culture, 56(3):770–772, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/593084.

Jorgensen:2010:CWK

[Jør10] Finn Arne Jørgensen. Cold War Kitchen: Americanization, Technology, and European Users (review). Technology and Culture, 51(2):515–516, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381282.

Jorgensen:2013:FON

[Jør13] Dolly Jørgensen. Finding Oil: The Nature of Petroleum Geology, 1859–1920 by Brian Frehner (review). Technology and Culture, 54 (1):202–203, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500812.

Jorgensen:2017:GAE

[Jør17] Finn Arne Jørgensen. The Great Acceleration: An Environmental History of the Anthropocene since 1945 by J. R. McNeill and Peter Engelke (review). Technology and Culture, 58(2):623–624, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/663016.

Josephson:2011:PPP

[Jos11] Paul Josephson. Petrostate: Putin, Power, and the New Russia (review). Technology and Culture, 52(4):855–857, October 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/454566.

Josephson:2013:MSE

[Jos13] Paul Josephson. A Martian Stranded on Earth: Alexander Bogdanov, Blood Transfusions, and Proletarian Science by Nikolai Krementsov (review). Technology and Culture, 54(1):197–198, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/500809.

Josephson:2016:RMN

[Jos16a] Paul Josephson. Rivers, Memory and Nation-Building: A History of the Volga and Mississippi Rivers by Dorothy Zeisler-Vralsted (review). Technology and Culture, 57(3):682–684, July 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/630943.

Josephson:2016:VSS

[Jos16b] Paul Josephson. Voices of the Soviet Space Program: Cosmonauts, Soldiers, and Engineers Who Took the USSR into Space by Slava Gerovitch (review). Technology and Culture, 57(2):492–493, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/619077.

Josephson:2019:BCT

[Jos19] Paul Josephson. Bad Call: Technology’s Attack on Referees and Umpires and How to Fix It by Harry Collins, Robert Evans, and Christopher Higgins (review). Technology and Culture, 60(3):929– 931, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
731454.

Kaiserfeld:2010:WSA

[Kai10] Thomas Kaiserfeld. Walter Schottky, Atomtheoretiker und Elektrotechniker: Sein Leben und Werk bis ins Jahr 1941 (review). Technology and Culture, 51(3):754–756, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390857.

Kaijser:2011:TTN

[Kai11] Arne Kaijser. The trail from trail: New challenges for historians of technology. Technology and Culture, 52(1):131–142, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416410.

Kaijser:2012:IIN

[Kai12] Arne Kaijser. Iceland Imagined: Nature, Culture, and Storytelling in the North Atlantic (review). Technology and Culture, 53(3):702– 703, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481916.

Kaijser:2014:THI

[Kai14] Arne Kaijser. Tom Hughes — international scholar. Technology and Culture, 55(4):953–957, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562838.

Kaiserfeld:2017:RIU

[Kai17] Thomas Kaiserfeld. Revolutionizing Innovation: Users, Communities, and Open Innovation ed. by Dietmar Harhoff and Karim R. Lakhani (review). Technology and Culture, 58(2):613–614, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/663010.

Kalba:2014:CRR

[Kal14] Laura Kalba. The Color Revolution by Regina Blaszczyk (review). Technology and Culture, 55(1):261–263, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/538928.

Kang:2013:AMS

[Kan13] Minsoo Kang. Automata and Mimesis on the Stage of Theatre History by Kara Reilly (review). Technology and Culture, 54(4):

971–973, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530523.

Kang:2016:MRM

[Kan16] Minsoo Kang. Medieval Robots: Mechanism, Magic, Nature, and Art by E. R. Truitt (review). Technology and Culture, 57(4): 1007–1008, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637939.

Kapoor:2019:WSW

[Kap19] Nathan Kapoor. “Who has seen the wind”: Imagining wind power for the generation of electricity in Victorian Britain. Technology and Culture, 60(2):467–493, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726934.

Karvonen:2017:HSM

[Kar17] Andrew Karvonen. A House in the Sun: Modern Architecture and Solar Energy in the Cold War by Daniel A. Barber (review). Technology and Culture, 58(4):1094–1095, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679837.

Karl:2018:DTA

[Kar18] Brian Karl. Digital Tradition: Arrangement and Labor in Istanbul’s Recording Studio Culture by Eliot Bates (review). Technology and Culture, 59(4):1000–1002, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711791.

Kasson:2017:WSV

[Kas17] John F. Kasson. Work Sights: The Visual Culture of Industry in Nineteenth-Century America by Vanessa Meikle Schulman (review). Technology and Culture, 58(1):269–271, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648265.

Katzir:2010:WPR

[Kat10] Shaul Katzir. War and peacetime research on the road to crystal frequency control. Technology and Culture, 51(1):99–125, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375781.

Katz:2011:CWH

[Kat11] Barry M. Katz. Cold War on the Home Front: The Soft Power of Midcentury Design (review). Technology and Culture, 52(1): 221–223, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416439.

Katz:2014:UME

[Kat14] Mark Katz. Ubiquitous Musics: The Everyday Sounds That We Don’t Always Notice Edited by Marta García Quiñones, Anahid Kassabian, and Elena Boschi (review). Technology and Culture, 55(3):737–739, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551617.

Keller:2010:IFM

[Kel10] Alex Keller. Ingenium: Five Machines that Changed the World (review). Technology and Culture, 51(3):777–778, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390870.

Keller:2011:SCS

[Kel11] Charles Keller. Shop Class as Soulcraft: An Inquiry into the Value of Work (review). Technology and Culture, 52(1):209–210, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416431.

Kemp:2012:EMH

[Kem12] R. Scott Kemp. The end of Manhattan: How the gas centrifuge changed the quest for nuclear weapons. Technology and Culture, 53(2):272–305, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476806.

Kenney:2014:CMU

[Ken14] Martin Kenney. Creating a Market University: How Academic Science Became an Economic Engine by Elizabeth Popp Berman (review). Technology and Culture, 55(4):1032–1033, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/562865.

Kenny:2019:AIU

[Ken19] Nicolas Kenny. American Illuminations: Urban Lighting, 1800– 1920 by David E. Nye (review). Technology and Culture, 60(2):633– 635, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
726957.

Ketabgian:2016:IMC

[Ket16] Tamara Ketabgian. Imperial Media: Colonial Networks and Information Technologies in the British Literary Imagination, 1857– 1918 by Aaron Worth (review). Technology and Culture, 57(1): 250–252, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611815.

Kilgore:2015:SIA

[Kil15] De Witt Douglas Kilgore. Spacefarers: Images of Astronauts and Cosmonauts in the Heroic Era of Spaceflight ed. by Michael J. Neufeld (review). Technology and Culture, 56(1):288–289, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/572137.

Kim:2012:GLF

[Kim12] Tae-Ho Kim. Gourmets in the Land of Famine: The Culture and Politics of Rice in Modern Canton (review). Technology and Culture, 53(1):225–226, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/466795.

Kim:2013:ISD

[Kim13] Mi Gyung Kim. Invention as a social drama: From an ascending machine to the aerostatic globe. Technology and Culture, 54 (4):853–887, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530514.

Kim:2014:TGH

[Kim14] Nanny Kim. Technology, Gender and History in Imperial China: Great Transformations Reconsidered by Francesca Bray (review). Technology and Culture, 55(3):740–742, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/551619.

Kim:2015:FVT

[Kim15] Daniel Kim. Future visions: Technology and citizenship in underwater dreams. Technology and Culture, 56(2):530–532, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582322.

Kim:2019:ATA

[Kim19] Mi Gyung Kim. Atmospheric Things: On the Allure of Elemental Development by Derek P. McCormack (review). Technology and Culture, 60(1):332–333, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719963.

Kinchy:2010:SSN

[Kin10] Abby J. Kinchy. States of Suspense: The Nuclear Age, Postmodernism, and United States Fiction and Prose (review). Technology and Culture, 51(1):282–284, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375800.

Kinchy:2011:RRS

[Kin11] Abby Kinchy. Race, Rigor, and Selectivity in U.S. Engineering: The History of an Occupational Color Line (review). Technology and Culture, 52(2):401–402, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/431376.

Kinney:2014:BSM

[Kin14] Jeremy R. Kinney. Blue Sky Metropolis: The Aerospace Century in Southern California ed. by Peter J. Westwick (review). Technology and Culture, 55(4):1023–1024, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/562860.

Kirk:2012:STM

[Kir12] Robert G. W. Kirk. “Standardization through mechanization”: Germ-free life and the engineering of the ideal laboratory animal. Technology and Culture, 53(1):61–93, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466778/pdf.

Kita:2012:ATW

[Kit12] Chigusa Kita. Alone Together: Why We Expect More from Technology and Less from Each Other (review). Technology and Culture, 53(4):960–961, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490758.

Kaplan:2019:AEE

[KL19] Judith Kaplan and Rebecca Lemov. Archiving endangerment, endangered archives: Journeys through the sound archives of Americanist anthropology and linguistics, 1911–2016. Technology and Culture, 60(2S):S161–S187, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/727302.

Kleiman:2010:RAD

[Kle10] Jordan Kleiman. The rise of agribusiness and the demise of the New Deal order: Shane Hamilton, Trucking Country: The Road to America’s Wal-Mart Economy. Technology and Culture, 51 (1):216–219, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375776.

Klein:2014:DA

[Kle14] Ursula Klein. Depersonalizing the arcanum. Technology and Culture, 55(3):591–621, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/551607.

Klein:2016:LMC

[Kle16] Ursula Klein. The Limits of Matter: Chemistry, Mining & Enlightenment by Hjalmar Fors (review). Technology and Culture, 57(3):668–670, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630935.

Kline:2013:FS

[Kli13] Ronald Kline. Foundational stories. Technology and Culture, 54 (1):117–129, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500790.

Kline:2017:HM

[Kli17] Ronald R. Kline. Humans and machines. Technology and Culture, 58(3):835–845, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669470.

Kline:2019:UNH

[Kli19] Ron Kline. The Unreliable Nation: Hostile Nature and Technological Failure in the Cold War by Edward Jones-Imhotep (review). Technology and Culture, 60(3):921–922, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731456.

Karafantis:2014:FSI

[KLŁ14] Layne Karafantis, John Laprise, and Sławomir Łotysz. Fortieth symposium of the International Committee for the History of Technology: ICOHTEC at the ICHSTM, Manchester, UK, 21–28 July 2013. Technology and Culture, 55(1):211–219, January 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/538914.

Knight:2013:ERC

[Kni13] David Knight. Early Railway Chemistry and Its Legacy by Colin A. Russell and John A. Hudson (review). Technology and Culture, 54(4):978–980, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530527.

Knight:2018:EVH

[Kni18] H. Jackson Knight. Engineering Victory: How Technology Won the Civil War by Thomas F. Army (review). Technology and Culture, 59(4):970–972, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711773.

Knobel:2014:SRR

[Kno14a] Cory Knobel. Standards: Recipes for Reality by Lawrence Busch (review). Technology and Culture, 55(1):251–253, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538922.

Knowles:2014:LDH

[Kno14b] Scott Gabriel Knowles. Learning from disaster?: The history of technology and the future of disaster research. Technology and Culture, 55(4):773–784, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562831.

Knowles:2016:RNS

[Kno16] Scott Knowles. Risk: Negotiating Safety in American Society by Arwen Mohun (review). Technology and Culture, 57(1):245–246, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
611812.

Katzir:2016:FSS

[KNSZ16] Shaul Katzir, Christopher Neumaier, M. Luísa Sousa, and David Zimmerman. Forty-second Symposium of the International Committee for the History of Technology and Fourth Meeting of the History of Electrical Technology: “History of High-Technologies and Their Socio-Cultural Contexts,” Tel Aviv, Israel, 16–20 August 2015. Technology and Culture, 57(1):216–224, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611806.

Koloski-Ostrow:2012:WRA

[KO12] Ann Olga Koloski-Ostrow. The Waters of Rome: Aqueducts, Fountains, and the Birth of the Baroque City (review). Technology and Culture, 53(3):698–700, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481914.

Koloski-Ostrow:2013:RWU

[KO13] Ann Olga Koloski-Ostrow. Reflections on water, its use, and its meaning for society from antiquity to the early modern period. Technology and Culture, 54(1):175–179, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500800.

Koloski-Ostrow:2014:TWP

[KO14] Ann Olga Koloski-Ostrow. Thirst: Water and Power in the Ancient World by Steven Mithen (review). Technology and Culture, 55 (4):997–998, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562845.

Kochetkova:2016:SFT

[Koc16] Elena Kochetkova. Seeing the forest and the trees: Western forestry systems and Soviet engineers, 1955–1964. Technology and Culture, 57(3):586–611, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630930.

Koeppel:2018:DGW

[Koe18] Gerard Koeppel. Designing Gotham: West Point Engineers and the Rise of Modern New York, 1817–1898 by Jon Scott Logel (review). Technology and Culture, 59(3):796–797, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/703153.

Konvitz:2010:WWP

[Kon10] Josef Konvitz. What’s wrong with this picture? Technology and Culture, 51(2):474–476, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381272.

Konvitz:2011:RM

[Kon11] Josef Konvitz. Response to Misa. Technology and Culture, 52 (1):232, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416445. See [Mis11c].

Korhonen:2018:TIF

[Kor18] Janne M. Korhonen. Tolerating the intolerable: Flash smelting of copper and the construction of technological constraints. Technology and Culture, 59(2):338–362, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/697235.

Koukal:2010:ASM

[Kou10] D. R. Koukal. “All that is solid melts into air”: Frank Donoghue, the last professors: The corporate university and the fate of the humanities. Technology and Culture, 51(1):227–231, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375773.

Kranakis:2010:LMT

[Kra10] Eda Kranakis. Looking into the mirror of time: Reflections on the life and work of Edwin T. Layton Jr., 1928–2009. Technology and Culture, 51(2):543–560, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/381254.

Kranakis:2011:IIG

[Kra11] Eda Kranakis. “And it is, it is a glorious thing /to be a pirate king!”: Adrian Johns, Piracy[: The Intellectual Property Wars from Gutenberg to Gates]. Technology and Culture, 52(4):796–800, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454533.

Kragh:2014:SSH

[Kra14] Helge Kragh. A Single Sky: How an International Community Forged the Science of Radio Astronomy by David P. D. Munns (review). Technology and Culture, 55(2):506–507, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546888.

Krapp:2017:DSI

[Kra17] Peter Krapp. Digital Signatures: The Impact of Digitization on Popular Music Sound by Ragnhild Brøvig–Hanssen and Anne Danielsen (review). Technology and Culture, 58(4):1110–1111, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
679846.

Krapp:2019:FST

[Kra19] Peter Krapp. Future Sounds: The Temporality of Noise by Stephen Kennedy (review). Technology and Culture, 60(3):926–927, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/731458.

Krebs:2010:VMF

[Kre10] Stefan Krebs. Das Versprechen mobiler Freiheit: Zur Kulturund Technikgeschichte von Kofferradio, Walkman, und Handy (review). Technology and Culture, 51(3):765–766, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390863.

Krebs:2014:MDS

[Kre14a] Stefan Krebs. Das Motorrad: Ein deutscher Sonderweg in die automobile Gesellschaft by Frank Steinbeck (review). Technology and Culture, 55(1):273–274, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/538934.

Krebs:2014:DGV

[Kre14b] Stefan Krebs. “Dial gauge versus senses 1–0”: German car mechanics and the introduction of new diagnostic equipment, 1950– 1980. Technology and Culture, 55(2):354–389, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546869.

Kreilkamp:2014:MTL

[Kre14c] Ivan Kreilkamp. Media, Technology and Literature in the Nineteenth Century: Image, Sound, Touch ed. by Colette Colligan and Margaret Linley (review). Technology and Culture, 55(4): 1002–1003, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562848.

Krige:2019:DEH

[Kri19a] John Krige. Democracy in Exile: Hans Speier and the Rise of the Defense Intellectual by Daniel Bessner (review). Technology and Culture, 60(3):919–920, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/731449.

Krige:2019:RIK

[Kri19b] John Krige. Regulating international knowledge exchange: The national security state and the American research university from the 1950s to today. Technology and Culture, 60(1):252–277, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/719952.

Krupar:2010:ASN

[Kru10] Jason Krupar. Atmospheric Science at NASA: A History (review). Technology and Culture, 51(3):751–752, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390855.

Krupar:2012:AFD

[Kru12] Jason Krupar. Atomic Frontier Days: Hanford and the American West (review). Technology and Culture, 53(3):724–725, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/481929.

Krupar:2013:FCN

[Kru13a] Jason Krupar. Fuel Cycle to Nowhere: U.S. Law and Policy on Nuclear Waste by Richard Burleson Stewart and Jane Bloom Stewart (review). Technology and Culture, 54(1):216–218, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/500820.

Krupar:2013:OOW

[Kru13b] Jason Krupar. Our Own Worst Enemy? Institutional Interests and the Proliferation of Nuclear Weapons Expertise by Sharon K. Weiner (review). Technology and Culture, 54(3):694–695, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520309.

Krupar:2014:PPP

[Kru14] Jason Krupar. Project Plowshare: The Peaceful Use of Nuclear Explosives in Cold War America by Scott Kaufman (review). Technology and Culture, 55(3):763–765, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/551632.

Krupar:2015:ROL

[Kru15] Jason Krupar. Robert Oppenheimer: A Life Inside the Center by Ray Monk (review). Technology and Culture, 56(3):774–775, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593086.

Krupar:2016:PIP

[Kru16] Jason Krupar. The Politics of Invisibility: Public Knowledge about Radiation Health Effects after Chernobyl by Olga Kuchinskaya (review). Technology and Culture, 57(2):490–492, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619076.

Kuijpers:2019:MSH

[Kui19] Maikel H. G. Kuijpers. Materials and skills in the history of knowledge: An archaeological perspective from the “Non-Asian” field. Technology and Culture, 60(2):604–615, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/726942.

Kumar:2012:NNS

[Kum12] Deepak Kumar. Nucleus and Nation: Scientists, International Networks, and Power in India (review). Technology and Culture, 53(3):720–721, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481927.

Kumar:2016:OSI

[Kum16] Deepak Kumar. The Outsourcer: The Story of India’s IT Revolution by Dinesh C. Sharma (review). Technology and Culture, 57(3):704–706, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630956.

Kumar:2019:BMW

[Kum19a] Prakash Kumar. “A big machine not working properly”: Elite narratives of India’s community projects, 1952–58. Technology and Culture, 60(4):1027–1058, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/741382.

Kumar:2019:I

[Kum19b] Prakash Kumar. Introduction. Technology and Culture, 60(4): 933–952, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741378.

Kurlinkus:2019:TCL

[Kur19a] Will Kurlinkus. Track Changes: A Literary History of Word Processing by Matthew G. Kirschenbaum (review). Technology and Culture, 60(1):340–342, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719968.

Kursell:2019:LMT

[Kur19b] Julia Kursell. Listening to more than sounds: Carl Stumpf and the experimental recordings of the Berliner Phonogramm-Archiv. Technology and Culture, 60(2S):S39–S63, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/727298.

Lacey:2016:SMH

[Lac16] Kate Lacey. Sounds of Modern History: Auditory Cultures in 19thand 20th-Century Europe ed. by Daniel Morat (review). Technology and Culture, 57(3):675–677, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/630939.

Ladd:2014:PCG

[Lad14] Brian Ladd. The People’s Car: A Global History of the Volkswagen Beetle by Bernhard Rieger (review). Technology and Culture, 55 (1):275–276, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538935.

Laemmli:2015:CPR

[Lae15] Whitney E. Laemmli. A case in pointe: Romance and regimentation at the New York City Ballet. Technology and Culture, 56(1):1–27, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572109.

Laemmli:2016:EMR

[Lae16] Whitney Laemmli. Embodiment and Mechanisation: Reciprocal Understandings of Body and Machine from the Renaissance to the Present by Daniel Black (review). Technology and Culture, 57(4): 1008–1010, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637940.

Lagendijk:2011:PMG

[Lag11] Vincent Lagendijk. Prime Movers of Globalization: The History and Impact of Diesel Engines and Gas Turbines (review). Technology and Culture, 52(4):852–853, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/454565.

Lagendijk:2013:RGR

[Lag13] Vincent Lagendijk. Red Gas: Russia and the Origins of European Energy Dependence by Per Högselius (review). Technology and Culture, 54(4):992–993, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530536.

Lagendijk:2015:TEP

[Lag15] Vincent Lagendijk. La traversée électrique des Pyrénées: Histoire de l’interconnexion entre la France et l’Espagne (Histoire de l’énergie no. 5) by Renan Viguié (review). Technology and Culture, 56(3):769–770, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/593083.

Lagendijk:2017:CRL

[Lag17] Vincent Lagendijk. Concrete Revolution: Large Dams, Cold War Geopolitics, and the U.S. Bureau of Reclamation by Christopher Sneddon (review). Technology and Culture, 58(4):1096–1097, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
679838.

Laird:2012:PBA

[Lai12] Pamela Walker Laird. Peddling Bicycles to America: The Rise of an Industry (review). Technology and Culture, 53(2):505–507, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476826.

Lakwete:2013:TIS

[Lak13] Angela Lakwete. Transition to an Industrial South: Athens, Georgia, 1830–1870 by Michael J. Gagnon (review). Technology and Culture, 54(3):663–664, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/520292.

Langins:2010:VAM

[Lan10] Janis Langins. Vauban, architecte de la modernité? (review). Technology and Culture, 51(1):240–242, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/375799.

Lankton:2011:IOT

[Lan11] Larry Lankton. Iron Ore Transport on the Great Lakes: The Development of a Delivery System to Feed American Industry (review). Technology and Culture, 52(3):640–641, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447799.

Lanthier:2013:RRT

[Lan13] Pierre Lanthier. Rivers run through it. Technology and Culture, 54 (1):180–183, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500801.

Lankton:2015:SUM

[Lan15] Larry Lankton. Seeing Underground: Maps, Models, and Mining Engineering in America by Eric C. Nystrom (review). Technology and Culture, 56(4):986–988, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601486.

Lane:2017:SLR

[Lan17] Maria Lane. Seeing like a Rover: How Robots, Teams, and Images Craft Knowledge of Mars by Janet Vertesi (review). Technology and Culture, 58(1):298–299, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/648282.

Laney:2018:SSE

[Lan18a] Monique Laney. Spaceflight in the Shuttle Era and Beyond: Redefining Humanity’s Purpose in Space by Valerie Neal (review). Technology and Culture, 59(3):813–815, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/703163.

Lanzarotta:2018:NSB

[Lan18b] Tess Lanzarotta. Network Sovereignty: Building the Internet across Indian Country by Marisa Elena Duarte (review). Technology and Culture, 59(3):808–809, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/703160.

Laprise:2012:SSR

[Lap12] John Laprise. Surveillance or Security? The Risks Posed by New Wiretapping Technologies (review). Technology and Culture, 53 (3):742–743, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481955.

Lassman:2018:LSA

[Las18] Thomas C. Lassman. The Long Space Age: The Economic Origins of Space Exploration from Colonial America to the Cold War by Alexander MacDonald (review). Technology and Culture, 59 (2):484–485, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/697247.

Lauer:2011:MRA

[Lau11] Josh Lauer. The Market Revolution in America: Liberty, Ambition, and the Eclipse of the Common Good (review). Technology and Culture, 52(1):197–198, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/416424.

Launius:2013:IOS

[Lau13] Roger D. Launius. Imagining Outer Space: European Astroculture in the Twentieth Century ed. by Alexander C. T. Geppert (review). Technology and Culture, 54(3):689–690, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/520306.

Launius:2017:FHR

[Lau17] Roger D. Launius. From Fishing Hamlet to Red Planet: India’s Space Journey ed. by P. V. Manoranjan Rao et al. (review). Technology and Culture, 58(1):278–279, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/648270.

LeCain:2011:HGC

[LeC11] Timothy LeCain. Hollowed Ground: Copper Mining and Community Building on Lake Superior, 1840s–1990s (review). Technology and Culture, 52(3):638–640, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/447798.

LeCain:2017:ALE

[LeC17] Timothy James LeCain. The Anti-Landscape ed. by David E. Nye and Sarah Elkind (review). Technology and Culture, 58(4): 1101–1103, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679841.

Lee:2013:ISK

[Lee13] Jung Lee. Invention without science: “Korean edisons” and the changing understanding of technology in colonial Korea. Technology and Culture, 54(4):782–814, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/530511.

Lee:2014:ADM

[Lee14] Kah-Wee Lee. Addiction by Design: Machine Gambling in Las Vegas by Natasha Dow Schüll (review). Technology and Culture, 55(1):278–280, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538937.

Leggett:2011:SWC

[Leg11] Don Leggett. Spectacle and witnessing: Constructing readings of Charles Parsons’s marine turbine. Technology and Culture, 52 (2):287–309, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431362.

Lerman:2010:CDC

[Ler10] Nina E. Lerman. Categories of difference, categories of power: Bringing gender and race to the history of technology. Technology and Culture, 51(4):893–918, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403271.

Leslie:2010:SCW

[Les10] Thomas Leslie. The Skyscraper and the City: The Woolworth Building and the Making of Modern New York (review). Technology and Culture, 51(2):506–507, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/381294.

Lesanu:2019:RGS

[Les19] Alexandru Lesanu. Removing a German sugar factory to the Soviet borderlands: A case study of the postwar technological transfer in Soviet Moldavia (1946–52). Technology and Culture, 60(3):866– 885, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
731446.

Leung:2014:ICM

[Leu14] Angela Ki Che Leung. Imperial Contagions: Medicine, Hygiene, and Cultures of Planning in Asia ed. by Robert Peckham and David M. Pomfret (review). Technology and Culture, 55(4):1016–1017, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562856.

Levere:2011:OCO

[Lev11] Trevor Levere. Aux origins de la chimie organique: Méthodes et pratiques des pharmaciens et des chimistes (1785–1835) (review). Technology and Culture, 52(3):629–631, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447793.

Levitt:2013:LET

[Lev13a] Theresa Levitt. Liberty, equality, technology: Virtuous inventors and base profiteers in the French industrial revolution and beyond. Technology and Culture, 54(4):815–819, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530512.

Levitt:2013:STB

[Lev13b] Theresa Levitt. Science and technology beyond the barricades. Technology and Culture, 54(2):377–381, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/506076.

Levitt:2017:SMM

[Lev17] Theresa Levitt. Les savoirs-mondes: Mobilités et circulation des savoirs depuis le Moyen Âge ed. by Pilar González Bernaldo, Liliane Hilaire–Peréz (review). Technology and Culture, 58(3):870– 872, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669476.

Li:2016:AER

[Li16] Jie Jack Li. The Antibiotic Era: Reform, Resistance, and the Pursuit of a Rational Therapeutics by Scott H. Podolsky (review). Technology and Culture, 57(3):698–699, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/630952.

Lieberman:2018:HPB

[Lie18] Jennifer L. Lieberman. Human Programming: Brainwashing, Automatons, and American Unfreedom by Scott Selisker (review). Technology and Culture, 59(3):810–811, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/703161.

Lifset:2018:SPE

[Lif18] Robert Lifset. Selling Power: Economics, Policy, and Electric Utilities Before 1940 by John L. Neufeld (review). Technology and Culture, 59(4):976–978, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711777.

Light:2011:DAC

[Lig11] Jennifer Light. Discriminating appraisals: Cartography, computation, and access to Federal mortgage insurance in the 1930s. Technology and Culture, 52(3):485–522, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/447779.

Light:2014:CDM

[Lig14] Jennifer S. Light. Close Up at a Distance: Mapping, Technology and Politics by Laura Kurgan (review). Technology and Culture, 55(3):769–770, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551635.

Lim:2019:PFE

[Lim19] Samson Lim. Photography and forgery in early capitalist Siam. Technology and Culture, 60(3):795–815, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731443.

Lindsay:2010:WUM

[Lin10] Jon R. Lindsay. “War upon the map”: User innovation in American military software. Technology and Culture, 51(3):619–651, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/390839.

Lindqvist:2011:HEA

[Lin11] Svante Lindqvist. How to establish an academic discipline: A Swedish experience. Technology and Culture, 52(2):233–245, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/431359.

Lindner:2013:SIN

[Lin13] Stephan H. Lindner. Der Schuh im Nationalsozialismus: Eine Produktgeschichte im deutsch-britisch-amerikanischen Vergleich by Anne Sudrow (review). Technology and Culture, 54(4):992–993, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
530535.

Lamont:2015:AAR

[LL15] Mark Lamont and Rebekah Lee. Arrive alive: Road safety in Kenya and South Africa. Technology and Culture, 56(2):464–488, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582317.

Li:2019:VFV

[LM19] Xiaochang Li and Mara Mills. Vocal features: From voice identification to speech recognition by machine. Technology and Culture, 60(2S):S129–S160, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/727301.

Long:2010:CPE

[Lon10] Pamela O. Long. The craft of premodern European history of technology: Past and future practice. Technology and Culture, 51(3):698–714, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390843.

Long:2011:SSM

[Lon11] Pamela O. Long. Ships and Shipping in Medieval Manuscripts (review). Technology and Culture, 52(3):618–619, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447787.

Long:2012:RWP

[Lon12a] Pamela O. Long. Rome’s World: The Peutinger Map Reconsidered (review). Technology and Culture, 53(2):479–480, April 2012. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/476827.

Longo:2012:KMR

[Lon12b] Bernadette Longo. From A to <A>: Keywords of Markup (review). Technology and Culture, 53(1):250–251, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466809.

Long:2015:CP

[Lon15] Pamela O. Long. A crooked path. Technology and Culture, 56 (2):498–507, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582320.

Lotysz:2015:EER

[Łot15] Sławomir Łotysz. Eastern European Railways in Transition: Nineteenth to Twenty-First Centuries ed. by Ralf Roth and Henry Jacolin (review). Technology and Culture, 56(2):549–551, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582332.

Lotysz:2016:UMI

[Łot16] Sławomir Łotysz. Unified Military Industries of the Soviet Bloc: Hungary and the Division of Labor in Military Production by Pál Germuska (review). Technology and Culture, 57(4):1017–1018, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
637945.

Loviglio:2011:EFR

[Lov11] Jason Loviglio. Early FM Radio: Incremental Technology in Twentieth-Century America (review). Technology and Culture, 52(3):652–654, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447805.

Lowengard:2010:LAP

[Low10] Sarah Lowengard. The Late Age of Print: Everyday Book Culture from Consumerism to Control (review). Technology and Culture, 51 (4):1046–1048, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403298.

Lowengard:2013:CDL

[Low13] Sarah Lowengard. California Design, 1930–1965: Living in a Modern Way ed. by Wendy Kaplan (review). Technology and Culture, 54(2):421–422, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506095.

Lowengard:2014:AAA

[Low14] Sarah Lowengard. Analyzing Art and Aesthetics Edited by Anne Collins Goodyear and Margaret Weitekamp (review). Technology and Culture, 55(3):739–740, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551618.

Lowen:2016:RUG

[Low16a] Rebecca Lowen. The Recombinant University: Genetic Engineering and the Emergence of Stanford Biotechnology by Doogab Yi (review). Technology and Culture, 57(4):1037–1039, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/637957.

Lowood:2016:RCA

[Low16b] Henry Lowood. Re-Collection: Art, New Media, and Social Memory by Richard Rinehart and Jon Ippolito (review). Technology and Culture, 57(2):501–502, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/619082.

Lowengard:2018:JVI

[Low18] Sarah Lowengard. John Vassos: Industrial Design for Modern Life by Danielle Shapiro (review). Technology and Culture, 59(2):490– 492, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
697251.

Lubar:2018:MHH

[Lub18] Steven Lubar. The Matter of History: How Things Create the Past by Timothy J. LeCain (review). Technology and Culture, 59 (4):963–964, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711769.

Luckin:2010:APH

[Luc10] Bill Luckin. Accident Prone: A History of Technology, Psychology, and Misfits of the Machine Age (review). Technology and Culture, 51(4):1024–1026, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/403285.

Lucsko:2011:SHV

[Luc11] David N. Lucsko. The Speed Handbook: Velocity, Pleasure, Modernism (review). Technology and Culture, 52(1):213–215, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416434.

Lucsko:2012:CCI

[Luc12] David Lucsko. Colin Chapman: Inside the Innovator (review). Technology and Culture, 53(1):216–218, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466790.

Luckin:2013:DEM

[Luc13] Bill Luckin. The Disaster Experts: Mastering Risk in Modern America by Scott Gabriel Knowles (review). Technology and Culture, 54(1):204–205, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/500813.

Lucsko:2014:CCA

[Luc14] David N. Lucsko. Of clunkers and camaros: Accelerated vehicle retirement programs and the automobile enthusiast, 1990– 2009. Technology and Culture, 55(2):390–428, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546870.

Luckin:2018:ANB

[Luc18a] Bill Luckin. The Age of Noise in Britain: Hearing Modernity by James G. Mansell (review). Technology and Culture, 59(4):980–981, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
711779.

Lucsko:2018:MYA

[Luc18b] David Lucsko. Machines of Youth: America’s Car Obsession by Gary S. Cross (review). Technology and Culture, 59(4):983–985, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
711781.

Maat:2016:DEP

[Maa16] Harro Maat. Development and Environmental Politics Unmasked: Authority, Participation and Equity in East Timor by Christopher Shephard (review). Technology and Culture, 57(4):1032–1034, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
637954.

Macrakis:2013:HSC

[Mac13] Kristie Macrakis. Hide and Seek: Camouflage, Photography, and the Media of Reconnaissance by Hanna Rose Shell (review). Technology and Culture, 54(3):696–697, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520310.

Macrakis:2014:SIS

[Mac14] Kristie Macrakis. Supervision: An Introduction to the Surveillance Society by John Gilliom and Torin Monahan (review). Technology and Culture, 55(2):515–516, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/546893.

MacKenzie:2015:MMC

[Mac15] Donald MacKenzie. Mechanizing the merc: The Chicago mercantile exchange and the rise of high-frequency trading. Technology and Culture, 56(3):646–675, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/593067.

Macedo:2016:SCT

[Mac16a] Marta Macedo. Standard cocoa: Transnational networks and technoscientific regimes in West African plantations. Technology and Culture, 57(3):557–585, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/630929.

Macfarlane:2016:EWE

[Mac16b] Daniel Macfarlane. Empire of Water: An Environmental and Political History of the New York City Water Supply by David Soll (review). Technology and Culture, 57(1):266–267, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611823.

Macedo:2017:OMG

[Mac17] Marta Macedo. Objects in Motion: Globalizing Technology ed. by Nina Möllers, Bryan Dewalt (review). Technology and Culture, 58(3):902–903, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669494.

Macola:2018:WCS

[Mac18] Giacomo Macola. Weapon of Choice: Small Arms and the Culture of Military Innovation by Matthew Ford (review). Technology and Culture, 59(2):487–489, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/697249.

MacDougall:2019:SPK

[Mac19] Robert MacDougall. Sympathetic physics: The Keely motor and the laws of thermodynamics in Nineteenth-Century culture. Technology and Culture, 60(2):438–466, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/726933.

Madureira:2019:RPG

[Mad19] Nuno Luís Madureira. Reckless proliferation and guardianship proliferation: The fast breeder nuclear reactor and the plutonium economy. Technology and Culture, 60(3):833–865, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731445.

Maerker:2013:ATD

[Mae13] Anna Maerker. Anatomizing the trade: Designing and marketing anatomical models as medical technologies, ca. 1700–1900. Technology and Culture, 54(3):531–562, July 2013. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520281.

Maines:2010:DIG

[Mai10] Rachel Maines. Defending the Indefensible: The Global Asbestos Industry and Its Fight for Survival (review). Technology and Culture, 51(3):748–750, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390853.

Maines:2012:TTP

[Mai12] Rachel Maines. Toys and Tools in Pink: Cultural Narratives of Gender, Science and Technology (review). Technology and Culture, 53(2):513–514, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476828.

Mains:2016:ULW

[Mai16] Daniel Mains. Urban Life-Worlds in Motion: African Perspectives ed. by Hans Peter Hahn, Kristin Kastner (review). Technology and Culture, 57(4):1030–1032, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/637953.

Malo:2017:DOC

[Mal17] Gabriel Malo. Designing Online Communities: How Designers, Developers, Community Managers, and Software Structure Discourse and Knowledge Production on the Web by Trevor Owens (review). Technology and Culture, 58(1):302–303, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648284.

Manuel:2013:MTE

[Man13] Jeffrey T. Manuel. Mr. Taconite: Edward W. Davis and the promotion of low-grade iron ore, 1913–1955. Technology and Culture, 54(2):317–345, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506072.

Manuel:2015:MMC

[Man15] Jeffrey T. Manuel. Mercury and the Making of California: Mining, Landscape, and Race, 1840–1890 by Andrew Scott Johnston (review). Technology and Culture, 56(3):758–759, July 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/593077.

Manganiello:2018:SWP

[Man18] Christopher J. Manganiello. The Slain Wood: Papermaking and Its Environmental Consequences in the American South by William Boyd (review). Technology and Culture, 59(1):171–173, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/692174.

Manuel:2019:PNP

[Man19] Jeffrey T. Manuel. The Price of Nuclear Power: Uranium Communities and Environmental Justice by Stephanie A. Malin (review). Technology and Culture, 60(2):644–645, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726955.

Marx:2010:TEH

[Mar10] Leo Marx. Technology: The emergence of a hazardous concept. Technology and Culture, 51(3):561–577, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390836.

Marks:2011:HFG

[Mar11a] Robert Marks. The Human Footprint: A Global Environmental History (review). Technology and Culture, 52(1):183–185, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416416.

Marrs:2011:FFU

[Mar11b] Aaron W. Marrs. The Fragile Fabric of Union: Cotton, Federal Politics, and the Global Origins of the Civil War (review). Technology and Culture, 52(1):201–202, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416426.

Marsh:2013:RFY

[Mar13] Allison C. Marsh. Revolution: The first 2,000 years of computing: The Computer History Museum, Mountain View, California. Technology and Culture, 54(3):640–649, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520285.

Mark:2016:E

[Mar16] Christopher Mark. To the Editor. Technology and Culture, 57 (3):714–717, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630961.

Marsh:2017:PIE

[Mar17a] Allison Marsh. Places of Invention ed. by Arthur P. Molella and Anna Karvellas (review). Technology and Culture, 58(2):617–618, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
663012.

Marsh:2017:PIR

[Mar17b] Allison Marsh. Places of Invention (review). Technology and Culture, 58(3):856–861, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/669472.

Margolis:2018:PMC

[Mar18] Emily A. Margolis. Photographs, Museums, Collections: Between Art and Information by Elizabeth Edwards, Christopher Morton (review). Technology and Culture, 59(1):170–171, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/692173.

Marklund:2019:TWM

[Mar19] Andreas Marklund. Trawling the wires: Mass surveillance of border-crossing communication in Denmark during World War II. Technology and Culture, 60(3):770–794, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731442.

Mas:2017:SWP

[Mas17] Catherine Mas. She wears the pants: The reform dress as technology in Nineteenth-Century America. Technology and Culture, 58(1):35–66, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648249.

Mateos:2017:MMG

[Mat17a] Gisela Mateos. The Mayan in the Mall: Globalization, Development, and the Making of Modern Guatemala by J. T. Way (review). Technology and Culture, 58(1):283–284, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648273.

Matsumoto:2017:ECP

[Mat17b] Mallory E. Matsumoto. The Evolution of Ceramic Production Organization in a Maya Community by Dean E. Arnold (review). Technology and Culture, 58(1):281–283, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/648272.

Matthews:2017:MIN

[Mat17c] Glenna Matthews. Make It New: A History of Silicon Valley Design by Barry M. Katz (review). Technology and Culture, 58(4): 1108–1110, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679845.

Matala:2019:FFC

[Mat19] Saara Matala. Flashy flagships of Cold War cooperation: The Finnish–Soviet Nuclear Icebreaker Project. Technology and Culture, 60(2):347–377, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/726930.

Mauskopf:2011:IEM

[Mau11] Seymour Mauskopf. Igniting early modern science through pyrotechnics. Technology and Culture, 52(3):606–609, July 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/447784.

Mavhunga:2010:GRP

[Mav10] Clapperton Mavhunga. Guns, Race, and Power in Colonial South Africa (review). Technology and Culture, 51(1):237–238, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375798.

Mamidipudi:2018:IIH

[MB18] Annapurna Mamidipudi and Wiebe E. Bijker. Innovation in Indian handloom weaving. Technology and Culture, 59(3):509–545, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703139.

McAleer:2014:PVD

[McA14] John McAleer. Panoramic visions: Denise Blake oleksijczuk, the first panoramas; erkki huhtamo, illusions in motion. Technology and Culture, 55(1):237–240, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/538918.

McCrossen:2011:TSH

[McC11] Alexis McCrossen. A Tenth of a Second: A History (review). Technology and Culture, 52(1):212–213, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/416433.

McClellan:2012:HD

[McC12] James E. McClellan III. Harold Dorn, 1928–2011. Technology and Culture, 53(3):671–677, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481909.

McCray:2013:DFD

[McC13] W. Patrick McCray. The Digital Flood: The Diffusion of Information Technology across the U.S., Europe, and Asia by James W. Cortada (review). Technology and Culture, 54(2):432–433, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/506101.

McComb:2014:IWA

[McC14a] Erinn McComb. Integrating Women into the Astronaut Corps: Politics and Logistics at NASA, 1972–2004 by Amy E. Foster (review). Technology and Culture, 55(2):507–509, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546889.

McCray:2014:HAD

[McC14b] W. Patrick McCray. How astronomers digitized the sky. Technology and Culture, 55(4):908–944, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/562836.

McComb:2015:TTW

[McC15] Erinn McComb. Texas Takes Wing: A Century of Flight in the Lone Star State by Barbara Ganson (review). Technology and Culture, 56(2):553–554, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582334.

McComb:2017:WPT

[McC17] Erinn McComb. Weekend Pilots: Technology, Masculinity, and Private Aviation in Postwar America by Alan Meyer (review). Technology and Culture, 58(1):288–290, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/648276.

McDonald:2010:CLC

[McD10] Chris McDonald. The Cultural Logic of Computation (review). Technology and Culture, 51(4):1056–1058, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/403304.

McDermott:2012:E

[McD12] John F. M. McDermott. To the Editor. Technology and Culture, 53(2):525–529, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/ 1/article/476847. See response [Cow12].

McDonald:2015:AIL

[McD15] Kate McDonald. Asymmetrical integration: Lessons from a railway empire. Technology and Culture, 56(1):115–149, January 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/572113.

McDonald:2017:AUI

[McD17] Kate McDonald. Airport Urbanism: Infrastructure and Mobility in Asia by Max Hirsh (review). Technology and Culture, 58(3):886– 888, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669485.

Mckellar:2013:BHT

[Mck13a] Shelley Mckellar. Broken Hearts: The Tangled History of Cardiac Care by David S. Jones (review). Technology and Culture, 54(4): 1003–1005, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530541.

McKellar:2013:NRF

[McK13b] Shelley McKellar. Negotiating risk: The failed development of atomic hearts in America, 1967–1977. Technology and Culture, 54(1):1–39, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/500795.

McLaren:2017:GTA

[McL17] Brian L. McLaren. A Genealogy of Tropical Architecture: Colonial Networks, Nature and Technoscience by Jiat-Hwee Chang (review). Technology and Culture, 58(4):1079–1081, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679829.

McNally:2014:LET

[McN14] Lou McNally. Lake Effect: Tales of Large Lakes, Arctic Winds, and Recurrent Snows by Mark Monmonier (review). Technology and Culture, 55(4):1014–1015, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/562855.

Medina:2018:FIA

[Med18] Eden Medina. Forensic identification in the aftermath of human rights crimes in Chile: A decentered computer history. Technology and Culture, 59(4S):S100–S133, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/712115.

Mehos:2010:SML

[Meh10] Donna C. Mehos. Sweet memories: The Laupahoehoe Train Museum, Hawai’i. Technology and Culture, 51(2):462–469, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381295.

Meikle:2010:ACS

[Mei10] Jeffrey L. Meikle. American Consumer Society, 1865–2005: From Hearth to HDTV (review). Technology and Culture, 51(2):525–527, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381296.

Meikle:2013:SSS

[Mei13] Jeffrey L. Meikle. Signs, Streets, and Storefronts: A History of Architecture and Graphics Along America’s Commercial Corridors by Martin Treu (review). Technology and Culture, 54(3):672–674, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
520297.

Mei:2019:SRJ

[Mei19] Jianjun Mei. Some reflections on Joseph Needham’s intellectual heritage. Technology and Culture, 60(2):594–603, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726941.

Mercelis:2012:LBT

[Mer12] Joris Mercelis. Leo Baekeland’s transatlantic struggle for bakelite: Patenting inside and outside of America. Technology and Culture, 53(2):366–400, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476809.

Mercelis:2017:EMC

[Mer17a] Joris Mercelis. The Eureka Myth: Creators, Innovators, and Everyday Intellectual Property by Jessica Silbey (review). Technology and Culture, 58(2):615–616, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/663011.

Merleaux:2017:BSR

[Mer17b] April Merleaux. Blood Sugar: Racial Pharmacology and Food Justice in Black America by Anthony Ryan Hatch (review). Technology and Culture, 58(3):893–895, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/669489.

Messeri:2016:EEJ

[Mes16] Lisa Messeri. Exploration and Engineering: The Jet Propulsion Laboratory and the Quest for Mars by Erik M. Conway (review). Technology and Culture, 57(3):693–694, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/630949.

Meyer:2012:AAR

[Mey12] Alan D. Meyer. The Age of Airpower (review). Technology and Culture, 53(2):497–498, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/476829.

Meyer:2015:NFA

[Mey15] Alan D. Meyer. NASA’s First A: Aeronautics from 1958 to 2008 by Robert G. Ferguson (review). Technology and Culture, 56(4): 997–998, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601492.

Meyer:2016:GDS

[Mey16] Alan D. Meyer. GPS Declassified: From Smart Bombs to Smartphones by Richard D. Easton and Eric F. Frazier (review). Technology and Culture, 57(1):276–278, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/611829.

Meyer:2017:OSR

[Mey17] Alan D. Meyer. The Other Space Race: Eisenhower and the Quest for Aerospace Security by Nicholas Michael Sambaluk (review). Technology and Culture, 58(2):597–598, April 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/663001.

Meyer:2018:PPH

[Mey18] Alan D. Meyer. The Problem with Pilots: How Physicians, Engineers, and Airpower Enthusiasts Redefined Flight by Timothy P. Schultz (review). Technology and Culture, 59(4):994–995, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/711787.

Meyer:2019:AWB

[Mey19] Alan D. Meyer. The American Way of Bombing: Changing Ethical and Legal Norms, From Flying Fortresses to Drones ed. by Matthew Evangelista and Henry Shue (review). Technology and Culture, 60(2):642–643, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726961.

Mannisto-Funk:2011:CTT

[MF11] Tiina Männistö-Funk. The crossroads of technology and tradition: Vernacular bicycles in rural Finland, 1880–1910. Technology and Culture, 52(4):733–756, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/454531.

Mannisto-Funk:2016:CHS

[MF16] Tiina Männistö-Funk. Cultural Histories of Sociabilities, Spaces and Mobilities ed. by Colin Divall (review). Technology and Culture, 57(2):466–467, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/619062.

Mannisto-Funk:2017:CBW

[MF17] Tiina Männistö-Funk. Claiming the Bicycle: Women, Rhetoric, and Technology in Nineteenth-Century America by Sarah Hallenbeck (review). Technology and Culture, 58(2):586–588, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662995.

Miller:2011:HB

[Mil11] David Philip Miller. History from between. Technology and Culture, 52(3):610–613, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447785.

Millard:2012:WWR

[Mil12a] Andre Millard. Widescreen Worldwide (review). Technology and Culture, 53(4):952–954, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/490754.

Miller:2012:SSL

[Mil12b] Christopher Miller. Second Suburb: Levittown, Pennsylvania (review). Technology and Culture, 53(2):504–505, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/476830.

Millard:2013:OHS

[Mil13a] Andre Millard. The Oxford Handbook of Sound Studies ed. by Trevor Pinch and Karin Bijsterveld (review). Technology and Culture, 54(2):393–394, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/506079.

Miller:2013:ADB

[Mil13b] Jonson Miller. Alexander Dallas Bache: Building the American Nation through Science and Education in the Nineteenth Century by Axel Jansen (review). Technology and Culture, 54(1):191–193, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500806.

Miller:2019:ASM

[Mil19] Daniel Miller. Anthropological studies of mobile phones. Technology and Culture, 60(4):1093–1097, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741386.

Mirowski:2012:MCG

[Mir12] Philip Mirowski. Minding the cybernetic gap. Technology and Culture, 53(1):192–195, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/466770.

Miranti:2014:GLF

[Mir14] Paul J. Miranti. A Great Leap Forward: 1930s Depression and U.S. Economic Growth by Alexander J. Field (review). Technology and Culture, 55(3):759–761, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551630.

Miranti:2019:NLT

[Mir19] Paul Miranti. Noise: Living and Trading in Electronic Finance by Alex Preda (review). Technology and Culture, 60(2):650–652, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
726958.

Misa:2011:WLW

[Mis11a] Thomas J. Misa. When the Lights Went Out: A History of Blackouts in America (review). Technology and Culture, 52(2):404–406, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431378.

Misa:2011:DFN

[Mis11b] Thomas J. Misa. David F. Noble, 22 July 1945 to 27 December 2010. Technology and Culture, 52(2):360–372, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/431365.

Misa:2011:E

[Mis11c] Thomas J. Misa. To the Editor. Technology and Culture, 52 (1):230–232, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/ 1/article/416444. See response [Kon11].

Mitcham:2011:PP

[Mit11] Carl Mitcham. Petroski’s policy. Technology and Culture, 52 (2):380–384, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431367.

Mitman:2012:WWT

[Mit12] Gregg Mitman. Wired Wilderness: Technologies of Tracking and the Making of Modern Wildlife (review). Technology and Culture, 53(1):242–243, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466804.

Mittlefehldt:2018:WWR

[Mit18] Sarah Mittlefehldt. Wood waste and race: The industrialization of biomass energy technologies and environmental justice. Technology and Culture, 59(4):875–898, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/711765.

Mitcham:2019:TAT

[Mit19] Carl Mitcham. Thinking about Technology: How the Technological Mind Misreads Reality by Gil Germain (review). Technology and Culture, 60(1):343–344, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719970.

Mizelle:2011:PPP

[Miz11] Richard Mizelle. Perilous Place, Powerful Storms: Hurricane Protection in Coastal Louisiana (review). Technology and Culture, 52(3):647–649, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447803.

Markovski:2011:ELD

[MM11] Petar Markovski and Suzanne Moon. Everette Lee DeGolyer and geology students mapping in the Arbuckle Mountains, Oklahoma, 1905. Technology and Culture, 52(1):127–130, January 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/416409.

Mody:2010:FPF

[Mod10] Cyrus C. M. Mody. Fermilab: Physics, the Frontier, and Megascience (review). Technology and Culture, 51(1):279–280, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375797.

Mody:2011:NBE

[Mod11] Cyrus C. M. Mody. Nanoethics: Big Ethical Issues with Small Technology (review). Technology and Culture, 52(2):436–437, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/431396.

Mody:2017:AJM

[Mod17] Cyrus C. M. Mody. Ann Johnson, 28 May 1965 to 11 December 2016. Technology and Culture, 58(2):570–577, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/662990.

Mohun:2012:FAC

[Moh12] Arwen Mohun. The Food Axis: Cooking, Eating, and the Architecture of American Houses (review). Technology and Culture, 53 (1):211–212, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466787.

Mohun:2015:VTC

[Moh15] Arwen P. Mohun. Vulnerability in Technological Cultures: New Directions in Research and Governance ed. by Anique Hommels, Jessica Mesman, and Wiebe E. Bijker (review). Technology and Culture, 56(2):575–576, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582346.

Mohsin:2016:EAH

[Moh16] Anto Mohsin. Epic audiovisual history of technology course by a historian of innovation: W. Bernard Carlson, “Understanding the inventions that changed the world”. Technology and Culture, 57(3):657–660, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630933.

Molotch:2015:BBP

[Mol15] Harvey Molotch. Bathroom by Barbara Penner (review). Technology and Culture, 56(3):787–788, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/593094.

Mom:2014:OAT

[Mom14] Gijs Mom. Orchestrating automobile technology: Comfort, mobility culture, and the construction of the “Family touring car,” 1917– 1940. Technology and Culture, 55(2):299–325, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546867.

Moon:2011:AB

[Moo11] Suzanne Moon. Accepting the baton. Technology and Culture, 52(1):1–5, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416403.

Moore:2015:EWP

[Moo15] Aaron Stephen Moore. Engineering War and Peace in Modern Japan, 1868–1964 by Takashi Nishiyama (review). Technology and Culture, 56(2):545–547, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582330.

Moon:2016:WPP

[Moo16a] David Moon. Works in Progress: Plans and Realities on Soviet Farms, 1930–1963 by Jenny Leigh Smith (review). Technology and Culture, 57(3):689–690, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630947.

Moon:2016:RBB

[Moo16b] Manyong Moon. Reconstructing Bodies: Biomedicine, Health, and Nation-building in South Korea since 1945 by John P. DiMoia (review). Technology and Culture, 57(2):487–488, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/619074.

Morrow:2010:CHS

[Mor10a] John H. Morrow, Jr. Conquest: How Societies Overwhelm Others (review). Technology and Culture, 51(1):235–237, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375796.

Morse:2010:CAM

[Mor10b] Victoria Morse. Cartography in Antiquity and the Middle Ages: Fresh Perspectives, New Methods (review). Technology and Culture, 51(1):239–240, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375795.

Morris:2012:WDM

[Mor12a] Peter J. T. Morris. Wiederaufbau: Das Deutsches Museum, 1945– 1970 (review). Technology and Culture, 53(4):939–940, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490746.

Morrow:2012:FWW

[Mor12b] John H. Morrow, Jr. The First World War: A Concise Global History (review). Technology and Culture, 53(3):712–713, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/481922.

Morris:2014:CBI

[Mor14a] David Z. Morris. Cars with the boom: Identity and territory in American postwar automobile sound. Technology and Culture, 55 (2):326–353, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/546868.

Morse:2014:SMP

[Mor14b] Victoria Morse. Ships on Maps: Pictures of Power in Renaissance Europe by Richard Unger (review). Technology and Culture, 55 (1):253–254, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538923.

Morton:2014:MNB

[Mor14c] David Morton. Making Noise: From Babel to the Big Bang & Beyond by Hillel Schwartz (review). Technology and Culture, 55 (2):486–487, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/546876.

Moraglio:2015:KDI

[Mor15] Massimo Moraglio. Knights of death: Introducing bicycles and motor vehicles to Turin, 1890–1907. Technology and Culture, 56 (2):370–393, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582313.

Morton:2017:JGR

[Mor17] David Morton. Junkyards, Gearheads, and Rust: Salvaging the Automotive Past by David N. Lucsko (review). Technology and Culture, 58(1):285–286, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/648274.

Moraglio:2018:CDA

[Mor18] Massimo Moraglio. La costruzione delle Alpi: Il Novecento e il modernismo alpino (1917–2017) [Building the Alps: The Twentieth Century and International Style Architecture] by Antonio De Rossi (review). Technology and Culture, 59(1):173–174, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/692175.

Mouat:2010:JMS

[Mou10] Jeremy Mouat. John Mackay: Silver King in the Gilded Age (review). Technology and Culture, 51(4):1030–1031, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/403288.

Musiani:2016:GTG

[MS16] Francesca Musiani and Valérie Schafer. Le gouvernement des technosciences: Gouverner le progrès et ses dégâts depuis 1945 ed. by Dominique Pestre (review). Technology and Culture, 57(2):485– 486, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
619073.

Mergoupi-Savaidou:2016:PST

[MSPLC16] Eirini Mergoupi-Savaidou, Faidra Papane-Lopoulou, and Ana Carneiro. Popularization of science, technology, and medicine in the “Periphery”: A step further? Technology and Culture, 57 (4):966–977, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637934.

Mueller:2015:GWI

[Mue15] Milton Mueller. The Global War for Internet Governance by Laura DeNardis (review). Technology and Culture, 56(3):784–785, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593092.

Mukerji:2012:AI

[Muk12] Chandra Mukerji. Analyzing infrastructure. Technology and Culture, 53(3):692–695, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481912.

Mukharji:2019:ASM

[Muk19] Projit Bihari Mukharji. Akarnan: The stethoscope and making of modern Ayurveda, Bengal, c. 1894–1952. Technology and Culture, 60(4):953–978, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741379.

Mullaney:2012:MTH

[Mul12a] Thomas S. Mullaney. The moveable typewriter: How Chinese typists developed predictive text during the height of Maoism. Technology and Culture, 53(4):777–814, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/490728.

Mullen:2012:DSM

[Mul12b] Megan Mullen. Demystifying some momentous changes. Technology and Culture, 53(1):185–188, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/466772.

Muller:2016:CAW

[Mül16a] Simone M. Müller. From cabling the Atlantic to wiring the world: A review essay on the 150th anniversary of the Atlantic telegraph cable of 1866. Technology and Culture, 57(3):507–526, July 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/630927.

Mulry:2016:TME

[Mul16b] Kate Mulry. Taming Manhattan: Environmental Battles in the Antebellum City by Catherine McNeur (review). Technology and Culture, 57(3):670–671, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630936.

Mullaney:2018:QCC

[Mul18a] Thomas S. Mullaney. QWERTY in China: Chinese computing and the radical alphabet. Technology and Culture, 59(4S):S34–S65, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
712113.

Mullaney:2018:SCC

[Mul18b] Thomas S. Mullaney. Shift CTRL: Computing and new media as global, cultural, sociopolitical, and ecological. Technology and Culture, 59(4S):S1–S6, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/712111.

Munns:2014:PSR

[Mun14] David P. Munns. Probing the Sky with Radio Waves: From Wireless Technology to the Development of Atmospheric Science by Chen-Pang Yeang (review). Technology and Culture, 55(3):767– 769, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
551634.

Munday:2016:MCM

[Mun16] Pat Munday. Mining cultures and Mary cults: Where the sacred and profane meet. Technology and Culture, 57(1):1–23, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611798.

Munday:2019:CAI

[Mun19] Pat Munday. The City That Ate Itself: Butte, Montana and Its Expanding Berkeley Pit by Brian James Leech (review). Technology and Culture, 60(3):906–907, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/731459.

Murphy:2012:RNA

[Mur12] Edward Murphy. The Ruins of the New Argentina: Peronism and the Remaking of San Juan after the 1944 Earthquake (review). Technology and Culture, 53(1):229–230, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466797.

Messeri:2015:GMN

[MV15] Lisa Messeri and Janet Vertesi. The greatest missions never flown: Anticipatory discourse and the “Projectory” in technological communities. Technology and Culture, 56(1):54–85, January 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/572111.

Macedo:2016:TNL

[MVÁ16] Marta Macedo and Jaume Valentines-Álvarez. Technology and nation: Learning from the periphery. Technology and Culture, 57 (4):989–997, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637936.

Nahum:2013:BWM

[Nah13] Andrew Nahum. Britain’s War Machine: Weapons, Resources, and Experts in the Second World War by David Edgerton (review). Technology and Culture, 54(1):209–210, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500816.

Nan:2012:LAR

[Nan12] Wang Nan. Lessons amid the Rubble: An Introduction to Post-Disaster Engineering and Ethics (review). Technology and Culture, 53(4):963–964, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490760.

Needell:2017:JWT

[Nee17] Allan A. Needell. James E. Webb, technocracy, and the New Deal roots of “Space Age Management”. Technology and Culture, 58(3):790–814, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669468.

Needham:2018:PHS

[Nee18] Andrew Needham. Power on the Hudson: Storm King Mountain and the Emergence of Modern American Environmentalism by Robert Lifset (review). Technology and Culture, 59(1):179–181, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692179.

Nelson:2016:WWC

[Nel16] Lisa Nelson. This Is Why We Can’t Have Nice Things by Whitney Phillips (review). Technology and Culture, 57(3):710–711, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/630959.

Nemer:2014:CPS

[Nem14] David Nemer. Coding Places: Software Practice in a South American City by Yuri Takhteyev (review). Technology and Culture, 55(2):517–518, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/546894.

Neufeld:2011:PPR

[Neu11] Michael J. Neufeld. Physics and Politics: Research and Research Support in Twentieth-Century Germany in International Perspective (review). Technology and Culture, 52(3):649–650, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/448143.

Neufeld:2013:ESM

[Neu13] Michael J. Neufeld. Eisenhower’s Sputnik Moment: The Race for Space and International Prestige by Yanek Mieczkowski (review). Technology and Culture, 54(4):998–999, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530538.

Neufeld:2014:NWF

[Neu14a] Michael J. Neufeld. NASA in the World: Fifty Years of International Collaboration in Space by John Krige, Angelina Long Callahan, and Ashok Maharaj (review). Technology and Culture, 55(3):770–772, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551636.

Neumaier:2014:EFV

[Neu14b] Christopher Neumaier. Eco-friendly versus cancer-causing: Perceptions of diesel cars in West Germany and the United States, 1970– 1990. Technology and Culture, 55(2):429–460, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546871.

Newfield:2010:SLM

[New10] Christopher Newfield. The Scientific Life: A Moral History of a Late Modern Vocation (review). Technology and Culture, 51(4): 1058–1060, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403305.

Newman:2018:FID

[New18] Andrew Newman. Fatal Isolation: The Devastating Paris Heat Wave of 2003 by Richard C. Keller (review). Technology and Culture, 59(1):178–179, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/692178.

Niemann:2012:IWC

[Nie12] Linda Grant Niemann. Iron Will: Cleveland-Cliffs and the Mining of Iron Ore, 1847–2006 (review). Technology and Culture, 53 (1):220–221, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466792.

Nikiforova:2015:CTR

[Nik15] Natalia Nikiforova. The concept of technology and the Russian cultural research tradition. Technology and Culture, 56(1):184–203, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572115.

Nikiforova:2017:HNN

[Nik17] Natalia Nikiforova. How Not to Network a Nation: The Uneasy History of the Soviet Internet by Benjamin Peters (review). Technology and Culture, 58(1):300–301, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/648283.

Nofre:2015:IBB

[Nof15] David Nofre. It Began with Babbage: The Genesis of Computer Science by Subrata Dasgupta (review). Technology and Culture, 56(2):537–538, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582325.

Nofre:2018:WAW

[Nof18] David Nofre. What Algorithms Want: Imagination in the Age of Computing by Ed Finn (review). Technology and Culture, 59 (3):811–813, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/703162.

Nordqvist:2015:BRI

[Nor15a] Petra Nordqvist. Biological Relatives: IVF, Stem Cells, and the Future of Kinship by Sarah Franklin (review). Technology and Culture, 56(1):303–305, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/572146.

Norton:2015:FPT

[Nor15b] Peter Norton. Four paradigms: Traffic safety in the Twentieth-Century United States. Technology and Culture, 56(2):319–334, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582311.

November:2011:WWW

[Nov11] Joseph November. When women were computers: LeAnn Erickson, Top Secret Rosies. Technology and Culture, 52(4):788–791, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454535.

November:2014:FWH

[Nov14] Joseph November. The Future Was Here: The Commodore Amiga by Jimmy Maher (review). Technology and Culture, 55(2):513–515, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546892.

Nofre:2014:WTB

[NPA14] David Nofre, Mark Priestley, and Gerard Alberts. When technology became language: The origins of the linguistic conception of computer programming, 1950–1960. Technology and Culture, 55(1):40–75, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538908.

Nappi:2019:RNN

[NW19] Carla Nappi and McKenzie Wark. Reading Needham now. Technology and Culture, 60(2):562–573, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/726938.

Nye:2011:DTA

[Nye11] David Nye. Designing Tomorrow: America’s World’s Fairs of the 1930s (review). Technology and Culture, 52(4):835–837, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454555.

Nye:2013:CCE

[Nye13] David E. Nye. Car Country: An Environmental History by Christopher W. Wells (review). Technology and Culture, 54(4):987–988, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
530532.

Nye:2014:FLH

[Nye14] David E. Nye. Flickering Light: A History of Neon by Christoph Ribbat (review). Technology and Culture, 55(2):502–504, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/546886.

Nye:2015:UBH

[Nye15] David E. Nye. Utopias: A Brief History from Ancient Writings to Virtual Communities by Howard P. Segal (review). Technology and Culture, 56(4):972–973, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/601477.

Nye:2017:WIR

[Nye17] David E. Nye. The Wheel: Inventions & Reinventions by Richard W. Bulliet (review). Technology and Culture, 58(3):869–870, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/669475.

Oakes:2019:PIE

[Oak19] Tim Oakes. The Promise of Infrastructure ed. by Nikhil Anand, Akhil Gupta, and Hannah Appel (review). Technology and Culture, 60(4):1112–1114, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/741392.

Oatsvall:2016:MEC

[Oat16] Neil Oatsvall. Militarizing the Environment: Climate Change and the Security State by Robert P. Marzec (review). Technology and Culture, 57(4):1034–1035, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/637955.

Oberdiek:2014:ETP

[Obe14] Hans Oberdiek. Ethics in Technology: A Philosophical Study by Topi Heikkerö (review). Technology and Culture, 55(4):995–996, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562844.

ODonnell:2012:BFG

[O’D12] Casey O’Donnell. On the back of a flying gryphon: Soaring over/ through the global game industry. Technology and Culture, 53 (1):196–199, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466769.

Oddy:2015:HD

[Odd15] Nicholas Oddy. This Hill is dangerous. Technology and Culture, 56(2):335–369, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582312.

OReagan:2018:FBN

[OF18] Douglas O’Reagan and Lee Fleming. The FinFET breakthrough and networks of innovation in the semiconductor industry, 1980– 2005: Applying digital tools to the history of technology. Technology and Culture, 59(2):251–288, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/697232.

OHar:2010:YYM

[O’H10] George O’Har. The Yankee Yorkshireman: Migration Lived and Imagined (review). Technology and Culture, 51(4):1028–1030, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403287.

Okamoto:2017:PUP

[Oka17a] Takuji Okamoto. The Postwar U.S. Policy toward Japan and the Revival of Japanese Technology: The Introduction of TV Broadcasting and Atomic Power in Japan and Hidetoshi Shibata by Kenzo Okuda (review). Technology and Culture, 58(3):885–886, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669484.

Okamoto:2017:RDL

[Oka17b] Takuji Okamoto. Reorganization, deregulation, and liberalization: Postwar development of the Japanese electric power industry and its change after 11 March 2011. Technology and Culture, 58(1): 182–193, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648256.

Okamoto:2018:LDI

[Oka18] Takuji Okamoto. Learning from a Disaster: Improving Nuclear Safety and Security after Fukushima ed. by Edward D. Blandford, Scott D. Sagan (review). Technology and Culture, 59(1):192–193, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692186.

Orain:2017:SVP

[OL17] Arnaud Orain and Sylvain Laubé. Scholars versus practitioners?: Anchor proof testing and the birth of a mixed culture in Eighteenth-Century France. Technology and Culture, 58(1):1–34, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/648248.

Olegario:2017:CKC

[Ole17a] Rowena Olegario. The Cotton Kings: Capitalism and Corruption in Turn-of-the-Century New York and New Orleans by Bruce E. Baker and Barbara Hahn (review). Technology and Culture, 58 (2):588–589, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/662996.

Olesko:2017:HPW

[Ole17b] Kathryn Olesko. Humboldts Preußen: Wissenschaft und Technik im Aufbruch by Ursula Klein (review). Technology and Culture, 58 (4):1085–1086, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679832.

Olwell:2010:TVC

[Olw10] Russell Olwell. Two views of Cassandra: Lawrence Badash, A Nuclear Winter’s Tale, John Mueller, Atomic Obsession. Technology and Culture, 51(4):1002–1005, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403276.

Olwell:2013:NCN

[Olw13] Russell Olwell. The Neutron’s Children: Nuclear Engineers and the Shaping of Identity by Sean F. Johnston (review). Technology and Culture, 54(3):692–694, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/520308.

Olwell:2014:NSA

[Olw14] Russell Olwell. Nuclear syntheses: Audra J. Wolfe, Competing with the Soviets; Martin V. Melosi, Atomic Age America. Technology and Culture, 55(3):723–725, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551612.

Olwage:2018:PRM

[Olw18] Grant Olwage. Paul Robeson’s microphone voice and the technologies of easy singing. Technology and Culture, 59(4):823–849, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
711763.

Onaga:2017:ISH

[OM17] Lisa Onaga and Aaron S. Moore. Introduction: Searching for the historical roots of 11 March 2011. Technology and Culture, 58 (1):154–158, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648253.

Onaga:2014:GBB

[Ona14] Lisa Onaga. Genentech: The Beginnings of Biotech by Sally Smith Hughes (review). Technology and Culture, 55(2):511–512, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/546891.

Onaga:2017:RLN

[Ona17] Lisa Onaga. Reconstructing the linear no-threshold model in Japan: A historical perspective on the technics of evaluating radiation exposure. Technology and Culture, 58(1):194–205, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/648257.

Onaga:2018:RBM

[Ona18] Lisa Onaga. Radiation Brain Moms and Citizen Scientists: The Gender Politics of Food Contamination after Fukushima by Aya Hirata Kimura (review). Technology and Culture, 59(1):194–195, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
692187.

OReagan:2017:KHP

[O’R17] Douglas O’Reagan. Know-how in postwar business and law. Technology and Culture, 58(1):121–153, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/648252.

OReagan:2019:DID

[O’R19] Douglas O’Reagan. Designing an Internet by David D. Clark (review). Technology and Culture, 60(4):1124–1126, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/741394.

Orain:2015:PGA

[Ora15] Arnaud Orain. La pièce et le geste: Artisans, marchands et savoir technique à Londres au XVIII^{e} siècle by Liliane Hilaire-Pérez (review). Technology and Culture, 56(3):752–754, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/593074.

Ore:2011:MHS

[Ore11] Janet Ore. Mobile home syndrome: Engineered woods and the making of a new domestic ecology in the Post–World War II Era. Technology and Culture, 52(2):260–286, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/431361.

Ore:2012:BWB

[Ore12] Janet Ore. Being and Well-Being: Health and the Working Bodies of Silicon Valley (review). Technology and Culture, 53(1):238–239, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466802.

Ore:2013:NND

[Ore13] Janet Ore. Nature Next Door: Cities and Trees in the American Northeast by Ellen Stroud (review). Technology and Culture, 54(3):676–677, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520299.

Oldenziel:2012:KJF

[OS12] Ruth Oldenziel and Johan Schot. Karen Johnson Freeze, 22 October 1945–19 March 2009: A tribute. Technology and Culture, 53(2):435–441, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476811.

Onaga:2016:DHD

[OS16] Lisa Onaga and Hanna Rose Shell. Digital histories of disasters: History of technology through social media. Technology and Culture, 57(1):225–230, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/611807.

Oslund:2010:LHW

[Osl10] Karen Oslund. Leviathan: The History of Whaling in America (review). Technology and Culture, 51(3):745–746, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390851.

Pacey:2018:HAH

[Pac18] Arnold Pacey. History and the art of hand-pump maintenance — reflections on my working life. Technology and Culture, 59(3):770– 780, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703146.

Paju:2016:HEC

[Paj16] Petri Paju. Hacking Europe: From Computer Cultures to Demoscenes ed. by Gerard Alberts, Ruth Oldenziel (review). Technology and Culture, 57(4):1041–1042, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/637959.

Pannabecker:2010:PPT

[Pan10] John Pannabecker. La Presse et les périodiques techniques en Europe, 1750–1950 (review). Technology and Culture, 51(3):763–764, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
390862.

Pandora:2012:SCT

[Pan12] Katherine Pandora. Switching Codes: Thinking Through Digital Technology in the Humanities and the Arts (review). Technology and Culture, 53(4):961–962, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/490759.

Pannabecker:2015:WWE

[Pan15] John Pannabecker. Women and Work in Eighteenth-Century France ed. by Daryl M. Hafter and Nina Kushner (review). Technology and Culture, 56(4):978–980, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601481.

Panzer:2016:ESP

[Pan16] Mary Panzer. At the Edge of Sight: Photography and the Unseen by Shawn Michelle Smith (review). Technology and Culture, 57(2):478–479, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/619069.

Pannabecker:2018:DPT

[Pan18a] John Pannabecker. Dompter Prométhée: Technologies et socialismes à l’âge romantique, 1820–1870 ed. by François Jarrige (review). Technology and Culture, 59(1):196–197, January 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/692161.

Pannabecker:2018:LGR

[Pan18b] John Pannabecker. L’amphithéâtre, la galerie et le rail: Le Conservatoire des arts et métiers, ses collections et le chemin de fer au XIX^{e} siècle by Lionel Dufaux (review). Technology and Culture, 59(4):968–970, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/711772.

Pannabecker:2018:RII

[Pan18c] John Pannabecker. Research, invent, improve: A Dictionnaire Technologique for non-elites (1822–35). Technology and Culture, 59(3):546–589, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/703140.

Panzer:2018:BAL

[Pan18d] Mary Panzer. Berenice Abbott: A Life in Photography by Julia Van Haaften (review). Technology and Culture, 59(4):985–986, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/711782.

Paris:2016:DAI

[Par16] Ivan Paris. Domestic appliances and industrial design: The Italian white-goods industry during the 1950s and 1960s. Technology and Culture, 57(3):612–648, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630931.

Parks:2017:SPH

[Par17] Richard C. Parks. The Social Project: Housing Postwar France by Kenny Cupers (review). Technology and Culture, 58(2):593–595, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
662999.

Parrinello:2018:SPS

[Par18] Giacomo Parrinello. Systems of power — a spatial envirotechnical approach to water power and industrialization in the Po Valley of Italy, ca. 1880–1970. Technology and Culture, 59(3):652–688, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703143.

Paton:2011:CCC

[Pat11] Miranda Paton. Cotton, Climate, and Camels in Early Islamic Iran: A Moment in World History (review). Technology and Culture, 52(1):180–181, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/416414.

Patel:2015:HNG

[Pat15] Sujata Patel. House, But No Garden: Apartment Living in Bombay’s Suburbs, 1898–1964 by Nikhil Rao (review). Technology and Culture, 56(4):1003–1005, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/601496.

Patton:2019:WDW

[Pat19] Elizabeth Patton. Where does work belong? Home-based work and communication technology within the American middle-class postwar home. Technology and Culture, 60(2):523–552, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726936.

Penner:2018:CKH

[Pen18] Barbara Penner. The Cornell kitchen: Housing and design research in postwar America. Technology and Culture, 59(1):48–94, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/692167.

Perdue:2012:CTK

[Per12] Peter C. Perdue. The Crafting of the 10,000 Things: Knowledge and Technology in Seventeenth-Century China (review). Technology and Culture, 53(2):483–484, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/476831.

Petroski:2010:EPN

[Pet10] Henry Petroski. The Engineering Project: Its Nature, Ethics, and Promise (review). Technology and Culture, 51(2):540–542, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381264.

Peters:2011:OHR

[Pet11] Tom F. Peters. Obelisk: A History (review). Technology and Culture, 52(2):389–390, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/431369.

Peters:2012:KZE

[Pet12] Peter F. Peters. Klangmaschinen zwischen Experiment und Medientechnik (review). Technology and Culture, 53(1):209–211, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466786.

Petroski:2013:GGW

[Pet13] Henry Petroski. Good Guys, Wiseguys, and Putting Up Buildings: A Life in Construction by Samuel C. Florman (review). Technology and Culture, 54(2):422–424, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506096.

Petersen:2014:SSO

[Pet14] Leif Inge Ree Petersen. Sultans of science: An opportunity lost. Technology and Culture, 55(4):973–987, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/562842.

Pettit:2016:CSA

[Pet16] Michael Pettit. The Classification of Sex: Alfred Kinsey and the Organization of Knowledge by Donna J. Drucker (review). Technology and Culture, 57(1):271–272, January 2016. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/611826.

Peters:2018:RMC

[Pet18] Benjamin Peters. Rise of the Machines: A Cybernetic History by Thomas Rid (review). Technology and Culture, 59(2):492–494, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
697252.

Petchey:2019:NZT

[Pet19] Peter Petchey. New Zealand’s technological participation in the international goldfields: The example of the Stamper battery. Technology and Culture, 60(2):494–522, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/726935.

Pfaffenberger:2013:BBW

[Pfa13] Bryan Pfaffenberger. Broken Ballots: Will Your Votes Count? by Douglas W. Jones and Barbara Simons (review). Technology and Culture, 54(4):1001–1003, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530540.

Philpotts:2014:CDN

[Phi14] Trey Philpotts. Charles Dickens’s Networks: Public Transportation and the Novel by Jonathan H. Grossman (review). Technology and Culture, 55(3):744–745, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/551621.

Pickering:2014:CFS

[Pic14] Andrew Pickering. Cybernetic futures: Stanislaw Lem, Summa Technologiae. Technology and Culture, 55(1):245–248, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538920.

Pike:2015:UMM

[Pik15] David Pike. Underground Movements: Modern Culture on the New York City Subway by Sunny Stalter-Pace (review). Technology and Culture, 56(1):278–279, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/572131.

Pinch:2010:ITG

[Pin10] Trevor Pinch. The invisible technologies of Goffman’s sociology from the merry-go-round to the Internet. Technology and Culture, 51(2):409–424, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381286.

Platt:2010:MEE

[Pla10] Rutherford H. Platt. Megalopolis: an enduring enigma. Technology and Culture, 51(1):223–226, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375774.

Plaiss:2016:NMP

[Pla16] Adam Plaiss. From natural monopoly to public utility: Technological determinism and the political economy of infrastructure in Progressive-Era America. Technology and Culture, 57(4):806–830, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
637926.

Plantin:2017:TII

[Pla17] Jean-Christophe Plantin. The Turn to Infrastructure in Internet Governance ed. by F. Musiani, et al. (review). Technology and Culture, 58(3):900–902, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/669493.

Plotnick:2012:ICE

[Plo12] Rachel Plotnick. At the interface: The case of the electric push button, 1880–1923. Technology and Culture, 53(4):815–845, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490729.

Pomeroy:2010:TLR

[Pom10] Steven Pomeroy. Truth, Lies, and O-Rings: Inside the Space Shuttle Challenger Disaster (review). Technology and Culture, 51(4): 1038–1040, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403293.

Popp:2011:MAC

[Pop11] Richard Popp. Machine-age communication: Media, transportation, and contact in the Interwar United States. Technology and Culture, 52(3):459–484, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/447778.

Popp:2013:SCA

[Pop13a] Richard K. Popp. The Sounds of Capitalism: Advertising, Music, and the Conquest of Culture by Timothy D. Taylor (review). Technology and Culture, 54(3):679–681, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520301.

Popplow:2013:MSP

[Pop13b] Marcus Popplow. Die moderne Strasse Planung, Bau und Verkehr vom 18. bis zum 20. Jahrhundert ed. by Hans-Liudger Dienel and Hans-Ulrich Schiedt (review). Technology and Culture, 54(2):403– 405, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506085.

Popp:2016:FRF

[Pop16] Richard K. Popp. FAXED: The Rise and Fall of the Fax Machine by Jonathan Coopersmith (review). Technology and Culture, 57 (1):278–280, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611830.

Pope:2019:CAI

[Pop19] Norris Pope. Color in the Age of Impressionism: Commerce, Technology, and Art by Laura Anne Kalba (review). Technology and Culture, 60(1):330–331, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719962.

Post:2010:BSH

[Pos10] Robert C. Post. Back at the start: History and technology and culture. Technology and Culture, 51(4):961–994, October 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/403274.

Post:2015:AAL

[Pos15] Bob Post. America’s Assembly Line by David E. Nye (review). Technology and Culture, 56(4):991–993, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601489.

Patiniotis:2016:BFG

[PR16] Manolis Patiniotis and Pedro M. P. Raposo. Beyond fixed geographies: Moving localities and the making of knowledge. Technology and Culture, 57(4):930–939, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/637931.

Prakash:2012:TMI

[Pra12] Gyan Prakash. Tracking Modernity: India’s Railway Culture and the Culture of Mobility (review). Technology and Culture, 53 (3):719–720, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481926.

Pratcher:2013:DVM

[Pra13] Anthony Pratcher. Desert Visions and the Making of Phoenix, 1860–2009 by Philip VanderMeer (review). Technology and Culture, 54(3):674–675, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520298.

Prasad:2019:BSR

[Pra19] Amit Prasad. Burdens of the scientific revolution: Euro/West-Centrism, black boxed machines, and the (post) colonial present. Technology and Culture, 60(4):1059–1082, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741383.

PresasIPuig:2010:TSB

[Pre10a] Albert Presas I Puig. Technoscientific synergies between Germany and Spain in the Twentieth Century: Continuity amid radical change. Technology and Culture, 51(1):80–98, January 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/375782.

Pretzer:2010:SCM

[Pre10b] William S. Pretzer. Structures of Change in the Mechanical Age: Technological Innovation in the United States, 1790–1865 (review). Technology and Culture, 51(4):1026–1028, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/403286.

Prescott:2013:PFS

[Pre13] Heather Prescott. The pill at fifty: Scientific commemoration and the politics of American memory. Technology and Culture, 54(4):735–745, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530509.

Prescott:2019:WMA

[Pre19] Heather Munro Prescott. We must not allow a contraception gap: Planned parenthood’s campaign for new birth control and feminist health activism in the 1990s. Technology and Culture, 60(3):816– 832, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
731444.

Priestley:2019:FPH

[Pri19] Mark Priestley. For Fun and Profit: A History of the Free and Open Source Software Revolution by Christopher J. Tozzi (review). Technology and Culture, 60(2):655–657, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/726960.

Pardo-Tomas:2013:TMV

[PT13] José Pardo-Tomás. Tintorería y Medicina en la Valencia del Siglo XV: El Manual de Joanot Valero by Lluís Cifuentes i Comamala and Ricardo Córdoba de la Llave (review). Technology and Culture, 54(4):968–969, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530521.

Puaksom:2017:SND

[Pua17] Davisakd Puaksom. Siam’s New Detectives: Visualizing Crime and Conspiracy in Modern Thailand by Samson W. Lim (review). Technology and Culture, 58(3):883–884, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/669483.

Pultz:2016:CDR

[Pul16] John Pultz. The Camera Does the Rest: How Polaroid Changed Photography by Peter Buse (review). Technology and Culture, 57 (4):1025–1027, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637950.

Pursell:2010:TCP

[Pur10] Carroll W. Pursell. Technologies as cultural practice and production. Technology and Culture, 51(3):715–722, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390844.

Purcell:2017:SPE

[Pur17] Darren Purcell. Spatializing Politics: Essays on Power and Place ed. by Delia Duong Ba Wendel and Fallon Samuels Aidoo (review). Technology and Culture, 58(4):1074–1076, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679826.

Quivik:2017:SPA

[Qui17] Fredric Quivik. Steam and power: Andreas Malm, Fossil Capital. Technology and Culture, 58(3):866–868, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669474.

Rabier:2013:SBI

[Rab13a] Christelle Rabier. Stays and Body Image in London: The Staymaking Trade, 1680–1810 by Lynn Sorge-English (review). Technology and Culture, 54(2):398–399, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506082.

Rabier:2013:ICM

[Rab13b] Christelle Rabier. Introduction: The crafting of medicine in the Early Industrial Age. Technology and Culture, 54(3):437–459, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520278.

Rae:2016:DMP

[Rae16] Gavin Rae. Design, Mediation, and the Posthuman ed. by Dennis M. Weiss, Amy D. Propen, and Colbey Emmerson Reid (review).

Technology and Culture, 57(1):283–284, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611833.

Russell:2011:NPS

[RAF^{+}11] Edmund Russell, James Allison, Thomas Finger, John K. Brown, Brian Balogh, and W. Bernard Carlson. The nature of power: Synthesizing the history of technology and environmental history. Technology and Culture, 52(2):246–259, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/431360.

Rampling:2010:AAH

[Ram10] Jennifer Rampling. Alchemy and Authority in the Holy Roman Empire (review). Technology and Culture, 51(2):497–498, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381289.

Rampling:2011:AAH

[Ram11] Jennifer Rampling. Alchemy and Authority in the Holy Roman Empire (review). Technology and Culture, 52(1):185–187, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416417.

Ramana:2014:ASB

[Ram14] M. V. Ramana. Atomic State: Big Science in Twentieth-Century India by Jahnavi Phalkey (review). Technology and Culture, 55 (3):762–763, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551631.

Randolph:2011:RNA

[Ran11] Scott Randolph. Railroad Noir: The American West and the End of the Twentieth Century (review). Technology and Culture, 52(3):637–638, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447797.

Rankin:2014:GRP

[Ran14] William Rankin. The geography of radionavigation and the politics of intangible artifacts. Technology and Culture, 55(3):622–674, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/551608.

Rankin:2015:HGR

[Ran15] William Rankin. Historical GIS Research in Canada ed. by Jennifer Bonnell and Marcel Fortin (review). Technology and Culture, 56 (4):1006–1008, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601498.

Ratcliff:2012:ACC

[Rat12] Jessica Ratcliff. Art to cheat the common-weale: Inventors, projectors, and patentees in English satire, ca.1630–70. Technology and Culture, 53(2):337–365, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/476808.

Reeds:2013:ATF

[Ree13] Karen Reeds. All the tools fit to print. Technology and Culture, 54(2):373–376, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506075.

Reed:2016:UNM

[Ree16] Christopher A. Reed. Unearthing the Nation: Modern Geology and Nationalism in Republican China by Grace Yen Shen (review). Technology and Culture, 57(1):260–262, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611820.

Rees:2018:CMI

[Ree18] Jonathan Rees. Cornering the Market: Independent Grocers and Innovation in American Small Business by Susan Spellman (review). Technology and Culture, 59(4):972–973, October 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/711774.

Rees:2019:RSC

[Ree19] Jonathan Rees. Refrigerator: The Story of Cool in the Kitchen by Helen Peavitt (review). Technology and Culture, 60(4):1109–1110, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
741408.

Regele:2018:WBC

[Reg18] Lindsay Schakenbach Regele. The world’s best carpets: Erastus Bigelow and the financing of Antebellum innovation. Technology and Culture, 59(1):126–151, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/692169.

Reilly:2012:SDL

[Rei12] Kara Reilly. Sublime Dreams of Living Machines: The Automaton in the European Imagination (review). Technology and Culture, 53(3):704–705, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/481917.

Rettberg:2016:GPW

[Ret16] Jill Walker Rettberg. Garments of Paradise: Wearable Discourse in the Digital Age by Susan Elizabeth Ryan (review). Technology and Culture, 57(2):502–504, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/619083.

Reuss:2014:RDM

[Reu14] Martin Reuss. A revolting development: Matthew wisnioski, engineers for change. Technology and Culture, 55(3):726–729, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/551613.

Reynard:2012:LPI

[Rey12] Pierre Claude Reynard. Le Laboratoire des pollutions industrielles: Paris, 1770–1830 (review). Technology and Culture, 53(3):705– 707, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481918.

Rhees:2014:PTE

[Rhe14] David Rhees. The Papers of Thomas A. Edison: Losses and Loyalties, April 1883–December 1884, Edited by Paul B. Israel et al. (review). Technology and Culture, 55(3):749–750, July 2014. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/551624.

Richter:2013:GCE

[Ric13] Amy G. Richter. Grand Central’s Engineer: William J. Wilgus and the Planning of Modern Manhattan by Kurt C. Schlichting (review). Technology and Culture, 54(3):670–672, July 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/520296.

Rieger:2016:AAE

[Rie16] Bernhard Rieger. Atlantic Automobilism: Emergence and Persistence of the Car 1895–1940 by Gijs Mom (review). Technology and Culture, 57(3):680–682, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/630942.

Riordan:2011:MMD

[Rio11] Michael Riordan. Makers of the Microchip: A Documentary History of Fairchild Semiconductor (review). Technology and Culture, 52(3):660–662, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447809.

Rizk:2010:BAA

[Riz10] Mysoon Rizk. The Big Archive: Art from Bureaucracy (review). Technology and Culture, 51(3):773–775, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/390868.

Roberts:2010:HTA

[Rob10a] Jonathan Roberts. Healing Traditions: African Medicine, Cultural Exchange, and Competition in South Africa, 1820–1948 (review). Technology and Culture, 51(2):533–535, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/381291.

Robertson:2010:NCC

[Rob10b] Frances Robertson. A Nation of Counterfeiters: Capitalists, Con Men, and the Making of the United States (review). Technology and Culture, 51(2):522–523, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/381298.

Robertson:2010:WAA

[Rob10c] Frances Robertson. The Work of Art in the Age of Its Technological Reproducibility, and Other Writings on Media (review). Technology and Culture, 51(1):275–277, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375794.

Roberts:2012:MDH

[Rob12] Peder Roberts. Merchants of Doubt: How a Handful of Scientists Obscured the Truth on Issues from Tobacco Smoke to Global Warming (review). Technology and Culture, 53(1):245–247, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
466806.

Roberts:2013:AIC

[Rob13] Lissa Roberts. Agency and industry: Charles C. Gillispie’s “The Natural History of Industry,” Then and Now. Technology and Culture, 54(4):922–941, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/530516.

Robertson:2017:LFR

[Rob17] Craig Robertson. Learning to file: Reconfiguring information and information work in the Early Twentieth Century. Technology and Culture, 58(4):955–981, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679820.

Rockman:2017:I

[Roc17] Seth Rockman. Introduction. Technology and Culture, 58(2):487– 505, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
662985.

Rohde:2019:LFP

[Roh19] Joy Rohde. Looking Forward: Prediction and Uncertainty in Modern America by Jamie L. Pietruska (review). Technology and Culture, 60(3):922–924, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/731462.

Roland:2010:TAW

[Rol10a] Alex Roland. Technology and the American Way of War since 1945 (review). Technology and Culture, 51(3):739–741, July 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/390848.

Roland:2010:WSM

[Rol10b] Alex Roland. On the water: Stories from Maritime America: National Museum of American History, Washington, D.C. Technology and Culture, 51(3):723–727, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/390845.

Roland:2010:SSH

[Rol10c] Alex Roland. The state of space history: Steven J. Dick and Roger D. Launius, eds., Societal Impact of Spaceflight. Technology and Culture, 51(1):220–222, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375775.

Roland:2010:WNA

[Rol10d] Alex Roland. Was the nuclear arms race deterministic? Technology and Culture, 51(2):444–461, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/381267.

Roland:2016:AII

[Rol16] Alex Roland. America Inc.? Innovation and Enterprise in the National Security State by Linda Weiss (review). Technology and Culture, 57(3):694–696, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630950.

Roland:2017:MJE

[Rol17] Alex Roland. Making Jet Engines in World War II: Britain, Germany, and the United States by Hermione Giffard (review). Technology and Culture, 58(3):878–879, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/669480.

Roland:2018:RPA

[Rol18] Alex Roland. Reinventing the Propeller: Aeronautical Specialty and the Triumph of the Modern Airplane by Jeremy R. Kinney (review).

Technology and Culture, 59(4):987–988, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711783.

Rome:2010:NCE

[Rom10] Adam Rome. The Nature of Cities: Ecological Visions and the American Urban Professions, 1920–1960 (review). Technology and Culture, 51(4):1043–1044, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/403296.

Rostgaard:2010:WSM

[Ros10] Marianne Rostgaard. Wool and Society: Manufacturing Policy, Economic Thought, and Local Production in 18th-Century Iceland (review). Technology and Culture, 51(1):244–246, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/375793.

Rosenband:2011:MRI

[Ros11] Leonard N. Rosenband. Midnight Ride, Industrial Dawn: Paul Revere and the Growth of American Enterprise (review). Technology and Culture, 52(4):823–824, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/454547.

Rosen:2012:OLC

[Ros12a] Paul Rosen. One Less Car: Bicycling and the Politics of Automobility (review). Technology and Culture, 53(2):507–508, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476834.

Rosenband:2012:IWL

[Ros12b] Leonard N. Rosenband. Industrial Work and Life: An Anthropological Reader. Vol. 78: London School of Economics Monographs on Social Anthropology (review). Technology and Culture, 53(2):477– 479, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476835.

Rosner:2013:RFH

[Ros13] David Rosner. The Rise and Fall of the Healthy Factory: The Politics of Industrial Health in Britain, 1914–60 by Vicky Long (review). Technology and Culture, 54(3):683–685, July 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/520303.

Rosenthal:2017:NIE

[Ros17] Caitlin Rosenthal. Numbers for the innumerate: Everyday arithmetic and Atlantic capitalism. Technology and Culture, 58(2):529– 544, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
662987.

Rothschild:2016:DAM

[Rot16] Rachel Rothschild. Détente from the air: Monitoring air pollution during the Cold War. Technology and Culture, 57(4):831–865, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
637927.

Russell:2014:SAI

[RS14] Andrew L. Russell and Valérie Schafer. In the shadow of ARPANET and Internet: Louis Pouzin and the Cyclades network in the 1970s. Technology and Culture, 55(4):880–907, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/562835.

Rubinson:2018:DCH

[Rub18] Paul Rubinson. Doomed to Cooperate: How American and Russian Scientists Joined Forces to Avert Some of the Greatest Post-Cold War Nuclear Dangers ed. by Siegfried Hecker (review). Technology and Culture, 59(1):187–188, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/692183.

Rudolph:2012:TMF

[Rud12] John L. Rudolph. Teaching materials and the fate of dynamic biology in American classrooms after Sputnik. Technology and Culture, 53(1):1–36, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466780.

Rudolph:2013:LDE

[Rud13] John L. Rudolph. Learning by Doing: Experiments and Instruments in the History of Science Teaching ed. by Peter Heering and Roland Wittje (review). Technology and Culture, 54(1):225–227, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500825.

Rudolph:2018:IEC

[Rud18] John L. Rudolph. Innocent Experiments: Childhood and the Culture of Popular Science in the United States by Rebecca Onion (review). Technology and Culture, 59(1):197–199, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/692162.

Ruskin:2011:HEO

[Rus11a] Steven Ruskin. The Heavens on Earth: Observatories and Astronomy in Nineteenth-Century Science and Culture (review). Technology and Culture, 52(3):628–629, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/447792.

Russo:2011:CAC

[Rus11b] Alexander Russo. Calling All Cars: Radio Dragnets and the Technology of Policing (review). Technology and Culture, 52(4):837–838, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454554.

Russo:2012:RUP

[Rus12] Alexander Russo. Radio Utopia: Postwar Audio Documentary in the Public Interest (review). Technology and Culture, 53(2):500– 502, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476836.

Ruuska:2011:GDI

[Ruu11] Alex Ruuska. Ghost dancing and the iron horse: Surviving through tradition and technology. Technology and Culture, 52(3):574–597, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
447782.

Russell:2018:AIT

[RV18] Andrew L. Russell and Lee Vinsel. After innovation, turn to maintenance. Technology and Culture, 59(1):1–25, January 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/692165.

Rydell:2011:MGT

[Ryd11] Robert W. Rydell. Mr. Gatling’s Terrible Marvel: The Gun that Changed Everything and the Misunderstood Genius Who Invented It (review). Technology and Culture, 52(2):397–399, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/431374.

Saha:2013:CCC

[Sah13a] Madhumita Saha. Cultivating the Colonies: Colonial States and Their Environmental Legacies ed. by Christina Folke Ax et al. (review). Technology and Culture, 54(1):184–185, January 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/500802.

Saha:2013:FSF

[Sah13b] Madhumita Saha. Food for soil, food for people: Research on food crops, fertilizers, and the making of “Modern” Indian agriculture. Technology and Culture, 54(2):289–316, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/506071.

Saha:2014:IPS

[Sah14] Madhumita Saha. Indigo Plantations and Science in Colonial India by Prakash Kumar (review). Technology and Culture, 55(2):492– 494, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546880.

Sambrook:2013:MGM

[Sam13] Stephen Sambrook. Making a Green Machine: The Infrastructure of Beverage Container Recycling by Finn Arne Jørgensen (review). Technology and Culture, 54(1):215–216, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500819.

Sanabria:2011:CRA

[San11] Sergio L. Sanabria. Compass and Rule: Architecture as Mathematical Practice in England 1500–1750 (review). Technology and Culture, 52(2):391–392, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/431370.

Sanabria:2012:WLS

[San12] Sergio Sanabria. Working with Limestone: The Science, Technology and Art of Medieval Limestone Monuments (review). Technology and Culture, 53(2):481–482, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/476837.

Sanford:2017:BBM

[San17] Ezelle Sanford III. Backwater Blues: The Mississippi Flood of 1927 in the African American Imagination by Richard M. Mizelle, Jr. (review). Technology and Culture, 58(2):590–591, April 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/662997.

Satia:2010:WWE

[Sat10a] Priya Satia. War, wireless, and empire: Marconi and the British warfare state, 1896–1903. Technology and Culture, 51(4):829–853, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403267.

Satya:2010:RRB

[Sat10b] Laxman D. Satya. Reproductive Restraints: Birth Control in India, 1877–1947 (review). Technology and Culture, 51(2):530–531, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381283.

Satchell:2013:WFT

[Sat13a] Veront M. Satchell. At work in the fields of their lords. Technology and Culture, 54(2):382–388, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506077.

Satia:2013:BMW

[Sat13b] Priya Satia. The Battle for the Mind: War and Peace in the Era of Mass Communication by Gary S. Messinger (review). Technology and Culture, 54(1):218–219, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/500821.

Saunders:2011:HCW

[Sau11] Richard Saunders. Historic Canals and Waterways of South Carolina (review). Technology and Culture, 52(4):824–826, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454546.

Saraiva:2017:TNL

[SC17] Tiago Saraiva and Ana Cardoso De Matos. Technological nocturne: The Lisbon Industrial Institute and romantic engineering (1849– 1888). Technology and Culture, 58(2):422–458, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/662983.

Scarpino:2011:AFH

[Sca11] Philip Scarpino. America’s Fires: A Historical Context for Policy and Practice (review). Technology and Culture, 52(3):646–647, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447802.

Scales:2014:SUP

[Sca14] Rebecca Scales. Soundscapes of the Urban Past: Staged Sound as Mediated Cultural Heritage ed. by Karin Bijsterveld (review). Technology and Culture, 55(1):265–267, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/538930.

Schultz:2010:MCS

[Sch10a] Joe Schultz. Michigan central station, Detroit, 2010. Technology and Culture, 51(4):889–892, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403270.

Schultz:2010:FSB

[Sch10b] Joseph Schultz. Froth! The Science of Beer (review). Technology and Culture, 51(3):759–760, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/390860.

Schwantes:2010:SEL

[Sch10c] Benjamin Schwantes. Stairway to Empire: Lockport, the Erie Canal, and the Shaping of America (review). Technology and Culture, 51(2):504–506, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381253.

Schiffer:2011:TCR

[Sch11a] Michael Brian Schiffer. Technology and Culture (review). Technology and Culture, 52(4):816–817, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/454544.

Schleifer:2011:EPS

[Sch11b] David Schleifer. Empty Pleasures: The Story of Artificial Sweeteners from Saccharin to Splenda (review). Technology and Culture, 52(4):849–850, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454562.

Schneider:2011:GWW

[Sch11c] Eric Schneider. Gender, Work and Wages in Industrial Revolution Britain (review). Technology and Culture, 52(1):195–196, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/416423.

Schneider:2011:SCT

[Sch11d] Jen Schneider. Scientific Cultures, Technological Challenges: A Transatlantic Perspective (review). Technology and Culture, 52 (1):228–229, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416443.

Schandler:2012:KOG

[Sch12a] Matthew Schandler. Korea’s Online Gaming Empire (review). Technology and Culture, 53(2):523–524, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/476838.

Schleifer:2012:PSH

[Sch12b] David Schleifer. The perfect solution: How trans fats became the healthy replacement for saturated fats. Technology and Culture, 53(1):94–119, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466777.

Schneider:2012:LCH

[Sch12c] Jen Schneider. Lab Coats in Hollywood: Science, Scientists, and Cinema (review). Technology and Culture, 53(1):252–253, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466782.

Schneider:2012:TNT

[Sch12d] Jen Schneider. Tech-Noir: A Theory of the Development of Popular Genres (review). Technology and Culture, 53(4):954–955, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/490755.

Schwantes:2012:SDA

[Sch12e] Carlos A. Schwantes. Out of Steam: Dieselization and American Railroads, 1920–1960 (review). Technology and Culture, 53 (1):218–219, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466791.

Scheffler:2013:BCD

[Sch13a] Robin Wolfe Scheffler. Biomedical Computing: Digitizing Life in the United States by Joseph Adam November (review). Technology and Culture, 54(4):1005–1006, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/530542.

Schleifer:2013:SSA

[Sch13b] David Schleifer. Sweet Stuff: An American History of Sweeteners from Sugar to Sucralose by Deborah Jean Warner (review). Technology and Culture, 54(1):205–207, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/500814.

Schmucki:2013:RRL

[Sch13c] Barbara Schmucki. Routes, Roads and Landscapes ed. by Brita Brenna et al. (review). Technology and Culture, 54(1):186–188, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500803.

Schneider:2013:SAT

[Sch13d] Jen Schneider. Science on American Television: A History by Marcel Chotkowski LaFollette (review). Technology and Culture, 54 (4):999–1001, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530539.

Schwartz:2013:DSN

[Sch13e] Hillel Schwartz. Discord: The Story of Noise by Mike Goldsmith (review). Technology and Culture, 54(3):656–657, July 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/520288.

Schaefer:2014:ICC

[Sch14a] Dagmar Schaefer. Imagined Civilizations: China, the West, and Their First Encounters by Roger Hart (review). Technology and Culture, 55(4):998–1000, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562846.

Schloss:2014:SCL

[Sch14b] Joseph Schloss. Sound Clash: Listening to American Studies Edited by Kara Keeling and Josh Kun (review). Technology and Culture, 55(3):736–737, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/551616.

Schramm:2014:PRV

[Sch14c] Jeff Schramm. The Pennsylvania Railroad. Volume 1: Building an Empire, 1846–1917 by Albert J. Churella (review). Technology and Culture, 55(1):258–259, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/538926.

Schlombs:2015:GCT

[Sch15a] Corinna Schlombs. Girls Coming to Tech!: A History of American Engineering Education for Women by Amy Sue Bix (review). Technology and Culture, 56(4):995–996, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601491.

Schneider:2015:SFT

[Sch15b] Jen Schneider. Science Fiction TV by J. P. Telotte (review). Technology and Culture, 56(3):779–780, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/593089.

Schot:2016:CSD

[Sch16] Johan Schot. Confronting the second deep transition through the historical imagination. Technology and Culture, 57(2):445–456, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
619058.

Schafer:2019:LME

[Sch19] Valérie Schafer. L’invention de la mémoire: Écrire, enregistrer, numériser [The invention of memory: writing, recording, digitizing] by Michel Laguës, Denis Beaudouin and Georges Chapoutier (review). Technology and Culture, 60(2):649–650, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726953.

Scranton:2012:EHU

[Scr12] Philip Scranton. Evolutionary History: Uniting History and Biology to Understand Life on Earth (review). Technology and Culture, 53(2):472–473, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476839.

Soens:2019:SWE

[SDJ19] Tim Soens, Greet De Block, and Iason Jongepier. Seawalls at work: Envirotech and labor on the North Sea Coast before 1800. Technology and Culture, 60(3):688–725, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/731440.

Seavers:2011:AGC

[Sea11] Stephanie Seavers. The Alchemy of Glass: Counterfeit, Imitation, and Transmutation in Ancient Glassmaking (review). Technology and Culture, 52(4):813–814, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/454539.

Seely:2015:WWA

[See15a] Bruce Seely. Wind Wizard: Alan G. Davenport and the Art of Wind Engineering by Siobhan Roberts (review). Technology and Culture, 56(1):281–283, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/572133.

Seely:2015:SSO

[See15b] Bruce E. Seely. The state of SHOT: Organizational capacity in a volunteer-based organization. Technology and Culture, 56(2):508– 529, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582321.

Segal:2010:GBF

[Seg10] Howard Segal. Getting Bucky Fuller right: Loretta Lorance, becoming Bucky Fuller. Technology and Culture, 51(4):1006–1009, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403277.

Segal:2014:GAH

[Seg14] Howard P. Segal. The Great American Hall of Wonders, Smithsonian American Art Museum, Washington, D.C., 15 July 2011– 8 January 2012. Technology and Culture, 55(1):227–236, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538917.

Senchyne:2017:RMP

[Sen17] Jonathan Senchyne. Rags make paper, paper makes money: Material texts and metaphors of capital. Technology and Culture, 58(2):545–555, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/662988.

Seth:2016:AEE

[Set16] Suman Seth. Asia, Europe, and the Emergence of Modern Science: Knowledge Crossing Boundaries ed. by Arun Bala (review). Technology and Culture, 57(2):471–472, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/619065.

Simon:2016:ET

[SGB16] Josep Simon and Antonio García-Belmar. Education and textbooks. Technology and Culture, 57(4):940–950, October 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/637932.

Smith-Howard:2015:HFH

[SH15] Kendra Smith-Howard. Home Fires: How Americans Kept Warm in the 19th Century by Sean Patrick Adams (review). Technology and Culture, 56(4):980–981, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601482.

Smith-Howard:2017:HAA

[SH17] Kendra Smith-Howard. Healing animals in an Antibiotic Age: Veterinary drugs and the professionalism crisis, 1945–1970. Technology and Culture, 58(3):722–748, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/669466.

Smith-Howard:2019:BRH

[SH19a] Kendra Smith-Howard. Before the Refrigerator: How We Used to Get Ice by Jonathan Rees (review). Technology and Culture, 60 (4):1107–1108, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741409.

Smith-Howard:2019:SCU

[SH19b] Kendra Smith-Howard. Slaughterhouse: Chicago’s Union Stock Yard and the World It Made by Dominic A. Pacyga (review). Technology and Culture, 60(1):321–322, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/719957.

Shallat:2010:RHP

[Sha10] Todd Shallat. Rivers in History: Perspectives on Waterways in Europe and North America (review). Technology and Culture, 51 (2):502–504, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381292.

Shah:2017:BBI

[Sha17] Esha Shah. BITS of Belonging: Information Technology, Water, and Neoliberal Governance in India by Simanti Dasgupta (review). Technology and Culture, 58(3):898–900, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669492.

Shell:2014:FAH

[She14a] Hanna Rose Shell. Films in the archive: Hollywood in Detroit. Technology and Culture, 55(3):711–715, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/551610.

Shell:2014:IE

[She14b] Hanna Rose Shell. Introduction from the Editor. Technology and Culture, 55(2):477–478, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/546873.

Shields:2012:TST

[Shi12] Mark A. Shields. Technology and Social Theory (review). Technology and Culture, 53(4):918–920, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/490734.

Shindell:2017:AAR

[Shi17] Matthew Shindell. After Apollo? Richard Nixon and the American Space Program by John M. Logsdon (review). Technology and Culture, 58(2):598–600, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/663002.

Shulman:2011:RHU

[Shu11] Peter A. Shulman. The richest hills unearthed: Timothy J. LeCain, Mass Destruction. Technology and Culture, 52(4):801–804, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454538.

Shulman:2015:ECL

[Shu15] Peter A. Shulman. Energy Capitals: Local Impact, Global Influence ed. by Joseph A. Pratt etal. (review). Technology and Culture, 56 (4):1000–1001, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601494.

Shulman:2019:GWS

[Shu19] Chaim Shulman. The groundbreaking water supply systems of central and Eastern European cities, 1300–1580. Technology and Culture, 60(3):726–769, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/731441.

Siddiqi:2010:CTN

[Sid10] Asif A. Siddiqi. Competing technologies, national(ist) narratives, and universal claims: Toward a global history of space exploration. Technology and Culture, 51(2):425–443, April 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/381261.

Siddiqi:2012:LLZ

[Sid12] Asif A. Siddiqi. Lenin’s Laureate: Zhores Alferov’s Life in Communist Science (review). Technology and Culture, 53(2):511–512, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476840.

Sideri:2018:PPL

[Sid18] Katerina Sideri. Patent Politics: Life Forms, Markets and the Public Interest in the United States and Europe by Shobita Parthasarathy (review). Technology and Culture, 59(3):817–818, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
703165.

Siebeneichner:2014:EHS

[Sie14] Tilmann Siebeneichner. Exploring the heavens: Space technology and religious implications. Technology and Culture, 55(4):988–994, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
562843.

Simon:2010:MCM

[Sim10] Roger D. Simon. The machine in context: Merritt Roe Smith’s Harpers Ferry Armory and the New Technology: The Challenge of Change. Technology and Culture, 51(4):1010–1017, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/403278.

Simons:2017:CCL

[Sim17a] Greg Simons. Carnage and Connectivity: Landmarks in the Decline of Conventional Military Power by David Betz (review). Technology and Culture, 58(1):290–291, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648277.

Simons:2017:BPH

[Sim17b] Peter Simons. Barnstorming the Prairies: How Aerial Vision Shaped the Midwest by Jason Weems (review). Technology and Culture, 58(2):591–593, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/662998.

Simonsen:2018:IAA

[Sim18] Dorthe Gert Simonsen. Island in the air — powered aircraft and the early formation of British airspace. Technology and Culture, 59(3):590–619, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/703141.

Singh:2012:MTH

[Sin12] J. P. Singh. Mobile Telecommunications in a High-Speed World: Industry Structure, Strategic Behaviour and Socio-Economic Status (review). Technology and Culture, 53(1):248–250, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466808.

Singh:2015:GIP

[Sin15] J. P. Singh. The Great Indian Phone Book: How the Cheap Cell Phone Changes Business, Politics, and Daily Life by Assa Doron and Robin Jeffrey (review). Technology and Culture, 56(2):563–565, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
582340.

Sismondo:2012:TZI

[Sis12] Sergio Sismondo. Trading Zones and Interactional Expertise: Creating New Kinds of Collaboration (review). Technology and Culture, 53(3):696–697, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481913.

Sjoblom:2011:DGL

[Sjö11] Gustav Sjöblom. Driving Germany: The Landscape of the German Autobahn, 1930–1970 (review). Technology and Culture, 52 (1):215–216, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416435.

Sa:2016:TCU

[SK16] Creso M. Sá and Andrew Kretz. Technology commercialization as university mission: Early historical developments at the University of Toronto. Technology and Culture, 57(1):119–143, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/611802.

Slade:2010:TNW

[Sla10] Joseph W. Slade. The Technology of the Novel: Writing and Narrative in British Fiction (review). Technology and Culture, 51(4): 1048–1050, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403299.

Slade:2011:WMC

[Sla11a] Joseph W. Slade. We Are the Machine: The Computer, the Internet, and Information in Contemporary German Literature (review). Technology and Culture, 52(2):427–429, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/431391.

Slaton:2011:UCS

[Sla11b] Amy E. Slaton. The Uses of Context: Steen Hyldgaard Christensen, Bernard Delahousse, and Martin Meganck, eds., engineering in context. Technology and Culture, 52(4):805–808, October 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/454537.

Slayton:2011:DRL

[Sla11c] Rebecca Slayton. From death rays to light sabers: Making laser weapons surgically precise. Technology and Culture, 52(1):45–74, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
416406.

Slayton:2012:CDE

[Sla12a] Rebecca Slayton. Continental Defense in the Eisenhower Era: Nuclear Antiaircraft Arms and the Cold War (review). Technology and Culture, 53(4):947–948, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/490751.

Slayton:2012:JET

[Sla12b] Rebecca Slayton. Judging Edward Teller: A Closer Look at One of the Most Influential Scientists of the Twentieth Century (review). Technology and Culture, 53(2):509–510, April 2012. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/476841.

Slayton:2014:OMI

[Sla14] Rebecca Slayton. The Optical Munitions Industry in Great Britain, 1888–1923 by Stephen C Sambrook (review). Technology and Culture, 55(2):497–499, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/546883.

Slaton:2015:CCM

[Sla15] Amy E. Slaton. Concrete and Culture: A Material History by Adrian Forty (review). Technology and Culture, 56(1):279–281, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
572132.

Slayton:2016:SWE

[Sla16] Rebecca Slayton. Scientists at War: The Ethics of Cold War Weapons Research by Sarah Bridger (review). Technology and Culture, 57(3):696–698, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/630951.

Slayton:2018:FMS

[Sla18] Rebecca Slayton. FastLane: Managing Science in the Internet World by Thomas J. Misa, Jeffrey R. Yost (review). Technology and Culture, 59(1):190–192, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/692185.

Slayton:2019:TWI

[Sla19] Rebecca Slayton. A Telephone for the World: Iridium, Motorola, and the Making of a Global Age by Martin Collins (review). Technology and Culture, 60(3):928–929, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/731450.

Shell:2017:DI

[SM17] Hanna Rose Shell and Gregg Mitman. The documentary impulse. Technology and Culture, 58(3):846–855, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669471.

Smith:2010:BFB

[Smi10a] Jenny Leigh Smith. Brezhnev’s Folly: The Building of BAM and Late Soviet Socialism (review). Technology and Culture, 51(2):513– 514, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381284.

Smith:2010:WDL

[Smi10b] Merritt Roe Smith. W. David Lewis, 1931–2007. Technology and Culture, 51(1):290–292, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375771.

Smith:2010:EAD

[Smi10c] Pamela H. Smith. “Extremely apt at doing things by all sorts of tools”. Technology and Culture, 51(2):470–473, April 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/381259.

Smith:2011:CCC

[Smi11a] John K. Smith. The catalyst club: Contentious chemistry and confounding innovation. Technology and Culture, 52(2):310–334, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431363.

Smith:2011:AIA

[Smi11b] Pamela H. Smith. The Arts of Industry in the Age of the Enlightenment (review). Technology and Culture, 52(3):624–626, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447790.

Smith:2013:CSE

[Smi13] Jenny Leigh Smith. Into the Cosmos: Space Exploration and Soviet Culture ed. by James T. Andrews and Asif A. Siddiqi (review). Technology and Culture, 54(2):419–420, April 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/506094.

Smith:2014:WPJ

[Smi14a] Crosbie Smith. Witnessing power: John Elder and the making of the marine compound engine, 1850–1858. Technology and Culture, 55(1):76–106, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538909.

Smith:2014:NNJ

[Smi14b] Jenny Leigh Smith. New Natures: Joining Environmental History with Science and Technology Studies Edited by Dolly Jørgensen, Finn Arne Jørgensen, and Sara B. Pritchard (review). Technology and Culture, 55(3):734–735, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/551615.

Smith:2017:STT

[Smi17] Mark C. Smith. Scientism and Technocracy in the Twentieth Century: The Legacy of Scientific Management by Richard G. Olson (review). Technology and Culture, 58(4):1087–1088, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/679833.

Snyder:2017:NTE

[Sny17] Hunter Snyder. The Nonhuman Turn ed. by Richard Grusin (review). Technology and Culture, 58(1):292–293, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648278.

SHT:2012:CMN

[Soc12] Society for the History of Technology. The Cleveland Meeting, 3–5 November 2011. Technology and Culture, 53(1):161–176, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466774.

Soluri:2011:SSA

[Sol11] John Soluri. In the Shadow of Slavery: Africa’s Botanical Legacy in the Atlantic World (review). Technology and Culture, 52(4): 819–821, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454543.

Sondergaard:2012:TMD

[Søn12] Morten Karnøe Søndergaard. Technology museums in Denmark. Technology and Culture, 53(3):678–687, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/481910.

Songster:2017:YRP

[Son17] E. Elena Songster. The Yellow River: The Problem of Water in Modern China by David A Pietz (review). Technology and Culture, 58(2):621–623, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/663015.

Soppelsa:2011:ITM

[Sop11] Peter Soppelsa. Intersections: Technology, mobility, and geography. Technology and Culture, 52(4):673–677, October 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/454529.

Soppelsa:2014:GBC

[Sop14] Peter Soppelsa. Greening Berlin: The Co-Production of Science, Politics, and Urban Nature by Jens Lachmund (review). Technology and Culture, 55(1):276–278, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/538936.

Soppelsa:2015:RAT

[Sop15] Peter Soppelsa. Reviving Animals as Technology: Ann Norton Greene, Horses at Work (review). Technology and Culture, 56 (1):252–260, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572121.

Sovacool:2016:CHC

[Sov16] Benjamin K. Sovacool. Coal and the historical complexity of energy security: Peter A. Shulman, Coal and Empire. Technology and Culture, 57(2):460–463, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/619060.

Stalter-Pace:2017:VCC

[SP17] Sunny Stalter-Pace. Vehicles: Cars, Canoes, and Other Metaphors of Moral Imagination ed. by David Lipset and Richard Handler (review). Technology and Culture, 58(1):286–288, January 2017. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/648275.

Speich:2011:DLR

[Spe11] Daniel Speich. Digitale Landschaften (review). Technology and Culture, 52(2):424–425, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/431389.

Spielmann:2015:VRH

[Spi15] Yvonne Spielmann. Video Revolutions. On the History of a Medium by Michael Z. Newman (review). Technology and Culture, 56(4): 1010–1011, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601500.

Sprenger:2019:EHI

[Spr19] Florian Sprenger. The Environment: A History of the Idea by Paul Warde, Libby Robin, and Sverker Sörlin (review). Technology and Culture, 60(3):900–902, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/731465.

Spude:2011:GGH

[Spu11] Robert Spude. Going for Gold: The History of Newmont Mining Corporation (review). Technology and Culture, 52(3):642–644, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447800.

Spude:2013:SRC

[Spu13] Robert L. Spude. Santa Rita del Cobre: A Copper Mining Community in New Mexico by Christopher J. Huggard and Terrence M. Humble (review). Technology and Culture, 54(2):410–412, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/506089.

Soppelsa:2019:OF

[SR19] Peter Soppelsa and Amy S. Rodgers. Origins of the flyswatter. Technology and Culture, 60(3):886–895, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731447.

Staudenmaier:2010:I

[SS10] John M. Staudenmaier, S.J. and Bruce Seely. In this issue. Technology and Culture, 51(1):a–e, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375770.

Soppelsa:2013:SDB

[SS13] Peter Soppelsa and Blair Stein. Santos-dumont’s blimp passes the Eiffel tower. Technology and Culture, 54(4):942–946, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/530517.

Suominen:2018:SMH

[SSH18] Jaakko Suominen, Antti Silvast, and Tuomas Harviainen. Smelling machine history: Olfactory experiences of information technology. Technology and Culture, 59(2):313–337, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/697234.

Stapleton:2010:WDB

[Sta10a] Darwin H. Stapleton. War and Disease: Biomedical Research on Malaria in the Twentieth Century (review). Technology and Culture, 51(2):535–536, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381274.

Staudenmaier:2010:PB

[Sta10b] John M. Staudenmaier, S.J. Passing the baton. Technology and Culture, 51(4):781–785, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/403264.

StaudenmaierSJ:2012:WLW

[Sta12] John M. Staudenmaier, S.J. Welcoming in a lovely world. Technology and Culture, 53(2):259–271, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/476805.

Stapleton:2014:EHH

[Sta14a] Darwin Stapleton. Environment, Health and History ed. by Virginia Berridge and Martin Gorsky (review). Technology and Culture, 55(2):487–488, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/546877.

Staudenmaier:2014:NTH

[Sta14b] John Staudenmaier, S.J. Notes for Tom Hughes’s funeral homily: University of Virginia Chapel, 6 April 2014. Technology and Culture, 55(4):970–972, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/562841.

Stafford:2017:ESO

[Sta17] Barbara Maria Stafford. Engaging Smithsonian Objects through Science, History, and the Arts ed. by Mary Jo Arnoldi (review). Technology and Culture, 58(2):618–620, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/663013.

Stearns:2010:BIR

[Ste10a] Peter N. Stearns. The British Industrial Revolution in Global Perspective (review). Technology and Culture, 51(4):1018–1019, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403281.

Stephens:2010:RBG

[Ste10b] Carlene E. Stephens. Ruth Belville: The Greenwich Time Lady (review). Technology and Culture, 51(1):248–249, January 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/375792.

Sterling:2011:CRC

[Ste11] Christopher H. Sterling. Communication Revolution: Critical Junctures and the Future of Media (review). Technology and Culture, 52(2):422–424, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/431388.

Sterling:2014:ARC

[Ste14] Christopher H. Sterling. American Radio in China: International Encounters with Technology and Communications, 1919–41 by Michael Krysko (review). Technology and Culture, 55(2):504– 505, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546887.

Stevens:2015:NBO

[Ste15] Hallam Stevens. Networking biology: The origins of sequencesharing practices in genomics. Technology and Culture, 56(4): 839–867, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601470.

Stein:2016:MMC

[Ste16a] Jesse Adams Stein. Masculinity and material culture in technological transitions: From letterpress to offset lithography, 1960s– 1980s. Technology and Culture, 57(1):24–53, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611799.

Stevens:2016:BDL

[Ste16b] Hallam Stevens. Big Data, Little Data, No Data: Scholarship in the Networked World by Christine L. Borgman (review). Technology and Culture, 57(3):706–708, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/630957.

Stein:2019:AAW

[Ste19] Blair Stein. Aerial Aftermaths: Wartime from Above by Caren Kaplan (review). Technology and Culture, 60(3):915–917, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/731457.

Stine:2011:NNY

[Sti11] Jeffrey K. Stine. The Nature of New York: An Environmental History of the Empire State (review). Technology and Culture, 52 (4):838–840, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/454556.

Stine:2018:HGW

[Sti18] Jeffrey K. Stine. Heavy Ground: William Mulholland and the St. Francis Dam Disaster by Norris Hundley Jr. and Donald C. Jackson (review). Technology and Culture, 59(3):798–799, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/703154.

Stone:2012:PB

[Sto12] John Stone. The point of the bayonet. Technology and Culture, 53 (4):885–908, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490731.

Stone:2013:SSS

[Sto13a] Glenn Davis Stone. Seeds, Science, and Struggle: The Global Politics of Transgenic Crops by Abby Kinchy (review). Technology and Culture, 54(4):990–991, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/530534.

Storey:2013:PST

[Sto13b] William Kelleher Storey. The Postcolonial Science and Technology Studies Reader ed. by Sandra Harding (review). Technology and Culture, 54(3):654–655, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/520287.

Storey:2019:EGV

[Sto19] William K. Storey. Empire of Guns: The Violent Making of the Industrial Revolution by Priya Satia (review). Technology and Culture, 60(2):632–633, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/726959.

Stradling:2010:ASE

[Str10] David Stradling. The Age of Smoke: Environmental Policy in Germany and the United States, 1880–1970 (review). Technology and Culture, 51(1):254–255, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/375791.

Strasser:2011:UGN

[Str11] Susan Strasser. The Untilled Garden: Natural History and the Spirit of Conservation in America, 1740–1840 (review). Technology and Culture, 52(1):190–191, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416420.

Strate:2012:NER

[Str12] Lance Strate. The Net Effect: Romanticism, Capitalism, and the Internet (review). Technology and Culture, 53(2):520–521, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476842.

Strauss:2013:ILT

[Str13] Jonathan Strauss. Improper Life: Technology and Biopolitics from Heidegger to Agamben by Timothy C. Campbell (review). Technology and Culture, 54(3):697–699, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/520311.

SullivanDeOliveira:2019:IEB

[Sul19] Patrick Luiz Sullivan De Oliveira. The Imagined Empire: Balloon Enlightenments in Revolutionary Europe by Mi Gyung Kim (review). Technology and Culture, 60(2):628–630, April 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/726951.

Summitt:2017:UWS

[Sum17] April R. Summitt. Unruly Waters: A Social and Environmental History of the Brazos River by Kenna Lang Archer (review). Technology and Culture, 58(4):1097–1099, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/679839.

Suominen:2012:DCR

[Suo12] Jaakko Suominen. Digital Cultures (review). Technology and Culture, 53(2):521–523, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/476843.

Sutter:2012:ENW

[Sut12] Paul S. Sutter. Engineering Nature: Water, Development, and the Global Spread of American Environmental Expertise (review). Technology and Culture, 53(1):221–223, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/466793.

Sutter:2019:SLS

[Sut19] Paul S. Sutter. Sustaining Lake Superior: An Extraordinary Lake in a Changing World by Nancy Langston (review). Technology and Culture, 60(4):1116–1117, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/741404.

Soderholm:2019:PBE

[SV19] Kristina Söderholm and Roine Viklund. Policy and business efforts for the reduced impact of mining on nature: When historical studies have something to offer policy makers. Technology and Culture, 60(1):192–218, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/719950.

Shell:2015:THD

[SW15] Hanna Rose Shell and Alex Wellerstein. Technologist-historian: Data visualization meets the archive. Technology and Culture, 56 (1):204–208, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572116.

Swann:2010:ECV

[Swa10] Paul Swann. Edited Clean Version: Technology and the Culture of Control (review). Technology and Culture, 51(4):1053–1054, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
403302.

Swade:2013:CAT

[Swa13a] Doron Swade. Codebreaker: Alan Turing’s Life and Legacy: 21 June 2012–21 October 2013, Science Museum, London. Technology and Culture, 54(4):957–962, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/530519.

Swanson:2013:MTB

[Swa13b] Drew A. Swanson. Making Tobacco Bright: Creating an American Commodity, 1617–1937 by Barbara Hahn (review). Technology and Culture, 54(2):402–403, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/506084.

Swanson:2015:SSH

[Swa15] Drew Swanson. The Story of N: A Social History of the Nitrogen Cycle and the Challenge of Sustainability by Hugh S. Gorman (review). Technology and Culture, 56(4):970–971, October 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/601476.

Swanson:2018:FHF

[Swa18] Drew Swanson. Farmers Helping Farmers: The Rise of the Farm and Home Bureaus, 1914–1935 by Nancy K. Berlage (review). Technology and Culture, 59(4):978–979, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711778.

Swanson:2019:TME

[Swa19] Drew Swanson. Taming the Muskingum by Emory L. Kemp (review). Technology and Culture, 60(1):322–324, January 2019. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/719958.

Szajewski:2015:DMA

[Sza15] Michael Szajewski. Digital Memory and the Archive by Wolfgang Ernst (review). Technology and Culture, 56(2):560–562, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/582338.

Tucker:2018:DAR

[TAF^{+}18] Jennifer Tucker, Glenn Adamson, Jonathan S. Ferguson, Josh Garrett-Davis, Erik Goldstein, Ashley Hlebinsky, David D. Miller, and Susanne Slavick. Display of arms: a roundtable discussion about the public exhibition of firearms and their history. Technology and Culture, 59(3):719–769, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/703145.

Tarr:2014:TLE

[Tar14] Joel A. Tarr. Toxic legacy: The environmental impact of the manufactured gas industry in the United States. Technology and Culture, 55(1):107–147, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/538910.

Tarr:2015:RAC

[Tar15] Joel Tarr. Rethinking the American City: An International Dialogue ed. by Miles Orvell and Klaus Benesch (review). Technology and Culture, 56(4):1005–1006, October 2015. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601497.

Taylor:2010:WPS

[Tay10] Peter Taylor. Well-pressed stories. Technology and Culture, 51 (2):477–479, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/381277.

Taylor:2014:AEW

[Tay14] Ben Taylor. “Accidents and emergencies: Welfare and safety in Europe and North America, c.1750–2000”: 9–11 September 2013, Oxford, UK. Technology and Culture, 55(1):220–222, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538915.

Taylor:2017:PVB

[Tay17] Grant D. Taylor. Peripheral Vision: Bell Labs, the S-C 4020, and the Origins of Computer Art by Zabet Patterson (review). Technology and Culture, 58(1):296–298, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648281.

Teich:2011:CAR

[Tei11] Albert H. Teich. The Commodification of Academic Research: Science and the Modern University (review). Technology and Culture, 52(3):650–652, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/447804.

Telotte:2011:ASF

[Tel11] J. P. Telotte. American Science Fiction Film and Television (review). Technology and Culture, 52(3):658–660, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447808.

Telotte:2013:AWI

[Tel13] J. P. Telotte. Astounding Wonder: Imagining Science and Science Fiction in Interwar America by John Cheng (review). Technology and Culture, 54(2):417–419, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/506093.

Tenner:2014:FDU

[Ten14] Edward Tenner. To Forgive Design: Understanding Failure by Henry Petroski (review). Technology and Culture, 55(1):249–251, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
538921.

Tenner:2015:DET

[Ten15] Edward Tenner. The Design of Everyday Things by Donald Norman (review). Technology and Culture, 56(3):785–787, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593093.

Theodore:2010:MTC

[The10a] David Theodore. Medicine and Technology in Canada, 1900–1950 (review). Technology and Culture, 51(2):537–538, April 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/381279.

Theriot:2010:OBG

[The10b] Jason P. Theriot. Oil: A Beginner’s Guide (review). Technology and Culture, 51(1):256–257, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375790.

Thomas:2010:TDM

[Tho10a] Matt Thomas. TV by Design: Modern Art and the Rise of Network Television (review). Technology and Culture, 51(2):527–528, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/381276.

Thompson:2010:MLU

[Tho10b] Steven L. Thompson. Motorcycling and Leisure: Understanding the Recreational PTW Rider (review). Technology and Culture, 51(3):758–759, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390859.

Thomas:2015:VHG

[Tho15] William Thomas. The Visioneers: How a Group of Elite Scientists Pursued Space Colonies, Nanotechnologies, and a Limitless Future by W. Patrick McCray (review). Technology and Culture, 56(1):289–291, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/572138.

Thomas:2016:HGC

[Tho16] William Thomas. How to glean culture from an evolving Internet: Richard Rogers, Digital Methods. Technology and Culture, 57 (1):238–241, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/611809.

Tilley:2019:GSD

[Til19] Helen Tilley. A great (scientific) divergence: Synergies and fault lines in global histories of science. Technology and Culture, 60 (2):583–593, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/726940.

Timmermann:2017:LKM

[Tim17] Carsten Timmermann. The Lock and Key of Medicine: Monoclonal Antibodies and the Transformation of Healthcare by Lara V. Marks (review). Technology and Culture, 58(4):1105–1106, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/679843.

Timmermans:2018:GSE

[Tim18] Stefan Timmermans. Good Science: The Ethical Choreography of Stem Cell Research by Charis Thompson (review). Technology and Culture, 59(2):503–504, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/697258.

Tingvall:2015:HTS

[Tin15] Claes Tingvall. The history of traffic safety: Describing 100 years. Technology and Culture, 56(2):489–492, April 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/582318.

Tinn:2018:MCC

[Tin18] Honghong Tinn. Modeling computers and computer models: Manufacturing economic-planning projects in Cold War Taiwan, 1959– 1968. Technology and Culture, 59(4S):S66–S99, October 2018. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/712114.

Tinn:2019:PIH

[Tin19] Honghong Tinn. Programmed Inequality: How Britain Discarded Women Technologists and Lost Its Edge in Computing by Marie Hicks (review). Technology and Culture, 60(4):1121–1122, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/741401.

Tkaczyk:2019:ATA

[Tka19] Viktoria Tkaczyk. Archival traces of applied research: Language planning and psychotechnics in interwar Germany. Technology and Culture, 60(2S):S64–S95, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/727299.

Tobbell:2013:HLE

[Tob13] Dominique A. Tobbell. Hormones for Life: Endocrinology, the Pharmaceutical Industry, and the Dream of a Remedy for Sterility, 1930–1970 by Christer Nordlund (review). Technology and Culture, 54(2):428–430, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/506099.

Todd:2010:IRR

[Tod10] Edmund N. Todd. The Industrial Revolution (review). Technology and Culture, 51(1):251–252, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/375789.

Todd:2011:NSS

[Tod11] Edmund Todd. Vom “nationalen” zum “sozialistischen” Selbst: Zur Erfahrungs-geschichte deutscher Chemiker und Ingenieure im 20. Jahrhundert (review). Technology and Culture, 52(4):847–848, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
454561.

Tomory:2011:BFG

[Tom11] Leslie Tomory. Building the first gas network, 1812–1820. Technology and Culture, 52(1):75–102, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416407.

Tomory:2014:CWC

[Tom14] Leslie Tomory. City Water, City Life: Water and the Infrastructure of Ideas in Urbanizing Philadelphia, Boston, and Chicago by Carl Smith (review). Technology and Culture, 55(3):747–749, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/551623.

Tomory:2015:LWS

[Tom15] Leslie Tomory. London’s water supply before 1800 and the roots of the networked city. Technology and Culture, 56(3):704–737, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/593069.

Tomory:2017:LWM

[Tom17] Leslie Tomory. London: Water and the Making of the Modern City by John Broich (review). Technology and Culture, 58(1):264–265, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
648262.

Thebaud-Sorger:2012:SIB

[TS12] Marie Thébaud-Sorger. The Sublime Invention: Ballooning in Europe, 1783–1820 (review). Technology and Culture, 53(2):487–488, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476844.

Tsukahara:2012:TSS

[Tsu12] Togo Tsukahara. Techno-science and the state in the Japanese Empire: Yang Daqing, Technology of Empire, Janis Mimura, Planning for Empire. Technology and Culture, 53(2):461–464, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476813.

Tucker:2012:SMF

[Tuc12] Jennifer Tucker. The Sympathetic Medium: Feminine Channeling, The Occult, and Communication Technologies, 1859–1919 (review). Technology and Culture, 53(1):213–214, January 2012. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/466788.

Tucher:2016:CHA

[Tuc16] Andie Tucher. The Camera as Historian: Amateur Photographers and Historical Imagination, 1885–1918 by Elizabeth Edwards (review). Technology and Culture, 57(1):254–255, January 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/611817.

Turner:2010:EHM

[Tur10] Fred Turner. Earthrise: How Man First Saw the Earth (review). Technology and Culture, 51(1):272–274, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/375788.

Turner:2018:MTP

[Tur18] Fred Turner. Millenarian tinkering: The Puritan roots of the maker movement. Technology and Culture, 59(4S):S160–S182, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/712117.

Tyabji:2012:NND

[Tya12] Nasir Tyabji. Negotiating nonalignment: Dilemmas attendant on initiating pharmaceutical production in India. Technology and Culture, 53(1):37–60, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/466779.

Tyabji:2013:MMH

[Tya13] Nasir Tyabji. Market Menagerie: Health and Development in Late Industrial States by Smita Srinivas (review). Technology and Culture, 54(3):699–701, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/520312.

Usenyuk:2016:PDU

[UHW16] Svetlana Usenyuk, Sampsa Hyysalo, and Jack Whalen. Proximal design: Users as designers of mobility in the Russian North. Technology and Culture, 57(4):866–908, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/637928.

Unger:2014:MJT

[Ung14] David S. Unger. MythBusters and the joy of technology. Technology and Culture, 55(2):482–485, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/546875.

Unger:2015:BST

[Ung15] Richard W. Unger. Brewing Science: Technology and Print, 1700– 1880 by James Sumner (review). Technology and Culture, 56 (2):535–536, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582324.

Uyeki:2011:LE

[Uye11] Eugene Uyeki. Letter to the Editor. Technology and Culture, 52(2):438–439, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431398.

Uziel:2019:FMH

[Uzi19] Daniel Uziel. Flying Man: Hugo Junkers and the Dream of Aviation by Richard Byers (review). Technology and Culture, 60(2):637–638, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
726947.

Vaillant:2013:SST

[Vai13] Derek W. Vaillant. At the speed of sound: Techno-aesthetic paradigms in U.S.–French international broadcasting, 1925–1942. Technology and Culture, 54(4):888–921, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530515.

Vaillant:2019:CMM

[Vai19] Derek W. Vaillant. Colonialists in the machine: Maria Paula Diogo and Dirk van Laak, Europeans Globalizing. Technology and Culture, 60(2):625–627, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/726944.

Vallor:2010:TVE

[Val10] Shannon Vallor. Technology and Values: Essential Readings (review). Technology and Culture, 51(2):538–540, April 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/381275.

VanCour:2014:SSC

[Van14] Shawn VanCour. Selling Sounds: The Commercial Revolution in American Music by David Suisman (review). Technology and Culture, 55(3):758–759, July 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/551629.

VanLieshout:2016:DDG

[Van16] Carry Van Lieshout. Droughts and dragons: Geography, rainfall, and Eighteenth-Century London’s water systems. Technology and Culture, 57(4):780–805, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/637925.

Vandermeer:2018:BDC

[Van18] Philip Vandermeer. Bulldozer: Demotion and Clearance of the Postwar Landscape by Francesca Russelo Ammon (review). Technology and Culture, 59(3):801–803, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/703156.

Vardi:2011:CCP

[Var11] Itai Vardi. Crash: Cinema and the Politics of Speed and Stasis (review). Technology and Culture, 52(3):656–657, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447807.

Varel:2018:DDL

[Var18] David A. Varel. Database of Dreams: The Lost Quest to Catalog Humanity by Rebecca Lemov (review). Technology and Culture, 59(2):498–499, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/697255.

vanderHorst:2017:FFG

[vdH17] Dan van der Horst. Fact and Fiction in Global Energy Policy: 15 Contentious Questions by Benjamin K. Sovacool, Marilyn A. Brown, Scott V. Valentine (review). Technology and Culture, 58 (3):891–893, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/669488.

vanderHoogte:2013:JCJ

[vdHP13] Arjo Roersch van der Hoogte and Toine Pieters. From Javanese coca to Java coca: An exemplary product of Dutch colonial agroindustrialism, 1880–1920. Technology and Culture, 54(1):90–116, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500798.

vanderStraeten:2019:BIM

[vdS19] Jonas van der Straeten. Borderlands of industrial modernity: Explorations into the history of technology in central Asia, 1850– 2000. Technology and Culture, 60(3):659–687, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731439.

vanderVleuten:2012:SWW

[vdV12] Erik van der Vleuten. Science for Welfare and Warfare: Technology and State Initiative in Cold War Sweden (review). Technology and Culture, 53(1):231–232, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/466798.

vanderVleuten:2014:CCS

[vdV14] Erik van der Vleuten. Cosmopolitan Commons: Sharing Risks and Resources across Borders ed. by Nil Disco and Eda Kranakis (review). Technology and Culture, 55(2):518–520, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546895.

Veenendaal:2010:TAC

[Vee10] Augustus J. Veenendaal, Jr. Tracks across Continents, Paths through History: The Economic Dynamics of Standardization in Railway Gauge (review). Technology and Culture, 51(2):511–513, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381271.

Venugopal:2012:SWI

[Ven12] Prasad Venugopal. Science, War and Imperialism: India in the Second World War (review). Technology and Culture, 53(1): 227–229, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/466796.

Venkat:2019:VMS

[Ven19] Bharat Jayram Venkat. A vital mediation: The sanatorium, before and after antibiotics. Technology and Culture, 60(4):979–1003, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
741380.

Vermij:2012:BRG

[Ver12] Rienk Vermij. Between Raphael and Galileo: Mutio Oddi and the Mathematical Culture of Late Renaissance Italy (review). Technology and Culture, 53(3):700–701, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/481915.

Vertesi:2013:EST

[Ver13] Janet Vertesi. Down to Earth: Satellite Technologies, Industries, and Cultures ed. by Lisa Parks and James Schwoch (review). Technology and Culture, 54(4):1008–1010, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/530544.

Vermij:2015:WMK

[Ver15] Rienk Vermij. Ways of Making and Knowing: The Material Culture of Empirical Knowledge ed. by Pamela H. Smith, Amy R. W. Meyers, and Harold J. Cook (review). Technology and Culture, 56 (4):977–978, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/601480.

Vest:2019:VRA

[Ves19] J. Martin Vest. Vinyl Records and Analog Culture in the Digital Age: Pressing Matters by Paul E. Winters (review). Technology and Culture, 60(2):654–655, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/726962.

Vetter:2019:WNT

[Vet19] Jeremy Vetter. Wired into Nature: The Telegraph and the North American Frontier by James Schwoch (review). Technology and Culture, 60(1):333–335, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719964.

Vining:2012:GHI

[Vin12] Margaret Vining. Grace Hopper and the Invention of the Information Age (review). Technology and Culture, 53(2):516–517, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/476845.

Vinsel:2015:DTP

[Vin15] Lee Jared Vinsel. Designing to the test: Performance standards and technological change in the U.S. automobile after 1966. Technology and Culture, 56(4):868–894, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/601471.

Vitz:2017:TSP

[Vit17] Matthew Vitz. Technology and the Search for Progress in Modern Mexico by Edward Beatty (review). Technology and Culture, 58(2):584–586, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/662994.

vanMeer:2012:TIF

[vM12a] Elisabeth van Meer. Text, image, flusser. Technology and Culture, 53(4):913–917, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490733.

vanMeer:2012:TPW

[vM12b] Elisabeth van Meer. The transatlantic pursuit of a world engineering federation: For the profession, the nation, and international peace, 1918–48. Technology and Culture, 53(1):120–145, January 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/466776.

vanMeer:2016:WRE

[vM16] Elisabeth van Meer. Writing the Rules for Europe: Experts, Cartels, and International Organizations by Wolfram Kaiser and Johan Schot (review). Technology and Culture, 57(2):483–485, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/619072.

vanMunster:2015:PNE

[vMS15] Rens van Munster and Casper Sylvest. Pro-nuclear environmentalism: Should we learn to stop worrying and love nuclear energy?

Technology and Culture, 56(4):789–811, October 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/601468.

Volti:2010:FPG

[Vol10] Rudi Volti. Ferdinand Porsche — Genesis of Genius: Road, Racing, and Aviation Innovation, 1900 to 1933 (review). Technology and Culture, 51(4):1040–1041, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/403294.

Volti:2011:YRF

[Vol11] Rudi Volti. The Yugo: The Rise and Fall of the Worst Car in History (review). Technology and Culture, 52(4):854–855, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454564.

Volti:2017:AHB

[Vol17] Rudi Volti. An Alternative History of Bicycles and Motorcycles: Two-Wheeled Transportation and Material Culture by Steven E. Alford, Suzanne Ferriss (review). Technology and Culture, 58(3):874– 876, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669478.

Voskuhl:2010:DPS

[Vos10] Adelheid Voskuhl. The Disordered Police State: German Cameralism as Science and Practice (review). Technology and Culture, 51 (4):1020–1021, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/403282.

Voskuhl:2016:EPT

[Vos16] Adelheid Voskuhl. Engineering philosophy: Theories of technology, German idealism, and social order in high-industrial Germany. Technology and Culture, 57(4):721–752, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/637923.

vanRiper:2012:FAA

[vR12] A. Bowdoin van Riper. Flying Across America: The Airline Passenger Experience (review). Technology and Culture, 53(2):498– 500, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
476833.

Wachelder:2017:PPI

[Wac17] Joseph Wachelder. From Playgrounds to PlayStation: The Interaction of Technology and Play by Carroll Pursell (review). Technology and Culture, 58(1):303–305, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/648285.

Wade:2016:PPB

[Wad16] Peter Wade. Postgenomics: Perspectives on Biology after the Genome ed. by Sarah S. Richardson and Hallam Stevens (review). Technology and Culture, 57(3):701–702, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/630954.

Wakefield:2011:NHW

[Wak11] Andre Wakefield. The New Holy Wars: Economic Religion vs. Environmental Religion in Contemporary America (review). Technology and Culture, 52(3):671–672, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/447815.

Walsh:2010:APU

[Wal10] John C. Walsh. The Archive of Place: Unearthing the Pasts of the Chilcotin Plateau (review). Technology and Culture, 51(3):772– 773, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
390867.

Walton:2018:MNA

[Wal18] Steven A. Walton. The Medieval New: Ambivalence in an Age of Innovation by Patricia Claire Ingham (review). Technology and Culture, 59(4):965–966, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/711770.

Wang:2010:DSP

[Wan10] Hsien-Chun Wang. Discovering steam power in China, 1840s– 1860s. Technology and Culture, 51(1):31–54, January 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/375784.

Wang:2013:CKT

[Wan13] Hsien-Chung Wang. Cultures of Knowledge: Technology in Chinese History ed. by Dagmar Schäfer (review). Technology and Culture, 54(3):657–659, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/520289.

Ward:2010:MMM

[War10] James A. Ward. Maxwell Motor and the Making of the Chrysler Corporation (review). Technology and Culture, 51(2):518–519, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381263.

Warner:2014:PMM

[War14] Deborah Warner. Pure and Modern Milk: An Environmental History since 1900 by Kendra Smith-Howard (review). Technology and Culture, 55(4):1007–1009, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/562851.

Weart:2016:TON

[Wea16] Spencer Weart. The Theater of Operations: National Security Affect from the Cold War to the War on Terror by Joseph Masco (review). Technology and Culture, 57(2):494–495, April 2016. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/619078.

Weart:2018:CPA

[Wea18] Spencer Weart. Cultures of Prediction in Atmospheric and Climate Science: Epistemic and Cultural Shifts in Computer-based Modelling and Simulation ed. by Matthias Heymann, Gabriele Gramelsberger, and Martin Mahony (review). Technology and Culture, 59(2):505–506, April 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/697259.

Weber:2015:SES

[Web15] Donald Weber. Safety or efficiency?: Strategies and conflicting interests in Belgian road-safety policy, 1920–1940. Technology and Culture, 56(2):394–419, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582314.

Weitekamp:2013:SOW

[Wei13] Margaret A. Weitekamp. Space Oddities: Women and Outer Space in Popular Film and Culture, 1960–2000 by Marie Lathers (review). Technology and Culture, 54(1):220–221, January 2013. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/500822.

Weitekamp:2015:MMS

[Wei15] Margaret Weitekamp. Marketing the Moon: The Selling of the Apollo Lunar Program by David Meerman Scott, Richard Jurek (review). Technology and Culture, 56(3):777–778, July 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/593088.

Wells:2010:RDC

[Wel10] Christopher Wells. Republic of Drivers: A Cultural History of Automobility in America (review). Technology and Culture, 51 (1):263–265, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375787.

Wellock:2012:EUB

[Wel12] Thomas R. Wellock. Engineering uncertainty and bureaucratic crisis at the Atomic Energy Commission, 1964–1973. Technology and Culture, 53(4):846–884, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/490730.

Welskopp:2015:RFH

[Wel15] Thomas Welskopp. Ribbon of Fire: How Europe Adopted and Developed US Strip Mill Technology (1920–2000) ed. by Jonathan Aylen and Ruggero Ranieri (review). Technology and Culture, 56(2):556–558, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/582336.

Wellock:2017:FMQ

[Wel17] Thomas R. Wellock. A figure of merit: Quantifying the probability of a nuclear reactor accident. Technology and Culture, 58(3):678– 721, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669465.

Wenzlhuemer:2017:TCM

[Wen17] Roland Wenzlhuemer. The telegraph and the control of material movements: A micro-study about the detachment of communication from transport. Technology and Culture, 58(3):625–649, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/669463.

Werrett:2012:TST

[Wer12] Simon Werrett. Technology on the spot: The trials of the Congreve rocket in India in the Early Nineteenth Century. Technology and Culture, 53(3):598–624, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/481907.

Westermann:2010:MMC

[Wes10a] Andrea Westermann. Manufacturing mass consumption in the GDR: Eli Rubin, synthetic socialism. Technology and Culture, 51 (1):232–234, January 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/375772.

Westfall:2010:BNF

[Wes10b] Catherine Westfall. The Boundaries of the New Frontier: Rhetoric and Communication at Fermi National Accelerator Laboratory (review). Technology and Culture, 51(1):281–282, January 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/375786.

Westfall:2011:VRS

[Wes11] Catherine Westfall. Velvet Revolution at the Synchrotron: Biology, Physics, and Change in Science (review). Technology and Culture, 52(2):431–432, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/431393.

Westfall:2015:HZC

[Wes15] Catherine Westfall. Hybrid Zone: Computers and Science at Argonne National Laboratory, 1946–1992 by Charles N. Yood (review). Technology and Culture, 56(1):295–296, January 2015. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/572141.

Westfall:2016:TVR

[Wes16] Catherine Westfall. Tunnel Visions: The Rise and Fall of the Superconducting Super Collider by Michael Riordan, Lillian Hoddeson, Adrienne W. Kolb (review). Technology and Culture, 57(4): 1036–1037, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/637956.

Wetmore:2015:DAE

[Wet15] Jameson M. Wetmore. Delegating to the automobile: Experimenting with automotive restraints in the 1970s. Technology and Culture, 56(2):440–463, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582316.

Wiesner-Hanks:2011:LTW

[WH11] Merry Wiesner-Hanks. Louder than Words: Ways of Seeing Women Workers in Eighteenth-Century France (review). Technology and Culture, 52(1):193–195, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/416422.

Whitney:2010:LLD

[Whi10] Kristoffer Whitney. Living lawns, dying Waters: The suburban boom, nitrogenous fertilizers, and the nonpoint source pollution dilemma. Technology and Culture, 51(3):652–674, July 2010. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/390840.

White:2011:PPP

[Whi11a] Sam White. Plows, Plagues, and Petroleum: How Humans Took Control of Climate (review). Technology and Culture, 52(1): 182–183, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416415.

Whitehouse:2011:GEB

[Whi11b] David Whitehouse. Glass Exchange between Europe and China, 1550–1800: Diplomatic, Mercantile, and Technological Interactions (review). Technology and Culture, 52(2):393–394, April 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/431371.

Whitney:2012:CNP

[Whi12] Kristoffer Whitney. Case(s) in (non)Point. Technology and Culture, 53(4):909–912, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/490732.

White:2013:OAE

[Whi13a] Sam White. The Ottoman Age of Exploration by Giancarlo Casale (review). Technology and Culture, 54(4):970–971, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/530522.

Whitney:2013:CCS

[Whi13b] Kristoffer Whitney. Crabgrass Crucible: Suburban Nature and the Rise of Environmentalism in Twentieth-Century America by Christopher C. Sellers (review). Technology and Culture, 54(2):426– 428, April 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
506098.

White:2014:GOP

[Whi14] Sam White. Guano and the Opening of the Pacific World by Gregory Cushman (review). Technology and Culture, 55(2):494–495, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
546881.

Wijfjes:2012:TDM

[Wij12] Huub Wijfjes. Television as Digital Media (review). Technology and Culture, 53(4):956–957, October 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/490756.

Wijfjes:2014:SCS

[Wij14] Huub Wijfjes. Spellbinding and crooning: Sound amplification, radio, and political rhetoric in international comparative perspective, 1900–1945. Technology and Culture, 55(1):148–185, January 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/538911.

Williams:2010:WSA

[Wil10] James C. Williams. The Way of the Ship: America’s Maritime History Reenvisioned, 1600–2000 (review). Technology and Culture, 51(3):741–743, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390849.

Williams:2011:TSS

[Wil11] James Williams. Thirty-seventh symposium of the International Committee for the History of Technology: “Reusing the Industrial Past,” Tampere, Finland, 10–15 August 2010. Technology and Culture, 52(2):373–379, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/431366.

Williams:2013:FTA

[Wil13] Amrys O. Williams. FOOD: Transforming the American table, 1950–2000: National Museum of American History, Smithsonian Institution, Washington, D.C. Technology and Culture, 54(4): 947–956, October 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/530518.

Williams:2014:RNH

[Wil14a] Amrys O. Williams. Refrigeration Nation: A History of Ice, Appliances, and Enterprise in America by Jonathan Rees (review). Technology and Culture, 55(4):1006–1007, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/562850.

Williams:2014:PPE

[Wil14b] James C. Williams. Past and Present Energy Societies: How Energy Connects Politics, Technologies, and Cultures ed. by Nina Möllers and Karin Zachmann (review). Technology and Culture, 55 (4):1004–1005, October 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/562849.

Williams:2014:OTA

[Wil14c] Rosalind Williams. Our Technological Age, from the inside out. Technology and Culture, 55(2):461–476, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/546872.

Wilkenfeld:2017:PPC

[Wil17a] George Wilkenfeld. Palazzos of Power: Central Stations of the Philadelphia Electric Company 1900–1930 by Aaron V. Wunsch, Joseph E. B. Elliott (review). Technology and Culture, 58(3):872– 874, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
669477.

Williams:2017:OSM

[Wil17b] Amrys O. Williams. Organic Struggle: The Movement for Sustainable Agriculture in the United States by Brian K. Obach (review). Technology and Culture, 58(3):895–896, July 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/669490.

Williford:2017:SPR

[Wil17c] Daniel Williford. Seismic politics: Risk and reconstruction after the 1960 earthquake in Agadir, Morocco. Technology and Culture, 58 (4):982–1016, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/679821.

Williford:2019:DMS

[Wil19] Daniel Williford. Dreamscapes of Modernity: Sociotechnical Imaginaries and the Fabrication of Power ed. by Sheila Jasanoff and Sang-Hyun Kim (review). Technology and Culture, 60(4): 1110–1112, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/741402.

Wisnioski:2011:BMD

[Wis11] Matt Wisnioski. Becoming MIT: Moments of Decision (review). Technology and Culture, 52(4):859–860, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/454567.

Wisnioski:2017:TCW

[Wis17] Matthew Wisnioski. Technology Choices: Why Occupations Differ in Their Embrace of New Technology by Diane E. Bailey and Paul M. Leonardi (review). Technology and Culture, 58(2):609–611, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
663008.

Williams:2011:OMM

[WM11] Rosalind H. Williams and Charles W. McFarland. Once more to the mountain. Technology and Culture, 52(3):441–458, July 2011. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/447777.

Wolfe:2010:RHV

[Wol10] Audra J. Wolfe. Radio’s Hidden Voice: The Origins of Public Broadcasting in the United States (review). Technology and Culture, 51(4):1050–1051, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/403300.

Wolters:2011:NCW

[Wol11a] Timothy S. Wolters. Network-Centric Warfare: How Navies Learned to Fight Smarter through Three World Wars (review). Technology and Culture, 52(3):665–667, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/447812.

Wolters:2011:RFM

[Wol11b] Timothy S. Wolters. Recapitalizing the fleet: A material analysis of Late-Nineteenth-Century U.S. Naval power. Technology and Culture, 52(1):103–126, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/416408.

Wolters:2014:RIS

[Wol14] Timothy Wolters. Re-inventing the Ship: Science, Technology, and the Maritime World, 1800–1918 ed. by Don Leggett and Richard Dunn (review). Technology and Culture, 55(2):496–497, April 2014. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/546882.

Wolfson:2015:EEJ

[Wol15] Richard Wolfson. Energy & Ethics: Justice and the Global Energy Challenge by Benjamin K. Sovacool (review). Technology and Culture, 56(2):569–570, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/582343.

Wood:2019:DWM

[Woo19] Nathaniel D. Wood. The Devil’s Wheels: Men and Motorcycling in the Weimar Republic by Sasha Disko (review). Technology and Culture, 60(2):635–637, April 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/726949.

Wormbs:2010:RSA

[Wor10] Nina Wormbs. Remembering the Space Age: Proceedings of the 50th Anniversary Conference (review). Technology and Culture, 51(3):750–751, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/390854.

Wormbs:2016:PPP

[Wor16a] Nina Wormbs. In Pursuit of a Promise: Perspectives on the Political Process to Establish the European Spallation Source (ESS) in Lund, Sweden ed. by Olof Hallonsten (review). Technology and Culture, 57(2):497–498, April 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/619080.

Worth:2016:WPG

[Wor16b] Aaron Worth. World Projects: Global Information before World War I by Markus Krajewski (review). Technology and Culture, 57(3):673–675, July 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/630938.

Wosk:2010:M

[Wos10a] Julie Wosk. Metropolis. Technology and Culture, 51(2):403–408, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
381287.

Wosk:2010:UFM

[Wos10b] Julie Wosk. Update on the film Metropolis. Technology and Culture, 51(4):1061–1062, October 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/403279.

Wosk:2017:MXM

[Wos17] Julie Wosk. Manus x Machina: Fashion in an Age of Technology by Metropolitan Museum of Art (review). Technology and Culture, 58(1):257–263, January 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/648261.

Wright:2012:HDL

[Wri12a] Gwendolyn Wright. Hotel Dreams: Luxury, Technology, and Urban Ambition in America, 1829–1929 (review). Technology and Culture, 53(2):490–491, April 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/476846.

Wright:2012:THV

[Wri12b] Terence Wright. Technologies of History: Visual Media and the Eccentricity of the Past (review). Technology and Culture, 53(3):717– 718, July 2012. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
481925.

Wrobel:2010:HCW

[Wro10] David Wrobel. How Cities Won the West: Four Centuries of Urban Change in Western North America (review). Technology and Culture, 51(2):508–509, April 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/381281.

Wu:2011:ERB

[Wu11] Xiujie Wu. Eating Rice from Bamboo Roots: The Social History of a Community of Handicraft Papermakers in Rural Sichuan, 1920– 2000 (review). Technology and Culture, 52(4):845–847, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/454558.

Wu:2019:NRN

[Wu19] Shellen Wu. Natural Resources and the New Frontier: Constructing Modern China’s Borderlands by Judd C. Kinzley (review). Technology and Culture, 60(4):1114–1115, October 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/741403.

Yang:2013:E

[Yan13] Daqing Yang. To the Editor. Technology and Culture, 54(1):228, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
500826.

Yang:2019:NSC

[Yan19] Daqing Yang. Navigating Semi-Colonialism: Shipping, Sovereignty, and Nation-Building in China, 1860–1937 by Anne Reinhardt (review). Technology and Culture, 60(3):910–911, July 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/731463.

Yates:2016:CCF

[Yat16] Alexia Yates. Le crédit à la consommation en France, 1947–1965: De la stigmatisation à la réglementation by Sabine Effosse (review). Technology and Culture, 57(1):273–274, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/611827.

Yost:2016:MCA

[Yos16] Jeffrey R. Yost. Managing Cyber Attacks in International Law, Business, and Relations: In Search of Cyber Peace by Scott J. Shackelford (review). Technology and Culture, 57(1):280–281, January 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
611831.

Zakariya:2015:EC

[Zak15] Nasser Zakariya. Exhibiting Cosmos. Technology and Culture, 56(3):738–744, July 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/593070.

Zanetti:2013:CMM

[Zan13] François Zanetti. Curing with machines: Medical electricity in Eighteenth-Century Paris. Technology and Culture, 54(3):503–530, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
520280.

Zarecor:2013:SCA

[Zar13] Kimberly Elman Zarecor. The Socialist Car: Automobility in the Eastern Bloc ed. by Lewis H. Siegelbaum (review). Technology and Culture, 54(1):213–214, January 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/500818.

Zarecor:2017:HPF

[Zar17] Kimberly Elman Zarecor. At Home in Postwar France: Modern Mass Housing and the Right to Comfort by Nicole C. Rudolph (review). Technology and Culture, 58(4):1099–1101, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/679840.

Zeller:2011:DSC

[Zel11] Thomas Zeller. Driving Spaces: A Cultural-Historical Geography of England’s M1 Motorway (review). Technology and Culture, 52 (1):217–218, January 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/416436.

Ziauddin:2017:SUS

[Zia17] Silvia Berger Ziauddin. Superpower underground: Switzerland’s rise to global bunker expertise in the Atomic Age. Technology and Culture, 58(4):921–954, October 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/679819.

Ziauddin:2018:SKS

[Zia18] Silvia Berger Ziauddin. Sharing Knowledge, Shaping Europe: U.S. Technological Collaboration and Nonproliferation by John Krige (review). Technology and Culture, 59(1):188–190, January 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/692184.

Zimmerman:2010:AEB

[Zim10] David Zimmerman. Arms, Economics, and British Strategy: From Dreadnoughts to Hydrogen Bombs (review). Technology and Culture, 51(3):738–739, July 2010. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/390847.

Zimring:2011:CEL

[Zim11] Carl A. Zimring. The complex environmental legacy of the automobile shredder. Technology and Culture, 52(3):523–547, July 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/447780.

Zimring:2013:HDW

[Zim13] Carl A. Zimring. Histories of the Dustheap: Waste, Material Cultures, Social Justice ed. by Stephanie Foote and Elizabeth Mazzolini (review). Technology and Culture, 54(3):703–704, July 2013. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/520314.

Zimmer:2017:BPS

[Zim17] Zac Zimmer. Bitcoin and Potosí Silver: Historical perspectives on cryptocurrency. Technology and Culture, 58(2):307–334, April 2017. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/662979.

Zizzamia:2019:UCC

[Ziz19] Daniel Zizzamia. Undermined in Coal Country: On the Measures in a Working Land by Bill Conlogue (review). Technology and Culture, 60(1):344–346, January 2019. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/719971.

Zagorski-Thomas:2016:SIS

[ZT16] Simon Zagorski-Thomas. The Sound of Innovation: Stanford and the Computer Music Revolution by Andrew J. Nelson (review). Technology and Culture, 57(4):1039–1041, October 2016. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/637958.

Zagorski-Thomas:2018:DMM

[ZT18] Simon Zagorski-Thomas. Directions in music by miles Davis: Using the ecological approach to perception and embodied cognition to analyze the creative use of recording technology in bitches brew. Technology and Culture, 59(4):850–874, October 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/711764.

Zylstra:2011:RAA

[Zyl11a] Geoff D. Zylstra. Railroads in the African American Experience: A Photographic Journey (review). Technology and Culture, 52(2):399– 400, April 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
431375.

Zylstra:2011:WFT

[Zyl11b] Geoff D. Zylstra. Whiteness, freedom, and technology: The racial struggle over Philadelphia’s streetcars, 1859–1867. Technology and Culture, 52(4):678–702, October 2011. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/454528.

Zylstra:2015:EPS

[Zyl15] Geoff D. Zylstra. Engineering Philadelphia: The Sellers Family and the Industrial Metropolis by Domenic Vitiello (review). Technology and Culture, 56(1):272–274, January 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/572128.